### v1

# Circuits and Backdoors: Five Shades of the SETH

Abstract
The SETH is a hypothesis of fundamental importance to (fine-grained) parameterized complexity theory and many important tight lower bounds are based on it. This situation is somewhat problematic, because the validity of the SETH is not universally believed and because in some senses the SETH seems to be “too strong” a hypothesis for the considered lower bounds. Motivated by this, we consider a number of reasonable weakenings of the SETH that render it more plausible, with sources ranging from circuit complexity, to backdoors for SAT-solving, to graph width parameters, to weighted satisfiability problems. Despite the diversity of the different formulations, we are able to uncover several non-obvious connections using tools from classical complexity theory. This leads us to a hierarchy of five main equivalence classes of hypotheses, with some of the highlights being the following:
  1. 1.
    We show that beating brute force search for SAT parameterized by a modulator to a graph of bounded pathwidth, or bounded treewidth, or logarithmic tree-depth, is actually the same question, despite the structural differences between these parameters. In fact, we show that beating brute force for any of these parameters is possible if and only if one can beat brute force for circuits of depth $\varepsilon n$ (a weaker version of the so-called NC-SETH). We also show that all these questions are equivalent to the existence of an $n^{(1-\varepsilon)k}$ algorithm for the weighted satisfiability problem associated with W[SAT]. A main tool we use here is Barrington’s theorem.
  2. 2.
    On the other hand, modulators to logarithmic pathwidth seem to belong to a different class. We show that beating brute force for this parameter is equivalent to beating brute force for SAT when given a strong 2-SAT backdoor. To prove this we rely on ideas from the classical Immerman–Szelepcsényi theorem.
  3. 3.
    Identifying an even harder class of problems, we show that beating brute force for SAT given a strong Horn backdoor is equivalent to the SETH for general circuits (without depth restrictions) and to beating brute force for the weighted satisfiability problem of the class W[P].
  4. 4.
    Going in the other direction, we revisit recent work of Esmer, Focke, Marx, and Rzazewski (ICALP ’24), and uncover some nuances in their lower bounds. In particular, we show that beating brute force for SAT parameterized by a modulator to constant tree-depth is equivalent to falsifying the standard form of the SETH. This allows us to show that their lower bound for $q$-Coloring (for parameter $(\sigma,\delta)$-hub) is actually not just an implication, but a reformulation of the SETH. On the other hand, we show that their lower bound for Max-Cut for the same parameter is equivalent to the maximization version of the SETH, which at the moment appears to be a weaker hypothesis.
For each of these classes we give at least one (non-SAT) problem which is equivalent to the class as a characteristic example application. Our work is a step towards a more nuanced meta-complexity theory where we advance from proving one-way implications of the SETH, to establishing a classification of interesting questions into equivalence classes. A somewhat unusual aspect is that our classification cuts across traditional time complexity classes, as we are often able to establish equivalences between lower bounds for FPT problems (of the form $(c-\varepsilon)^{k}n^{O(1)}$) and non-FPT problems (of the form $n^{(1-\varepsilon)k}$).
keywords:
SETH, SAT Backdoors, Parameterized Complexity, Fine-grained Complexity
category:
[…]
1 Introduction
1.1 Background and Motivation
The Strong Exponential Time Hypothesis (SETH [25]) states that the satisfiability of CNF formulas cannot be decided significantly faster than by brute force, that is, in time $(2-\varepsilon)^{n}$, assuming the formulas have large enough arity [^1] To ease presentation, in this section we use informal statements such as this one to describe all complexity hypotheses. In case of doubt, we refer the reader to Section A.1 where precise definitions are given.. Since its introduction, the SETH has played a major role as an indispensable tool in the development of fine-grained complexity theory. In this paper we focus on its influence on fine-grained parameterized complexity.
One of the most important techniques in the parameterized algorithms toolbox is the use of dynamic programming algorithms over tree decompositions, and the question of whether such algorithms can be improved is of central importance. It was therefore a very exciting development for the field when Lokshtanov, Marx, and Saurabh [33] pioneered a line of work which allowed to provide tight lower bounds for a large class of algorithms of this type. For instance, one of their results was that the standard algorithm for 3-Coloring, which has complexity $3^{\textrm{tw}}n^{O(1)}$, cannot be improved to $(3-\varepsilon)^{\textrm{tw}}n^{O(1)}$, where tw denotes the treewidth. Similar tight results were given in [33] and many subsequent works for many other problems, and also for other parameters (for example clique-width, cutwidth) [6, 8, 10, 13, 14, 16, 17, 18, 21, 22, 23, 27, 28, 29, 32, 35, 38]. An important common point of all these lower bounds is that they are generally shown to hold assuming the SETH.
The SETH is too strong (i).
The line of work summarized above can be fairly considered quite successful: for many of the most famous problems, not only do we know how to solve them in time $c^{k}n^{O(1)}$ for some constant $c$ and parameter $k$, but we know the precise value of the optimal constant $c$, assuming the SETH. This is great, if one believes the SETH, but it is important to keep in mind that the SETH is not universally believed and indeed, there seem to be serious reasons to doubt its validity, much more than that of more standard assumptions (we refer the reader to the argumentation given by Williams [43] who, somewhat provocatively, assigns the SETH a rather low probability of being true; see also [44]). Although SETH-based lower bounds are still useful (as pointed out in [33] they show that before improving, for instance, the 3-Coloring algorithm we mentioned, we should first try to disprove the SETH), it is somewhat problematic if SETH-based arguments are the only evidence of hardness we have for the most important problems.
This line of reasoning has in fact already been well-developed in the (non-parameterized) fine-grained complexity community. In particular, there are several works which reprove (and improve) lower bounds based on the SETH using a more plausible assumption, such as the NC-SETH, which states the same lower bound but for circuits of poly-logarithmic depth, rather than CNF formulas. We refer the reader to the work of Abboud, Hansen, Vassilevska Williams, and Williams [3] for a more extensive discussion of why the NC-SETH is a preferable hypothesis, or equivalently why falsifying the NC-SETH would have far more surprising consequences than disproving the SETH. The advantages of the NC-SETH over the SETH are taken into account in several other works in this area, for example [1, 9, 41]. In contrast, in parameterized complexity the SETH is often referred to as an “established” hypothesis, though in fairness this is mostly done in comparison to (even more) wide open hypotheses, such as the Set Cover Conjecture [15]. Given the importance of such lower bounds for the wider field, it would be useful if some more convincing evidence could be provided.
The SETH is too strong (ii).
The discussion above explains that the SETH is not an ideal starting point for proving lower bounds, as its validity is still subject to some doubt. There is, however, another, more counter-intuitive reason to argue that the SETH is not ideal, and that is that the SETH is “too good” in comparison with the lower bounds we are interested in. This is a line of reasoning recently explored by Esmer, Focke, Marx, and Rzazewski [15], who observed that for many of the problems considered in [33] (including $q$-Coloring, Independent Set, and Max-Cut), a more refined reduction shows that the SETH implies a much stronger lower bound: Not only are current algorithms optimal for treewidth, but they are still optimal for a novel, extremely restricted parameter, which they call $(\sigma,\delta)$-hub. A $(\sigma,\delta)$-hub, (for $\sigma,\delta$ to be considered absolute constants) is a set of vertices (a modulator) whose removal leaves a collection of bounded-size (size $\sigma$) components with bounded degree (degree $\delta$) in the modulator. The class of graphs with a $(\sigma,\delta)$-hub of size $k$ is a tiny fraction of the class of graphs of treewidth $k$, so this is a strong indication that the lower bounds for the more interesting parameters should be achievable by a much weaker assumption.
Equivalences are better than implications.
[…]
Coming back to our setting, the above mean that beyond the practical reason we are not satisfied by a lower bound implied by the SETH (we are not sufficiently confident that the SETH is true), there is a philosophical reason not to be happy, which is that it is natural to believe that most of the problems considered are in fact considerably harder than the SETH. It is therefore a natural complexity-theoretic goal to attempt to classify them in a hierarchy of increasing hardness, not just for the added confidence we would gain in the intractability of questions that end up being higher in the hierarchy, but for the extra insight that we are likely to gain from this classification.
1.2 Summary of Results
One way to explain the main direction of our results is to come back to the recent work of Esmer, Focke, Marx, and Rzazewski [15] we mentioned. We share their initial motivation, which is the observation that the SETH does not just imply tight treewidth-based lower bounds, such as those of [33], but in fact much more. However, where their next step is to say “Let us then try to extract the strongest implication of the SETH”, we say “Let us then explore how we can replace the SETH by a more solid assumption, while still extracting interesting lower bounds”. We will therefore study numerous variants of the SETH, coming from several different areas, including: (i) circuit versions, similar to the NC-SETH (ii) backdoor versions, such as “If we have a SAT formula with a 2-SAT backdoor, can we avoid brute-forcing the backdoor”? (a backdoor is a set of variables such that setting them simplifies the formula) (iii) modulator to graph width versions, such as “If we have a SAT formula with a modulator to a formula of constant treewidth, can we avoid brute-forcing the modulator?” (iv) weighted versions where the question is whether a satisfying assignment of weight $k$ exists (v) maximization versions where the objective is to maximize the number of satisfied clauses.
At this point it is worth pausing and trying to reassure a worried reader. It can be said that one of the weak points of fine-grained complexity theory is the proliferation of reasonable-sounding complexity hypotheses, with no obvious connection to each other. It is natural to think that our decision to study a relatively large and diverse collection of variants of the SETH can lead to nothing more than a chaotic mess, worsening this problem. Thankfully, the main take-home message of our results is that, even though we consider quite a few variations, we are able to identify only five equivalence classes (one of which contains the SETH itself), which align neatly with classes from classical complexity theory. In a sense, the conclusion we draw is optimistic as our results indicate that there is a great deal of complexity-theoretic structure waiting to be discovered here.
Let us then summarize our results, referring also to Figure 1. We identify the following equivalence classes of complexity hypotheses:
  1. 1.
    Coming back to the NC-SETH, we see that even though the hypothesis is stated for circuits of poly-logarithmic depth in [3], in practice it is often understood to apply to circuits of depth $n^{o(1)}$ (e.g. [9]). We take this reasoning to its logical conclusion and formulate this hypothesis for circuits of depth $\Omega(n)$, or more precisely for depth at most $\varepsilon n$ (we call this LD-C-SETH, for Linear-Depth Circuits). We discover that this hypothesis is equivalent to stating that brute force is unavoidable for CNF formulas supplied with a modulator to one of three standard widths: (i) constant pathwidth (ii) constant treewidth or (iii) logarithmic tree-depth. We also show that disproving the LD-C-SETH is equivalent to avoiding brute force for the W[SAT]-complete problem of finding a weight-$k$ satisfying assignment to a given (general, non-CNF) formula. This class corresponds to the middle box of Figure 1.
  2. 2.
    Going further, we consider the SETH for circuits without any restriction (Circuit-SETH), an assumption which seems much more plausible than the SETH. We show that avoiding brute force in this case can be done if and only if we can avoid brute force in one of two problems (i) deciding satisfiability for a CNF supplied with a Horn backdoor (ii) deciding weight-$k$ satisfiability for an arbitrary circuit, a problem complete for W[P]. This is the top-most class of Figure 1.
  3. 3.
    Coming back to the LD-C-SETH, we consider the graph width parameters which are closest the ones already considered. For modulators to logarithmic pathwidth (which would be the logical next step after logarithmic tree-depth) we show that avoiding brute-force enumeration is possible if and only if we can avoid it for 2-SAT backdoors.
  4. 4.
    In the other direction, the logical next step below modulators to constant pathwidth is modulators to constant tree-depth. Here we show that the hypothesis that brute-force algorithms are optimal is in fact equivalent to the standard SETH. The parameter $(\sigma,\delta)$-hub introduced by [15] is therefore included here, as it is a restriction of bounded tree-depth modulators. This is the bottom class of Figure 1. In a similar spirit we show that for Max-SAT, the standard version of the hypothesis (for the number of variables) and the version for the size of a $(\sigma,\delta)$-hub are equivalent (this is the second class from the bottom in Figure 1).
[…]
Figure 1: Summary of results. We identify five main equivalence classes, with arrows indicating implication, that is, pointing to weaker hypotheses. From bottom to top the five classes are problems equivalent to the SETH; the SETH for Max-SAT; the SETH for circuits of linear depth; the SETH for 2-SAT backdoors; and the SETH for arbitrary circuits. The figure also shows some connections with related hypotheses. Full definitions are given in Section A.1.
We have therefore identified five natural equivalence classes of hypotheses. In order to make sense of these results, let us attempt to draw an informal parallel with classical complexity theory. All the hypotheses we consider concern problems where the obvious algorithm is a brute-force “guess-and-check”; what differs is the difficulty of the verification procedure. If the verification procedure is P-complete (for example Horn-SAT or evaluating an arbitrary circuit) we end up in the top-most class of Figure 1. Similarly, if it is NL-complete (equivalent to a reachability problem) we end up in the class right below, while the class of the LD-C-SETH seems to correspond to NC1. Finally, the class immediately above the SETH seems somewhat similar to TC0, since in Max-SAT what we need to do if given an assignment is count the number of satisfied clauses and compare this number with a target – this corresponds to the functioning of a threshold gate. Armed with this intuition, we can identify some further complexity questions as characteristic examples of each class.
Applications and Implications
The initial motivation we gave for this investigation was that we would like to increase our confidence in SETH-based lower bounds for problems parameterized by treewidth. One implication of our results is that we have managed to do so for many problems. Indeed, we show that improving upon the trivial algorithm for $q$-Coloring for parameters modulator size to constant pathwidth, constant treewidth, or logarithmic tree-depth is equivalent to the LD-C-SETH. This implies that improving upon the standard algorithm for pathwidth or treewidth is at least as hard as falsifying the LD-C-SETH (and hence the NC-SETH). As explained, this seems significantly more solid than a result based on the SETH. Furthermore, recent work by Lampis [31] showed that a number of fine-grained lower bounds for problems parameterized by pathwidth (including Independent Set, Dominating Set, Set Cover and others) are equivalent to the pw-SETH, the hypothesis that SAT cannot be solved in time $(2-\varepsilon)^{\textrm{pw}}|\phi|^{O(1)}$, where pw is the input formula’s primal pathwidth. Since the pw-SETH is implied by the same hypothesis for parameter modulator to logarithmic pathwidth, all the hypotheses we consider here except those of the top-most class of Figure 1 also imply the lower bounds of [31]. We have thus succeeded in giving further evidence that known lower bounds for pathwidth and treewidth hold. For instance, even though it was previously known that no $(3-\varepsilon)^{\textrm{pw}}n^{O(1)}$ algorithm is possible for Dominating Set under the SETH, our results (together with those of [31]) establish this under the NC-SETH.
Beyond this, we give a characteristic example problem for each class, based on the intuition linking the classes of Figure 1 with traditional complexity classes. In particular, we show that the Circuit-SETH is equivalent to beating brute-force for the problem of deleting $k$ vertices from a graph, so that it becomes $r$-degenerate, for any fixed $r\geq 2$; this problem is known to be P-complete. Similarly, we give a reachability-type problem ($k$-Neighborhood-Cut) that is equivalent to the 2SatB-SETH, which is natural, since reachability-type problems are usually NL-complete.
These applications are of particular interest for the two bottom classes in light of the results of [15]. In this work it was shown that if the SETH is true, it is impossible to avoid brute force for (among others) $q$-Coloring and Max-Cut, parameterized by $(\sigma,\delta)$-hub size. This perhaps creates the impression that we should have the same amount of confidence in the two lower bounds. Contrary to this, we show that their lower bound for $q$-Coloring is in fact equivalent to the SETH, while their lower bound for Max-Cut is equivalent to the MaxSat-SETH. If it were possible to show that these lower bounds are equivalent to each other, this would also establish that the Max-3-SAT-SETH (the restriction of MaxSat-SETH to CNFs of arity $3$) implies the SETH. As pointed out in [15], this is a major open problem. Our results therefore uncover some (perhaps unforeseen) nuance in the lower bounds of [15].
Meta-complexity.
Is is worth stressing that, even though we present Figure 1 as a collection of “complexity classes”, and indeed we draw a parallel with classical such classes, the results we present are in a sense orthogonal to normal notions of time complexity. What we mean by this is that, as can be seen in Figure 1, we often show equivalence for problems which (are widely believed to) belong in distinct classes, namely, problems which are FPT on the one hand and W[SAT]- or W[P]-complete on the other. In other words, the equivalence classes of Figure 1 do not indicate that all the problems in each class have the same time complexity; but rather that if we could improve the running time of the best algorithm for one, we could improve it for the others, albeit obtaining different types of running times for each problem. We remark that a similar situation holds for the pw-SETH [31].
1.3 Overview of Techniques
Let us now sketch the main ideas used to obtain the results mentioned above. In all cases, a large part of the intuition needed is obtained by the parallels we drew between these classes and traditional complexity classes. More precisely:
  1. 1.
    For the hypotheses which turn out to be equivalent to the LD-C-SETH, one of the main technical tools we rely on is Barrington’s theorem [5], which allows us to transform a circuit on $n$ inputs and depth $\epsilon n$ into a branching program of constant width and length exponential in $\epsilon n$. This program can then be simulated by a formula with constant pathwidth and a modulator representing the inputs. For the converse direction we need to show that an algorithm disproving the LD-C-SETH can be used to decide the satisfiability of a CNF $\phi$ with a modulator $M$ of size $m$, such that deleting $M$ produces a formula of tree-depth at most $O(\log|\phi|)$. For this, we need to eliminate all the variables of the formula except those of $M$. We show how to do this in a way that is exponential in the tree-depth (and hence polynomial in the size of the formula).
  2. 2.
[…]
  3. 3.
    For the equivalences related to Circuit-SETH we rely on the intuition that computing the value of a circuit and deciding the satisfiability of a Horn formula are both P-complete problems.
  4. 4.
    For the equivalences given in the two bottom classes, the hard part is showing that (Max-)SAT parameterized by a modulator (to constant tree-depth, or a $(\sigma,\delta)$-hub) can be reduced to the same problem parameterized by the number of variables, that is, that we can eliminate the extra variables. For SAT this is done by an exhaustive application of the standard resolution rule, starting from the leaves of the rooted trees, while for Max-SAT parameterized by $(\sigma,\delta)$-hub, this is done by enumerating all assignments to each (bounded-size) component and replacing it with new constraints to its (bounded-size) neighborhood.
1.4 Previous Work
As mentioned, SETH-based lower bounds are an area of very active research in parameterized complexity. Beyond the results following [33], which establish running time lower bounds of the form $(c-\varepsilon)^{k}n^{O(1)}$, for some parameter $k$ and fixed constant $c$, the SETH has also been used to obtain lower bounds for non-FPT problems. The most famous example is perhaps the result of Patrascu and Williams that $k$-Dominating Set cannot be solved in time $n^{k-\varepsilon}$, under the SETH [40].
The question of whether the SETH is the “right” assumption has appeared before in the literature. In general fine-grained complexity this manifests itself via works which attempt to replace the SETH with more plausible assumptions, such as the ones mentioned above, or which attempt to prove the same lower bound based on several hypotheses which have no known relation [2, 4]. Furthermore, the question of which problems are exactly equivalent to the SETH has already been posed, though few such problems are known [11]. We also note that the connection between the NC-SETH and Barrington’s theorem was already used in [9].
In the area of parameterized complexity, investigations on whether the SETH is the right complexity assumption have mostly focused on whether, assuming the SETH we can obtain the same lower bounds but for more restricted cases than for treewidth ([15, 27]). However, a strongly related recent work [31] investigated the so-called pw-SETH, that is, the hypothesis that SAT cannot be solved in time $(2-\varepsilon)^{\textrm{pw}}|\phi|^{O(1)}$. Even though this is closely related to our motivation, the pw-SETH turns out to cover a completely different algorithmic paradigm, as it essentially speaks about the optimality of dynamic programming algorithms, while the hypotheses we consider here refer to the optimality of branching algorithms under different regimes. As a result, most of the hypotheses we consider here imply the pw-SETH, with the exception of Circuit-SETH, for which this remains an interesting open problem.
1.5 Discussion and Directions for Further Work
We believe that the results of this paper, together with recent related works ([15, 31]) point the way to a new and interesting research direction investigating the interplay between fine-grained complexity theory and parameterized input structure. The envisioned goal here is a theory where lower bound statements form a comprehensible hierarchy where we can eventually pinpoint the reasons that a lower bound is believed to hold; for example we have seen that the LD-C-SETH is essentially equivalent to improving upon a generic “guess-and-check” algorithm with an NC1 verifier. For which other algorithmic paradigms can we identify such natural complete problems? Equally importantly, how do different paradigms relate to each other? In particular, the pw-SETH informally captures the difficulty of improving upon a generic DP algorithm over a linear structure, and the Circuit-SETH the difficulty of improving upon a generic “guess-and-check” algorithm with a polynomial-time verifier. Discovering a connection between these two hypothesis would allow us to better understand which paradigm is more powerful; conversely the two may be incomparable, and it would be interesting to see if this can be expressed via the corresponding hypotheses.
For some more concrete research directions, we mention that in the interest of keeping the presentation manageable we have only given a limited number of examples for each equivalence class. It would of course be interesting to discover some more such problems. In particular, one obvious question is to consider the remaining lower bounds given in [15], especially their lower bound for Independent Set parameterized by $(\sigma,\delta)$-hub. Breaking this bound is shown to be at least as hard as the SETH in [15], while conversely the MaxSat-SETH can be shown to be at least as hard, via the equivalence we show with the $(\sigma,\delta)H$-MaxSat-SETH. In other words, Independent Set probably belongs in one of the two bottom classes of Figure 1 for parameter $(\sigma,\delta)$-hub, so it would be interesting to determine which one. Another question to consider is finding a natural characterization for SAT parameterized by a modulator to logarithmic treewidth, as we have already considered constant treewidth, pathwidth, and tree-depth, as well as logarithmic pathwidth and tree-depth.
1.6 Paper Organization
We first remark that we have tried to use a self-explanatory naming scheme, where X-SETH means the hypothesis that it is impossible to improve upon brute-force search for SAT parameterized by a parameter X, or when X is a complexity class, for the complete problem of X. Nevertheless, the definitions of all hypotheses used in this paper (and shown in Figure 1) and the most basic relations between them are given in Appendix A. Even though Appendix A serves as a useful reference, we have made an effort to give precise statements in all theorems, making sure that potentially non-trivial details in the formulations of various hypothesis, such as the order of quantifications of some constants, are clearly spelled out. For the rest of the paper, after some preliminaries, we present the basic connections for the two bottom classes of Figure 1 in Section 3, and then present one class in each of Section 4, Section 5, and Section 6. In Section 7 we present one further characteristic example for each class.
2 Preliminaries
We assume the reader is familiar with the basics of parameterized complexity as given for example in [12]. We refer the reader there for the definitions of treewidth and pathwidth. A graph $G$ has tree-depth $k$ if there is a bijective mapping of the vertices of $G$ to the vertices of a rooted tree of depth $k$ so that for every edge $uv$, the images of $u,v$ have an ancestor-descendant relation. Following [15], for integers $\sigma,\delta$, a $(\sigma,\delta)$-hub of a graph $G$ is a set of vertices $M$ such that each component of $G-M$ has size at most $\sigma$ and at most $\delta$ neighbors in $M$.
A modulator of a graph is a set of vertices such that removing it leaves a graph with a desired property. When $G$ is a graph and $M$ a modulator we will often write $G-M$ to denote the graph obtained after removing the modulator.
We deal with several variations of SAT, the problem of deciding if a given formula has a satisfying assignment. We write $k$-SAT for the restriction to clauses of arity at most $k$ and Max-SAT (or Max-$k$-SAT) for the more general problem of determining the assignment that satisfies the maximum number of clauses. For these problems we allow repeated clauses in the input, or equivalently we allow the input to assign (unary) weights to the clauses. An assignment of weight $k$ is an assignment that sets exactly $k$ variables to True.
For a CNF formula we define its primal graph as the graph that has a vertex for each variable and edges between variables that appear in the same clause. When $\phi$ is a CNF formula and $M$ is a set of variables, we will write $\phi-M$ to denote the formula obtained by deleting every clause that contains a variable of $M$ and, slightly abusing notation, we will also use $\phi-M$ to refer to the resulting primal graph.
We will also consider satisfiability problems for general formulas and circuits. A circuit is a DAG where vertices of positive in-degree are labeled as disjunctions ($\lor$), conjunctions ($\land$), or negations ($\neg$) and there is a unique sink, called the output gate. Vertices of zero in-degree are called input gates and the satisfiability problem asks if there is a truth assignment to the inputs that makes the output gate evaluate to True. A formula is a circuit where the maximum out-degree of any non-input gate is $1$. Unless otherwise stated we will assume that the fan-in (in-degree) of all circuits is bounded, but in most cases unbounded fan-in gates can be replaced by trees of bounded fan-in gates. The depth of a circuit is the largest distance from an input gate to the output. The size of a circuit is the number of arcs of the corresponding DAG.
Given a CNF formula $\phi$, a strong backdoor set is a set of variables such that for each assignment to these variables we obtain a formula that belongs to a base class. We will only deal with strong backdoors, so we will simply refer to the as backdoors. Furthermore, we will deal with 2-SAT and Horn backdoors, where a formula is Horn if each clause contains at most a positive literal. Recall that 2-SAT is NL-complete and Horn-SAT is P-complete, therefore both problems are solvable in polynomial time. This means that for a given backdoor the satisfiability of $\phi$ can be decided in time $2^{b}|\phi|^{O(1)}$ by trying all assignments to the backdoor. The term backdoor was introduced in [45]. Note that backdoors and modulators are closely related concepts and indeed the notion of modulator to constant treewidth which we use in this paper has also been called a backdoor to constant treewidth [20]. For more information on backdoors see [19].
In all cases where it is relevant, we will assume that the modulator or backdoor related to the problem at hand is given in the input. This is because we want to concentrate on the complexity of solving SAT given the modulator, rather than the complexity of finding the modulator. This is standard in parameterized lower bounds, see e.g. [33].
We give a detailed list of the complexity hypotheses that appear in this paper as well as simple observations about their relations in Appendix A.
3 Problems Equivalent to (Max-)CNF-SAT
In this section we deal with two hypotheses: the standard SETH and its natural weakening for Max-SAT, which we refer to as the MaxSat-SETH. One of our main motivations is to take a more careful look at the results of [15]. Recall that in this recent work it was pointed out that the SETH is “too strong”, if one aims to obtain lower bound results regarding standard width parameters, such as treewidth. For this reason, [15] focuses on a parameter which they call $(\sigma,\delta)$-hub, which is a severe restriction of treewidth, pathwidth, and tree-depth modulators. Their main result is that many of the standard lower bounds go through (assuming the SETH) even for this very restricted parameter.
In this paper we offer a more nuanced view that clarifies a bit more the picture presented in [15]. In particular, we show that for some of their lower bound results, not only is the SETH not too strong, but actually the lower bounds they obtain are equivalent to the SETH. We achieve this mainly via Theorem 3.1 below, which shows that the SETH is equivalent to the tdM-SETH, that is the hypothesis that it is impossible to avoid a brute-force enumeration for a given modulator to a graph of bounded tree-depth. Since $(\sigma,\delta)$-hub is a parameter that is a restriction of modulators to bounded tree-depth, we obtain that avoiding brute force enumeration for this parameter is also equivalent to the SETH (Corollary 3.3). In other words, Theorem 3.1 collapses three seemingly distinct parameters (the number of variables, the size of a $(\sigma,\delta)$-hub, and the size of a constant tree-depth modulator) into a single hypothesis. As we show in Section 7, these results establish that the lower bound of [15] for Coloring is actually equivalent to the SETH, and this remains true even if we parameterize by tree-depth modulator rather than $(\sigma,\delta)$-hub.
Despite the above, our results do not indicate that the SETH is the “correct” assumption for all the lower bounds of [15]. In the second part of this section we deal with the MaxSat-SETH, the hypothesis stating the same bound as the SETH but for the problem of maximizing the number of satisfied clauses of a CNF formula. Since this is clearly a problem at least as hard as deciding if a CNF formula can be satisfied, this assumption is at least as plausible as the SETH. Indeed, this assumption is also implied by the so-called Max-3Sat-Hypothesis, which states the same time bound but for CNFs of maximum arity $3$. As mentioned in [15], the relation between the SETH and the Max-3Sat-Hypothesis is currently unknown, which can perhaps be seen as a reason to believe the MaxSat-SETH as more plausible than either of the other two hypothesis. Our second result in this section is that the MaxSat-SETH remains equivalent if we parameterize not by the number of variables but by $(\sigma,\delta)$-hub size (Theorem 3.5). This allows us to establish in Section 7 that lower bounds concerning optimization problems, such as the one on Max Cut parameterized by $(\sigma,\delta)$-hub in [15], are more likely to be equivalent to MaxSat-SETH, hence perhaps more believable, than lower bounds equivalent only to the SETH.
Before we go on, let us mention that the results of this section rely on a precise formulation of the SETH (and related hypotheses) which assumes hardness for $k$-CNF formulas, for sufficiently large $k$. We note that this is important in this section, because we are dealing with extremely restrictive graph parameters. Once we move on to Section 4, it will immediately become possible to break down large clauses into smaller ones by introducing new variables, so the arity of the considered CNFs will no longer be of material importance (see Section A.2).
3.1 Equivalent Formulations of the SETH
Theorem 3.1.
The following statements are equivalent:
  1. 1.
    (SETH is false) There is an $\epsilon>0$ such that for all $k$ there is an algorithm that solves $k$-SAT in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$.
  2. 2.
    (tdM-SETH is false) There is an $\epsilon>0$ such that for all $c,k$ there is an algorithm that takes as input a $k$-CNF formula $\phi$ and a modulator $M$ of $\phi$ of size $d$, such that $\phi-M$ has tree-depth at most $c$, and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{d}|\phi|^{O(1)}$.
Proof 3.2.
The second statement clearly implies the first, as the set of all variables is a modulator to tree-depth 00. The interesting direction is to show that if we have a $k$-SAT algorithm running in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$, then this implies a faster algorithm parameterized by constant tree-depth modulator.
We are given a $k$-CNF formula $\phi$ and a modulator $M$ such that removing $M$ from the primal graph produces a graph of tree-depth $c$, for some fixed $c$. Assume that for the resulting graph after we remove $M$ we have a forest of rooted trees, where each tree has height at most $c$, and a bijection of the variables of $\phi$ (outside $M$) to the vertices of this forest so that two variables appear in the same clause if and only if one is an ancestor of the other.
Our strategy will be to use the standard resolution rule to eliminate from $\phi$ all the variables except those of the modulator. We will then apply the assumed algorithm for $k^{\prime}$-SAT, for some $k^{\prime}$ depending on $c$ and $k$.
Recall that resolution is the following rule: when a CNF formula contains two clauses $C,C^{\prime}$ such that for some variable $x$, $C$ contains $x$ and $C^{\prime}$ contains $\neg x$, then we can add to the formula the clause the contains all literals of $C\cup C^{\prime}$ except $x,\neg x$ without affecting satisfiability. We will say that we exhaustively resolve over variable $x$ if we (i) apply this rule to each pair of clauses $C,C^{\prime}$ containing $x,\neg x$ (ii) after doing so remove from $\phi$ all clauses that contain $x$ or $\neg x$.
Claim 1.
Exhaustively resolving over a variable $x$ does not affect the satisfiability of $\phi$.
{claimproof}
It is easy to see that the clauses added to $\phi$ do not affect satisfiability, as they can be inferred from existing clauses. To see that deleting all clauses that contain the variable $x$ does not affect satisfiability, we take two cases. If the new formula has a satisfying assignment $\sigma$ such that for all clauses $C$ which contain $x$ positive in $\phi$, a literal of $C$ other than $x$ is set to True by $\sigma$, then we can satisfy $\phi$ by extending $\sigma$ by setting $x$ to False. Otherwise, we claim that all clauses $C^{\prime}$ that contain $\neg x$ in $\phi$ have a literal other than $\neg x$ set to True by $\sigma$, so we can set $x$ to False and satisfy $\phi$. Indeed, we are now in a case where some clause $C$ which contains $x$ positive has all its literals set to False by $\sigma$, but this clause has been paired with all $C^{\prime}$ which contained $\neg x$. Since the new formula is satisfied by $\sigma$, each $C^{\prime}$ must contain a True literal.
Our algorithm is to apply for each $i\in[c]$, starting from $i=c$ and descending, exhaustive resolution to all the variables which correspond to vertices of the rooted trees which are at distance $i$ from the root. After each iteration we have eliminated a layer of leaves and decreased the tree-depth of $\phi-M$ by one, without affecting the satisfiability of the formula. We now show by induction that the result of this process after $c$ iterations is a formula with maximum arity $k^{\prime}\leq 2^{c}k$ whose number of clauses is at most $m^{2^{c}}$, where $m$ is the number of clauses of $\phi$.
We show this claim by induction on $c$. For $c=1$, $M$ is in fact a vertex cover of the primal graph. For each variable $x$ of the independent set we apply exhaustive resolution. Each new clause is made up of two old clauses, so has arity at most $2k$. Furthermore, if $x$ appears in $d(x)$ clauses, we will construct at most $d^{2}(x)/2$ clauses when resolving over $x$. The total number of new clauses is then at most $\sum_{x\not\in M}d^{2}(x)/2\leq m^{2}/2$, where we use the fact that $m\geq\sum_{x\not\in M}d(x)$, because the sets of clauses that contain two distinct variables not in $M$ are disjoint, and because of standard properties of the quadratic function. Since we can assume that $m$ is large enough, the total number of clauses we now have is at most $m+\frac{m^{2}}{2}\leq m^{2}$.
We now proceed with the inductive step. Suppose that the claim is true for smaller values of $c$ and we have an instance of maximum arity $k$ and $m$ clauses. By the same arguments as before, applying exhaustive resolution to all leaves produces an instance where $\phi-M$ has tree-depth $c-1$, the maximum arity is at most $k^{\prime}\leq 2k$ and the number of clauses is at most $m^{\prime}=m^{2}$. By inductive hypothesis, completing the procedure will produce a formula with arity at most $2^{c-1}k^{\prime}\leq 2^{c}k$ and with at most $(m^{\prime})^{2^{c-1}}\leq m^{2^{c}}$ clauses, as desired.
In the end we have a formula with $n=|M|$ variables, arity $k^{\prime}\leq 2^{c}k$ and size $|\phi^{\prime}|\leq|\phi|^{2^{c}}$. Running the presumed satisfiability algorithm on this formula implies the theorem.
Corollary 3.3.
There is an $\epsilon>0$ such that for all $k,\sigma,\delta>0$ there is an algorithm which takes as input a $k$-CNF formula $\phi$ and a $(\sigma,\delta)$-hub of size $m$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$ if and only if the SETH is false.
Proof 3.4.
The proof consists of observing that a $(\sigma,\delta)$-hub is a modulator to tree-depth at most $\sigma$ and invoking Theorem 3.1.
3.2 Equivalent Formulation of the SETH for Max-SAT
Theorem 3.5.
The following are equivalent:
  1. 1.
    (MaxSat-SETH is false) There exists $\varepsilon>0$ such that for all $k$ there is an algorithm that takes as input a $k$-CNF $\phi$ on $n$ variables and computes an assignment satisfying the maximum number of clauses in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$.
  2. 2.
    ($(\sigma,\delta)H$-MaxSat-SETH is false) There exists $\varepsilon>0$ such that for all $k,\sigma,\delta$ there is an algorithm that takes as input a $k$-CNF $\phi$ and a $(\sigma,\delta)$-hub $M$ of size $m$ and computes an assignment satisfying the maximum number of clauses in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
Proof 3.6.
It is easy to see that the second statement implies the first, as the set of all variables is a $(\sigma,\delta)$-hub. Recall that we allow clauses to be repeated in the input, so the running time of statement 1 cannot be written as $O((2-\varepsilon)^{n})$, because even for $k$-CNF, for $k$ fixed, the size of the formula cannot be bounded by a function of $n$ alone.
For the interesting direction, we consider one by one the components of $\phi-M$, each of which has size at most $\sigma$ and at most $\delta$ neighbors in $M$. Without loss of generality we assume that each component has size exactly $\sigma$ and exactly $\delta$ neighbors, by adding dummy variables or neighbors as needed.
Take such a component $C$ and for each assignment $s$ to its $\delta$ neighbor variables in $M$, calculate the assignment to the variables of $C$ that maximizes the number of clauses involving variables of $C$ which are satisfied (this can be done by going through the $2^{\sigma}$ possible assignments to the variables of the component). Let $w_{s}$ be the number of such clauses. Our goal is to remove the variables of $C$ from $\phi$ and add some clauses involving the variables of $N(C)$ such that for each assignment $s$ the total weight of satisfied new clauses is proportional to $w_{s}$. More precisely, we do the following: for each assignment $s$ to the variables of $N(C)$ we construct a clause falsified by this assignment (and only this assignment) to $N(C)$, and give it weight $\sum_{s^{\prime}\neq s}w_{s^{\prime}}$, that is give it weight that is equal to the sum of the weights corresponding to all other assignments to $N(C)$. We now observe that if we use the assignment $s$ for $N(C)$, the total weight we obtain from the new clauses is $w_{s}+(2^{\delta}-2)\sum_{s^{\prime}}w_{s^{\prime}}$, that is, the contribution of each assignment (including $s$) is counted $22^{\delta}-$ 2 times, but $w_{s}$ is counted once more, as a $w_{s}$ term appears in the weight of all satisfied clauses. We increase the target value of the Max-SAT instance by $(2^{\delta}-2)\sum_{s^{\prime}}w_{s^{\prime}}$ and implement the weights by taking copies of the constructed clauses. After performing this process exhaustively we are left with a Max-SAT instance with maximum arity $\max\{k,\delta\}$ and size at most $O(2^{\sigma+\delta}|\phi|)$, so executing the supposed algorithm gives the lemma.
4 Modulators to Constant Width and Linear Depth Circuits
In this section we present a sequence of reductions which show that several seemingly distinct fine-grained complexity hypotheses are actually the same. Our results our summarized in Theorem 4.1 below.
Theorem 4.1.
The following statements are equivalent:
  1. 1.
    (LD-C-SETH is false) There is an $\epsilon>0$ and an algorithm which takes as input a boolean circuit of size $s$ on $n$ inputs with depth at most $\epsilon n$ and decides if the circuit is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
  2. 2.
    (pwM-SETH is false) There is an $\epsilon>0$ such that for all fixed $c>0$ we have an algorithm which takes as input a CNF formula $\phi$ and a size $m$ modulator of $\phi$ to pathwidth at most $c$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
  3. 3.
    (twM-SETH is false) The same as the previous statement but for a modulator to treewidth at most $c$.
  4. 4.
[…]
  5. 5.
    (W[SAT]-SETH is false) There is an $\epsilon>0$ such that there is an algorithm which takes as input a general Boolean formula $\phi$ on $n$ variables and an integer $k$ and decides if $\phi$ has a satisfying assignment that sets exactly $k$ variables to True in time $n^{(1-\varepsilon)k}|\phi|^{O(1)}$.
In other words, we make connections between three types of hypotheses: one involving the difficulty of satisfying a circuit (with a mild restriction on its depth); three different hypotheses involving structural graph parameters; and a hypothesis on the complexity of the standard complete problem for the class W[SAT]. Before we proceed to the proofs, let us give some high-level intuition regarding these connections.
Among the five statements of Theorem 4.1 three have an obvious chain of implications: solving SAT parameterized by the size of a modulator to logarithmic tree-depth is at least as hard as solving it for a modulator to constant treewidth, which is at least as hard as for a modulator to constant pathwidth. These relations follow readily from the fact that $\textrm{pw}\leq tw$ and $\textrm{td}\leq\textrm{tw}\log n$ for all graphs. What is perhaps a bit surprising is that these three parameters turn out to be of equivalent complexity.
In order to add the remaining chains of implications to complete Theorem 4.1 we need several steps. The one that is conceptually the least obvious is perhaps the implication 2$\Rightarrow$1. Here we need to prove that the satisfiability problem for any circuit (with the stated restriction on its depth) can be reduced to SAT with a modulator to only constant pathwidth. The idea now is that the variables of the modulator will correspond to the inputs of the circuit, so we need to encode the workings of the circuit via a formula of constant pathwidth. This sounds counter-intuitive, since evaluating a circuit is a P-complete problem, but constant pathwidth SAT instances can be solved in logarithmic space. This observation makes it clear why it is crucial that we have placed a restriction on the depth of the circuit. In order to achieve our goal we rely on a classical result from computational complexity theory known as Barrington’s theorem [5] which allows us to transform any boolean circuit of depth $d$ into a branching program of length $2^{O(d)}$ and constant width. The fact that the branching program has constant width means that its workings can be encoded via a constant pathwidth formula, while the fact that the depth of the input circuit is at most $\epsilon n$ means that the exponential blow-up in the size can be tolerated by carefully dealing with the involved constants.
The other interesting implication is 5$\Rightarrow$4. To better explain this, suppose we were instead trying to prove 1$\Rightarrow$4. In that case we would try to construct a circuit whose inputs are the variables of the modulator given to us in statement 4. For this, we would need to somehow eliminate all the remaining variables, using the fact that they form a graph of logarithmic tree-depth. For this we use an inductive construction which proceeds bottom-up, constructing for each node of the tree a formula that eliminates all variables in its sub-tree. To do this we try both values of the variable and construct two copies of the formulas we have for its children, one for each value. This construction is of course exponential in the tree-depth, but since the tree-depth is $O(\log|\phi|)$, the result is polynomial in the input formula. To obtain instead 5$\Rightarrow$4 we use standard tricks that encode an assignment to $n$ variables via a weight $k$ assignment to $2^{n/k}$ variables.
Finally, for 1$\Rightarrow$5 we can construct a circuit with $k\log n$ inputs, encoding the $k$ variables to be set to True in the sought weight-$k$ assignment. A subtle point here is that we need to ensure that the circuit we construct does not have too large depth. However, for this we can rely on the fact that the W[SAT]-complete problem of statement 5 deals with boolean formulas, that is, circuits of out-degree $1$, which always admit an equivalent reformulation with logarithmic depth. Hence, we construct an instance with $k\log n$ input and depth $O(\log n)$, which for $k$ large enough satisfies the requirements of statement 1.
Proof 4.2 (Proof of Theorem 4.1).
[…]
4.1 W[SAT] to Pathwidth, Treewidth, Logarithmic Tree-depth Modulators
Lemma 4.3.
Statement 5 of Theorem 4.1 implies statement 4.
Proof 4.4.
We suppose there exist constants $\epsilon>0,d>0$ and an algorithm that takes as input a Boolean formula $\psi$ (that is, a Boolean circuit of maximum out-degree $1$) with $N$ inputs and an integer $k$ and decides if $\psi$ has a satisfying assignment of weight $k$ in time $N^{(1-\varepsilon)k}|\psi|^{d}$.
Fix now a constant $c>0$ and we are given a CNF formula $\phi$ and a modulator $M$ of size $m$ such that removing all vertices of $M$ from the primal graph of $\phi$ results in a graph of tree-depth $h$ with $h\leq c\log|\phi|$. Suppose we are given a rooted tree of height $h$ such that the variables of $\phi$ (outside $M$) can be bijectively mapped to nodes of the tree in a way that if two variables appear in a clause together, then one is an ancestor of the other. We want to decide the satisfiability of $\phi$ in time $(2-\varepsilon^{\prime})^{m}|\phi|^{O(1)}$.
Our high-level plan is to construct from $\phi$ a Boolean formula $\psi$ and an integer $k$ satisfying the following:
  1. 1.
    $\phi$ is satisfiable if and only if $\psi$ has a satisfying assignment of weight $k$.
  2. 2.
    $\psi$ has $N=O(2^{m/k})$ input variables and can be constructed in time $2^{m/k}|\phi|^{O(1)}$.
In particular, we will set $k=\lceil\frac{2d}{\varepsilon}\rceil$. Then $N^{(1-\varepsilon)k}|\psi|^{d}$ can be upper bounded as follows: $N^{(1-\varepsilon)k}=O(2^{(1-\varepsilon)m})$ and $|\psi|^{d}=2^{dm/k}|\phi|^{O(1)}$. But $\frac{d}{k}\leq\frac{\varepsilon}{2}$, therefore the algorithm runs in $2^{(1-\frac{\varepsilon}{2})m}|\phi|^{O(1)}$.
We now describe how to construct $\psi$ from $\phi$, assuming we have the rooted tree of height $h\leq c\log|\phi|$ we mentioned covering the primal graph of $\phi$ outside of $M$. We first describe the construction of an auxiliary formula $\psi_{r}$, whose input variables are only the variables of $M$, where $r$ is the root of the tree of height $h$. The formula will be built inductively, by constructing a formula $\psi_{b}$ for each node $b$ of the tree, starting from the leaves. In the remainder we use, for a node $b$ of the tree, $h_{b}$ to denote the height of $b$ in the tree, that is its maximum distance from any of its descendants, and $\phi_{b}$ to denote the formula made up of the clauses of $\phi$ that contain a variable associated with $b$ or one of its descendants. We can assume without loss of generality that all clauses of $\phi$ contain a variable outside of $M$ (if not, replace $C$ with $(C\lor x),(C\lor\neg x)$, where $x$ is a new variable outside $M$), hence we have $\phi_{r}=\phi$ and also if $b,b^{\prime}$ are nodes such that $h_{b}=h_{b^{\prime}}$, then $\phi_{b}$ and $\phi_{b^{\prime}}$ are made up of disjoint sets of clauses of $\phi$. To see this, if there were a clause that contains a variable associated with $b$ (or one of its descendants) and $b^{\prime}$ (or one of its descendants) this would add to the primal graph an edge that is not between an ancestor and a descendant, contradiction.
The formula $\psi_{b}$ will have the following properties:
  1. 1.
    The input variables of $\psi_{b}$ are the variables of $M$ and all the variables associated with ancestors of $b$ in the tree.
  2. 2.
    $\psi_{b}$ is satisfied by an assignment to its input variables if and only if it is possible to extend this assignment to the variables associated with $b$ and its descendants, so that all clauses of $\phi_{b}$ are satisfied.
  3. 3.
    $|\psi_{b}|=O(2^{h_{b}}|\phi_{b}|)$.
As mentioned, $\psi_{r}$ will be the formula obtained by this procedure for the root of the tree and will only have $M$ as input variables.
In order to construct $\psi_{b}$ we proceed by induction on $h_{b}$. If $h_{b}=0$ then $\psi_{b}$ is constructed by taking $\phi_{b}$, and obtaining two formulas $\phi_{b}^{0},\phi_{b}^{1}$ by setting the variable associated with $b$ to False, True, respectively. We set $\psi_{b}=\phi_{b}^{0}\lor\phi_{b}^{1}$. Clearly, the three properties are satisfied.
Now suppose $h_{b}>0$ and let $b_{1},\ldots,b_{s}$ be the children of $b$ in the tree. Let $x$ be the variable associated with $b$. We have already constructed formulas $\psi_{b_{1}},\ldots,\psi_{b_{s}}$. For each $i\in[s]$ let $\psi_{b_{i}}^{0}$ (respectively $\psi_{b_{i}}^{1}$) be the formula obtained from $\psi_{b_{i}}$ if we set $x$ to False (respectively True) and simplify the formula accordingly. Let $C_{b}$ be the set of clauses of $\phi$ containing $x$ but no variable associated with a descendant of $b$. Let $C_{b}^{0}$ (respectively $C_{b}^{1}$) be the CNF formula made up of the clauses of $C_{b}$ we obtain if we set $x$ to False (respectively True), that is $C_{b}^{0}$ is the set of clauses containing $x$ positive, but with $x$ removed (similarly for $C_{b}^{1}$). We define $\psi_{b}$ as follows:
$\psi_{b}=\left(C_{b}^{0}\land\bigwedge_{i\in[s]}\psi_{b_{i}}^{0}\right)\lor\left(C_{b}^{1}\land\bigwedge_{i\in[s]}\psi_{b_{i}}^{1}\right)$
Let us argue why $\psi_{b}$ satisfies the three properties. First, $\psi_{b}$ has the right set of variables, as we have eliminated $x$ from all formulas.
Second, suppose $\psi_{b}$ is satisfied by some assignment to its input variables and suppose without loss of generality that the first term of the disjunction is satisfied (the other case is symmetric), therefore $C_{b}^{0}$ is satisfied. We set $x$ to False, and thus satisfy all of $C_{b}$. For all $i\in[s]$ we have that $\psi_{b_{i}}^{0}$ is satisfied, therefore $\psi_{b_{i}}$ is satisfied by the current assignment to $M$ and the ancestors of $b$, extended with $x$ set to False. By induction, this can be extended to an assignment satisfying $\phi_{b_{i}}$. Since we have satisfied $C_{b}$ and $\phi_{b_{i}}$ for all $i\in[s]$, we have satisfied $\phi_{b}$. For the converse direction, if it is possible to extend some assignment to $M$ and the ancestors of $b$ so that $\phi_{b}$ is satisfied, then this assignment satisfies $\psi_{b}$. Indeed, suppose that in the assumed extension $x$ is set to False (again the other case is symmetric). This implies that all the clauses of $C_{b}^{0}$ must be satisfied, as they contain $x$ positive. Furthermore, by induction, since we can satisfy $\phi_{b_{i}}$ (for all $i\in[s]$) by setting $x$ to False, we also satisfy $\psi_{b_{i}}$ by setting $x$ to False, hence $\psi_{b_{i}}^{0}$ evaluates to True. Therefore, the first term of the disjunction of $\psi_{b}$ evaluates to True.
Third, to bound the size of $\psi_{b}$ we have $|C_{b}^{0}|+|C_{b}^{1}|\leq 2|\phi_{b}|$ and we also have $2\sum_{i\in[s]}\psi_{b_{i}}=O(2^{b_{h}}\sum_{i\in[s]}|\phi_{b_{i}}|)=O(2^{b_{h}}|\phi_{b}|)$, (where we used the fact that the $\phi_{b_{i}}$ are disjoint) giving the desired bound. We therefore have $|\psi_{r}|=|\phi|^{O(c)}=|\phi|^{O(1)}$.
We now transform $\psi_{r}$ into a formula $\psi$ such that $\psi$ has a satisfying assignment of weight $k$ if and only if $\psi_{r}$ has a satisfying assignment. To do so, we partition the $m$ input variables of $\psi_{r}$ into $k$ sets $V_{i}$, $i\in[k]$, each of size at most $\lceil m/k\rceil$. For each assignment $\sigma$ to the variables $V_{i}$ we construct a new variable $x_{i,\sigma}$. For each $i\in[k]$ we add a clause $\left(\bigvee_{\sigma}x_{i,\sigma}\right)$. For each $y\in V_{i}$ let $\sigma^{1}(y)$ (respectively $\sigma^{0}(y)$) be the set of assignments to $V_{i}$ which set $y$ to True (respectively False). Then, we replace every appearance of $y$ in $\psi_{r}$ with the clause $\bigvee_{\sigma\in\sigma^{1}(y)}x_{i,\sigma}$ and every appearance of $\neg y$ with the clause $\bigvee_{\sigma\in\sigma^{0}(y)}x_{i,\sigma}$. Note that now we have $|\psi|=O(2^{m/k}|\psi_{r}|)=2^{m/k}|\phi|^{O(1)}$, as promised. Furthermore, it is not hard to see that $\psi$ can only be satisfied by an assignment of weight $k$ if we select for each $i\in[k]$ a unique variable $x_{i,\sigma}$ to set to True, thus creating a one-to-one correspondence between satisfying assignments to $\phi$ and weight-$k$ satisfying assignments to $\psi$.
4.2 Linear-depth Circuits and Constant Pathwidth Modulators
In this section we present Lemma 4.7 which, as mentioned will rely on the transformation of Boolean circuits to bounded-width branching programs given by Barrington’s theorem. Before we begin let us recall some definitions and background.
Definition 4.5.
A layered branching program of width $w$ and length $\ell$ with input variables $x_{1},\ldots,x_{n}$ is a DAG with the following properties:
  1. 1.
    Vertices are partitioned into $\ell$ sets $V_{1},\ldots,V_{\ell}$ (the layers), such that each arc goes from a vertex in $V_{i}$ to a vertex in $V_{i+1}$. For each $i\in[\ell]$ we have $|V_{i}|\leq w$.
  2. 2.
    There is a unique source (the start vertex), while $V_{\ell}$ contains two vertices (the accept and reject vertices). All other vertices have strictly positive in- and out-degree.
  3. 3.
    Each arc is labeled True or False and each vertex with positive out-degree is labeled with an input variable.
We say that a branching program accepts an assignment $\sigma$ to its variables if when we start at the source and we follow a path where for each vertex labeled $x_{i}$ we follow the out-going arc labeled $\sigma(x_{i})$, we reach the accept vertex.
We now recall Barrington’s classical result.
Theorem 4.6.
([5]) There is an algorithm which, given a boolean circuit on $n$ inputs with depth $d$, produces an equivalent branching program on the same inputs with width $5$ and length $O(4^{d})$ in time $O(4^{d})$.
Lemma 4.7.
Statement 2 of Theorem 4.1 implies statement 1.
Proof 4.8.
The proof relies heavily on Theorem 4.6: given an arbitrary Boolean circuit we will first construct an equivalent branching program and then convert this branching program into a CNF formula. The formula will have a modulator to small pathwidth because the branching program has constant width.
More precisely, suppose that for some $\varepsilon>0$ there exists an algorithm which for all $c>0$ takes as input a CNF formula $\phi$ and a modulator $M$ of its primal to pathwidth $c$, with $|M|=m$, and decides if $\phi$ is satisfiable in time $2^{(1-\varepsilon)m}|\phi|^{c^{\prime}}$, for some $c^{\prime}>0$. Let $\varepsilon^{\prime}=\frac{\varepsilon}{4c^{\prime}}$. We claim that we can obtain an algorithm which, given an $n$-input circuit $C$ of size $s$ and depth $\epsilon^{\prime}n$ decides if $C$ is satisfiable in time $2^{(1-\varepsilon^{\prime})n}s^{O(1)}$.
We first invoke Theorem 4.6 and obtain a branching program of length $O(4^{\epsilon^{\prime}n})$ and width $5$. We then construct a CNF formula whose variables are the $n$ variables $x_{1},\ldots,x_{n}$ representing the inputs to the branching program, plus the $3\ell$ variables $y_{1,t},y_{2,t},y_{3,t}$, for $t\in[\ell]$. The intuitive meaning of these variables is that the assignment to $y_{1,t},y_{2,t},y_{3,t}$ should encode the vertex which the execution of the branching program will reach in layer $t$. We make an injective mapping from the vertices of $V_{t}$ to the assignments of $y_{1,t},y_{2,t},y_{3,t}$ and add clauses to exclude assignments not associated with a vertex of $V_{t}$. We then add clauses to ensure that arcs are followed: for each assignment to $y_{1,t},y_{2,t},y_{3,t},y_{1,t+1},y_{2,t+1},y_{3,t+1}$ as well as the at most five input variables used as labels for the vertices of $V_{t}$, we construct clauses ensuring the transitions conform to the arcs of the branching program. Finally, we add clauses ensuring that in layer $1$ we begin at the start vertex and in the last layer we reach that accept vertex. It is not hard to see that the constructed CNF formula is satisfiable if and only if there exists an assignment to the input variables accepted by the branching program, and hence the circuit.
We observe that in the formula we have constructed there is a modulator of size $n$ such that removing those variables results in a graph of pathwidth at most 5 (we can use a sequence of bags containing $\{y_{1,t},y_{2,t},y_{3,t},y_{1,t+1},y_{2,t+1},y_{3,t+1}\}$ for all $t$). If we execute the supposed algorithm on this instance the running time will be at most $2^{(1-\varepsilon)n}|\phi|^{c^{\prime}}$. We have $|\phi|=O(4^{\epsilon^{\prime}n})=O(2^{\frac{\varepsilon n}{2c^{\prime}}})$. The algorithm therefore runs in $O(2^{(1-\frac{\varepsilon}{2})n})=O(2^{(1-\varepsilon^{\prime})n})$.
4.3 Linear-Depth Circuits and Weighted SAT
Lemma 4.9.
Statement 1 of Theorem 4.1 implies statement 5.
Proof 4.10.
Suppose there exist $\varepsilon>0,c>0$ such that satisfiability of $n$-input circuits of depth $\epsilon n$ and size $s$ can be decided in time $2^{(1-\varepsilon)n}s^{c}$.
We are given an $N$-variable formula $\phi$ and an integer $k$ and are asked to decide if $\phi$ has a weight-$k$ satisfying assignment in time $N^{(1-\varepsilon^{\prime})k}|\phi|^{O(1)}$. We can assume without loss of generality that $|\phi|\leq N^{\varepsilon k/6}$, otherwise the problem can be solved in $N^{k}|\phi|^{O(1)}=|\phi|^{O(1/\varepsilon)}=|\phi|^{O(1)}$ time by trying out all assignments. Therefore, $\log|\phi|\leq\frac{\varepsilon k\log N}{6}$.
We now consider $\phi$ as a boolean circuit. Using standard techniques, we can balance $\phi$ to obtain an equivalent circuit of depth at most $3\log|\phi|\leq\frac{\varepsilon k\log N}{2}$ (see e.g. [7]) while maintaining a size that is polynomial in $|\phi|$. We add to $\phi$ $k\lceil\log N\rceil$ new inputs, partitioned into $k$ groups of $\lceil\log N\rceil$ inputs, whose intuitive meaning is that they encode the indices of the $k$ inputs to $\phi$ which are meant to be set to True in a satisfying assignment. For each group we enumerate all $O(N)$ assignments and associate each with an input of $\phi$. We transform the $N$ input gates of $\phi$ into disjunction gates of fan-in $k$, whose inputs are conjunctions of fan-in $\lceil\log N\rceil$ connected to the new inputs in one of the $k$ groups via appropriate $\neg$ gates so that an input of $\phi$ is set to True if and only if its index is the assignment to one of the $k$ groups of new inputs. This construction can be made to have depth $O(\log\log N+\log k)$ by replacing high fan-in gates with binary trees. It is now not hard to see that the new circuit has a satisfying assignment that gives distinct assignments to each of the $k$ groups if and only if $\phi$ has a weight-$k$ satisfying assignment. What remains is to add for each distinct $i,i^{\prime}\in[k]$ some gadget that ensures that the assignments to groups $i,i^{\prime}$ are not identical (and hence we select an assignment of weight exactly $k$ for $\phi$). This can easily be done with $O(k^{2})$ formulas of size $O(\log N)$ so the total depth remains at most the depth of $\phi$ plus $O(\log\log N+\log k)$. We can therefore assume that the depth of the circuit we have now constructed is at most $\varepsilon k\log N$. The circuit has at most $k\log N+k$ inputs and has been constructed in polynomial time, so running the supposed satisfiability algorithm takes $2^{(1-\varepsilon)k\log N}2^{k}|\phi|^{O(1)}$ which for $N>2^{2/\varepsilon}$ is upper-bounded by $N^{(1-\varepsilon/2)k}|\phi|^{O(1)}$.
5 Modulators to Logarithmic Pathwidth and 2-SAT Backdoors
In this section we show an equivalence between the following two hypotheses.
Theorem 5.1.
The following are equivalent:
  1. 1.
    (2SatB-SETH is false) There is an algorithm that takes as input a CNF formula $\phi$ and a strong 2-SAT backdoor of size $b$, and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{b}|\phi|^{O(1)}$.
  2. 2.
[…]
Before we proceed, let us give some intuition. One can think of both problems that we are dealing with here as reachability problems, in the following sense: if we are given an assignment to a 2-SAT backdoor, we can verify if it can be extended to a satisfying assignment by solving 2-SAT, which is known to be complete for the class NL, that is, equivalent to reachability. On the other hand, if we have an instance of logarithmic pathwidth, the intuition of the class EPNL defined in [26] (see also [31]) is that solving such an instance corresponds exactly to deciding if a non-deterministic machine with space $O(\log n)$ will accept a certain input, that is, this problem is again complete for NL. This intuition leads us to believe that modulators to logarithmic pathwidth and 2-SAT backdoors are equivalent.
There is, however, one complication in the above reasoning. What we have established is that once someone supplies us with an assignment to the 2-SAT backdoor, checking if this assignment is correct has the same complexity as checking whether an assignment to the logarithmic pathwidth modulator is correct. However, what we really care about is whether an algorithm which avoids enumerating all assignments in one case can be used to avoid enumerating the assignments in the other. Hence, an argument which assumes that the assignment is known cannot be used as a black box. What we need is a reduction that transfers the answer to all possible assignments for a backdoor to the possible assignments to a modulator, without going through the assignments (ideally running in polynomial time).
Given this, it is natural to revisit the reasons for which both 2-SAT and SAT on graphs of logarithmic pathwidth are essentially equivalent to reachability. We then run into the following obstacle: while SAT on graphs of logarithmic pathwidth easily reduces to and from reachability, 2-SAT is actually only easily seen to be equivalent to the complement of reachability. The reason 2-SAT ends up being complete for NL is the famous Immerman–Szelepcsényi theorem, which established that NL=coNL [24].
[…]
Definition 5.2.
Suppose we are given a DAG $G=(V,E)$, a designated source $s$ and sink $t$ of $G$, a set $M$ of $m$ boolean variables $x_{1},\ldots,x_{m}$, and an annotation function which assigns to some of the arcs of $G$ a literal (that is, a variable or its negation) from $M$.
The Ann-s-t-Reach problem is the following: decide if there exists an assignment to $M$ such that if we only keep in the graph arcs with no annotation or with an annotation set to True by the assignment, then there is a directed path from $s$ to $t$.
The Ann-s-t-Non-Reach problem is the following: decide if there exists an assignment to $M$ such that if we keep the same arcs as above, then there is no directed path from $s$ to $t$.
Intuitively, Ann-s-t-Reach allows us to encode $2^{m}$ instances of reachability simultaneously: for each assignment to the variables of $M$ we obtain a distinct graph. It is important to note here that Ann-s-t-Reach and Ann-s-t-Non-Reach are not complementary problems. In both problems the outer quantification is existential: we ask if there exists an assignment to $M$ such that, in one case a path exists, and in the other no path exists.
[…]
Proof 5.3 (Proof of Theorem 5.1).
We split the proof into three parts. In Lemma 5.4 we show an equivalence between 2-SAT backdoors and Ann-s-t-Non-Reach; in Lemma 5.10 we show an equivalence between SAT with a modulator to logarithmic pathwidth and Ann-s-t-Reach; and in Lemma 5.12 we show that Ann-s-t-Reach and Ann-s-t-Non-Reach have the same complexity by revisiting the proof of the Immerman–Szelepcsényi theorem.
5.1 2-SAT and Non-Reachability
Lemma 5.4.
There is an algorithm solving Ann-s-t-Non-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$, for some $\varepsilon>0$ if and only if the 2SatB-SETH is false.
Proof 5.5.
The proof is broken down into Lemma 5.6 and Lemma 5.8.
Lemma 5.6.
If there is an algorithm falsifying 2SatB-SETH, then there is an algorithm solving Ann-s-t-Non-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$.
Proof 5.7.
We use a standard reduction of non-reachability to 2-SAT, except we have to take into account the arc annotations. Suppose we are given a Ann-s-t-Non-Reach instance. Let $G=(V,E)$ be the given DAG on $n$ vertices. We construct a CNF formula with variables the variables of $M$, given with the Ann-s-t-Non-Reach instance, to which we add a variable $x_{v}$ for each $v\in V$. We will set $M$ to be the backdoor set. Now, for each arc $(v_{1},v_{2})\in E$ we add the clause ($(\neg x_{v_{1}}\lor x_{v_{2}})$ to the formula if the arc is not annotated, otherwise we add the clause ($(\neg x_{v_{1}}\lor x_{v_{2}}\lor\neg\ell)$ if the arc is annotated with the literal $\ell$. We also add the unit clauses $(x_{s})$ and ($(\neg x_{t})$. This completes the construction.
We now claim that if the formula we constructed is satisfiable, using the same assignment to $M$ yields a digraph where there is no path from $s$ to $t$. Indeed, suppose that for the same assignment there is a path from $s$ to $t$, meaning that the assignment sets to True all the literals used as annotations to the edges of the path. This implies that in the corresponding closes, the backdoor literals we have added are set to False by the satisfying assignment. But then this would create a chain of implications given $x_{s}\to x_{t}$, which in conjunction with the two unit clauses would falsify the formula.
For the converse direction, suppose that there is an assignment to $M$ that makes $t$ unreachable from $s$. We use the same assignment in the formula and extend it by setting $x_{v}$ to True if and only if $v$ is reachable from $s$ under the same assignment to $M$. It is not hard to see that this assignment will validate the formula: the only possibility of failure is for a clause representing the arc $(v_{1},v_{2})$, where $v_{1}$ is reachable from $s$ and $v_{2}$ is not. In such a case, the arc must be annotated by a literal that is set to False by the assignment to $M$, so the corresponding clause is satisfied.
Finally, we observe that the construction is polynomial and the size of the backdoor is exactly $m$, as each clause contains at most $2$ of the new variables.
Lemma 5.8.
If there is an algorithm solving Ann-s-t-Non-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$, for some $\varepsilon>0$, then the 2SatB-SETH is false.
Proof 5.9.
We want to reduce 2-SAT to Ann-s-t-Non-Reach while using the annotation to encode the backdoor variables. We are given a CNF formula $\phi$ with a set $B=\{y_{1},\ldots,y_{b}\}$ of $b$ backdoor variables and $n$ non-backdoor variables $X=\{x_{1},\ldots,x_{n}\}$. We assume without loss of generality that each clause contains exactly $2$ non-backdoor variables: indeed, no clause can contain three or more such variables (otherwise $B$ would not be a strong backdoor), and if $C$ contains fewer than $2$ such variables, we can replace it with the pair of clauses $(C\lor x_{i}),(C\lor\neg x_{i})$, for some $x_{i}$ that does not appear in $C$.
We now describe a basic implication digraph: for each variable $x_{i}\in X$ we construct two vertices $x_{i},\neg x_{i}$. For each clause $C$, let $\ell_{1},\ell_{2}$ be the literals of $C$ made up of variables of $X$. We construct an arc from $\neg\ell_{1}$ to $\ell_{2}$ and an arc from $\neg\ell_{2}$ to $\ell_{1}$. If $C$ has size $2$ we do not annotate these arcs. Otherwise, if $C$ contains $b^{\prime}\leq b$ other literals, we subdivide these arcs so that they become directed paths of length $b^{\prime}$ and annotate each arc of the resulting path with the negation of a distinct literal of $C$ involving a variable of $B$. The intuitive meaning of this construction is that we have a path from $\neg\ell_{1}$ to $\ell_{2}$ and from $\neg\ell_{2}$ to $\ell_{1}$ if and only if all the backdoor literals of $C$ are set to False.
The digraph above is not necessarily a DAG, but we can construct an equivalent DAG as follows: take $2n$ copies of the graph above, number them $1,\ldots,2n$, for each vertex $\ell$ in copy $i$ add non-annotated arcs to the same vertex in all later copies, and for each edge $(\ell_{1},\ell_{2})$ and each $i\in[2n-1]$ add an arc from the vertex $\ell_{1}$ in copy $i$ to the vertex $\ell_{2}$ in copy $i+1$ maintaining the annotation of the arc. We now have the property that there is a path from $\ell_{1}$ to $\ell_{2}$ in the original graph if and only if there is a path from $\ell_{1}$ in the first copy to $\ell_{2}$ in the last copy of the vertex set in the resulting DAG. This equivalence continues to hold taking into account the arc annotations. Call the DAG we have constructed the implication DAG and denote it by $D$.
Now our construction is as follows: start with two vertices $s,t$ and for each $i\in[n]$ construct $2$ copies of the implication DAG, call them $D_{i,0},D_{i,1}$ for $i\in[n]$. For $i\in[n]$ put an arc from $s$ to the first vertex $x_{i}$ in $D_{i,0}$ from the last $\neg x_{i}$ in $D_{i,0}$ to the first $\neg x_{i}$ in $D_{i,1}$ and from the last $x_{i}$ in $D_{i,1}$ to $t$. We set $M=B$ and this completes the construction which can clearly be carried out in polynomial time. What remains is to argue for correctness.
Suppose now that the original CNF is satisfiable. We want to show that by using the same assignment to $M$ we eliminate all $s\to t$ paths. Indeed, suppose we find such a path, which for some $i\in[n]$ must go through $D_{i,0},D_{i,1}$. In particular, this means we have a path from $x_{i}$ to $\neg x_{i}$ and a path from $\neg x_{i}$ to $x_{i}$ in $D$ while respecting the annotations. However, recall that for each clause we have added a path of annotated arcs containing the negations of all backdoor literals of the clause, meaning that if we can follow an arc, the assignment to $M$ must be setting all the backdoor literals to False. We conclude that the assignment cannot be satisfying to the formula, as we can infer the implications $(x_{i}\to\neg x_{i})$ and ($(\neg x_{i}\to x_{i})$.
For the converse direction, suppose there is an assignment to $B$ which eliminates all $s\to t$ paths. We claim the same assignment can be extended to a satisfying assignment of the original formula. Suppose for contradiction that this is not the case and suppose that we are working with a maximal counter-example, that is, for the given $n,b$ we have started our reduction from a formula $\phi$ that satisfies the following: (i) $\phi$ is unsatisfiable (ii) there is an assignment to $B$ that eliminates all $s\to t$ paths (iii) every other $\phi^{\prime}$ with the same backdoor size $b$ that satisfies (i) and (ii) either has $n^{\prime}>n$ non-backdoor variables or the Ann-s-t-Non-Reach instance constructed for $\phi^{\prime}$ has at most as many arcs as that for $\phi$. In other words, $\phi$ is the counter-example with smallest $n$ and among these has the maximum number of arcs in the resulting Ann-s-t-Non-Reach instance. Assuming these we will reach a contradiction, which will show that a counter-example cannot exist.
We will attempt to find a good assignment to $\phi$. Consider an $i\in[n]$ and since there is no path from $s$ to $t$ it must be the case that either there is no path from (the first) $x_{i}$ to (the last) $\neg x_{i}$ in $D$, or that there is no path from (the first) $\neg x_{i}$ to (the last) $x_{i}$ in $D$, or both.
Suppose that there is no path from $x_{i}$ to $\neg x_{i}$, but that there is a path from $\neg x_{i}$ to $x_{i}$, for some $i\in[n]$. We then set $x_{i}$ to True. Symmetrically, suppose that there is no path from $\neg x_{i}$ to $x_{i}$, but that there is a path from $x_{i}$ to $\neg x_{i}$. Then we set $x_{i}$ to False. Do this for all variables that satisfy one of these conditions.
We claim that this assignment so far satisfies all clauses containing a variable that has been set. Indeed, suppose $x_{i}$ has been set to True but there is a clause such that all its backdoor literals are false and its other non-backdoor literal is $\ell$. The interesting case is if $C$ contains $x_{i}$ negated (otherwise $C$ is satisfied by $x_{i}$), so we have the directed paths $x_{i}\to\ell$ and $\neg\ell\to\neg x_{i}$, which can be followed under the current assignment to $M$. Then, there is a path $\neg\ell\to\neg x_{i}\to x_{i}\to\ell$, which implies that the variable involved in $\ell$ has been set to a value that satisfies $C$. The reasoning is the same if $x_{i}$ has been set to False.
At this point we have an assignment to $B$ and an assignment to $X^{\prime}\subseteq X$ which together satisfy all clauses involving variables of $X^{\prime}$. If $X^{\prime}\neq\emptyset$ then we are done, because we can give the assignment we have calculated to $X^{\prime}$ and remove from $\phi$ all clauses containing a variable from $X^{\prime}$. The new formula $\phi^{\prime}$ would be a smaller counter-example, contradicting the selection of $\phi$. Indeed, the Ann-s-t-Non-Reach instance for $\phi^{\prime}$ is a sub-graph of the one for $\phi$, so it cannot contain any $s\to t$ path, while $\phi^{\prime}$ must be unsatisfiable, because if it were satisfiable, using the assignment we calculated for $X^{\prime}$ would satisfy all of $\phi$.
We then must have $X^{\prime}=\emptyset$. This implies that for all $i\in[n]$ we have neither a path from $x_{i}$ to $\neg x_{i}$ nor from $\neg x_{i}$ to $x_{i}$ in $D$. We claim that setting any $x_{i}$ to True and simplifying the formula also yields a smaller counter-example. Indeed, it will be more convenient to do this by adding the clause $(x_{i}\lor x_{i})$ to $\phi$, obtaining an Ann-s-t-Non-Reach instance with more arcs (this clause cannot have already existed, as then we would have had $X^{\prime}\neq\emptyset$). The new formula is clearly also unsatisfiable (we added a clause to $\phi$ which was assumed unsatisfiable). Furthermore, there is still no $s\to t$ path in the Ann-s-t-Non-Reach instance for $\phi^{\prime}$. To see this, suppose that there is now such a path $s\to x_{j}\to\neg x_{j}\to\neg x_{j}\to x_{j}\to t$. Previously, neither the $x_{j}\to\neg x_{j}$ path (in $D_{j,0}$) nor the $\neg x_{j}\to x_{j}$ path (in $D_{j,1}$) existed, so both paths must be using the arc $\neg x_{i}\to x_{i}$ which is the only newly added arc to $D$. So we already had paths $x_{i}\to\neg x_{j}$ and $\neg x_{j}\to\neg x_{i}$, which means we already had a path $x_{i}\to\neg x_{i}$, contradicting the assumption that $X^{\prime}=\emptyset$. We conclude that $\phi^{\prime}$ is a counter-example with the same number of variables but strictly more arcs in the corresponding Ann-s-t-Non-Reach instance, contradicting the selection of $\phi$. Hence, no counter-example exists.
5.2 Log-Pathwidth and Reachability
Lemma 5.10.
[…]
Proof 5.11.
For one direction, assume that there is an algorithm solving Ann-s-t-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$ and we are given a CNF formula $\phi$ and a modulator $M$ of size $m$ such that removing $M$ from $\phi$ results in a graph of pathwidth $O(\log|\phi|)$. Suppose we are given a path decomposition of the resulting graph, and the bags are numbered $B_{1},\ldots,B_{r}$. Recall that each clause of $\phi$ induces a clique on the primal graph and by standard properties of tree decompositions for each clause $C$ there is a bag that contains all variables of $C$ (outside of $M$). Repeating bags as necessary we assume that there is an injective mapping from the clauses to the bags so that each clause is mapped to a bag that contains all its variables outside of $M$. Note that if a clause contains only variables of $M$, then it can be mapped to any bag.
We construct an Ann-s-t-Reach instance as follows. For each $i\in[r]$ and for each assignment $\sigma$ to the variables of $B_{i}$ we construct two vertices $v_{i,\sigma,1},v_{i,\sigma,2}$. For each $i\in[r-1]$, for each assignment $\sigma$ to $B_{i}$ and $\sigma^{\prime}$ to $B_{i+1}$ such that $\sigma,\sigma^{\prime}$ agree on $B_{i}\cap B_{i+1}$, we add an arc $(v_{i,\sigma,2},v_{i+1,\sigma^{\prime},1})$ without annotation.
For each $i\in[r]$ and assignment $\sigma$ to $B_{i}$ we do the following: if $\sigma$ satisfies the (at most one) clause $C$ mapped to $B_{i}$ or if no clause is mapped to $B_{i}$, then add an arc $(v_{i,\sigma,1},v_{i,\sigma,2})$ without annotation. If there is a clause $C$ mapped to $B_{i}$ but $C$ is not satisfied by $\sigma$, we add an arc $(v_{i,\sigma,1},v_{i,\sigma,2})$ annotated with each literal of $C$ from $M$. In particular, if $C$ contains no literals from $M$ we do not add an arc $(v_{i,\sigma,1},v_{i,\sigma,2})$, while if it contains several, we add several parallel arcs (we can then subdivide them to avoid constructing a multi-graph), each annotated with a distinct literal. We add a vertex $s$ and arcs $(s,v_{1,\sigma,1})$ for all $\sigma$, as well as a vertex $t$ and arcs $(v_{r,\sigma,2},t)$. We retain the set $M$ between the two instances.
It is now not hard to see that there is a one-to-one correspondence between satisfying assignments to $\phi$ and $s\to t$ paths. Fix an assignment to $M$. If there is an $s\to t$ path, it must use for each $i$ an arc $(v_{i,\sigma,1},v_{i,\sigma,2})$, which allows us to infer an assignment $\sigma$ for $B_{i}$. This assignment is consistent because we only added arcs between bags for consistent pairs of assignments. It is satisfying because if a clause $C$ is mapped to $B_{i}$ we can only use the arc $(v_{i,\sigma,1},v_{i,\sigma,2})$ if $\sigma$ satisfies the clause or if the assignment to $M$ agrees with one of the literals of the clause; furthermore, all clauses are mapped to some bag. For the converse direction, if there is an assignment we construct a path by using in each bag the arc $(v_{i,\sigma,1},v_{i,\sigma,2})$ corresponding to the assignment $\sigma$ used for $B_{i}$.
Let us now move to the other half of the lemma. Suppose there is an algorithm deciding satisfiability with a modulator $M$ to pathwidth $O(\log|\phi|)$ and we want to use this algorithm to solve Ann-s-t-Reach. We construct a formula which retains the set $M$ of annotation variables as the modulator. Given a digraph $G=(V,E)$ with $n$ vertices we assume without loss of generality that $n$ is a power of $2$ (otherwise add some dummy vertices). We construct a formula with $n\log n$ extra variables, called $x_{i,j}$, $i\in[n],j\in[\log n]$. The intended meaning is that the $\log n$ variables $x_{i,1},\ldots,x_{i,\log n}$ encode the index of the $i$-th vertex we would visit in an $s\to t$ path. To enforce this meaning we assume the vertices of $G$ are numbered from 00 to $n-1$ and with each vertex we associate a distinct truth assignment to $\log n$ variables by reading these variables as a binary number. We add clauses enforcing the variables $x_{1,j},j\in[\log n]$ to take the value corresponding to $s$ and the variables $x_{n,j},j\in[\log n]$ to take the value corresponding to $t$. For each $i\in[n-1]$ we add at most $n^{2}$ clauses to the formula as follows: for each $(u,v)\not\in E$ such that $u\neq v$, we add a clause ensuring that if the $x_{i,j}$ variables encode $u$ and the $x_{i+1,j}$ variables encode $v$, then the formula is not satisfied; for each $(u,v)\in E$ which is annotated with a literal $\ell$ we add the same clause (which is falsified if $x_{i,j},x_{i+1,j}$ encode $u,v$), but we add to this clause the literal $\ell$. If $(u,v)\in E$ is not annotated or if $u=v$, we do not add a clause. It is not hard to see that, if we remove the variables of $M$, this formula has pathwidth $O(\log n)$, because we can make $n-1$ bags, each containing $x_{i,j},x_{i+1,j},j\in[\log n]$ and this covers all edges.
To argue for correctness, if there is an $s\to t$ path in the original instance for some assignment to $M$ we use the same assignment in $\phi$ and extend it to a satisfying assignment as follows: if the path visits $r\leq n$ vertices in total, we set for each $i\in[r]$ the assignment to $x_{i,j},j\in[\log n]$ to encode the $i$-th vertex of the path; all remaining groups encode $t$. This satisfies the formula. For the converse direction, we can infer from a satisfying assignment to the formula a path of length at most $n-1$ by interpreting the assignment to $x_{i,j},j\in[\log n]$ as the encoding of the $i$-th vertex of the path, and the clauses we have added ensure that the path is valid, that is we are only followed to stay in the same vertex, follow an arc without annotation, or follow an arc whose annotation is set to True by the assignment to $M$.
5.3 Reachability and Non-Reachability
Lemma 5.12.
For all $\varepsilon>0$, there is an algorithm solving Ann-s-t-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$ if and only if there is an algorithm solving Ann-s-t-Non-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$.
Proof 5.13.
We describe a polynomial-time algorithm which takes as input a DAG $G_{1}=(V_{1},E_{1})$, with $|V_{1}|=n$, a source $s$, a sink $t$, with the arcs annotated with variables from a set $M$. The algorithm produces a new DAG $G_{2}=(V_{2},E_{2})$, a source $s_{2}$, a sink $t_{2}$, with arcs annotated from the same set $M$. It guarantees the following: for each assignment to $M$, $G_{2}$ has an $s_{2}\to t_{2}$ path respecting the assignment if and only if $G_{1}$ does not have any $s\to t$ path respecting the assignment. Observe that if we achieve this, we immediately obtain reductions from Ann-s-t-Reach to Ann-s-t-Non-Reach in both directions, implying the lemma.
We will reuse the ideas of the famous Immerman–Szelepcsényi theorem, so the reader may think of $G_{2}$ as a graph that encodes the working of a non-deterministic machine which is verifying that $G_{1}$ has no $s\to t$ path. The construction is of course made more complicated because we also need to take into account the arc annotations of $G_{1}$.
Before we begin, let us slightly edit $G_{1}$. Take $n=|V_{1}|$ copies of $V_{1}$, constructing for each $v\in V_{1}$ the vertices $v^{1},\ldots,v^{n}$. For each $i\in[n-1]$ and $v\in V_{1}$ we add the arc $(v^{i},v^{i+1})$ without annotation; for each arc $(u,v)$ and each $i\in[n-1]$ we add the arc $(u^{i},v^{i+1})$, maintaining the annotation of $(u,v)$. Finally, for the vertex $s^{n-1}$ and all $v\in V_{i}\setminus\{t\}$ we add the non-annotated arc $(s^{n-1},v^{n})$. This transformation has achieved the following: for each assignment to $M$, an $s\to t$ path exists in $G_{1}$ if and only if an $s^{1}\to t^{n}$ path exists in the new graph; there is a path from $s_{1}$ to $v^{n}$ for all $v\in V_{1}\setminus\{t\}$. As a result the question of whether an $s\to t$ path exists in $G_{1}$ is equivalent to asking whether the number of vertices reachable from $s^{1}$ in $\{v^{n}\ |\ v\in V_{1}\}$ is equal to $n$. Call the new graph $G_{1}^{\prime}$ and we want to construct $G_{2}$ such that for all assignments to $M$, there is a path $s_{2}\to t_{2}$ in $G_{2}$ if and only if $n-1$ vertices of $\{v^{n}\ |\ v\in V_{1}\}$ are reachable from $s^{1}$ in $G_{1}^{\prime}$. In the remainder, when we refer to the $i$-th layer of $G_{1}^{\prime}$ we will mean the set $\{v^{i}\ |v\in V_{1}\}$. Observe that $G_{1}^{\prime}$ has the property that all arcs go from a layer $i$ to a layer $i+1$.
We begin the construction of $G_{2}$ with $n^{2}$ vertices $s_{i,j}$ for $i,j\in[n]$, plus a source $s_{2}$ and a sink $t_{2}$. The intended meaning is that (for all assignments to $M$) $s_{i,j}$ will be reachable from $s_{2}$ if and only if there are exactly $j$ vertices of $G_{1}^{\prime}$ in layer $i$ which are reachable from $s^{1}$. We add a non-annotated arc $(s_{n,n-1},t_{2})$ and a non-annotated arc $(s_{2},s_{1,1})$. For each $i,j\in[n]$ we want to construct a checker gadget $C_{i,j}$ such that there will be an arc from $s_{i,j}$ to $C_{i,j}$ and arcs from $C_{i,j}$ to $s_{i+1,j^{\prime}}$ for some $j^{\prime}\in[n]$. We want to achieve the following property for all assignments $\sigma$ to $M$. Suppose that under some assignment $\sigma$, there are exactly $j$ vertices of layer $i$ of $G_{1}^{\prime}$ reachable from $s^{1}$. Then, there will be a path from $s_{i,j}$ to $s_{i+1,j^{\prime}}$ through $C_{i,j}$ if and only if the number of vertices of layer $i+1$ of $G_{1}^{\prime}$ reachable from $s^{1}$ is $j^{\prime}$. If we achieve this property we are done, as any path starts from $s_{1,1}$ and it is the case that exactly $1$ vertex of layer $1$ is reachable from $s^{1}$ ($s^{1}$ itself), so we can proceed by induction and we see that there is a path to $s_{n,n-1}$ (hence to $t_{2}$) if and only if $n-1$ vertices of the last layer of $G_{1}^{\prime}$ are reachable. Hence, there will be a $s_{2}\to t_{2}$ path in $G_{2}$ if and only if there is no $s_{1}\to t_{1}$ path in $G_{1}$.
Let us now describe the construction of the checker $C_{i,j}$. For each $\alpha,\beta\in\{0,\ldots,n\}$ we construct a vertex $x_{i,j,\alpha,\beta}$. We place an incoming non-annotated arc to $x_{i,j,0,0}$ from $s_{i,j}$. The intended meaning of these vertices is the following: assuming that $j$ is correct (that is, there are exactly $j$ vertices of layer $i$ reachable from $s^{1}$ in $G_{1}^{\prime}$), there should exist a path from $s_{i,j}$ to $x_{i,j,\alpha,\beta}$ if and only if among the $\alpha$ first vertices of layer $i+1$ of $G_{1}^{\prime}$ there are exactly $\beta$ which are reachable from $s^{1}$. Clearly, the property holds for $x_{i,j,0,0}$. Given this intended meaning we place a non-annotated arc from $x_{i,j,n,\beta}$ to $s_{i+1,\beta}$ for all $\beta\in[n]$. Now if we establish that the vertices $x_{i,j,\alpha,\beta}$ satisfy the property we stated, then the checker gadget functions correctly. What remains then is to add some gadgets that imply this property.
For each $\alpha\in\{0,\ldots,n-1\}$ and $\beta\in\{0,\ldots,n-1\}$ we will construct a gadget called $\mathrm{Adj}_{i,j,\alpha,\beta}$ and a gadget called $\mathrm{NonAdj}_{i,j,\alpha,\beta}$, each with a single input and output vertex. The intended meaning is that a path should exist from the input to the output of the $\mathrm{Adj}$ gadget if and only if the $(\alpha+1)$-th vertex of layer $i+1$ of $G_{1}^{\prime}$ is reachable from $s^{1}$; such a path should exist in the $\mathrm{NonAdj}$ gadget if and only if the same vertex is not reachable from $s^{1}$. Given this meaning we place a non-annotated arc from $x_{i,j,\alpha,\beta}$ to the input of $Adj_{i,j,\alpha,\beta}$ and from the output of that gadget to $x_{i,j,\alpha+1,\beta+1}$; and an annotated arc from $x_{i,j,\alpha,\beta}$ to the input of $NonAdj_{i,j,\alpha,\beta}$ and from the output of that gadget to $x_{i,j,\alpha+1,\beta}$. What remains is to describe these gadgets and prove they work as expected.
The $\mathrm{Adj}_{i,j,\alpha,\beta}$ gadget is conceptually simple. Suppose that $v$ is the $(\alpha+1)$-th vertex of $V_{1}$ and we are testing if $v^{i+1}$ is reachable from $s^{1}$ in $G_{1}^{\prime}$. The gadget is simply a copy of the first $i+1$ layers of $G_{1}^{\prime}$, with input vertex $s^{1}$ and output vertex $v^{i+1}$. Clearly, there is a path from the input to the output of the gadget if and only if $v^{i+1}$ is reachable from $s^{1}$ in $G_{1}^{\prime}$.
The $\mathrm{NonAdj}_{i,j,\alpha,\beta}$ gadget is more complicated as it relies on the key insight of the Immerman–Szelepcsényi theorem which is that we can certify that a vertex $v$ is not reachable from $s$ at distance $i+1$ if we know how many vertices are reachable at distance $i$, because we can list all such vertices, verify that their cardinality is correct, and verify that none of them have an arc to $v$. For each $\gamma\in\{0,\ldots,n\}$ and $\delta\in\{0,\ldots,j\}$ we construct a vertex $y_{i,j,\alpha,\beta,\gamma,\delta}$, whose intended role is the following: assuming $j$ is correct, there is a path from the input of the gadget to $y_{i,j,\alpha,\beta,\gamma,\delta}$ if and only if among the first $\gamma$ vertices of the $i$-th layer of $G_{i}$ there are at least $\delta$ vertices which are reachable from $s^{1}$ but which do not have an arc that leads to $v$ under the current assignment, where again $v$ is the $(\alpha+1)$-th vertex of $V_{1}$. The input vertex of the gadget is $y_{i,j,\alpha,\beta,0,0}$ and the output is $y_{i,j,\alpha,\beta,n,j}$. For all $\gamma\in\{0,\ldots,n-1\}$ and $\delta\in\{0,\ldots,j\}$ we add a non-annotated arc from $y_{i,j,\alpha,\beta,\gamma,\delta}$ to $y_{i,j,\alpha,\beta,\gamma+1,\delta}$. The intended meaning is that if among the first $\gamma$ vertices of the $i$-th layer there are $\delta$ satisfying the property above, then there certainly are at least $\delta$ vertices satisfying this property among the first $\gamma+1$ vertices. What remains is to add a gadget so that we can transition from $y_{i,j,\alpha,\beta,\gamma,\delta}$ to $y_{i,j,\alpha,\beta,\gamma+1,\delta+1}$ if and only if the $(\gamma+1)$-th vertex for the $i$-th layer, call it $u^{i}$, satisfies the property. If there is a non-annotated arc from $u^{i}$ to $v^{i+1}$, we do nothing. If there is no such arc or the arc is annotated, we need to check if $u^{i}$ is reachable from $s^{1}$. We therefore add an $\mathrm{Adj}$ gadget checking this, and add a non-annotated arc from $y_{i,j,\alpha,\beta,\gamma,\delta}$ to the input of that gadget. From the output of that gadget we add an arc to $y_{i,j,\alpha,\beta,\gamma+1,\delta+1}$ annotated as follows: if $(u^{i},v^{i+1})$ is not an arc of $G_{1}^{\prime}$, then we add no annotation; otherwise $(u^{i},v^{i+1})$ is annotated with a literal $\ell$, and we use the annotation $\neg\ell$ for the new arc. This completes the construction.
Correctness now follows from the description of the gadgets. It is worth noting that a vertex $u^{i}$ can be used as part of the certificate that $v^{i+1}$ is not reachable only when the (possible) arc $(u^{i},v^{i+1})$ cannot be used either because it does not exist or because of its annotation. Since we have added the same annotation negated to the output arc of the $\mathrm{NonAdj}$ gadget, this correctly checks if $u^{i}$ is a valid part of the certificate that $v^{i+1}$ is not reachable. The rest of the proof follows along the lines of the classical Immerman–Szelepcsényi theorem [24].
6 Circuit-SAT and Horn Backdoors
Our goal in this section is to prove the following theorem:
Theorem 6.1.
The following statements are equivalent:
  1. 1.
    (Circuit-SETH is false) There exists $\varepsilon>0$ and an algorithm that takes as input a Boolean circuit of size $s$ on $n$ inputs and decides if the circuit is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
  2. 2.
    (HornB-SETH is false) There exists $\varepsilon>0$ and an algorithm that takes as input a CNF formula $\phi$ together with a strong Horn backdoor of size $b$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{b}|\phi|^{O(1)}$.
  3. 3.
    (W[P]-SETH is false, weak version) There exist $\varepsilon>0,k_{0}>0$ and an algorithm which for all $k>k_{0}$ takes as input a monotone Boolean circuit of size $s$ and decides if the circuit can be satisfied by setting at most $k$ inputs to True in time $O(s^{(1-\varepsilon)k})$.
  4. 4.
    (W[P]-SETH is false, strong version) There exists $\varepsilon>0$ and an algorithm which takes as input a Boolean circuit of size $s$ on $n$ inputs and an integer $k$ and decides if the circuit can be satisfied by setting exactly $k$ of its inputs to True in time $n^{(1-\varepsilon)k}s^{O(1)}$.
Before we go on let us give some intuition. Similarly to Section 5, it will be helpful to recall some basic facts from computational complexity theory. For the first two statements, we are dealing with the satisfiability problem for circuits and SAT with a Horn backdoor. Intuitively, it is natural to expect these two problems to be equivalent because once we fix either the input to a given circuit or the assignment to the Horn backdoor, deciding if the circuit is satisfied or if the backdoor assignment can be extended to a satisfying assignment are problems with the same classical complexity, that is, they are P-complete [39]. As in Section 5 this intuition is not in itself sufficient to imply the equivalence, because what we want is not to claim that once we have an assignment, checking it is of the same complexity in both cases, but rather that if we can avoid checking all assignments in one case, we can avoid it in both. Nevertheless, the classical results on the P-completeness of these two problems are indeed the main ingredient we need, and our proof mostly consists of verifying that we can extend them to the new setting. We then go on to consider the prototypical W[P]-complete problem of the third statement. It is natural to expect that since the first and third statement refer to circuit satisfiability problems, without any particular restriction on the circuits, it should be possible to transform one to the other, and indeed we are able to do so using standard techniques, similar to what we did in Theorem 4.1.
Proof 6.2 (Proof of Theorem 6.1).
The proof is given in the four lemmas below which show the implications $1\Rightarrow 2$ (Lemma 6.3), $2\Rightarrow 1$ (Lemma 6.5), $3\Rightarrow 1$ (Lemma 6.7), $1\Rightarrow 4$ (Lemma 6.9).
We note that the implication $4\Rightarrow 3$ is easy to see. First, statement 3 talks about monotone circuits, while statement 4 about general circuits. Second, if we have an algorithm running in time $n^{(1-\varepsilon)k}s^{O(1)}$ we can see that $n\leq s$ (as the input gates are part of the circuit) so the running time is at most $O(s^{(1-\varepsilon)k+c})$ for some constant $c$. We can now set $k_{0}=\frac{2c}{\varepsilon}$ and we then have $\varepsilon k>\frac{\varepsilon k}{2}+c$ if $k>k_{0}$, therefore the running time is at most $O(s^{(1-\frac{\varepsilon}{2})k})$.
6.1 Circuit-SAT and Horn Backdoors
Lemma 6.3.
Statement 1 of Theorem 6.1 implies statement 2.
Proof 6.4.
Suppose that we have an algorithm solving Circuit SAT in time $(2-\varepsilon)^{n}s^{O(1)}$, for some fixed $\varepsilon>0$. We are given a CNF formula $\phi$ and a strong Horn backdoor of size $b$ of $\phi$. Our plan is to construct a Boolean circuit $C$ satisfying the following:
  1. 1.
    $C$ is satisfiable if and only if $\phi$ is.
  2. 2.
    $C$ has $b$ inputs and can be constructed in time $|\phi|^{O(1)}$.
Clearly, if $C$ can be constructed in time polynomial in $|\phi|$, then its size is also polynomial in $|\phi|$, so the supposed algorithm for Circuit SAT gives us the promised algorithm for SAT with a Horn backdoor.
The high-level idea now is to implement a circuit which takes as input the values of the $b$ backdoor variables of $\phi$ and then runs the standard polynomial-time algorithm to check satisfiability of Horn formulas. Recall that this algorithm starts with the all-False assignment to all (non-backdoor) variables, and then iterates, at each step setting a variable to True if this is forced. This will result in either a satisfying assignment or a contradiction after at most $n$ iterations, where $n$ is the number of (non-backdoor) variables.
Suppose that the variables of $\phi$ that belong to the backdoor are $y_{1},\ldots,y_{b}$ and the remaining variables are $x_{1},\ldots,x_{n}$. Assume $\phi$ has $m$ clauses, numbered $c_{1},\ldots,c_{m}$.
We construct a circuit with the following gates:
  1. 1.
    $b$ input gates, identified with the variables $y_{i},i\in[b]$.
  2. 2.
    For each non-backdoor variable $x_{i},i\in[n]$, we construct $n+2$ disjunction ($\lor$) gates $x_{i,t},t\in\{0,\ldots,n+1\}$. The intuitive meaning is that $x_{i,t}$ should evaluate to True if the Horn satisfiability algorithm would set variable $x_{i}$ to True after $t$ iterations of the main loop. One of the inputs of each $x_{i,t}$ is the gate $x_{i,t-1}$, for all $i,t\in[n+1]$. We specify the remaining inputs below.
  3. 3.
    For each clause $c_{j},j\in[m]$, we construct a disjunction ($\lor$) gate using the variables $x_{i,n+1}$ and the input gates as inputs. In particular, if $x_{i}$ appears positive in $c_{j}$ we connect $x_{i,n+1}$ to the gate representing $c_{j}$, and if it appears negative, we connect $x_{i,n+1}$ to a negation ($\neg$) gate whose output is sent to the gate representing $c_{j}$, and similarly for input gates. We route the output of all gates representing clauses to a $\land$ gate which is the output of the whole circuit. Intuitively, the circuit contains a copy of $\phi$, with the $x_{i}$ variables replaced by the gates $x_{i,n+1}$.
To complete the construction we describe the gates we need to add so that the $x_{i,t}$ gates play their intuitive role. For each variable $x_{i},i\in[n]$, for each $t\in[n+1]$ and each clause $c_{j},j\in[m]$, such that $x_{i}$ appears positive in $c_{j}$ we add a conjunction ($\land$) gate $c_{i,j,t}$ whose output is sent to $x_{i,t}$. The inputs to $c_{i,j,t}$ are all gates $x_{i^{\prime},t-1}$ such that $x_{i^{\prime}}$ appears (negative) in $c_{j}$, all gates $y_{i^{\prime\prime}}$ such that $y_{i^{\prime\prime}}$ appears negative in $c_{j}$, and the negation of each gate $y_{i^{\prime\prime}}$, such that $y_{i^{\prime\prime}}$ appears positive in $c_{j}$. Finally, set all $x_{i,0}$ gates, for $i\in[n]$ to False and simplify the circuit accordingly. This completes the construction, which can clearly be carried out in polynomial time. We have described a circuit with unbounded fan-in, but it is easy to transform it into a circuit of bounded fan-in without significantly increasing its size, in polynomial time. Note also that we have used the fact that if $x_{i}$ appears positive in $c_{j}$ every other non-backdoor variable $x_{i^{\prime}}$ must appear negative in $c_{j}$; this follows from the fact that if two variables appear positive in a clause, at least one of them must be in the backdoor [37].
It is easy to see that if the circuit is satisfiable then so is the formula, because we can extract an assignment from the values of the input gates and the gates $x_{i,n+1}$. This assignment satisfies the formula since we have added a copy of $\phi$ at the end of the circuit, with inputs the $y_{i}$ and $x_{i,n+1}$ gates.
For the converse direction, we first make some easy observations: for each $i\in[n]$ and $t\in[n+1]$ we have that if $x_{i,t-1}$ evaluates to True, then $x_{i,t}$ evaluates to True, since $x_{i,t-1}$ is an input to $x_{i,t}$. As a result, as $t$ increases, the set of gates $x_{i,t}$ which evaluate to True can only monotonically increase. Furthermore, if for some $t\in[n+1]$ we have for all $i\in[n]$ that $x_{i,t-1}$ and $x_{i,t}$ evaluate to the same value, then for all $t^{\prime}>t$ and for all $i$ we have $x_{i,t^{\prime}}$ evaluates to the same value as $x_{i,t}$. This is because the $c_{i,j,t}$ gates and the $c_{i,j,t+1}$ gates will evaluate to the same values, forcing $x_{i,t+1}$ to have the same value as $x_{i,t}$, and we can proceed by induction. It now follows that $x_{i,n}$ evaluates to the same value as $x_{i,n+1}$ for all $i\in[n]$. Indeed, there are $n+2$ possible values of $t$ (time steps), at each step the set of True gates can only monotonically increase, so we have at most $n+1$ distinct sets of True gates. As a result, by pigeonhole principle, there exists $t$ such that $x_{i,t}$ evaluates to the same value as $x_{i,t+1}$ for all $i\in[n]$, which implies the same for $t=n$.
Suppose now that $\phi$ is satisfiable and fix a satisfying assignment. Using the same assignment restricted to the backdoor variables as input to our circuit we claim that we obtain a satisfying assignment to the circuit. To see this, we claim that if for some $t\in[n+1]$ we have that $x_{i,t}$ evaluates to True, then the satisfying assignment to $\phi$ must also set $x_{i}$ to True. For the sake of contradiction, suppose that $t$ is minimum such that $x_{i,t}$ evaluates to True but the satisfying assignment sets $x_{i}$ to False. Since $x_{i,t}$ is True and $t$ is minimum, one of the $c_{i,j,t}$ gates must evaluate to True. But this means that the satisfying assignment sets all literals of the clause $c_{j}$ to False, except $x_{i}$, because by the minimality of $t$ we know that for each $x_{i^{\prime},t-1}$ which evaluates to True, $x_{i^{\prime}}$ is True in the satisfying assignment. Hence, the satisfying assignment must set $x_{i}$ to True.
We can now conclude that if $x_{i,n+1}$ evaluates to True, then the satisfying assignment to $\phi$ sets $x_{i}$ to True. Suppose then for contradiction that the circuit outputs False, which implies that there is a gate added in the third step, representing $c_{j}$, which evaluates to False. It must be the case that $c_{j}$ has as input an $x_{i,n+1}$ which evaluates to False while $x_{i}$ is set to True in the satisfying assignment, because in all other cases the inputs to $c_{j}$ would be the same as those dictated by the satisfying assignment to $\phi$ and $c_{j}$ would evaluate to True. Furthermore, $x_{i}$ must appear positive in $c_{j}$ and is therefore the only non-backdoor variable which appears positive in $c_{j}$. We now claim that the gate $c_{i,j,n+1}$ must evaluate to True, reaching a contradiction. Indeed, because $x_{i^{\prime},n}$ and $x_{i^{\prime},n+1}$ evaluate to the same value for all $i^{\prime}\in[n]$, all non-backdoor variables of $c_{j}$ except $x_{i}$ (which all appear negative in $c_{j}$) are set to True in $x_{i,n+1}$, hence in $x_{i,n}$, and backdoor variables are shared between the gates $c_{j},c_{i,j,n+1}$, so all inputs to $c_{i,j,n+1}$ evaluate to True, contradiction. Hence, we have an assignment satisfying the circuit.
Lemma 6.5.
Statement 2 of Theorem 6.1 implies statement 1.
Proof 6.6.
Suppose we have an algorithm solving SAT, given a Horn backdoor of size $b$, in time $(2-\varepsilon)^{b}|\phi|^{O(1)}$. We are now given a circuit $C$ with $n$ inputs and size $s$. Our plan is to construct from $C$ a CNF formula $\phi$ and a Horn backdoor such that:
  • $\phi$ is satisfiable if and only if $C$ is.
  • The backdoor has size $n$ and $\phi$ can be constructed in time $s^{O(1)}$.
Clearly, if $\phi$ can be constructed in time $s^{O(1)}$, it also has size polynomial in $s$, so the supposed algorithm for SAT with a Horn backdoor would give us an algorithm for Circuit-SAT.
The high-level idea is now to revisit the standard proof that Horn-SAT is P-complete, which relies on a reduction from the Circuit Value Problem. More precisely, assume that the given circuit is represented by a DAG that contains only $\land$,$\lor$ , ∨, and $\neg$ gates of maximum fan-in $2$. We first transform the circuit (in polynomial time) so that $\neg$ gates only appear directly connected to the inputs. Since the circuit is a DAG, we can assume that the gates are topologically sorted. We construct $\phi$ as follows:
  1. 1.
    For each input gate of $C$ we construct a variable $x_{i}$, $i\in[n]$. These variables will form the Horn backdoor.
  2. 2.
    For each input, $\land$, and $\lor$ gate $c$ of $C$, we construct two variables: $c^{p}$ and $c^{n}$. Intuitively, in circuit assignments that set $c$ to True, $c^{p}$ must be set to True and $c^{n}$ to False, while the opposite must happen for circuit assignments that set $c$ to False. Thus, for each such gate we add the clause ($(\neg c^{p}\lor\neg c^{n})$.
  3. 3.
    For each input gate $c$ corresponding to variable $x_{i}$ we add the clauses $(x_{i}\leftrightarrow c^{p})$ and $(x_{i}\leftrightarrow\neg c^{n})$.
  4. 4.
    For each negation gate $c$, whose input corresponds to the variable $x_{i}$, we add the clauses $(x_{i}\leftrightarrow\neg c^{p})$ and $(x_{i}\leftrightarrow c^{n})$.
  5. 5.
    For each conjunction gate $c$, whose inputs are gates $c_{1},c_{2}$ we add the clauses ($(\neg c_{1}^{p}\lor\neg c_{2}^{p}\lor c^{p}),(\neg c_{1}^{n}\lor c^{n}),(\neg c_{2}^{n}\lor c^{n}),(\neg c^{p}\lor c_{1}^{p}),(\neg c^{p}\lor c_{2}^{p})$
  6. 6.
    For each disjunction gate $c$, whose inputs are gates $c_{1},c_{2}$ we add the clauses ($(\neg c_{1}^{n}\lor\neg c_{2}^{n}\lor c^{n}),(\neg c_{1}^{p}\lor c^{p}),(\neg c_{2}^{p}\lor c^{p}),(\neg c^{n}\lor c_{1}^{n}),(\neg c^{n}\lor c_{2}^{n})$
  7. 7.
    For the unique output gate $c$ of the circuit we add the unit clauses $c^{p}$ and $\neg c^{n}$.
The construction can be performed in polynomial time. Furthermore, all clauses except those of the third and fourth steps are Horn clauses. The clauses of the third and fourth steps all have size $2$ and contain a variable from the backdoor, so we have a valid strong Horn backdoor of size $n$, as promised. It is now not hard to see that if $C$ has a satisfying assignment, then we can obtain a satisfying assignment to $\phi$ by giving to the variables of the backdoor the values of the satisfying assignment to $C$ and then for each gate $c$ of the circuit setting $c^{p},c^{n}$ to True, False respectively if $c$ is set to True and to False, True respectively otherwise.
For the converse direction, suppose we have a satisfying assignment to $\phi$. We claim that using the assignment to the backdoor variables as input to $C$ satisfies the circuit. To see this, we prove by induction that for each gate $c$, exactly one of $c^{p},c^{n}$ must be set to True in the satisfying assignment, and $c^{p}$ is set to True if and only if gate $c$ has value True in the circuit $C$ for the given input assignment. Note that it is clear that at most one of $c^{p},c^{n}$ is set to True, due to the clauses of the second step. Furthermore, for the base case we observe that input and negation gates satisfy the property, due to the clauses of the third and fourth steps. Consider then a gate $c$ with inputs $c_{1},c_{2}$, which are assumed to satisfy the property. We have four cases:
  1. 1.
    $c$ is a conjunction that has value True in $C$. Then $c_{1},c_{2}$ must evaluate to True in $C$, since these gates satisfy the property we have $c_{1}^{p},c_{2}^{p}$ set to True, so the clause ($(\neg c_{1}^{p}\lor\neg c_{2}^{p}\lor c^{p})$ ensures that $c^{p}$ is True (and $c^{n}$ is False) as desired.
  2. 2.
    $c$ is a conjunction that has value False in $C$. Then one of $c_{1},c_{2}$ must evaluate to False in $C$, say without loss of generality $c_{1}$, and since these gates satisfy the property we have $c_{1}^{n}$ set to True. So, the clause ($(\neg c_{1}^{n}\lor c^{n})$ ensures that $c^{n}$ is True (and $c^{p}$ is False) as desired.
  3. 3.
    $c$ is a disjunction that has value True in $C$. Then one of $c_{1},c_{2}$ must evaluate to True in $C$, say without loss of generality $c_{1}$, and since these gates satisfy the property we have $c_{1}^{p}$ set to True. So, the clause ($(\neg c_{1}^{p}\lor c^{p})$ ensures that $c^{p}$ is True (and $c^{n}$ is False) as desired.
  4. 4.
    $c$ is a disjunction that has value False in $C$. Then $c_{1},c_{2}$ must evaluate to False in $C$, since these gates satisfy the property we have $c_{1}^{n},c_{2}^{n}$ set to True, so the clause ($(\neg c_{1}^{n}\lor\neg c_{2}^{n}\lor c^{n})$ ensures that $c^{n}$ is True (and $c^{p}$ is False) as desired.
Therefore, $\phi$ is satisfiable if and only if $C$ is.
6.2 Circuit-SAT and Weighted Circuit-SAT
Lemma 6.7.
Statement 3 of Theorem 6.1 implies statement 1.
Proof 6.8.
Suppose we have an algorithm which takes as input a monotone boolean circuit $C^{\prime}$ of size $S$ and $N$ inputs and an integer $k$ and decides if $C^{\prime}$ has a satisfying assignment setting $k$ inputs to True in time $O(S^{(1-\varepsilon)k})$, for some $\varepsilon>0$, assuming $k>k_{0}$ for some fixed $k_{0}$. We will use this supposed algorithm to solve Circuit-SAT. We are given a circuit $C$ with $n$ inputs and size $s$. We will assume that $s<2^{\frac{\varepsilon}{2k_{0}}n}$, because otherwise $2^{n}\leq s^{\frac{2k_{0}}{\varepsilon}}=s^{O(1)}$, so we immediately have a polynomial time algorithm for such cases by trying out all input assignments.
We construct a monotone circuit $C^{\prime}$ and an integer $k=k_{0}$ with the following properties:
  1. 1.
    $C^{\prime}$ has a satisfying assignment setting $k$ of its inputs to True if and only if $C$ is satisfiable.
  2. 2.
    $C^{\prime}$ can be constructed in time $O(2^{n/k}s)$.
  3. 3.
    $C^{\prime}$ has $N=O(2^{n/k})$ inputs and size $S=O(Ns)=O(2^{\frac{n}{k}+\frac{\varepsilon n}{2k}})$.
If we achieve the above, then we obtain the lemma as follows: we can decide satisfiability of $C$ by running the reduction and then the supposed algorithm on $C^{\prime}$. The running time is dominated by the second step which takes time $O(S^{(1-\varepsilon)k})=O(2^{(1-\varepsilon)(1+\frac{\varepsilon}{2})n})=O(2^{(1-\varepsilon^{\prime})n})$ for some appropriate $\varepsilon^{\prime}$, as desired.
Let us then describe the construction of $C^{\prime}$ in which we start from $C$ and add some gadgets. We assume that after running standard transformations $C$ only contains negation gates at the first layer, that is, every input of a negation is an input to $C$. Partition the inputs of $C$ arbitrarily into $k$ groups of (almost) equal size. Each group contains at most $\lceil\frac{n}{k}\rceil$ inputs of $C$. Consider a group $I_{i}$ of input gates of $C$, for $i\in[k]$ and enumerate all $2^{|I_{i}|}=O(2^{n/k})$ possible assignments to the inputs of $I_{i}$. For each such assignment $\sigma$ we construct a new input gate $g_{i,\sigma}$. We convert the gates of $I_{i}$ from input gates to disjunction ($\lor$) gates and for each assignment $\sigma$ we connect $g_{i,\sigma}$ to all gates of $I_{i}$ such that $\sigma$ assigns them value True. Similarly, for each negation gate whose input is in $I_{i}$ we convert it into a disjunction gate and for each $\sigma$ we connect $g_{i,\sigma}$ to all gates which were previously negations such that $\sigma$ was assigning to their input the value False. Since we have converted all negations to disjunctions the circuit is now monotone.
At this point it is not hard to see that the new circuit has $N=O(k2^{n/k})=O(2^{n/k})$ inputs (as $k$ is a fixed constant), its size is at most $O(Ns)$ (in the worst case we connected each new input to each old gate) and that a satisfying assignment to $C$ exists if and only if there exists a satisfying assignment to $C^{\prime}$ that sets exactly one new input gate from each of the $k$ groups to True and all others to False.
We now add some gadgets to ensure that a satisfying assignment of weight $k$ must set one input from each group to True. For each $i\in[k]$ add a disjunction ($\lor$) gate $t_{i}$ and connect all $g_{i,\sigma}$ to $t_{i}$. Add a conjunction gate that will be the new output of the circuit and feed it as input all $t_{i}$, for $i\in[k]$ as well as the output gate of $C$. This completes the construction. The total size is still $S=O(Ns)$ and the construction can be performed in linear time in the size of the output. We can also transform the circuit into an equivalent bounded fan-in circuit without increasing the size by more than a constant factor. With the extra gates we added in the last step we ensure that at least one input gate from each group is set to True, so since we are looking for an assignment of weight exactly $k$, exactly one input is set to True in each group.
Lemma 6.9.
Statement 1 of Theorem 6.1 implies statement 4.
Proof 6.10.
Suppose there is an algorithm that takes as input a boolean circuit $C^{\prime}$ of size $S$ with $N$ inputs and decides if $C^{\prime}$ is satisfiable in time $2^{(1-\varepsilon)N}S^{c}$, for some constants $\varepsilon>0,c>0$. We are given a circuit $C$ with $n$ inputs and size $s$ as well as an integer $k$ and we want to decide if $C$ has a satisfying assignment of weight exactly $k$ in time $n^{(1-\varepsilon^{\prime})k}s^{O(1)}$. We will do this by constructing a circuit $C^{\prime}$ with the following properties:
  1. 1.
    $C^{\prime}$ is satisfiable if and only if $C$ has a satisfying assignment of weight $k$.
  2. 2.
    $C^{\prime}$ has at most $N=k\log n+k$ input gates and can be constructed in time $O(sk^{2}n^{2})$.
If we achieve the above we obtain the desired algorithm because we can produce $C^{\prime}$ and run the supposed Circuit-SAT algorithm on it. The running time is dominated by the second step which takes $2^{(1-\varepsilon)N}S^{c}$. We have $2^{(1-\varepsilon)N}\leq n^{(1-\varepsilon)k}2^{k}$ and $S^{c}=O(s^{c}(kn)^{2c})$. Observe now that we can assume that $n$ and $k$ are sufficiently large. In particular, if $k<8c/\varepsilon$, then since without loss of generality $s>n$, the trivial $O(n^{k}s)$ time algorithm which checks all assignments of weight $k$ would run in time $s^{O(1)}$. But if $n,k$ are sufficiently large, then $2^{k}(kn)^{2c}<n^{\frac{\varepsilon k}{2}}$ and the total running time is at most $n^{(1-\frac{\varepsilon}{2})k}s^{O(1)}$ as desired.
We construct $C^{\prime}$ by adding some gadgets to $C$. First, number the input gates of $C$ as $c_{1},\ldots,c_{n}$. We transform them into disjunction gates. For each $i\in[n]$ we construct $k$ new gates $c_{i}^{1},\ldots,c_{i}^{k}$ and feed their output into $c_{i}$. The intuitive meaning is that $c_{i}^{j}$ will be True if $c_{i}$ is the $j$-th input set to True in a satisfying assignment of weight $k$. Each $c_{i}^{j}$, for $i\in[n],j\in[k]$ is a conjunction gate.
We now construct $k\lceil\log n\rceil\leq k\log n+k$ new input gates, partitioned into $k$ groups $I_{1},\ldots,I_{k}$, each containing $\lceil\log n\rceil$ input gates. The intuitive meaning is that the assignment to the gates of $I_{j}$ should encode the index of the $j$-th input of $C$ set to True in a satisfying assignment. For each $j\in[k]$ and each assignment $\sigma$ to the input gates of $I_{j}$ we interpret $\sigma$ as a binary number in $[n]$ and suppose that this number is $i$. We then connect the input gates of $I_{j}$ to $c_{i}^{j}$ via appropriate negations such that $c_{i}^{j}$ evaluates to True if and only if the gates of $I_{j}$ are given assignment $\sigma$.
At this point it is not hard to see that $C$ is satisfiable with an assignment of weight $k$ if and only if $C^{\prime}$ is satisfiable via an assignment which for each distinct $j,j^{\prime}\in[k]$ gives distinct assignments to $I_{j},I_{j^{\prime}}$. What remains is to add some gadgets so that $I_{j},I_{j^{\prime}}$ cannot have the same assignment, so that satisfying assignments to $C^{\prime}$ encode satisfying assignments of weight exactly $k$ to $C$. This is easily achieved by adding for each pair of distinct $j,j^{\prime}\in[k]$ a disjunction gate, which takes its input from $2\lceil\log n\rceil$ conjunction gates, where each conjunction gate checks if the assignments to $I_{j},I_{j^{\prime}}$ differ in a corresponding bit by giving False to the bit of $I_{j}$ and True to the bit of $I_{j^{\prime}}$ or vice-versa.
Let us analyze the properties of this construction. The number of inputs of $C^{\prime}$ is at most $k\lceil\log n\rceil\leq k\log n+k$. The number of wires we added is at most $O(kn\log n)$ for the wires incident on $c_{i}^{j}$ gates, and $O(k^{2}\log n)$ for the gadget that ensures we have an assignment of weight $k$. The construction can be performed in linear time in the size of the output and the total size is at most $O(sk^{2}n^{2})$.
7 Applications
In this section we list five characteristic example of our equivalence classes. In each case we highlight a problem and a lower bound on its time complexity such that breaking the bound is equivalent to falsifying the hypotheses of the corresponding class.
7.1 SETH-equivalence for Coloring
Theorem 7.1.
For all $q\geq 3$, there is an $\varepsilon>0$ and an algorithm solving $q$-Coloring in time $(q-\varepsilon)^{m}n^{O(1)}$, where $m$ is the size of a given modulator to constant tree-depth, if and only if the SETH is false. The same holds if $m$ is the size of a given $(\sigma,\delta)$-hub.
Proof 7.2.
It was already established in [15] that if the SETH is true, then there is no algorithm for $q$-Coloring with the stated running time, even for parameter the size of a given $(\sigma,\delta)$-hub. Therefore, if the SETH is true, no such algorithm exists for parameter size of modulator to constant tree-depth.
We establish that if the SETH is false, then we do get such a $q$-Coloring algorithm for parameter modulator to constant tree-depth. We are given an $n$-vertex graph $G$ and a modulator $M$ of size $m$ so that $G-M$ has tree-depth $c$. We will use Theorem 3.1 and reduce this instance to an instance of $k$-SAT with a modulator to constant tree-depth. According to Theorem 3.1, if the SETH is false, we can solve such an instance is time $2^{(1-\varepsilon)m^{\prime}}|\phi|^{O(1)}$, where $m^{\prime}$ is the size of the new modulator, for some $\varepsilon$. In particular, we will make sure that $m^{\prime}\leq(1+\frac{\varepsilon}{2})m\log q+O(1)$, and that $\phi$ has size polynomial in $n$. Then, the running time of the whole procedure will be at most $2^{(1-\varepsilon)(1+\frac{\varepsilon}{2})m\log q}n^{O(1)}=q^{(1-\varepsilon^{\prime})m}n^{O(1)}$ for some appropriate $\varepsilon^{\prime}$.
To begin our construction, we consider every vertex of $G-M$. For each such $v$, we define $q$ variables $x_{v}^{1},\ldots,x_{v}^{q}$, indicating informally that $v$ takes color $i\in[q]$. We add the clause $(x_{v}^{1}\lor\ldots\lor x_{v}^{q})$. For each $uv\in E$ with $u,v\not\in M$ we add for each $i\in[q]$ the clause ($(\neg x_{v}^{i}\lor\neg x_{u}^{i})$. It is not hard to see that the graph we have constructed so far has tree-depth at most $cq$. All the other variables we will construct will belong in the new modulator.
Consider now the vertices of $M$ and for some integer $\gamma$ to be defined later partition them into groups of size at most $\gamma$ as equitably as possible. We have then at most $\tau=\lceil\frac{m}{\gamma}\rceil\leq\frac{m}{\gamma}+1$ groups, $M_{1},\ldots,M_{\tau}$. For some integer $\rho$, also to be defined later, we replace each $M_{i}$ with a set of $\rho$ new variables $Y_{i}$. We select the integers $\gamma,\rho$ so that we have the following:
$2^{(1-\frac{\varepsilon}{2})\rho}\leq q^{\gamma}\leq 2^{\rho}$
Since $q^{\gamma}\leq 2^{\rho}$, we can map injectively all proper colorings of $G[M_{i}]$ to assignments of $Y_{i}$. For each assignment that is not an image of a proper coloring we add a clause ensuring that this assignment cannot be used in a satisfying assignment. Furthermore, for each $uv\in E$ such that $u,v$ belong in distinct sets $M_{i},M_{j}$ we consider all pairs of colorings of $M_{i},M_{j}$ that set $u,v$ to the same color. For each such pair, we add a clause ensuring that $Y_{i},Y_{j}$ cannot simultaneously take the assignments which are the images of these colorings. Finally, for each $u\in M_{i}$ and $v\not\in M$ such that $uv\in E$, for each color $j\in[q]$ we add for each coloring of $M_{i}$ that colors $u$ with $j$ a clause that contains $\neg x_{v}^{j}$ and literals which are all falsified if we select for $Y_{i}$ the image of this coloring of $M_{i}$.
This completes the construction and it is not hard to see that there is a one-to-one correspondence between proper colorings of $G$ and satisfying assignments of the CNF formula we have constructed. To bound the size of the new modulator we observe that we have $\frac{m}{\gamma}+1$ groups of $\rho$ variables, but $(1-\frac{\varepsilon}{2})\rho\leq\gamma\log q$, therefore, $\rho\leq(1+\frac{\varepsilon}{2})\gamma\log q$, so the total number of variables of the modulator is at most $(\frac{m}{\gamma}+1)\rho\leq(1+\frac{\varepsilon}{2})m\log q+O(1)$.
Finally, to argue that we can select appropriate $\gamma,\rho$, we observe that it is sufficient that $2^{\rho}>2^{(1-\frac{\varepsilon}{2})\rho}q$, because then there is an integer power of $q$ between $2^{\rho}$ and $2^{(1-\frac{\varepsilon}{2})\rho}$. This gives $\rho>\frac{4\log q}{\varepsilon}$, so selecting $\rho$ to be any integer at least this large also allows us to select an appropriate $\gamma$.
7.2 MaxSAT-equivalence for Max Cut
Theorem 7.3.
There is an $\varepsilon>0$ and an algorithm which takes as input an $n$-vertex graph $G$ and a $(\sigma,\delta)$-hub of size $h$ and computes a maximum cut of $G$ in time $(2-\varepsilon)^{h}n^{O(1)}$ if and only if the MaxSat-SETH is false.
Proof 7.4.
One direction is easy: given a Max Cut instance $G=(V,E)$, we can construct a CNF formula by constructing a variable $x_{v}$ for each vertex $v$ and for each edge $uv$ the clauses $(x_{u}\lor x_{v})$ and ($(\neg x_{u}\lor\neg x_{v})$. The primal graph of this formula is simply $G$, so the $(\sigma,\delta)$-hub is preserved. It is now not hard to see that the maximum number of clauses that it is possible to satisfy in the new instance is exactly $|E|+\textrm{max-cut}(G)$, so an algorithm as described in the second statement of Theorem 3.5 gives the desired algorithm for Max Cut.
For the converse direction, we are given a $k$-CNF formula $\phi$ with $n$ variables and $m$ clauses and a target value $t$ and are asked if we can satisfy at least $t$ clauses. We reuse the reduction of [33], except with smaller weights, so that we obtain a $(\sigma,\delta)$-hub. Let us give the relevant details. We will first describe a reduction to edge-weighted Max Cut and then remove the weights.
For each variable of $\phi$, $x_{1},\ldots,x_{n}$ we construct a vertex with the same name. We also construct a vertex $x_{0}$. For each clause $C_{j}$ of $\phi$ we construct a cycle of length $14k+$ 1, one of whose vertices is $x_{0}$, while the others are new vertices. Each edge of this cycle has weight $8k$. Number the vertices of the cycle $p_{j,1},\ldots,p_{j,4k}$, starting from a neighbor of $x_{0}$. Suppose that $C_{j}$ contains literals $\ell_{1},\ell_{2},\ldots,\ell_{k^{\prime}}$ for $k^{\prime}\leq k$. For each $i\in[k^{\prime}]$ if $\ell_{i}$ is a positive appearance of a variable $x_{r}$, then we connect $p_{j,4i+2},p_{j,4i+3}$ to $x_{r}$, otherwise if it is a negative appearance of $x_{r}$, then we connect $p_{j,4i+1},p_{j,4i+2}$ to $x_{r}$. These edges have weight $1$. Furthermore, we attach $k-k^{\prime}$ leaves to $p_{j,1}$ with edges of weight $1$. We now claim that there is an assignment satisfying $t$ clauses if and only if there is a cut of weight at least $32k^{2}m+km+t$.
For the forward direction, if there is such an assignment, place all the vertices representing False variables on the same side as $x_{0}$ in the cut and all the vertices representing True variables on the other side. For each unsatisfied clause, pick a bipartition of the corresponding odd cycle so that $4k$ of the $14k+$ 1 edges of the cycle are cut and the uncut edge is incident on $x_{0}$. This contributes $32k^{2}$ from the edges of the cycle. We also obtain at least $k^{\prime}$ from the edges connecting the cycle to the variable vertices and $k-k^{\prime}$ from the attached leaves by placing each leaf to the opposite side of its neighbor, so we get $32k^{2}+k$ for each unsatisfied clause. For each satisfied clause $C_{j}$, let $\ell_{i}$ be the first literal that is set to True by the assignment. For $i^{\prime}<i$ we set $p_{j,4i^{\prime}+2},p_{j,4i^{\prime}+4}$ on the same side as $x_{0}$ and $p_{j,4i^{\prime}+1},p_{j,4i^{\prime}+3}$ on the other side; while for $i^{\prime}>i$ we use the opposite partition. We now have that the vertex preceding $p_{j,4i+1}$ and the vertex after $p_{j,4i+4}$ in the cycle are both on the side of $x_{0}$. For the four remaining vertices, if $\ell_{i}$ is a variable that appears positive, we set $p_{j,4i+2},p_{j,4i+3}$ on the same side as $x_{0}$, cutting both edges connecting these vertices to the corresponding variable vertex, and $p_{j,4i+1},p_{j,4i+4}$ on the other side; while if $\ell_{i}$ is a variable that appears negative, we set $p_{j,4i+1},p_{j,4i+2},p_{j,4i+4}$ on the side opposite $x_{0}$ and $p_{j,4i+3}$ on the side of $x_{0}$, again cutting both edges to the corresponding variable vertex. We now observe that for a satisfied clause we have cut $132k^{2}+k+$ 1 edges, obtaining the bound on the total size of the cut.
For the converse direction we observe that an optimal cut must cut $4k$ of the $14k+$ 1 edges of the odd cycle constructed for each clause. Indeed, it is not possible to cut more (since the cycle is odd), while if we leave two edges uncut, we may replace the solution by a solution that cuts $4k$ of the edges. This will add $8k$ to the total weight, more than making up for the at most $2k$ lost edges connecting the cycle to the rest of the graph. Thus, for each cycle we obtain $32k^{2}$ from the edges inside the cycle. Furthermore, we can obtain at most $k+1$ from the remaining edges incident on the cycle. To see this, note that we have added $k^{\prime}$ pairs of edges (one pair for each literal of the clause) which are adjacent to consecutive vertices in the cycle. If we cut both edges in a pair, an edge of the cycle is uncut, so there is at most one pair of cut edges and from each other pair we cut at most one edge. Taking into account the $k-k^{\prime}$ incident on leaves (which are always cut), we have that each cycle contributes at most $132k^{2}+k+$ 1 to the cut. Hence, there must be at least $t$ cycles contributing exactly this much, which we claim correspond to satisfied clauses by the assignment we extract by setting to False variables whose corresponding vertices are on the same side as $x_{0}$. Indeed, take a cycle representing $C_{j}$ and suppose that $C_{j}$ is not satisfied by this assignment. However, there exists a literal $\ell_{i}$ in $C_{j}$ such that we cut both edges added for $\ell_{i}$. If $\ell_{i}$ is positive, then $p_{j,4i+2},p_{j,4i+3}$ have a common neighbor. If $C_{j}$ is not satisfied, this neighbor must be on the same side as $x_{0}$, so $p_{j,4i+2},p_{j,4i+3}$ are on the other side. It is now not hard to see via a parity argument that two edges of the cycle must be uncut, contradiction. The reasoning is similar if $\ell_{i}$ is negative, as then $p_{j,4i+1},p_{j,4i+2}$ must have a common neighbor, which must be on the side not containing $x_{0}$ (otherwise $C_{j}$ is satisfied), so $p_{j,4i+1},p_{j,4i+2}$ are on the same side as $x_{0}$, and again two edges are uncut. Hence, a cut of the specified size leads to an assignment satisfying at least $t$ clauses.
In order to remove weights, for each edge $uv$ of weight $w$ we construct $w$ parallel paths of length $3$ between $uv$ and increase the target by $2w$. It is not hard to see that any decision on $u,v$ can be extended in a way that cuts at least $2w$ of the new edges, and it is possible to cut all $3w$ new edges if and only if $u,v$ are on different sides.
After the transformation we have that the variable vertices plus $x_{0}$ form a hub of size $n+1$ such that its removal leaves components with $O(k^{2})$ vertices. Furthermore, each component is connected to at most $k$ vertices of the hub. We therefore have a $(\sigma,\delta)$-hub, for $\sigma=O(k^{2})$ and $\delta=k$.
7.3 Coloring and Pathwidth Modulators
Theorem 7.5.
For all $q\geq 3$ we have the following. There is an $\varepsilon>0$ such that for all $c>0$ there is an algorithm that takes as input an $n$-vertex graph $G$ and a modulator $M$ of size $m$ such that $G-M$ has pathwidth at most $c$ and decides if $G$ is $q$-colorable in time $(q-\varepsilon)^{m}n^{O(1)}$ if and only if the LD-C-SETH is false. The same is true if we replace “pathwidth $c$” with “tree-depth $c\log n$” in the preceding statement.
Proof 7.6.
Clearly, an algorithm parameterized by modulator to tree-depth $c\log n$ implies the existence of an algorithm parameterized by modulator to pathwidth $c$. We show two statements: if the LD-C-SETH is false, then there is an algorithm for parameter modulator to tree-depth $c\log n$; and if there is an algorithm for parameter modulator to pathwidth $c$, then the LD-C-SETH is false.
The first claim readily follows from the proof of Theorem 7.1. Recall that in that proof we showed how to transform an instance $G$ of $q$-Coloring with a modulator $M$ to tree-depth $c$ to an instance of SAT with a modulator of size essentially $|M|\log q$, while ensuring that in the new instance the removal of the modulator leaves a graph of tree-depth at most $cq$. We run the same reduction, except now the tree-depth is at most $cq\log n$. One of the statements of Theorem 4.1 states that we can avoid brute-force for SAT parameterized by a modulator to logarithmic tree-depth if and only if the LD-C-SETH is false.
For the second claim we need to reduce SAT with parameter modulator to pathwidth $c$ to $q$-Coloring. Suppose for some $q\geq 3$ there exists $\varepsilon>0$ such that for all $c^{\prime}$ there is a $q$-Coloring algorithm which takes as input a modulator to pathwidth $c^{\prime}$ of size $m^{\prime}$ and solves $q$-Coloring in time $q^{(1-\varepsilon)m^{\prime}}n^{O(1)}$.
We are given a CNF formula $\phi$ and a modulator $M$ of size $m$ such that $\phi-M$ has pathwidth $c$. Using Section A.2 we can assume that $\phi$ is a 3-CNF formula. Select integers $\gamma,\rho$ such that we have the following:
$q^{(1-\frac{\varepsilon}{2})\gamma}<2^{\rho}<q^{\gamma}$
For such integers to exist it is sufficient to have $q^{\gamma}>2q^{(1-\frac{\varepsilon}{2})\gamma}$, which gives $\gamma>\frac{2}{\varepsilon\log q}$. Selecting $\gamma=\lceil\frac{2}{\varepsilon\log q}\rceil$ then guarantees that an appropriate integer $\rho$ exists.
We will first reduce our 3-CNF formula to an instance of CSP with alphabet $[q]$ and arity at most $3\gamma$. In this problem we are given a list of variables which take values in $[q]$ and a list of constraints, each of which involves at most $3\gamma$ variables. Each constraint lists the assignment combinations to the involved variables which are acceptable. The goal is to select an assignment so that we have selected an acceptable combination in each constraint. Observe that $q$-Coloring is a special case of this problem where the arity is $2$ and all constraints state that the endpoints of an edge must receive distinct colors.
We now partition $M$ into groups of variables of size $\rho$ as equitably as possible, constructing at most $\frac{m}{\rho}+1$ groups. For each group we construct in our CSP instance a group of $\gamma$ variables. We construct an injective mapping from boolean assignments to a group of variables of the original instance to an $q$-ary assignment to the corresponding group of variables of the new instance. Since $2^{\rho}<q^{\gamma}$, this is always possible. For each variable of $\phi$ outside of $M$ we construct a variable whose acceptable values are only $\{1,2\}$ (representing True, False). For each clause $C$ of $\phi$ we construct a constraint with maximum arity $3\gamma$: if a variable of $C$ is part of a group $M_{i}$ of variables of $M$, we involve all the $\gamma$ variables representing this group in the constraint; otherwise we involve the single variable that represents the boolean variable outside of $M$ appearing in $C$. We now list all combinations of assignments to the at most $3\gamma$ involved variables and we set as unacceptable those combinations which are not an image of boolean assignments that satisfy the clause. This completes the construction and it is not hard to see that the answer is preserved. Furthermore, if we consider the primal graph of the CSP instance, the graph is unchanged once we remove the (new) modulator. The new modulator contains at most $(\frac{m}{\rho}+1)\gamma$ variables. But $(1-\frac{\varepsilon}{2})\gamma\log q<\rho$ so the size of the new modulator is at most $m^{\prime}\leq(1-\frac{\varepsilon}{2})\frac{m}{\log q}+O(1)$. If we had an algorithm solving a CSP instance $\psi$ with bounded arity and alphabet $[q]$ in time $q^{(1-\varepsilon)m^{\prime}}|\psi|^{O(1)}$, where $m^{\prime}$ is the modulator to constant pathwidth, this would give the desired algorithm for deciding our original CNF. We will therefore show how to obtain this from the supposed $q$-Coloring algorithm.
We first reduce the arity of the CSP instance to $2$. For each constraint $C$ involving more than $2$ and at most $3\gamma$ variables, let $\mathcal{S}_{C}$ be the set of acceptable assignments, which has cardinality at most $q^{3\gamma}=O(1)$. We construct $12|\mathcal{S}_{C}|+$ 1 new variables and connect them in an odd cycle with constraints of arity $2$, ensuring that all the variables which are involved in a constraint take distinct values, that $|\mathcal{S}_{C}|$ variables, which are pairwise non-adjacent, have possible values $\{1,2,3\}$, and that all the others have possible values $\{1,2\}$. The intuition here is that we have formed an odd cycle and we can use colors $\{1,2,3\}$ to color it. Since the cycle is odd, we must use color $3$ somewhere, and if we use color $3$ once we can 2-color the rest. The idea then is to associate each of the $|\mathcal{S}_{C}|$ variables that have three possible values with a satisfying assignment to $C$ and add constraints that ensure that if this variable takes color $3$ then the variables of $C$ must agree with the corresponding assignment. For this, for each of the $|\mathcal{S}_{C}|$ variables of the cycle that represents some assignment $\sigma$ we add constraints of arity $2$ involving this variable and each variable of $C$. The constraints state that if the variable of the cycle has value $1$ or $2$, then everything is acceptable, otherwise the other variable must agree with $\sigma$. We do this exhaustively and it is not hard to see that the answer is preserved. We have to argue that the pathwidth of the primal graph (after removing the modulator) is almost preserved. Indeed, for each constraint $C$, find a bag that contains all its variables (outside of $M$), and after this bag add a sequence of copies of it, where we add a path decomposition of the cycle we have constructed. Repeating this will not add to the pathwidth more than a small additive constant.
Finally, we want to reduce the CSP of arity $2$ and alphabet $q$ to $q$-Coloring. We will reduce to List Coloring where all lists are subsets of $[q]$. Adding a clique of size $q$ and appropriately connecting to the graph to simulate the lists will then complete the reduction. To reduce to List Coloring we can use the weak edge gadgets of [30]. For each variable of the CSP instance we construct a vertex. For each constraint over $u,v$ we consider every assignment to $u,v$ that is ruled out by this constraint. A weak edge is a gadget that we can add between $u,v$ that will rule out the corresponding coloring, but will always be colorable if we pick any other color combination for $u,v$. In fact, a weak edge is implemented using a path of length $3$ between $u,v$ (with appropriate lists) so using weak edges instead of edges does not affect the pathwidth of the graph by more than an additive constant. By adding weak edges for each forbidden combination of assignments we obtain an equivalent $q$-Coloring instance. The size of the modulator has been preserved and we obtain the theorem.
7.4 k-Neighborhood Cut
In this section we consider the following problem:
Definition 7.7.
In the $k$-Neighborhood-Cut problem we are given as input a DAG $G$ with two designated vertices $s,t$. The question is whether it is possible to select $k$ vertices of $G$, other than $s,t$, such that deleting all the out-neighbors of the selected vertices results in a graph where no path from $s$ to $t$ exists.
Before we proceed, let us note that it is trivial that $k$-Neighborhood-Cut admits an algorithm running in time $n^{k+O(1)}$: simply try out all solutions and for each solution check if there is a path from $s$ to $t$ avoiding the out-neighborhoods of selected vertices. Furthermore, it is not hard to see that the problem is W[2]-hard parameterized by $k$. Indeed, take an instance of $k$-Dominating Set $G=(V,E)$ and construct a DAG that consists of two copies of $V$, call them $V_{1},V_{2}$, such that there is an arc from each $u\in V_{1}$ to each $v\in V_{2}$ if $u=v$ or $uv\in E$. Add a vertex $s$ with arcs to all of $V_{2}$ and a vertex $t$ with arcs from all of $V_{2}$. Clearly, selecting a dominating set in $V_{1}$ is necessary and sufficient to eliminate all paths from $s$ to $t$. From the results of Patrascu and Williams [40], the above reduction implies that $k$-Neighborhood-Cut cannot be solved in time $n^{k-\varepsilon}$, assuming the SETH.
In this section we show the same lower bound but assuming the weaker assumption that SAT cannot be solved faster than brute-force for instances with a 2-SAT backdoor. More interestingly, we show that this is precisely the right assumption for this problem, in the sense that a better algorithm can be obtained if and only if the assumption is false.
Theorem 7.8.
There exist $\varepsilon>0,k\geq 2$ such that $k$-Neighborhood-Cut can be solved in time $n^{k-\varepsilon}$ if and only if the 2SatB-SETH is false.
Proof 7.9.
We will show reductions in both directions from Ann-s-t-Non-Reach which was established to be equivalent to 2SatB-SETH in Section 5. As a reminder, in Ann-s-t-Non-Reach we are given a DAG with its arcs annotated from a set of $m$ boolean variables. We are asked to select an assignment to the variables so that in the DAG resulting from keeping only non-annotated arcs or arcs whose annotation is set to True by the assignment there is no $s\to t$ path, for two given vertices $s,t$. We have shown that 2SatB-SETH is equivalent to assuming that it is impossible to solve Ann-s-t-Non-Reach without considering essentially all assignments to the $m$ boolean variables.
First, suppose we have an instance $G=(V,E)$ of Ann-s-t-Non-Reach on a DAG with $n$ vertices and a set $M$ of $m$ boolean variables. We construct a DAG $G^{\prime}$ as an instance of $k$-Neighborhood-Cut as follows. We begin with the same DAG $G$. We then partition the variables into $k$ sets, call them $M_{1},\ldots,M_{k}$, each of size at most $\frac{m}{k}+1$. For each set $M_{i}$ we enumerate all assignments $\sigma$, and for each assignment we construct a vertex $x_{i,\sigma}$. We construct for each $i\in[k]$ a vertex $y_{i}$ and add arcs $sy_{i}$, $y_{i}t$ as well as all arcs $x_{i,\sigma}y_{i}$. For each arc $e$ of $G$ that is annotated with a literal from a variable of $M_{i}$, we subdivide the arc and call the new vertex $z_{e}$. We add an arc to $z_{e}$ from each $x_{i,\sigma}$ such that $\sigma$ sets the literal of $e$ to False. This completes the construction, which can be carried out in time at most $2^{m/2}n^{O(1)}$ for all $k\geq 2$. The new graph has $N$ vertices with $N\leq 2^{m/k}+n^{O(1)}$.
We now observe that the reduction above preserves the solution. In particular, because of the vertices $y_{i}$ we are forced to select exactly one $x_{i,\sigma}$ for each $i\in[k]$, so there is a one-to-one correspondence between assignments to $M$ and potential solutions in the new instance. Furthermore, an annotated arc can be used if and only if we have selected the encoding of an assignment that sets the annotation to True.
Suppose that there is an algorithm for $k$-Neighborhood-Cut running in time $N^{k-\varepsilon}$ for some fixed $k\geq 2,\varepsilon>0$. We can now solve Ann-s-t-Non-Reach by running the reduction described above and then solving $k$-Neighborhood-Cut. The running time will be dominated by the second part, which would take $N^{k-\varepsilon}=2^{\frac{m}{k}(k-\varepsilon)}n^{O(1)}=2^{(1-\frac{\varepsilon}{k})m}n^{O(1)}$, which would falsify the 2SatB-SETH by the results of Section 5.
For the converse direction, suppose we are given an instance $G=(V,E)$ of $k$-Neighborhood-Cut on $n$ vertices. We will reduce it to Ann-s-t-Non-Reach as follows. We start with the same DAG and construct a set $M$ of $k\log n$ variables (assume without loss of generality that $n$ is a power of $2$, otherwise adding dummy vertices to $G$). Intuitively, the variables of $M$ are partitioned into $k$ sets $M_{1},\ldots,M_{k}$, where the assignment of $M_{i}$ is meant to encode the selection of the $i$-th vertex of the solution to $k$-Neighborhood-Cut.
We now edit $G$ as follows. We replace every vertex $u\in V$ with two vertices $u_{in},u_{out}$. Each arc $uv$ is replaced with the arc $u_{out},v_{in}$. We set $s_{out}$ as the start vertex and $t_{in}$ as the destination vertex. For each $u\in V$ we further add $k+1$ vertices $u_{0},u_{1},\ldots,u_{k}$. The idea is that we should be able to go from $u_{in}$ to $u_{i}$ if and only if the first $i$ vertices of the solution do not have any arc to $u$ in the original instance. We add an arc $u_{in}u_{0}$ and an arc $u_{k}u_{out}$.
Let us now describe the gadget we add between $u_{i-1}$ and $u_{i}$ for all $i\in[k]$. Intuitively, we should be able to go from $u_{i-1}$ to $u_{i}$ if and only if the assignment to $M_{i}$ does not encode a vertex that dominates $u$. For each $v\in V$ such that $vu\not\in E$ and $v\neq u$ we therefore construct a path from $u_{i-1}$ to $u_{i}$ of length $\log n$ going through new vertices. The arcs of the path are annotated with literals from $M_{i}$ ensuring that we can only follow all the arcs if the assignment to $M_{i}$ encodes vertex $v$.
This completes the construction. Observe that each arc of $G$ has been replaced by $k$ collections of at most $n$ parallel paths, each of length $\log n$, so the new graph has at most $n^{5}$ vertices. We observe that the answer is preserved. Indeed, a solution $S$ of $k$-Neighborhood-Cut can be translated to an assignment to $M$ in the natural way and we can see that if $u\in S$ dominates $v$, then it is impossible to go from $v_{in}$ to $v_{out}$ with the current assignment. In the other direction, if we have an assignment, we can extract a set $S$ in $G$ by interpreting the assignments to the $k$ groups as the indices of the selected vertices.
Suppose now that there is an algorithm solving Ann-s-t-Non-Reach in time $2^{(1-\varepsilon)m}N^{d}$, for some fixed $d$. The running time of executing the reduction and then the algorithm would be at most $2^{(1-\varepsilon)k\log n}n^{5d}=n^{k-\varepsilon k+5d}$. This implies that when $k\geq\frac{5d}{\varepsilon}+1$ the running time is at most $n^{k-\varepsilon}$, therefore there exist $k,\varepsilon$ for which we can obtain a fast enough algorithm for $k$-Neighborhood-Cut.
7.5 Degenerate Deletion
In this section we consider the problem of editing a graph by deleting vertices so that the result becomes $r$-degenerate. Recall that a graph is $r$-degenerate if all of its induced subgraphs contain a vertex of degree at most $r$. Equivalently, a graph is $r$-degenerate if there exists an ordering of its vertices $v_{1},\ldots,v_{n}$ such that each $v_{i}$ has at most $r$ neighbors among vertices with higher index.
The problem we are interested in is whether we can delete $k$ vertices of a given graph so that the graph becomes $r$-degenerate (where $k$ can be thought of as the parameter and $r\geq 2$ is an absolute constant). This problem was shown to be W[P]-complete by Mathieson [36]. By reusing essentially the same chain of reductions, we are able to show the following equivalence between this problem and Circuit-SETH.
Theorem 7.10.
For all $r\geq 2$ we have the following. There exists $\epsilon>0,k_{0}>0$ and an algorithm which for all $k>k_{0}$, if given as input an $n$-vertex graph, decides if $G$ can be made $r$-degenerate by deleting $k$ vertices in time $n^{(1-\varepsilon)k}$ if and only if the Circuit-SETH is false.
Proof 7.11.
For the direction that Circuit-SETH implies hardness for our problem we essentially reuse the reduction of [36], which relies on a previous reduction of [42], and observe that these reductions from the standard W[P]-complete problem we dealt with in Theorem 6.1 actually preserve the value of $k$ exactly and only increase the size of the instance by a constant factor, therefore the lower bound is preserved as well. For the other direction we will perform a direct reduction to Circuit-SAT.
Let us give the relevant details. For the first direction, we start with an instance of monotone weight-$k$ circuit satisfiability. By Theorem 6.1, if we can solve such an instance in time $s^{(1-\varepsilon)k}$ for monotone circuits of $n$ inputs and size $s$, assuming $k$ is a large enough constant, we will falsify the Circuit-SETH. As shown by Szeider [42], there is a simple reduction from this problem to the Cyclic Monotone Circuit Activation problem. In the latter problem we are given a monotone boolean circuit which is a general digraph (not a DAG) and are asked to select $k$ initial gates to activate (set to True). From that point on at each round a gate becomes True if at least one of its in-neighbors is True for disjunction ($\lor$) gates and if all its in-neighbors are True, for conjunction ($\land$) gates. The question is whether there is a selection of $k$ initial gates that eventually makes everything True.
Szeider observed that we can reduce weight-$k$ monotone circuit satisfiability to this problem as follows: take $k+1$ copies of the given circuit, identify their inputs, and then replace the inputs with conjunctions whose inputs are the $k+1$ outputs. If there is a weight-$k$ assignment to the original circuit we use the same activation set and it activates all outputs, therefore eventually all gates. If on the other hand there is a way to activate everything, because we have $k+1$ copies, there is at least one copy in which the output was activated using only the input gates, so from this we can extract a satisfying assignment to the original circuit. Mathieson [36] presents a reduction from this problem which, for each $r\geq 2$ produces an equivalent instance of $r$-Degenerate Deletion, without modifying $k$, by replacing each gate with a constant-size gadget. Executing these reductions starting from a circuit of size $s$ we obtain a graph on $N=O(s)$ vertices. If the problem could be solved in time $N^{(1-\varepsilon)k}$, we would have an algorithm deciding weight-$k$ satisfiability of the original instance in time $O(s^{(1-\varepsilon)k})$, disproving the Circuit-SETH.
For the converse direction, suppose that Circuit-SETH is false and that we have an instance of the $r$-Degenerate Deletion problem with budget $k$ on a graph with $n$ vertices. Assume without loss of generality that $n$ is a power of $2$ (otherwise add some dummy vertices). We will reduce this to satisfiability of a circuit with $k\log n$ inputs and size $n^{O(1)}$. Essentially, this circuit encodes the workings of a polynomial-time algorithm which checks if a solution is correct. The solution is encoded by the $k\log n$ inputs which are meant to be seen as $k$ groups of $\log n$ bits, encoding the indices of the selected vertices to delete. The algorithm we are trying to encode can be seen as follows: at each step we mark some active vertices (initially this set is the set of deleted vertices); then if a vertex has at most $r$ unmarked neighbors, we mark it; continue this exhaustively for $n$ rounds and in the end check if all vertices are marked. This algorithm can easily be encoded by a circuit of $n$ layers, each consisting of $n$ gates, where vertex $j$ of layer $i$ encodes whether vertex $j$ is marked after $i$ rounds.
If we had the supposed algorithm for checking circuit satisfiability, it would run in time $2^{(1-\varepsilon)k\log n}n^{O(1)}=n^{(1-\varepsilon)k+O(1)}$. Setting $k_{0}$ sufficiently large that $\epsilon k_{0}/2$ is larger than the constant additive term in the exponent gives the desired algorithm.
Appendix A Complexity Hypotheses
A.1 List of Hypotheses
We list here the complexity hypotheses that appear in this paper.
  • SAT-related
    1. 1.
      SETH: [25] For all $\varepsilon>0$ there exists $k$ such that there is no algorithm solving $k$-SAT for $n$-variable formulas $\phi$ in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$, where $n$ is the number of variables.
    2. 2.
      MaxSat-SETH: The same as the previous statement but for Max-SAT, that is, the problem of deciding if there is an assignment satisfying at least $t$ clauses, where $t$ is given in the input.
  • Circuits
    1. 3.
      Circuit-SETH: For all $\varepsilon>0$ there is no algorithm deciding if an $n$-input circuit of size $s$ is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
    2. 4.
      LD-C-SETH: The same as the previous statement, but restricted to circuits of depth at most $\varepsilon n$.
  • Backdoors
    1. 5.
      2SatB-SETH: For all $\varepsilon>0$ there is no algorithm deciding if a CNF formula $\phi$ with a given strong 2-SAT backdoor of size $b$ is satisfiable in time $(2-\varepsilon)^{b}|\phi|^{O(1)}$.
    2. 6.
      HornB-SETH: The same as the previous statement, but for a strong Horn backdoor.
  • Modulators
    1. 7.
      tdM-SETH: For all $\varepsilon>0$ there exist $c,k$ such that there is no algorithm that takes as input a $k$-CNF formula $\phi$ and a modulator $M$ of size $m$ such that $\phi-M$ has tree-depth at most $c$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
    2. 8.
      pwM-SETH: For all $\varepsilon>0$ there exists $c$ such that there is no algorithm that takes as input a CNF formula $\phi$ and a modulator $M$ of size $m$ such that $\phi-M$ has pathwidth at most $c$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
    3. 9.
      twM-SETH: The same as the previous statement but for treewidth at most $c$.
    4. 10.
[…]
    5. 11.
[…]
    6. 12.
      $(\sigma,\delta)H$-MaxSat-SETH: For all $\varepsilon>0$ there exist $k,\sigma,\delta$ such that there is no algorithm which takes as input a $k$-CNF $\phi$ and a $(\sigma,\delta)$-hub $M$ of size $m$ and computes an assignment satisfying the maximum number of clauses in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
    7. 13.
      $(\sigma,\delta)H$-SETH: The same as the previous statement but for deciding if the input formula is satisfiable.
  • Weighted SAT
    1. 14.
      W[SAT]-SETH: For all $\varepsilon>0$ there is no algorithm which takes as input a (general) Boolean formula $\phi$ of size $s$ on $n$ variables and an integer $k$ and decides if $\phi$ is satisfiable in time $n^{(1-\varepsilon)k}s^{O(1)}$.
    2. 15.
      W[P]-SETH (weak version): Same as the previous statement but for Boolean circuits.
    3. 16.
      W[P]-SETH (strong version): For all $\varepsilon>0,k_{0}>0$, there is no algorithm which for all $k>k_{0}$ takes as input a Boolean circuit of size $s$ and decides if the circuit has a weight-$k$ satisfying assignment in time $s^{(1-\varepsilon)k}$.
  • Others
    1. 17.
      Max-3-SAT-SETH: The same as MaxSat-SETH but for CNFs of maximum arity $3$.
    2. 18.
      NC-SETH: The same as LD-C-SETH but for circuits of depth at most $\log^{O(1)}n$.
    3. 19.
      pw-SETH: For all $\varepsilon>0$ there is no algorithm that solves 3-SAT in time $(2-\varepsilon)^{\textrm{pw}}n^{O(1)}$ for formulas on $n$ variables of primal pathwidth pw.
A.2 Relations between Classes of Hypotheses
Many of the relations between the hypotheses we consider can be inferred directly from the definitions or from the relations between graph widths. In particular, we have $\textrm{tw}\leq\textrm{pw}\leq\textrm{td}\leq\textrm{tw}\log n$ for any graph on $n$ vertices. This gives $1\Rightarrow 7\Rightarrow 8\Rightarrow 9\Rightarrow 10\Rightarrow 11\Rightarrow 13\Rightarrow 12$. Furthermore, it is clear that deciding if a general circuit is satisfiable is at least as hard as a circuit with depth $\varepsilon n$, which is at least as hard as a CNF formula, so $1\Rightarrow 4\Rightarrow 3$. We have $14\Rightarrow 15$ as formulas are a special case of circuits. Also $2\Rightarrow 12$ as the set of all variables is a $(\sigma,\delta)$-hub.
It is worth noting that the arity $k$ appears explicitly in statement 7, but not in statements 8 to 11. This is because of the following observation, which allows us to decrease the arity once the parameter we are dealing with allows the construction of long paths.
{observation}
There is a polynomial time algorithm which takes as input a CNF formula $\phi$ and a modulator $M$ and produces an equisatisfiable 3-CNF formula $\psi$ by adding variables to $\phi$ such that $M$ remains a modulator of $\psi$ and we have $\textrm{pw}(\psi-M)\leq\textrm{pw}(\phi-M)+2$, $\textrm{tw}(\psi-M)\leq\textrm{tw}(\phi-M)+2$, $\textrm{td}(\psi-M)\leq\textrm{td}(\phi-M)+O(\log|\phi|)$.
Proof A.1.
We first state an easy claim.
Claim 2.
Let $G=(V,E)$ be a graph and $\mathcal{K}\subseteq 2^{V}$ be a set of cliques of $G$. Let $G^{\prime}$ be a graph obtained from $G$ by adding for each $K\in\mathcal{K}$ a path going through at most $\ell$ new vertices and connecting some of the vertices of the path to vertices of $K$. Then $\textrm{pw}(G^{\prime})\leq\textrm{pw}(G)+2$, $\textrm{tw}(G^{\prime})\leq\textrm{tw}(G)+2$, and $\textrm{td}(G^{\prime})\leq\textrm{td}(G)+\log(\ell+1)$.
{claimproof}
Assume we start with a tree or path decomposition of $G$ or a rooted tree proving the tree-depth of $G$. For path and tree decompositions we know that there is a bag $B$ that contains the clique $K$. Take a path decomposition of the path that we connected to $K$ and add to all bags the contents of bag $B$ (the width has now been increased by $2$). In the case of treewidth, attach one bag of this decomposition to $B$; in the case of pathwidth insert this decomposition after $B$. For tree-depth, let $k\in K$ be the vertex that is furthest from the root. We observe that all vertices of $K\setminus\{k\}$ must be ancestors of $k$. We insert below $k$ a tree of depth at most $\log(\ell+1)$ containing the path we connected to $k$. Here we are using the fact that a path of $\ell$ vertices has tree-depth at most $\log(\ell+1)$ (which can easily be seen by induction).
We apply the following standard trick exhaustively: as long as $\phi$ contains a clause of arity $r>3$, say $C=(\ell_{1}\lor\ell_{2}\lor\ldots\lor\ell_{r})$, introduce $r-3$ new variables $z_{1},\ldots,z_{r-3}$ and replace $C$ with the clauses $(\ell_{1}\lor\ell_{2}\lor z_{1})\land(\neg z_{1}\lor\ell_{3}\lor z_{2})\land(\neg z_{2}\lor\ell_{4}\lor z_{3})\land\ldots\land(\neg z_{r-3}\lor\ell_{r-1}\lor\ell_{r})$. Observe now that the variables of $C$ form a clique in the primal graph, while the new variables induce a path whose neighborhood is contained in this clique. We can therefore apply the claim, with $\mathcal{K}$ being the set of all clauses on which we applied this transformation.
We note here two further easy observations justifying the organization of hypotheses we have made in Figure 1.
Theorem A.2.
(MaxSat-SETH $\Rightarrow$ LD-C-SETH) If there is an $\varepsilon>0$ and an algorithm that decides satisfiability of $n$-input circuits of size $s$ and depth at most $\varepsilon n$ in time $(2-\varepsilon)^{n}s^{O(1)}$, then there is an $\varepsilon^{\prime}>0$ and an algorithm that given as input an $n$-variable CNF formula $\phi$ finds the assignment satisfying the maximum number of clauses in time $(2-\varepsilon^{\prime})^{n}|\phi|^{O(1)}$.
Proof A.3.
Construct a circuit with $n$ inputs and a gate for each clause appropriately connected to the inputs through negations. This circuit can be made to have fan-in $2$ and depth $\log n$. For each target $t$ we now want to add some layers to the circuit so that it is satisfiable if and only if we can satisfy at least $t$ clauses of the original instance. Running the assumed algorithm for each $t$ allows us to compute the maximum number of satisfied clauses, while self-reducibility allows us to find an assignment that satisfies this many clauses.
[…]
Theorem A.4.
(2SatB-SETH $\Rightarrow$ Circuit-SETH) If there is an $\varepsilon>0$ and an algorithm that decides satisfiability of $n$-input circuits of size $s$ in time $(2-\varepsilon)^{n}s^{O(1)}$, then there is an $\varepsilon^{\prime}>0$ and an algorithm that given as input a CNF formula $\phi$ and a strong 2-SAT backdoor of size $b$ decides if $\phi$ is satisfiable in time $(2-\varepsilon^{\prime})^{b}|\phi|^{O(1)}$.
Proof A.5.
The idea is to construct a circuit implementing the standard polynomial-time algorithm for 2-SAT. Recall that one way to decide satisfiability for 2-SAT is to run resolution exhaustively, where resolution is the rule that, given two clauses $C,C^{\prime}$ that contain two opposite literals $x,\neg x$, adds to the formula a new clause containing all the other literals of $C,C^{\prime}$. It is known that a 2-CNF formula is satisfiable if and only if performing this exhaustively does not produce the empty formula.
We are given a CNF $\phi$ and a backdoor $B$ of size $b$. Let $x_{1},\ldots,x_{n}$ be the non-backdoor variables of $\phi$. We construct a circuit with $b$ inputs. Observe that there are at most $4n^{2}$ clauses we can construct with the variables $x_{1},\ldots,x_{n}$. For each $t\in[0,4n^{2}]$ and clause $C$ that can be constructed from these variables we will have in our circuit a gate $x_{t,C}$ whose meaning is that the algorithm has inferred $C$ after $t$ steps. The gates $x_{0,C}$ can be computed from the inputs: $x_{0,C}$ is the disjunction over all clauses of $\phi$ that contain $C$ of the conjunction of the negation of all their other literals. It is now easy to compute $x_{i+1,C}$ from layer $x_{i,C}$ and the output of the circuit is $x_{4n^{2},\emptyset}$. The size of the circuit is polynomial in $\phi$.