### v2

# Circuits and Backdoors: Five Shades of the SETH

Abstract
The Strong Exponential Time Hypothesis (SETH) is a standard assumption in (fine-grained) parameterized complexity and many tight lower bounds are based on it. We consider a number of reasonable weakenings of the SETH, with sources from (i) circuit complexity (ii) backdoors for SAT-solving (iii) graph width parameters and (iv) weighted satisfiability problems. Our goal is to arrive at formulations which are simultaneously more plausible as hypotheses, but also capture interesting and robust notions of complexity. Using several tools from classical complexity theory we are able to consolidate these numerous hypotheses into a hierarchy of five main equivalence classes of increasing solidity. This framework serves as a step towards structurally classifying a variety of SETH-based lower bounds into intermediate equivalence classes.
To illustrate the applicability of our framework, for each of our classes we give at least one (non-SAT) problem which is equivalent to the class as a characteristic example application. As our main showcase, we consider a natural parameterization of Independent Set by vertex deletion distance from several standard graph classes. Our framework allows us to make connections that would have previously been hard to see: for instance, we consider the question of whether (Weighted) Independent Set can be solved faster than $2^{k}n^{O(1)}$ on graphs which are $k$ vertices away from a class $\mathcal{C}$ where the problem is tractable, such as interval graphs or cographs. We provide precise characterizations of the difficulty of breaking such bounds, in particular proving that obtaining $(2-\varepsilon)^{k}n^{O(1)}$ time algorithms for Cograph+kv+kv or Block+kv+kv graphs is equivalent to obtaining a fast satisfiability algorithm for circuits of depth $\varepsilon n$; while solving the weighted version for Interval+kv+kv graphs is equivalent to the (seemingly) harder problem of obtaining a fast satisfiability algorithm for SAT parameterized by a 2-SAT backdoor.
keywords:
SETH, SAT Backdoors, Parameterized Complexity, Fine-grained Complexity
category:
[…]
1 Introduction
The objective of this paper is to study the structural relations between fine-grained questions in parameterized complexity theory. In order to avoid becoming too abstract and philosophical, let us begin with a concrete example that illustrates what we mean by that.
Consider the Independent Set problem and fix a class of graphs $\mathcal{C}$ on which Independent Set is in P and which satisfies the very mild condition of being hereditary (closed under vertex deletion). How hard is it to solve Independent Set on $\mathcal{C}+kv$, that is, the class of graphs which are $k$ vertex deletions away from being in $\mathcal{C}$? Assuming the $k$ vertices are marked in the input, the obvious time upper bound is $2^{k}n^{O(1)}$: guess which of the $k$ vertices to take in the solution, simplify the instance, and solve the rest in polynomial time. Notice that this is a very standard “distance to triviality” parameterization, of which numerous are considered in the parameterized complexity literature, and that our algorithmic argument is based on a trivial branching procedure.
Given the above, a very natural question to ask is for which classes $\mathcal{C}$ one could hope to do better than $2^{k}n^{O(1)}$. Let us cherry-pick a few example classes which satisfy our criteria, such as interval graphs, cluster graphs, cographs, and block graphs [^1] The precise definitions of these classes are recalled in Section 8, but are not necessary at this point.. Since cluster graphs are a subset of the other classes, one should clearly start there. However, a recent result of Esmer, Focke, Marx, and Rzazewski [19] shows that even for cluster graphs, we cannot achieve $(2-\varepsilon)^{k}n^{O(1)}$, unless the SETH is false [^2] As we explain later this is not explicitly stated in [19] but follows from their construction. We give details in Section 8.! It therefore appears that despite the fact that, say, interval graphs are a much wider class than cluster graphs, the complexity of all these problems (and many other similar cases, say $\mathcal{C}$ being chordal, claw-free, $P_{5}$-free, etc.) simply collapse to one identical case. Should we view this as cause for celebration (because we have completely resolved all these problems)? Or should we view this as cause for concern (are we missing something)?
Before we answer this, let us take a step back into classical complexity theory. Consider the following questions: we are given a CNF formula $\phi$ and are asked (i) is $\phi$ satisfiable? (ii) does $\phi$ have a unique satisfying assignment? (iii) is the majority of assignments satisfying for $\phi$? (iv) can we satisfy $\phi$ even if we alternate picking values for variables with an adversary who is trying to falsify it? A knowledgeable reader can easily recognize that the complexity of these questions is strongly believed to be very different, as they correspond to complete problems for the classes NP, $\Sigma_{2}^{p}$, PP, and PSPACE respectively. Nevertheless, a reader less familiar with complexity theory may see a similarity between this situation and the questions of the previous paragraph. Indeed, all the questions we asked about $\phi$ can be solved in exponential time (and polynomial space), while exponential time is necessary even for the easiest one (under standard hypotheses). We observe then that a large effort of classical complexity theory has been devoted into understanding the relative structural complexities of various problems, even (and perhaps especially) in cases where the problems are believed to have the same time and space complexity. This effort has been quite impactful as the different complexity nuances captured by all the intermediate classes have often turned out to be important in other aspects (for instance, kernelization lower bounds typically rely on the assumption that the polynomial hierarchy does not collapse, which would not make sense as a statement if we only cared about deterministic time complexity).
Given this history, the thesis of this paper is that if, in the context of fine-grained parameterized complexity, we settle just for SETH-based lower bounds, then we are indeed missing something. Going back to our example, we would like to replicate this success of classical complexity theory in our setting and obtain a more nuanced understanding of the complexity of our questions. In the same way that traditional complexity theory allows us to go from saying “this problem is likely intractable (not in P)” to saying “this problem is exactly as hard as $\Sigma_{2}^{p}$ or PSPACE”, we would like to be able to go from saying “brute-force is inevitable for Independent Set on the class $\mathcal{C}+kv$ (under SETH)” to “avoiding brute force for Independent Set on $\mathcal{C}+kv$ is exactly as hard as this class of problems”. The obstacle is of course that, at the moment, fine-grained parameterized theory is missing the framework and the vocabulary necessary to do this. In other words, we do not yet have a classification effort which would allow us to take questions which are all “SETH-hard” and order them into equivalence classes, according to their relative difficulty.
The goal of this paper is exactly to explore this new direction, start building the basic blocks of such a framework, and show that this is a fruitful and promising approach. We will do this by identifying five natural equivalence classes (Figure˜1), one of which contains the SETH, and giving natural complete problems for each as well as an informal characterization connecting them to classical complexity classes. To showcase the applicability of this framework we will also apply it to the Independent Set problem discussed above for the four classes we cherry-picked (Figure˜2). The framework will allow us to compare the relative complexities of cases which do not have an obvious graph-theoretic relation (for example cographs and interval graphs are incomparable classes), but also relate them with questions that initially seem completely unrelated, stemming from circuit complexity or backdoors for SAT-solving. We see this investigation as a first step that will allow us to more carefully classify known tight lower bounds from the literature into equivalence classes, with the added benefit of identifying the lower bounds which should be considered most plausible (because it must be taken into account that the SETH is not a universally believed assumption).
1.1 Background and Motivation
Having summarized the general direction of this paper, let us give a more detailed review of the relevant background. Recall that the Strong Exponential Time Hypothesis (SETH [30]) states that the satisfiability of CNF formulas cannot be decided significantly faster than by brute force, that is, in time $(2-\varepsilon)^{n}$, assuming the formulas have large enough arity [^3] To ease presentation, in this section we use informal statements such as this one to describe all complexity hypotheses. In case of doubt, we refer the reader to Appendix A, where precise definitions are given.. Since its introduction, the SETH has been an indispensable tool in the development of fine-grained complexity theory. In this paper we focus on its influence on fine-grained parameterized complexity.
One of the most important techniques in the parameterized algorithms toolbox is the use of dynamic programming algorithms over tree decompositions, and the question of whether such algorithms can be improved is of central importance. It was therefore a very exciting development for the field when Lokshtanov, Marx, and Saurabh [38] pioneered a line of work which allowed to provide tight lower bounds for a large class of algorithms of this type. For instance, one of their results was that the standard algorithm for 3-Coloring, which has complexity $3^{\textrm{tw}}n^{O(1)}$, cannot be improved to $(3-\varepsilon)^{\textrm{tw}}n^{O(1)}$, where tw denotes the treewidth. Similar tight results were given in [38] and many subsequent works for many other problems, and also for other parameters (for example clique-width, cutwidth) [7, 9, 14, 17, 18, 20, 21, 22, 25, 27, 28, 32, 33, 34, 37, 40, 43]. An important common point of all these lower bounds is that they are generally shown to hold assuming the SETH.
The SETH is too strong (i).
One of the motivations of this paper is to question the wisdom of basing all these numerous lower bound results on one-way implications of the SETH. Even though the SETH is mostly considered a reasonable hypothesis, it is important to keep in mind that it is not universally believed. Indeed, there seem to be serious reasons to doubt its validity, much more than that of more standard assumptions (we refer the reader to the argumentation given by Williams [49] who, somewhat provocatively, assigns the SETH only a one-in-four probability of being true; see also [50]). In other words, even if one believes that the SETH is probably true, because its status is sufficiently open one should not necessarily consider a SETH-based lower bound as the final word on a problem’s status.
This line of reasoning has already been well-developed in the (non-parameterized) fine-grained complexity community. In particular, there are several works which reprove (and improve) lower bounds based on the SETH using a more plausible assumption, such as the NC-SETH, which states the same lower bound but for circuits of poly-logarithmic depth, rather than CNF formulas. We refer the reader to the work of Abboud, Hansen, Vassilevska Williams, and Williams [3] for a more extensive discussion of why the NC-SETH is a preferable hypothesis, or equivalently why falsifying the NC-SETH would have far more surprising consequences than disproving the SETH; see also [1, 10, 47]. In the parameterized complexity setting, similar work was only recently initiated in [36] (more on this below).
The SETH is too strong (ii).
The discussion above explains that the SETH is not an ideal starting point for proving lower bounds, as its validity is still subject to some doubt. There is, however, another, more counter-intuitive reason to argue that the SETH is not ideal, and that is that the SETH is “too good” in comparison with the lower bounds we are interested in. This is a line of reasoning taken by several previous works [26, 32]. In particular, Jaffke and Jansen [32] showed that the correct base for $q$-Coloring is exactly $q$ not only when we parameterize by pathwidth or treewidth, but even when we parameterize by vertex deletion distance to path-forest (a significantly more restrictive parameter). More recently, this approach was further explored by Esmer, Focke, Marx, and Rzazewski [19], who observed that for many of the problems considered in [38] (including $q$-Coloring, Independent Set, and Max-Cut), a more refined reduction shows that the SETH implies a much stronger lower bound: Not only are current algorithms optimal for treewidth, but they are still optimal for a novel, extremely restricted parameter, which they call $(\sigma,\delta)$-hub. A $(\sigma,\delta)$-hub, (for $\sigma,\delta$ to be considered absolute constants) is a set of vertices (a modulator) whose removal leaves a collection of bounded-size (size $\sigma$) components with bounded degree (degree $\delta$) in the modulator. The class of graphs with a $(\sigma,\delta)$-hub of size $k$ is a tiny fraction of the class of graphs of treewidth $k$, so this is a strong indication that the lower bounds for the more interesting parameters should be achievable by a much weaker assumption.
Another way to see the above is that the SETH is talking about the hardness of a different class of problems than the width-based lower bounds that it implies. Indeed, observe that we have crucially moved from problems where the “obvious” algorithm is dynamic programming (parameterized by treewidth or other widths) to problems where the “obvious” algorithm is branching (for example, decide the color of the vertices of the modulator). This leads us to believe that there are several “complexity classes” lying between the SETH and the treewidth-based lower bounds discussed above and waiting to be discovered.
Equivalences are better than implications.
[…]
Much of the state of the art in fine-grained parameterized complexity can then be thought of as being close to the superficial understanding we sketched above: we show that our problems are harder than the SETH, and consider the case closed. The thesis of this paper is that there is something to be gained by more carefully comparing the relative complexities of our problems, and for this we need to organize them into equivalence classes. Interestingly, we will show that the equivalence classes we end up with can sometimes be orthogonal to traditional complexity classes.
[…]
Figure 1: Summary of results. We identify five main equivalence classes, with arrows indicating implication, that is, pointing to weaker hypotheses. From bottom to top the five classes are problems equivalent to the SETH; the SETH for Max-SAT; the SETH for circuits of linear depth; the SETH for 2-SAT backdoors; and the SETH for arbitrary circuits. The figure also shows some connections with related hypotheses. Full definitions are given in Appendix˜A.
1.2 Summary of Results
One way to explain the main direction of our results is to come back to the recent work of Esmer, Focke, Marx, and Rzazewski [19] we mentioned. We share their initial motivation, which is the observation that the SETH does not just imply tight treewidth-based lower bounds, such as those of [38], but in fact much more. However, where their next step is to say “Let us then try to extract the strongest implication of the SETH”, we say “Let us then explore how we can replace the SETH by a different (more solid) hypothesis, and try to understand what type of question we capture”. We will therefore study numerous variants of the SETH, coming from several different areas, including: (i) circuit versions, similar to the NC-SETH (ii) backdoor versions, such as “If we have a SAT formula with a 2-SAT backdoor, can we avoid brute-forcing the backdoor”? (a backdoor is a set of variables such that setting them simplifies the formula) (iii) modulator to graph width versions, such as “If we have a SAT formula with a modulator to a formula of constant treewidth, can we avoid brute-forcing the modulator?” (iv) weighted versions where the question is whether a satisfying assignment of weight $k$ exists (v) maximization versions where the objective is to maximize the number of satisfied clauses.
At this point it is worth pausing and trying to reassure a worried reader. One of the weak points of fine-grained complexity theory is the proliferation of reasonable-sounding complexity hypotheses, with no obvious connection to each other. It is natural to think that our decision to study a relatively large and diverse collection of variants of the SETH can only make things worse. Thankfully, the main take-home message of our results is that, even though we consider quite a few variations, we are able to identify only five equivalence classes (one of which contains the SETH itself), which align neatly with classes from classical complexity theory. In a sense, the conclusion we draw is optimistic as our results indicate that there is a great deal of complexity-theoretic structure waiting to be discovered here.
Let us then summarize our results, referring also to Figure˜1. We identify the following equivalence classes of complexity hypotheses:
  1. 1.
    Possibly the most interesting of the classes we consider is the middle one of Figure˜1, which is motivated by the NC-SETH. Even though the NC-SETH is stated for circuits of poly-logarithmic depth in [3], in practice it is often understood to apply to circuits of depth $n^{o(1)}$ (e.g. [10]). We take this reasoning to its logical conclusion and formulate this hypothesis for circuits of depth $\Omega(n)$, or more precisely for depth at most $\varepsilon n$ (we call this LD-C-SETH, for Linear-Depth Circuits). We discover that this hypothesis is equivalent to stating that brute force is unavoidable for CNF formulas supplied with a modulator to one of three standard widths: (i) constant pathwidth (ii) constant treewidth or (iii) logarithmic tree-depth. We find it somewhat surprising that these parameters, which are know to strictly generalize each other in the graph-theoretic sense, collapse to the same hypothesis. We also show that disproving the LD-C-SETH is equivalent to avoiding brute force for the W[SAT]-complete problem of finding a weight-$k$ satisfying assignment to a given (general, non-CNF) formula. This class corresponds to the middle box of Figure˜1. An important remark here is that we are able to extend these results to MaxSAT (see Section˜4.4), as this requires a foray into arithmetic circuits and is crucial later on to obtain the precise characterization of the complexity of Independent Set on graphs which are close to being cographs or block graphs. This class therefore seems to be quite robust and contains some very natural equivalent problems unrelated to SAT.
  2. 2.
    Going further, we consider the SETH for circuits without any restriction (Circuit-SETH), an assumption which seems much more plausible than the SETH. We show that avoiding brute force in this case can be done if and only if we can avoid brute force in one of two problems (i) deciding satisfiability for a CNF supplied with a Horn backdoor (ii) deciding weight-$k$ satisfiability for an arbitrary circuit, a problem complete for W[P]. This is the top-most class of Figure˜1.
  3. 3.
    Coming back to the LD-C-SETH, we consider the graph width parameters which are closest to the ones already considered. For modulators to logarithmic pathwidth (which would be the logical next step after logarithmic tree-depth) we show that avoiding brute-force enumeration is possible if and only if we can avoid it for 2-SAT backdoors. As we explain below, one of our main showcase results is to show that an equivalent characterization of this class can be given by Weighted Independent Set on $\textrm{Interval}+kv$ graphs, so this class also seems quite natural.
  4. 4.
    In the other direction, the logical next step below modulators to constant pathwidth is modulators to constant tree-depth. Here we show that the hypothesis that brute-force algorithms are optimal is in fact equivalent to the standard SETH. The parameter $(\sigma,\delta)$-hub introduced by [19] is therefore included here, as it is a restriction of bounded tree-depth modulators. In other words, our main result here is that hypotheses based on any parameter more restricted than “modulator to bounded tree-depth” all collapse to the normal SETH. This is the bottom class of Figure˜1. In a similar spirit we show that for Max-SAT, the standard version of the hypothesis (for the number of variables) and the version for the size of a $(\sigma,\delta)$-hub are equivalent (this is the second class from the bottom in Figure˜1).
Take-home message
We have identified five natural equivalence classes of hypotheses. In order to make sense of these results, let us attempt to draw an informal analogy with classical complexity theory. All the hypotheses we consider concern problems where the obvious algorithm is a brute-force “guess-and-check”; what differs is the difficulty of the verification procedure. If the verification procedure is P-complete (for example Horn-SAT or evaluating an arbitrary circuit) we end up in the top-most class of Figure˜1. Similarly, if it is NL-complete (equivalent to a reachability problem) we end up in the class right below, while the class of the LD-C-SETH seems to correspond to NC1. Finally, the class immediately above the SETH seems somewhat similar to (a restricted version of) TC0, since in Max-SAT what we need to do if given an assignment is count the number of satisfied clauses and compare this number with a target – this corresponds to the functioning of a threshold gate. What we are then essentially doing is reconstructing a parallel of a part of traditional complexity theory around fine-grained variations of the SETH and what remains is to populate these classes with further problems. Armed with this intuition, we can identify some further complexity questions as characteristic examples of each class.
Applications and Implications
One motivation we gave for this investigation was that we would like to increase our confidence in SETH-based lower bounds for problems parameterized by treewidth. One implication of our results is that we have managed to do so for many problems. Indeed, we show that improving upon the trivial algorithm for $q$-Coloring for parameters modulator size to constant pathwidth, constant treewidth, or logarithmic tree-depth is equivalent to the LD-C-SETH. This implies that improving upon the standard algorithm for pathwidth or treewidth is at least as hard as falsifying the LD-C-SETH (and hence the NC-SETH). As explained, this seems significantly more solid than a result based on the SETH. Furthermore, recent work by Lampis [36] showed that a number of fine-grained lower bounds for problems parameterized by pathwidth (including Independent Set, Dominating Set, Set Cover and others) are equivalent to the pw-SETH, the hypothesis that SAT cannot be solved in time $(2-\varepsilon)^{\textrm{pw}}|\phi|^{O(1)}$, where pw is the input formula’s primal pathwidth. Since the pw-SETH is implied by the same hypothesis for parameter modulator to logarithmic pathwidth, all the hypotheses we consider here except those of the top-most class of Figure˜1 also imply the lower bounds of [36]. We have thus succeeded in giving further evidence that known lower bounds for pathwidth and treewidth hold. For instance, even though it was previously known that no $(3-\varepsilon)^{\textrm{pw}}n^{O(1)}$ algorithm is possible for Dominating Set under the SETH, our results (together with those of [36]) establish this under the NC-SETH.
Beyond this, we give a characteristic example problem for each class, based on the intuition linking the classes of Figure˜1 with traditional complexity classes. In particular, we show that the Circuit-SETH is equivalent to beating brute-force for the problem of deleting $k$ vertices from a graph, so that it becomes $r$-degenerate, for any fixed $r\geq 2$; this problem is known to be P-complete. Similarly, we give a reachability-type problem ($k$-Neighborhood-Cut) that is equivalent to the 2SatB-SETH, which is natural, since reachability-type problems are usually NL-complete.
These applications are of particular interest for the two bottom classes in light of the results of [19]. In this work it was shown that if the SETH is true, it is impossible to avoid brute force for (among others) $q$-Coloring and Max-Cut, parameterized by $(\sigma,\delta)$-hub size. This perhaps creates the impression that we should have the same amount of confidence in the two lower bounds. Contrary to this, we show that their lower bound for $q$-Coloring is in fact equivalent to the SETH, while their lower bound for Max-Cut is equivalent to the MaxSat-SETH. If it were possible to show that these lower bounds are equivalent to each other, this would also establish that the Max-3-SAT-SETH (the restriction of MaxSat-SETH to CNFs of arity $3$) implies the SETH. As pointed out in [19], this is a major open problem. Our results therefore uncover some (perhaps unforeseen) nuance in the lower bounds of [19].
Relation with Time Complexity Classes.
It is worth stressing that, even though we present Figure˜1 as a collection of “complexity classes”, and indeed we draw a parallel with classical such classes, the results we present are in a sense orthogonal to normal notions of time complexity. What we mean by this is that, as can be seen in Figure˜1, we often show equivalence for problems which (are widely believed to) belong in distinct classes, namely, problems which are FPT on the one hand and W[SAT]- or W[P]-complete on the other. In other words, the equivalence classes of Figure˜1 do not indicate that all the problems in each class have the same time complexity; but rather that if we could improve the running time of the best algorithm for one, we could improve it for the others, albeit obtaining different types of running times for each problem. We remark that a similar situation holds for the pw-SETH [36].
logpw-SETHLD-C-SETHMax-SAT-SETHSETHWIS Interval+kvWIS Cograph+kvWIS Block+kvIS Cograph+kvIS Block+kvTheorem˜8.18BinWIS Cluster+kvWIS Cluster+kvIS Cluster+kvTheorem˜8.10Theorem˜8.12Theorem˜8.16Theorem˜8.14Theorem˜8.8Theorem˜8.6[19]
Figure 2: Summary of our results for (Weighted) Independent Set parameterized by vertex deletion distance to four graph classes: block, cluster, interval graphs, and cographs. Arrows indicate implication and a problem name indicates the hypothesis that brute-force is inevitable for this case. We consider unweighted versions (IS), unary weighted (WIS) and binary weighted (BinWIS).
A showcase application.
At this point we have established a framework based on five version of the SETH which are, we believe, quite natural and roughly corresponds to five complexity classes, endowed with nice complete problems. Nevertheless, it needs to be said that complexity classes can only be considered as interesting as the problems they contain or capture, or as Papadimitriou and Yannakakis put it, classes that contain interesting problems are discovered not invented [45]. Trying to verify that this is the case for our classes, we return to the motivating example this paper began with: (Weighted) Independent Set on $\mathcal{C}+kv$ graphs, where $\mathcal{C}$ are the graph classes we cherry-picked: cluster graphs, cographs, block graphs, and interval graphs. Our results are summarized in Figure˜2. We identify the LD-C-SETH as a dividing line for the complexity of this problem: for $\textrm{Cluster}+kv$ graphs the problem is at most as hard as the LD-C-SETH, even if binary weights are allowed, while for $\textrm{Cograph}+kv$ and $\textrm{Block}+kv$ the problem is exactly as hard as the LD-C-SETH, in both the unweighted case and the weighted case (with unary weights). However, Weighted Independent Set (again with unary weights) appears to be strictly harder for $\textrm{Interval}+kv$ graphs, as we show that this version is equivalent to the $\log\!\textrm{pw}$M-SETH. Observe that in this way the framework allows us to make comparisons which are not obvious graph-theoretically: cographs, block graphs, and interval graphs are incomparable classes, so a priori one would not have necessarily expected cographs and block graphs to be equivalent but interval graphs to be harder. Of course, a broader advantage is that integrating these questions in our framework allows us to compare these questions with many other, seemingly unrelated questions (such as whether $q$-Coloring can be solved faster than brute-force on graphs with a small modulator to constant pathwidth, which we discussed previously). We believe these results show that this is a direction that merits further investigation.
1.3 Overview of Techniques
Let us now sketch the main ideas used to obtain the results mentioned above. In all cases, a large part of the intuition needed is obtained by the parallels we drew between these classes and traditional complexity classes. More precisely:
  1. 1.
    For the hypotheses which turn out to be equivalent to the LD-C-SETH, one of the main technical tools we rely on is Barrington’s theorem [6], which allows us to transform a circuit on $n$ inputs and depth $\varepsilon n$ into a branching program of constant width and length exponential in $\varepsilon n$. This program can then be simulated by a formula with constant pathwidth and a modulator representing the inputs. For the converse direction we need to show that an algorithm disproving the LD-C-SETH can be used to decide the satisfiability of a CNF $\phi$ with a modulator $M$ of size $m$, such that deleting $M$ produces a formula of tree-depth at most $O(\log|\phi|)$. For this, we need to eliminate all the variables of the formula except those of $M$. We show how to do this in a way that is exponential in the tree-depth (and hence polynomial in the size of the formula). A particular technical challenge here is that, to generalize this result to MaxSAT (which will be crucial for later applications) we need to extend these arguments to arithmetic circuits and then move back to boolean circuits to obtain equivalence with the LD-C-SETH.
  2. 2.
    For the equivalence between 2SatB-SETH and $\log\!\textrm{pw}$M-SETH we rely on the intuition that both 2-SAT and satisfiability for SAT formulas of pathwidth $O(\log|\phi|)$ are essentially reachability problems. However, a closer look reveals that a direct reduction is non-obvious, as deciding 2-SAT is in fact more readily seen to be equivalent to non-reachability. We therefore need to transfer the classical Immerman–Szelepcsényi theorem, which established that NL=coNL, to our setting [29].
  3. 3.
    For the equivalences related to Circuit-SETH we rely on the intuition that computing the value of a circuit and deciding the satisfiability of a Horn formula are both P-complete problems.
  4. 4.
    For the equivalences given in the two bottom classes, the hard part is showing that (Max-)SAT parameterized by a modulator (to constant tree-depth, or a $(\sigma,\delta)$-hub) can be reduced to the same problem parameterized by the number of variables, that is, that we can eliminate the extra variables. For SAT this is done by an exhaustive application of the standard resolution rule, starting from the leaves of the rooted trees, while for Max-SAT parameterized by $(\sigma,\delta)$-hub, this is done by enumerating all assignments to each (bounded-size) component and replacing it with new constraints to its (bounded-size) neighborhood.
  5. 5.
    Finally, for our applications to Independent Set on graph classes, we need to rely on several unrelated ideas. In particular, we use standard but non-obvious facts from circuit complexity, such as the fact that $n$ $n$-bit integers can be summed by a bounded fan-in circuit of depth $O(\log n)$; we use the fact that the circuits of the LD-C-SETH can be considered acyclic, which in turn allows us to encode them using the tree structure hidden in cographs and block graphs; we use the fact that the natural DP algorithm for Weighted Independent Set on interval graphs can be translated to a reachability question to obtain equivalence with the $\log\!\textrm{pw}$M-SETH in this case.
1.4 Previous Work
[…]
The question of whether the SETH is the “right” assumption has appeared before in the literature. In general fine-grained complexity this manifests itself via works which attempt to replace the SETH with more plausible assumptions, such as the ones mentioned above, or which attempt to prove the same lower bound based on several hypotheses which have no known relation [2, 4]. Furthermore, the question of which problems are exactly equivalent to the SETH has already been posed, though few such problems are known [15]. We also note that the connection between the NC-SETH and Barrington’s theorem was already used in [10].
In the area of parameterized complexity, investigations on whether the SETH is the right complexity assumption have mostly focused on whether, assuming the SETH we can obtain the same lower bounds but for more restricted cases than for treewidth ([19, 32]). However, a strongly related recent work [36] investigated the so-called pw-SETH, that is, the hypothesis that SAT cannot be solved in time $(2-\varepsilon)^{\textrm{pw}}|\phi|^{O(1)}$. Even though this is closely related to our motivation, the pw-SETH turns out to cover a completely different algorithmic paradigm, as it essentially speaks about the optimality of dynamic programming algorithms, while the hypotheses we consider here refer to the optimality of branching algorithms under different regimes. As a result, most of the hypotheses we consider here imply the pw-SETH, with the exception of Circuit-SETH, for which this remains an interesting open problem.
1.5 Discussion and Directions for Further Work
We believe that the results of this paper, together with recent related works ([19, 36]) point the way to a new and interesting research direction investigating the interplay between fine-grained complexity theory and parameterized input structure. The envisioned goal here is a theory where lower bound statements form a comprehensible hierarchy where we can eventually pinpoint the reasons that a lower bound is believed to hold; for example we have seen that the LD-C-SETH is essentially equivalent to improving upon a generic “guess-and-check” algorithm with an NC1 verifier. For which other algorithmic paradigms can we identify such natural complete problems? Equally importantly, how do different paradigms relate to each other? In particular, the pw-SETH informally captures the difficulty of improving upon a generic DP algorithm over a linear structure, and the Circuit-SETH the difficulty of improving upon a generic “guess-and-check” algorithm with a polynomial-time verifier. Discovering a connection between these two hypothesis would allow us to better understand which paradigm is more powerful; conversely the two may be incomparable, and it would be interesting to see if this can be expressed via the corresponding hypotheses.
More broadly, an interesting research direction would of course be to further populate the classes introduced in this paper. Specific questions include:
  1. 1.
    Characterize the complexity of Independent Set parameterized by a modulator to an unweighted interval graph. In the weighted case, the problem is equivalent to the $\log\!\textrm{pw}$M-SETH; could it be easier (for instance, equivalent to the LD-C-SETH) for unweighted graphs?
  2. 2.
    As mentioned, Jaffke and Jansen [32] showed that brute force is unavoidable for $q$-Coloring parameterized by a modulator to a path-forest. Our result improves upon this in one sense (because we show equivalence to the LD-C-SETH, rather than SETH-hardness), but is weaker in another (because we parameterize by a modulator to a graph of constant pathwidth, rather than a path). Can this gap be closed? We conjecture that $q$-Coloring parameterized by feedback vertex set or modulator size to a path-forest is in fact easier than the LD-C-SETH. Evidence for such a claim could be established, for instance, by coming up with a reduction from this problem to CNF-SAT.
  3. 3.
    In a similar vein, a natural question to ask is the precise complexity of Independent Set parameterized by feedback vertex set. Again, our results can probably be used to show hardness (indeed LD-C-SETH-equivalence) when we have a modulator to a graph of constant treewidth, but not for a graph of treewidth $1$. What is the smallest treewidth $c$ for which parameterizing Independent Set by vertex-deletion distance to treewidth $c$ is LD-C-SETH-equivalent?
1.6 Paper Organization
We first remark that we have tried to use a self-explanatory naming scheme, where X-SETH means the hypothesis that it is impossible to improve upon brute-force search for SAT parameterized by a parameter X, or when X is a complexity class, for the complete problem of X. Nevertheless, the definitions of all hypotheses used in this paper (and shown in Figure˜1) and the most basic relations between them are given in Appendix˜A. Even though Appendix˜A serves as a useful reference, we have made an effort to give precise statements in all theorems, making sure that potentially non-trivial details in the formulations of various hypotheses, such as the order of quantifications of some constants, are clearly spelled out. For the rest of the paper, after some preliminaries, we present the basic connections for the two bottom classes of Figure˜1 in Section˜3, and then present one class in each of Section˜4, Section˜5, and Section˜6. In Section˜7 we present one further characteristic example for each class. Finally, in Section˜8 we present our results on Independent Set summarized in Figure˜2.
2 Preliminaries
We assume the reader is familiar with the basics of parameterized complexity as given for example in [16]. We refer the reader there for the definitions of treewidth and pathwidth. A graph $G$ has tree-depth $k$ if there is a bijective mapping of the vertices of $G$ to the vertices of a rooted tree of depth $k$ so that for every edge $uv$, the images of $u,v$ have an ancestor-descendant relation. We sometimes call such a mapping a tree-depth decomposition. Following [19], for integers $\sigma,\delta$, a $(\sigma,\delta)$-hub of a graph $G$ is a set of vertices $M$ such that each component of GMG-M has size at most $\sigma$ and at most $\delta$ neighbors in $M$.
A modulator of a graph is a set of vertices such that removing it leaves a graph with a desired property. When $G$ is a graph and $M$ a modulator we will often write GMG-M to denote the graph obtained after removing the modulator.
We deal with several variations of SAT, the problem of deciding if a given formula has a satisfying assignment. We write $k$-SAT for the restriction to clauses of arity at most $k$ and Max-SAT (or Max-$k$-SAT) for the more general problem of determining the assignment that satisfies the maximum number of clauses. For these problems we allow repeated clauses in the input, or equivalently we allow the input to assign (unary) weights to the clauses. An assignment of weight $k$ is an assignment that sets exactly $k$ variables to True.
For a CNF formula we define its primal graph as the graph that has a vertex for each variable and edges between variables that appear in the same clause. When $\phi$ is a CNF formula and $M$ is a set of variables, we will write $\phi-M$ to denote the formula obtained by deleting every clause that contains a variable of $M$ and, slightly abusing notation, we will also use $\phi-M$ to refer to the resulting primal graph.
We will also consider satisfiability problems for general formulas and circuits. A boolean circuit is a DAG where vertices of positive in-degree are labeled as disjunctions ($\lor$), conjunctions ($\land$), or negations ($\neg$) and there is a unique sink, called the output gate. Vertices of zero in-degree are called input gates and the satisfiability problem asks if there is a truth assignment to the inputs that makes the output gate evaluate to True. A formula is a circuit where the maximum out-degree of any non-input gate is $1$. Unless otherwise stated we will assume that the fan-in (in-degree) of all circuits is bounded, but in most cases unbounded fan-in gates can be replaced by trees of bounded fan-in gates. The depth of a circuit is the largest distance from an input gate to the output. The size of a circuit is the number of arcs of the corresponding DAG. We will also consider arithmetic circuits, where the value of each gate is an integer rather than a boolean and will mainly be concerned with circuits where the gates implement the addition (++) and maximum ($\max$) operations, that is, they output the sum or maximum of their inputs respectively. Boolean circuits can be interpreted as arithmetic circuits via the natural correspondence that False is 0 and True is $1$. To make the converse transition, we will make use of the following known facts (see e.g. [11]).
Theorem 2.1.
Given an integer $n$ it is possible to construct in time polynomial in $n$ bounded fan-in boolean circuits with $n^{2}$ inputs, $O(n)$ outputs, size $n^{O(1)}$ and depth $O(\log n)$ solving the following problems: (i) computing the maximum of $n$ integers of $n$ bits each (ii) computing the sum of $n$ integers of $n$ bits each. Furthermore, there exists a bounded fan-in boolean circuit with $2n$ inputs, encoding two $n$-bit binary numbers, that has depth $O(\log n)$ and decides if the first number is larger than the second.
Using Theorem˜2.1 we will be able to convert an arithmetic circuit which only uses $\max$ and ++ gates into a boolean circuit, without increasing the depth by much, even if the $\max$ and ++ gates do not have bounded fan-in (because Theorem˜2.1 talks about computing the sum or maximum of $n$ inputs).
Given a CNF formula $\phi$, a strong backdoor set is a set of variables such that for each assignment to these variables we obtain a formula that belongs to a base class. We will only deal with strong backdoors, so we will simply refer to them as backdoors. Furthermore, we will deal with 2-SAT and Horn backdoors, where a formula is Horn if each clause contains at most one positive literal. Recall that 2-SAT is NL-complete and Horn-SAT is P-complete, therefore both problems are solvable in polynomial time. This means that for a given backdoor the satisfiability of $\phi$ can be decided in time $2^{b}|\phi|^{O(1)}$ by trying all assignments to the backdoor. The term backdoor was introduced in [51]. Note that backdoors and modulators are closely related concepts and indeed the notion of modulator to constant treewidth which we use in this paper has also been called a backdoor to constant treewidth [24]. For more information on backdoors see [23].
In all cases where it is relevant, we will assume that the modulator or backdoor related to the problem at hand is given in the input. This is because we want to concentrate on the complexity of solving SAT given the modulator, rather than the complexity of finding the modulator. This is standard in parameterized lower bounds, see e.g. [38].
We give a detailed list of the complexity hypotheses that appear in this paper as well as simple observations about their relations in Appendix˜A.
3 Problems Equivalent to (Max-)CNF-SAT
In this section we deal with two hypotheses: the standard SETH and its natural weakening for Max-SAT, which we refer to as the MaxSat-SETH. One of our main motivations is to take a more careful look at the results of [19]. Recall that in this recent work it was pointed out that the SETH is “too strong”, if one aims to obtain lower bound results regarding standard width parameters, such as treewidth. For this reason, [19] focuses on a parameter which they call $(\sigma,\delta)$-hub, which is a severe restriction of treewidth, pathwidth, and tree-depth modulators. Their main result is that many of the standard lower bounds go through (assuming the SETH) even for this very restricted parameter.
In this paper we offer a more nuanced view that clarifies a bit more the picture presented in [19]. In particular, we show that for some of their lower bound results, not only is the SETH not too strong, but actually the lower bounds they obtain are equivalent to the SETH. We achieve this mainly via Theorem˜3.1 below, which shows that the SETH is equivalent to the tdM-SETH, that is the hypothesis that it is impossible to avoid a brute-force enumeration for a given modulator to a graph of bounded tree-depth. Since $(\sigma,\delta)$-hub is a parameter that is a restriction of modulators to bounded tree-depth, we obtain that avoiding brute force enumeration for this parameter is also equivalent to the SETH (Corollary˜3.3). In other words, Theorem˜3.1 collapses three seemingly distinct parameters (the number of variables, the size of a $(\sigma,\delta)$-hub, and the size of a constant tree-depth modulator) into a single hypothesis. As we show in Section˜7, these results establish that the lower bound of [19] for Coloring is actually equivalent to the SETH, and this remains true even if we parameterize by tree-depth modulator rather than $(\sigma,\delta)$-hub.
Despite the above, our results do not indicate that the SETH is the “correct” assumption for all the lower bounds of [19]. In the second part of this section we deal with the MaxSat-SETH, the hypothesis stating the same bound as the SETH but for the problem of maximizing the number of satisfied clauses of a CNF formula. Since this is clearly a problem at least as hard as deciding if a CNF formula can be satisfied, this assumption is at least as plausible as the SETH. Indeed, this assumption is also implied by the so-called Max-3Sat-Hypothesis, which states the same time bound but for CNFs of maximum arity $3$. As mentioned in [19], the relation between the SETH and the Max-3Sat-Hypothesis is currently unknown, which can perhaps be seen as a reason to believe the MaxSat-SETH as more plausible than either of the other two hypothesis. Our second result in this section is that the MaxSat-SETH remains equivalent if we parameterize not by the number of variables but by $(\sigma,\delta)$-hub size (Theorem˜3.5). This allows us to establish in Section˜7 that lower bounds concerning optimization problems, such as the one on Max Cut parameterized by $(\sigma,\delta)$-hub in [19], are more likely to be equivalent to MaxSat-SETH, hence perhaps more believable, than lower bounds equivalent only to the SETH.
Before we go on, let us mention that the results of this section rely on a precise formulation of the SETH (and related hypotheses) which assumes hardness for $k$-CNF formulas, for sufficiently large $k$. We note that this is important in this section, because we are dealing with extremely restrictive graph parameters. Once we move on to Section˜4, it will immediately become possible to break down large clauses into smaller ones by introducing new variables, so the arity of the considered CNFs will no longer be of material importance (see Section˜A.2).
3.1 Equivalent Formulations of the SETH
Theorem 3.1.
The following statements are equivalent:
  1. 1.
    (SETH is false) There is an $\epsilon>0$ such that for all $k$ there is an algorithm that solves $k$-SAT in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$.
  2. 2.
    (tdM-SETH is false) There is an $\epsilon>0$ such that for all $c,k$ there is an algorithm that takes as input a $k$-CNF formula $\phi$ and a modulator $M$ of $\phi$ of size $d$, such that $\phi-M$ has tree-depth at most $c$, and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{d}|\phi|^{O(1)}$.
Proof 3.2.
The second statement clearly implies the first, as the set of all variables is a modulator to tree-depth 0. The interesting direction is to show that if we have a $k$-SAT algorithm running in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$, then this implies a faster algorithm parameterized by constant tree-depth modulator.
We are given a $k$-CNF formula $\phi$ and a modulator $M$ such that removing $M$ from the primal graph produces a graph of tree-depth $c$, for some fixed $c$. Assume that for the resulting graph after we remove $M$ we have a forest of rooted trees, where each tree has height at most $c$, and a bijection of the variables of $\phi$ (outside $M$) to the vertices of this forest so that two variables appear in the same clause if and only if one is an ancestor of the other.
Our strategy will be to use the standard resolution rule to eliminate from $\phi$ all the variables except those of the modulator. We will then apply the assumed algorithm for $k^{\prime}$-SAT, for some $k^{\prime}$ depending on $c$ and $k$.
Recall that resolution is the following rule: when a CNF formula contains two clauses $C,C^{\prime}$ such that for some variable $x$, $C$ contains $x$ and $C^{\prime}$ contains $\neg x$, then we can add to the formula the clause that contains all literals of $C\cup C^{\prime}$ except $x,\neg x$ without affecting satisfiability. We will say that we exhaustively resolve over variable $x$ if we (i) apply this rule to each pair of clauses $C,C^{\prime}$ containing $x,\neg x$ (ii) after doing so remove from $\phi$ all clauses that contain $x$ or $\neg x$.
Claim 1.
Exhaustively resolving over a variable $x$ does not affect the satisfiability of $\phi$.
{claimproof}
It is easy to see that the clauses added to $\phi$ do not affect satisfiability, as they can be inferred from existing clauses. To see that deleting all clauses that contain the variable $x$ does not affect satisfiability, we take two cases. If the new formula has a satisfying assignment $\sigma$ such that for all clauses $C$ which contain $x$ positive in $\phi$, a literal of $C$ other than $x$ is set to True by $\sigma$, then we can satisfy $\phi$ by extending $\sigma$ by setting $x$ to False. Otherwise, we claim that all clauses $C^{\prime}$ that contain $\neg x$ in $\phi$ have a literal other than $\neg x$ set to True by $\sigma$, so we can set $x$ to True and satisfy $\phi$. Indeed, we are now in a case where some clause $C$ which contains $x$ positive has all its literals set to False by $\sigma$, but this clause has been paired with all $C^{\prime}$ which contained $\neg x$. Since the new formula is satisfied by $\sigma$, each $C^{\prime}$ must contain a True literal.
Our algorithm is to apply for each $i\in[c]$, starting from $i=c$ and descending, exhaustive resolution to all the variables which correspond to vertices of the rooted trees which are at distance $i$ from the root. After each iteration we have eliminated a layer of leaves and decreased the tree-depth of $\phi-M$ by one, without affecting the satisfiability of the formula. We now show by induction that the result of this process after $c$ iterations is a formula with maximum arity $k^{\prime}\leq 2^{c}k$ whose number of clauses is at most $m^{2^{c}}$, where $m$ is the number of clauses of $\phi$.
We show this claim by induction on $c$. For $c=1$, $M$ is in fact a vertex cover of the primal graph. For each variable $x$ of the independent set we apply exhaustive resolution. Each new clause is made up of two old clauses, so has arity at most $2k$. Furthermore, if $x$ appears in $d(x)$ clauses, we will construct at most $d^{2}(x)/2$ clauses when resolving over $x$. The total number of new clauses is then at most $\sum_{x\not\in M}d^{2}(x)/2\leq m^{2}/2$, where we use the fact that $m\geq\sum_{x\not\in M}d(x)$, because the sets of clauses that contain two distinct variables not in $M$ are disjoint, and because of standard properties of the quadratic function. Since we can assume that $m$ is large enough, the total number of clauses we now have is at most $m+\frac{m^{2}}{2}\leq m^{2}$.
We now proceed with the inductive step. Suppose that the claim is true for smaller values of $c$ and we have an instance of maximum arity $k$ and $m$ clauses. By the same arguments as before, applying exhaustive resolution to all leaves produces an instance where $\phi-M$ has tree-depth c1c-1, the maximum arity is at most $k^{\prime}\leq 2k$ and the number of clauses is at most $m^{\prime}=m^{2}$. By inductive hypothesis, completing the procedure will produce a formula with arity at most $2^{c-1}k^{\prime}\leq 2^{c}k$ and with at most $(m^{\prime})^{2^{c-1}}\leq m^{2^{c}}$ clauses, as desired.
In the end we have a formula with $n=|M|$ variables, arity $k^{\prime}\leq 2^{c}k$ and size $|\phi^{\prime}|\leq|\phi|^{2^{c}}$. Running the presumed satisfiability algorithm on this formula implies the theorem.
Corollary 3.3.
There is an $\epsilon>0$ such that for all $k,\sigma,\delta>0$ there is an algorithm which takes as input a $k$-CNF formula $\phi$ and a $(\sigma,\delta)$-hub of size $m$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$ if and only if the SETH is false.
Proof 3.4.
The proof consists of observing that a $(\sigma,\delta)$-hub is a modulator to tree-depth at most $\sigma$ and invoking Theorem˜3.1.
3.2 Equivalent Formulation of the SETH for Max-SAT
Theorem 3.5.
The following are equivalent:
  1. 1.
    (MaxSat-SETH is false) There exists $\varepsilon>0$ such that for all $k$ there is an algorithm that takes as input a $k$-CNF $\phi$ on $n$ variables and computes an assignment satisfying the maximum number of clauses in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$.
  2. 2.
    ($(\sigma,\delta)H$-MaxSat-SETH is false) There exists $\varepsilon>0$ such that for all $k,\sigma,\delta$ there is an algorithm that takes as input a $k$-CNF $\phi$ and a $(\sigma,\delta)$-hub $M$ of size $m$ and computes an assignment satisfying the maximum number of clauses in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
Proof 3.6.
It is easy to see that the second statement implies the first, as the set of all variables is a $(\sigma,\delta)$-hub. Recall that we allow clauses to be repeated in the input, so the running time of statement 1 cannot be written as $O((2-\varepsilon)^{n})$, because even for $k$-CNF, for $k$ fixed, the size of the formula cannot be bounded by a function of $n$ alone.
For the interesting direction, we consider one by one the components of $\phi-M$, each of which has size at most $\sigma$ and at most $\delta$ neighbors in $M$. Without loss of generality we assume that each component has size exactly $\sigma$ and exactly $\delta$ neighbors, by adding dummy variables or neighbors as needed.
Take such a component $C$ and for each assignment $s$ to its $\delta$ neighbor variables in $M$, calculate the assignment to the variables of $C$ that maximizes the number of clauses involving variables of $C$ which are satisfied (this can be done by going through the $2^{\sigma}$ possible assignments to the variables of the component). Let $w_{s}$ be the number of such clauses. Our goal is to remove the variables of $C$ from $\phi$ and add some clauses involving the variables of $N(C)$ such that for each assignment $s$ the total weight of satisfied new clauses is proportional to $w_{s}$. More precisely, we do the following: for each assignment $s$ to the variables of $N(C)$ we construct a clause falsified by this assignment (and only this assignment) to $N(C)$, and give it weight $\sum_{s^{\prime}\neq s}w_{s^{\prime}}$, that is give it weight that is equal to the sum of the weights corresponding to all other assignments to $N(C)$. We now observe that if we use the assignment $s$ for $N(C)$, the total weight we obtain from the new clauses is $w_{s}+(2^{\delta}-2)\sum_{s^{\prime}}w_{s^{\prime}}$, that is, the contribution of each assignment (including $s$) is counted $2^{\delta}-2$ times, but $w_{s}$ is counted once more, as a $w_{s}$ term appears in the weight of all satisfied clauses. We increase the target value of the Max-SAT instance by $(2^{\delta}-2)\sum_{s^{\prime}}w_{s^{\prime}}$ and implement the weights by taking copies of the constructed clauses. After performing this process exhaustively we are left with a Max-SAT instance with maximum arity $\max\{k,\delta\}$ and size at most $O(2^{\sigma+\delta}|\phi|)$, so executing the supposed algorithm gives the lemma.
4 Modulators to Constant Width and Linear Depth Circuits
In this section we present a sequence of reductions which show that several seemingly distinct fine-grained complexity hypotheses are actually the same. Our results are summarized in Theorem˜4.1 below, where we initially focus on the decision version of SAT (we discuss MaxSAT further below).
Theorem 4.1.
The following statements are equivalent:
  1. 1.
    (LD-C-SETH is false) There is an $\varepsilon>0$ and an algorithm which takes as input a (bounded fan-in) boolean circuit of size $s$ on $n$ inputs with depth at most $\varepsilon n$ and decides if the circuit is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
  2. 2.
    (pwM-SETH is false) There is an $\varepsilon>0$ such that for all fixed $c>0$ we have an algorithm which takes as input a CNF formula $\phi$ and a size $m$ modulator of $\phi$ to pathwidth at most $c$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
  3. 3.
    (twM-SETH is false) The same as the previous statement but for a modulator to treewidth at most $c$.
  4. 4.
    ($\log\!\textrm{td}$M-SETH is false) The same as the previous statement, but for a modulator to tree-depth at most $c\log|\phi|$.
  5. 5.
    (W[SAT]-SETH is false) There is an $\varepsilon>0$ such that there is an algorithm which takes as input a general Boolean formula $\phi$ on $n$ variables and an integer $k$ and decides if $\phi$ has a satisfying assignment that sets exactly $k$ variables to True in time $n^{(1-\varepsilon)k}|\phi|^{O(1)}$.
In other words, we make connections between three types of hypotheses: one involving the difficulty of satisfying a circuit (with a mild restriction on its depth); three different hypotheses involving structural graph parameters; and a hypothesis on the complexity of the standard complete problem for the class W[SAT]. Before we proceed to the proofs, let us give some high-level intuition regarding these connections.
Among the five statements of Theorem˜4.1 three have an obvious chain of implications: solving SAT parameterized by the size of a modulator to logarithmic tree-depth is at least as hard as solving it for a modulator to constant treewidth, which is at least as hard as for a modulator to constant pathwidth. These relations follow readily from the fact that $\textrm{pw}\leq tw$ and $\textrm{td}\leq\textrm{tw}\log n$ for all graphs. What is perhaps a bit surprising is that these three parameters turn out to be of equivalent complexity.
[…]
Proof 4.2 (Proof of Theorem˜4.1).
The implications $4\Rightarrow 3\Rightarrow 2$ follow immediately from the fact that on all $n$-vertex graphs we have $\textrm{tw}\leq\textrm{pw}$ and $\textrm{td}\leq\textrm{tw}\log n$. The implication $5\Rightarrow 4$ is given in Lemma˜4.3. The implication $2\Rightarrow 1$ is given in Lemma˜4.7. The implication $1\Rightarrow 5$ is given in Lemma˜4.9.
MaxSAT
In addition to the above we extend the results of Theorem˜4.1 to show equivalence also with MaxSAT parameterized by the size of a modulator to logarithmic tree-depth (hence also constant pathwidth and treewidth). We present this result in Section˜4.4. This extension is non-trivial, as it requires to generalize the basic idea used in Theorem˜4.1 (namely, eliminating variables using a tree-depth decomposition) from the realm of boolean to the realm of arithmetic circuits, because now to select the best assignment for a variable we have to calculate the number of satisfied clauses (this naturally corresponds to a circuit with $\max$ and ++ gates). The main challenge is to reduce such circuits to boolean ones without significantly increasing their depth, so that we can still obtain equivalence with the LD-C-SETH, and for this we need to rely on classical facts from circuit complexity such as the fact that computing the sum of $N$ integers on $N$ bits can be done with boolean circuits of depth only $O(\log N)$ (mentioned previously in Theorem˜2.1).
4.1 W[SAT] to Pathwidth, Treewidth, Logarithmic Tree-depth Modulators
Lemma 4.3.
Statement 5 of Theorem˜4.1 implies statement 4.
Proof 4.4.
We suppose there exist constants $\varepsilon>0,d>0$ and an algorithm that takes as input a Boolean formula $\psi$ (that is, a Boolean circuit of maximum out-degree $1$) with $N$ inputs and an integer $k$ and decides if $\psi$ has a satisfying assignment of weight $k$ in time $N^{(1-\varepsilon)k}|\psi|^{d}$.
Fix now a constant $c>0$ and we are given a CNF formula $\phi$ and a modulator $M$ of size $m$ such that removing all vertices of $M$ from the primal graph of $\phi$ results in a graph of tree-depth $h$ with $h\leq c\log|\phi|$. Suppose we are given a rooted tree of height $h$ such that the variables of $\phi$ (outside $M$) can be bijectively mapped to nodes of the tree in a way that if two variables appear in a clause together, then one is an ancestor of the other. We want to decide the satisfiability of $\phi$ in time $(2-\varepsilon^{\prime})^{m}|\phi|^{O(1)}$.
Our high-level plan is to construct from $\phi$ a Boolean formula $\psi$ and an integer $k$ satisfying the following:
  1. 1.
    $\phi$ is satisfiable if and only if $\psi$ has a satisfying assignment of weight $k$.
  2. 2.
    $\psi$ has $N=O(2^{m/k})$ input variables and can be constructed in time $2^{m/k}|\phi|^{O(1)}$.
In particular, we will set $k=\lceil\frac{2d}{\varepsilon}\rceil$. Then $N^{(1-\varepsilon)k}|\psi|^{d}$ can be upper bounded as follows: $N^{(1-\varepsilon)k}=O(2^{(1-\varepsilon)m})$ and $|\psi|^{d}=2^{dm/k}|\phi|^{O(1)}$. But $\frac{d}{k}\leq\frac{\varepsilon}{2}$, therefore the algorithm runs in $2^{(1-\frac{\varepsilon}{2})m}|\phi|^{O(1)}$.
We now describe how to construct $\psi$ from $\phi$, assuming we have the rooted tree of height $h\leq c\log|\phi|$ we mentioned covering the primal graph of $\phi$ outside of $M$. We first describe the construction of an auxiliary formula $\psi_{r}$, whose input variables are only the variables of $M$, where $r$ is the root of the tree of height $h$. The formula will be built inductively, by constructing a formula $\psi_{b}$ for each node $b$ of the tree, starting from the leaves. In the remainder we use, for a node $b$ of the tree, $h_{b}$ to denote the height of $b$ in the tree, that is its maximum distance from any of its descendants, and $\phi_{b}$ to denote the formula made up of the clauses of $\phi$ that contain a variable associated with $b$ or one of its descendants. We can assume without loss of generality that all clauses of $\phi$ contain a variable outside of $M$ (if not, replace $C$ with $(C\lor x),(C\lor\neg x)$, where $x$ is a new variable outside $M$), hence we have $\phi_{r}=\phi$ and also if $b,b^{\prime}$ are nodes such that $h_{b}=h_{b^{\prime}}$, then $\phi_{b}$ and $\phi_{b^{\prime}}$ are made up of disjoint sets of clauses of $\phi$. To see this, if there were a clause that contains a variable associated with $b$ (or one of its descendants) and $b^{\prime}$ (or one of its descendants) this would add to the primal graph an edge that is not between an ancestor and a descendant, contradiction.
The formula $\psi_{b}$ will have the following properties:
  1. 1.
    The input variables of $\psi_{b}$ are the variables of $M$ and all the variables associated with ancestors of $b$ in the tree.
  2. 2.
    $\psi_{b}$ is satisfied by an assignment to its input variables if and only if it is possible to extend this assignment to the variables associated with $b$ and its descendants, so that all clauses of $\phi_{b}$ are satisfied.
  3. 3.
    $|\psi_{b}|=O(2^{h_{b}}|\phi_{b}|)$.
As mentioned, $\psi_{r}$ will be the formula obtained by this procedure for the root of the tree and will only have $M$ as input variables.
In order to construct $\psi_{b}$ we proceed by induction on $h_{b}$. If $h_{b}=0$ then $\psi_{b}$ is constructed by taking $\phi_{b}$, and obtaining two formulas $\phi_{b}^{0},\phi_{b}^{1}$ by setting the variable associated with $b$ to False, True, respectively. We set $\psi_{b}=\phi_{b}^{0}\lor\phi_{b}^{1}$. Clearly, the three properties are satisfied.
Now suppose $h_{b}>0$ and let $b_{1},\ldots,b_{s}$ be the children of $b$ in the tree. Let $x$ be the variable associated with $b$. We have already constructed formulas $\psi_{b_{1}},\ldots,\psi_{b_{s}}$. For each $i\in[s]$ let $\psi_{b_{i}}^{0}$ (respectively $\psi_{b_{i}}^{1}$) be the formula obtained from $\psi_{b_{i}}$ if we set $x$ to False (respectively True) and simplify the formula accordingly. Let $C_{b}$ be the set of clauses of $\phi$ containing $x$ but no variable associated with a descendant of $b$. Let $C_{b}^{0}$ (respectively $C_{b}^{1}$) be the CNF formula made up of the clauses of $C_{b}$ we obtain if we set $x$ to False (respectively True), that is $C_{b}^{0}$ is the set of clauses containing $x$ positive, but with $x$ removed (similarly for $C_{b}^{1}$). We define $\psi_{b}$ as follows:
$\psi_{b}=\left(C_{b}^{0}\land\bigwedge_{i\in[s]}\psi_{b_{i}}^{0}\right)\lor\left(C_{b}^{1}\land\bigwedge_{i\in[s]}\psi_{b_{i}}^{1}\right)$
Let us argue why $\psi_{b}$ satisfies the three properties. First, $\psi_{b}$ has the right set of variables, as we have eliminated $x$ from all formulas.
Second, suppose $\psi_{b}$ is satisfied by some assignment to its input variables and suppose without loss of generality that the first term of the disjunction is satisfied (the other case is symmetric), therefore $C_{b}^{0}$ is satisfied. We set $x$ to False, and thus satisfy all of $C_{b}$. For all $i\in[s]$ we have that $\psi_{b_{i}}^{0}$ is satisfied, therefore $\psi_{b_{i}}$ is satisfied by the current assignment to $M$ and the ancestors of $b$, extended with $x$ set to False. By induction, this can be extended to an assignment satisfying $\phi_{b_{i}}$. Since we have satisfied $C_{b}$ and $\phi_{b_{i}}$ for all $i\in[s]$, we have satisfied $\phi_{b}$. For the converse direction, if it is possible to extend some assignment to $M$ and the ancestors of $b$ so that $\phi_{b}$ is satisfied, then this assignment satisfies $\psi_{b}$. Indeed, suppose that in the assumed extension $x$ is set to False (again the other case is symmetric). This implies that all the clauses of $C_{b}^{0}$ must be satisfied, as they contain $x$ positive. Furthermore, by induction, since we can satisfy $\phi_{b_{i}}$ (for all $i\in[s]$) by setting $x$ to False, we also satisfy $\psi_{b_{i}}$ by setting $x$ to False, hence $\psi_{b_{i}}^{0}$ evaluates to True. Therefore, the first term of the disjunction of $\psi_{b}$ evaluates to True.
Third, to bound the size of $\psi_{b}$ we have $|C_{b}^{0}|+|C_{b}^{1}|\leq 2|\phi_{b}|$ and we also have $2\sum_{i\in[s]}\psi_{b_{i}}=O(2^{b_{h}}\sum_{i\in[s]}|\phi_{b_{i}}|)=O(2^{b_{h}}|\phi_{b}|)$, (where we used the fact that the $\phi_{b_{i}}$ are disjoint) giving the desired bound. We therefore have $|\psi_{r}|=|\phi|^{O(c)}=|\phi|^{O(1)}$.
We now transform $\psi_{r}$ into a formula $\psi$ such that $\psi$ has a satisfying assignment of weight $k$ if and only if $\psi_{r}$ has a satisfying assignment. To do so, we partition the $m$ input variables of $\psi_{r}$ into $k$ sets $V_{i}$, $i\in[k]$, each of size at most $\lceil m/k\rceil$. For each assignment $\sigma$ to the variables $V_{i}$ we construct a new variable $x_{i,\sigma}$. For each $i\in[k]$ we add a clause $\left(\bigvee_{\sigma}x_{i,\sigma}\right)$. For each $y\in V_{i}$ let $\sigma^{1}(y)$ (respectively $\sigma^{0}(y)$) be the set of assignments to $V_{i}$ which set $y$ to True (respectively False). Then, we replace every appearance of $y$ in $\psi_{r}$ with the clause $\bigvee_{\sigma\in\sigma^{1}(y)}x_{i,\sigma}$ and every appearance of $\neg y$ with the clause $\bigvee_{\sigma\in\sigma^{0}(y)}x_{i,\sigma}$. Note that now we have $|\psi|=O(2^{m/k}|\psi_{r}|)=2^{m/k}|\phi|^{O(1)}$, as promised. Furthermore, it is not hard to see that $\psi$ can only be satisfied by an assignment of weight $k$ if we select for each $i\in[k]$ a unique variable $x_{i,\sigma}$ to set to True, thus creating a one-to-one correspondence between satisfying assignments to $\phi$ and weight-$k$ satisfying assignments to $\psi$.
4.2 Linear-depth Circuits and Constant Pathwidth Modulators
In this section we present Lemma˜4.7 which, as mentioned will rely on the transformation of Boolean circuits to bounded-width branching programs given by Barrington’s theorem. Before we begin let us recall some definitions and background.
Definition 4.5.
A layered branching program of width $w$ and length $\ell$ with input variables $x_{1},\ldots,x_{n}$ is a DAG with the following properties:
  1. 1.
    Vertices are partitioned into $\ell$ sets $V_{1},\ldots,V_{\ell}$ (the layers), such that each arc goes from a vertex in $V_{i}$ to a vertex in $V_{i+1}$. For each $i\in[\ell]$ we have $|V_{i}|\leq w$.
  2. 2.
    There is a unique source (the start vertex), while $V_{\ell}$ contains two vertices (the accept and reject vertices). All other vertices have strictly positive in- and out-degree.
  3. 3.
    Each arc is labeled True or False and each vertex with positive out-degree is labeled with an input variable.
We say that a branching program accepts an assignment $\sigma$ to its variables if when we start at the source and we follow a path where for each vertex labeled $x_{i}$ we follow the out-going arc labeled $\sigma(x_{i})$, we reach the accept vertex.
We now recall Barrington’s classical result.
Theorem 4.6.
([6]) There is an algorithm which, given a boolean circuit on $n$ inputs with depth $d$, produces an equivalent branching program on the same inputs with width $5$ and length $O(4^{d})$ in time $O(4^{d})$.
Lemma 4.7.
Statement 2 of Theorem˜4.1 implies statement 1.
Proof 4.8.
The proof relies heavily on Theorem˜4.6: given an arbitrary Boolean circuit we will first construct an equivalent branching program and then convert this branching program into a CNF formula. The formula will have a modulator to small pathwidth because the branching program has constant width.
More precisely, suppose that for some $\varepsilon>0$ there exists an algorithm which for all $c>0$ takes as input a CNF formula $\phi$ and a modulator $M$ of its primal graph to pathwidth $c$, with $|M|=m$, and decides if $\phi$ is satisfiable in time $2^{(1-\varepsilon)m}|\phi|^{c^{\prime}}$, for some $c^{\prime}>0$. Let $\varepsilon^{\prime}=\frac{\varepsilon}{4c^{\prime}}$. We claim that we can obtain an algorithm which, given an $n$-input circuit $C$ of size $s$ and depth $\varepsilon^{\prime}n$ decides if $C$ is satisfiable in time $2^{(1-\varepsilon^{\prime})n}s^{O(1)}$.
We first invoke Theorem˜4.6 and obtain a branching program of length $O(4^{\varepsilon^{\prime}n})$ and width $5$. We then construct a CNF formula whose variables are the $n$ variables $x_{1},\ldots,x_{n}$ representing the inputs to the branching program, plus the $3\ell$ variables $y_{1,t},y_{2,t},y_{3,t}$, for $t\in[\ell]$. The intuitive meaning of these variables is that the assignment to $y_{1,t},y_{2,t},y_{3,t}$ should encode the vertex which the execution of the branching program will reach in layer $t$. We make an injective mapping from the vertices of $V_{t}$ to the assignments of $y_{1,t},y_{2,t},y_{3,t}$ and add clauses to exclude assignments not associated with a vertex of $V_{t}$. We then add clauses to ensure that arcs are followed: for each assignment to $y_{1,t},y_{2,t},y_{3,t},y_{1,t+1},y_{2,t+1},y_{3,t+1}$ as well as the at most five input variables used as labels for the vertices of $V_{t}$, we construct clauses ensuring the transitions conform to the arcs of the branching program. Finally, we add clauses ensuring that in layer $1$ we begin at the start vertex and in the last layer we reach that accept vertex. It is not hard to see that the constructed CNF formula is satisfiable if and only if there exists an assignment to the input variables accepted by the branching program, and hence the circuit.
We observe that in the formula we have constructed there is a modulator of size $n$ such that removing those variables results in a graph of pathwidth at most 5 (we can use a sequence of bags containing $\{y_{1,t},y_{2,t},y_{3,t},y_{1,t+1},y_{2,t+1},y_{3,t+1}\}$ for all $t$). If we execute the supposed algorithm on this instance the running time will be at most $2^{(1-\varepsilon)n}|\phi|^{c^{\prime}}$. We have $|\phi|=O(4^{\varepsilon^{\prime}n})=O(2^{\frac{\varepsilon n}{2c^{\prime}}})$. The algorithm therefore runs in $O(2^{(1-\frac{\varepsilon}{2})n})=O(2^{(1-\varepsilon^{\prime})n})$.
4.3 Linear-Depth Circuits and Weighted SAT
Lemma 4.9.
Statement 1 of Theorem˜4.1 implies statement 5.
Proof 4.10.
Suppose there exist $\varepsilon>0,c>0$ such that satisfiability of $n$-input circuits of depth $\varepsilon n$ and size $s$ can be decided in time $2^{(1-\varepsilon)n}s^{c}$.
We are given an $N$-variable formula $\phi$ and an integer $k$ and are asked to decide if $\phi$ has a weight-$k$ satisfying assignment in time $N^{(1-\varepsilon^{\prime})k}|\phi|^{O(1)}$. We can assume without loss of generality that $|\phi|\leq N^{\varepsilon k/6}$, otherwise the problem can be solved in $N^{k}|\phi|^{O(1)}=|\phi|^{O(1/\varepsilon)}=|\phi|^{O(1)}$ time by trying out all assignments. Therefore, $\log|\phi|\leq\frac{\varepsilon k\log N}{6}$.
We now consider $\phi$ as a boolean circuit. Using standard techniques, we can balance $\phi$ to obtain an equivalent circuit of depth at most $3\log|\phi|\leq\frac{\varepsilon k\log N}{2}$ (see e.g. [8]) while maintaining a size that is polynomial in $|\phi|$. We add to $\phi$ $k\lceil\log N\rceil$ new inputs, partitioned into $k$ groups of $\lceil\log N\rceil$ inputs, whose intuitive meaning is that they encode the indices of the $k$ inputs to $\phi$ which are meant to be set to True in a satisfying assignment. For each group we enumerate all $O(N)$ assignments and associate each with an input of $\phi$. We transform the $N$ input gates of $\phi$ into disjunction gates of fan-in $k$, whose inputs are conjunctions of fan-in $\lceil\log N\rceil$ connected to the new inputs in one of the $k$ groups via appropriate $\neg$ gates so that an input of $\phi$ is set to True if and only if its index is the assignment to one of the $k$ groups of new inputs. This construction can be made to have depth $O(\log\log N+\log k)$ by replacing high fan-in gates with binary trees. It is now not hard to see that the new circuit has a satisfying assignment that gives distinct assignments to each of the $k$ groups if and only if $\phi$ has a weight-$k$ satisfying assignment. What remains is to add for each distinct $i,i^{\prime}\in[k]$ some gadget that ensures that the assignments to groups $i,i^{\prime}$ are not identical (and hence we select an assignment of weight exactly $k$ for $\phi$). This can easily be done with $O(k^{2})$ formulas of size $O(\log N)$ so the total depth remains at most the depth of $\phi$ plus $O(\log\log N+\log k)$. We can therefore assume that the depth of the circuit we have now constructed is at most $\varepsilon k\log N$. The circuit has at most $k\log N+k$ inputs and has been constructed in polynomial time, so running the supposed satisfiability algorithm takes $2^{(1-\varepsilon)k\log N}2^{k}|\phi|^{O(1)}$ which for $N>2^{2/\varepsilon}$ is upper-bounded by $N^{(1-\varepsilon/2)k}|\phi|^{O(1)}$.
4.4 MaxSAT and Modulators of Logarithmic Tree-depth
We are now ready to show that the results of Theorem˜4.1 remain valid if instead of SAT we consider the more general MaxSAT problem. In particular, a fast satisfiability algorithm for circuits of depth $\varepsilon n$ would be sufficient not only to decide satisfiability of formulas which are close to having logarithmic tree-depth, but also to decide the maximum number of satisfiable clauses of such formulas. This will become crucial later, because the versions of Independent Set we consider will be relatively straightforward to reduce to MaxSAT on formulas which have a small modulator to instances of constant treewidth (and hence logarithmic tree-depth).
Theorem 4.11.
The following two statements are equivalent:
  1. 1.
    (LD-C-SETH is false) There is an $\varepsilon>0$ and an algorithm which takes as input a (bounded fan-in) boolean circuit of size $s$ on $n$ inputs with depth at most $\varepsilon n$ and decides if the circuit is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
  2. 2.
    (MaxSAT for logarithmic tree-depth modulators) There is an $\varepsilon>0$ such that for all fixed $c>0$ we have an algorithm which takes as input a CNF formula $\phi$, a target value $t$, and a size $m$ modulator of $\phi$ to tree-depth at most $c\log|\phi|$ and decides if there exists an assignment that satisfies at least $t$ clauses of $\phi$ in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
Proof 4.12.
The implication $2\Rightarrow 1$ follows immediately from Theorem˜4.1, because an algorithm that can solve MaxSAT can also solve SAT, implying the fourth statement of Theorem˜4.1. We therefore focus on the implication $1\Rightarrow 2$. Fix an $\varepsilon>0$ for which we have an algorithm that can decide the satisfiability of boolean circuits of depth $\varepsilon n$ in time $(2-\varepsilon)^{n}s^{O(1)}$.
Fix a $c>0$ and suppose we are given a formula $\phi$ with a modulator $M$ of size $m$ such that $\phi-M$ has tree-depth at most $c\log|\phi|$. We want to find an assignment that satisfies as many clauses of $\phi$ as possible in time $(2-\varepsilon^{\prime})^{m}|\phi|^{O(1)}$, for some $\varepsilon^{\prime}>0$. We will assume without loss of generality that, for all fixed $\delta>0$ we have $|\phi|<2^{\delta m}$, as otherwise we can enumerate all assignments to $M$ in time $2^{m}<|\phi|^{1/\delta}=|\phi|^{O(1)}$ and for each such assignment compute its best extension to the rest of the instance in time $\phi^{O(1)}$ (using the fact that the remaining instance has logarithmic tree-depth). By choosing an appropriate $\delta$ we can therefore assume that any quantity that is $O(\log|\phi|)$ is upper-bounded by $\varepsilon m$.
Let us first give a high-level overview of our approach. We will begin by constructing an arithmetic circuit using bounded fan-in $\max$ and ++ gates and depth $O(\log|\phi|)$. The idea is that the gates of the circuit will allow us to eliminate all the variables of $\phi$ except those of $M$. In particular, for each variable of $\phi-M$ we have a $\max$ gate (we select the better assignment for the variable) whose inputs are ++ gates with two inputs: the number of clauses immediately satisfied by giving a specific value to the variable, and the circuit corresponding to the remaining formula. Hence, feeding an assignment to the variables of $M$ to the circuit will compute the maximum number of clauses that can be satisfied by extending this assignment. Once we have this arithmetic circuit we would like to convert it into a boolean circuit, however doing this in a naive way will increase the depth to super-logarithmic in $|\phi|$. We will therefore first “compress” the circuit: we will replace the bounded fan-in gates with gates of fan-in $\log\log|\phi|$, in the process reducing the depth to $\log|\phi|/\log\log|\phi|$. We then use the fact that the sum or maximum of $N$ numbers of $N$ bits can be computed by a boolean circuit of depth $O(\log N)$, or in our case, the sum of $\log|\phi|$ numbers on $\log|\phi|$ bits can be computed by a boolean circuit of depth $\log\log|\phi|$. Plugging this circuit in the place of high fan-in gates produces the final boolean circuit of the desired depth.
Let us proceed to the construction. We will construct a boolean circuit with $n=m$ inputs and depth at most $O(\log|\phi|)<\varepsilon m$, on which we can apply the supposed algorithm. The circuit will be satisfiable if and only if there exists an assignment satisfying $t$ clauses of $\phi$, so we will obtain the implication.
We first construct an arithmetic circuit. In such a circuit the value of each gate is an integer and we will use the operations ++ and $\max$. We describe the construction of such a circuit by induction.
Our arithmetic circuit will have as inputs the variables of $M$ with the natural correspondence between boolean and integer values, namely, False is 0 and True is $1$. In order to allow us to encode literals we will also use $\neg$ gates, which given input $x$ output 1x1-x; these will only be used in the layer immediately after the inputs, while the rest of the circuit will consist only of ++ and $\max$ gates of fan-in $2$, as well as integer constants.
We make the following inductive claim: we can produce from $\phi$ and $M$ an arithmetic circuit $C$ which satisfies the following:
  1. 1.
    For each assignment $\sigma$ to $M$, if we feed $\sigma$ to the circuit $C$ we obtain as output the number of clauses satisfied by the best assignment of $\phi$ that is consistent with $\sigma$.
  2. 2.
    If the primal graph of $\phi-M$ is connected, then the depth of $C$ is at most $3\textrm{td}(\phi-M)+2\log|\phi|+1$.
  3. 3.
    If the primal graph of $\phi-M$ is not connected, then the depth of $C$ is at most $3\textrm{td}(\phi-M)+2\log|\phi|+2$.
Claim 2.
An arithmetic circuit satisfying the above can be constructed in polynomial time, given $\phi$ and $M$ as well as a tree-depth decomposition of $\phi-M$.
{claimproof}
We show the claim by induction. For the base case, suppose $M$ contains all variables of $\phi$. Then, we can construct a circuit as follows: for each clause, construct a $\max$ gate, and feed it as input all the literals of the clause (directly or through $\neg$ gates); then add a ++ gate that computes the sum of the gates constructed for the clauses. This circuit uses ++ and $\max$ gates of unbounded fan-in, but can easily be converted to one that uses gates of fan-in $2$ and depth at most $2\log|\phi|+1$ as desired.
We now have two inductive cases, depending on whether $\phi-M$ is connected or not. First, suppose that $\phi-M$ is connected, there exists therefore a variable $x\not\in M$ such that $\textrm{td}(\phi-M-\{x\})=\textrm{td}(\phi-M)-1$. We compute two formulas $\phi_{0},\phi_{1}$, by setting $x$ to 0 or $1$ respectively and simplifying $\phi$ (that is, deleting satisfied clauses and falsified literals). Inductively construct circuits $C_{0},C_{1}$ for the two formulas. The circuit $C$ has a $\max$ gate as output; we feed into this gate two inputs, calculated by ++ gates: the first ++ gate takes as input the output of $C_{0}$ and an integer equal to the number of clauses that contain the literal $\neg x$, while the other ++ gate takes as input the output of $C_{1}$ and an integer equal to the number of clauses that contain $x$. Correctness of the new circuit can be established by induction. The depth is at most $2+\max\{d_{0},d_{1}\}$, where $d_{0},d_{1}$ are the depths of $C_{0},C_{1}$ respectively. We have $d_{0}\leq 3\textrm{td}(\phi-M-\{x\})+2\log|\phi|+2\Rightarrow 2+d_{0}\leq 3\textrm{td}(\phi-M)+2\log|\phi|+1$, as desired.
Suppose then that $\phi-M$ is disconnected and the connected components correspond to formulas $\phi_{1},\phi_{2},\ldots,\phi_{\ell}$. Intuitively, we would like to construct the circuits $C_{1},\ldots,C_{\ell}$ and then feed their output to a ++ gate, but this is slightly complicated from the fact that we want to keep the fan-in at $2$. If one of the formulas, say $\phi_{1}$ is large, that is, $|\phi_{1}|\geq|\phi|/2$, then we can proceed by building one circuit $C_{1}$ for $\phi_{1}$ and another $C_{2}$ for $\phi_{2}\cup\ldots\cup\phi_{\ell}$, and feeding their outputs to a ++ gate. The circuit $C_{1}$ has depth $1$ less than the desired bound (as $\phi_{1}$ has a connected primal graph), while $C_{2}$ has sufficiently smaller depth as its size has decreased by a factor of at least $2$, compared to $\phi$.
We are therefore left with the case where $\phi-M$ is disconnected into formulas $\phi_{1},\ldots,\phi_{\ell}$, and all these subformulas have size at most $|\phi|/2$. Partition the formulas $\phi_{1},\ldots,\phi_{\ell}$ into two sets $L,R$ by initially setting $L=\{\phi_{1},\ldots,\phi_{\ell}\}$ and $R=\emptyset$ and then repeatedly selecting an arbitrary formula from $L$ and moving it to $R$ as long as we maintain the invariant that the total size of formulas in $L$ is at least $|\phi|/2$. We now are at a situation where the total size of formulas in $R$ is at most $|\phi|/2$; $L$ is such that if we remove any formula it contains then its size becomes at most $|\phi|/2$; all formulas $\phi_{i}$ have size at most $|\phi|/2$. We now construct three circuits: a circuit $C_{R}$ representing all formulas of $R$; a circuit $C_{1}$ representing an arbitrarily chosen formula $\phi_{i}\in L$; and a circuit $C_{L}$ representing $L\setminus\{\phi_{i}\}$. We feed the outputs of $C_{1}$ and $C_{L}$ into a ++ gate, and the output of that gate with the output of $C_{R}$ into an ++ gate, which is the final output. We observe that because all three circuits represent formulas whose size is at least a factor of $2$ smaller than $\phi$, their depths are at least $3$ less than the desired bound, so the two extra layers we added allow us to stay within the target depth.
We now show how the circuit we have produced can be turned into a circuit of smaller depth by increasing the fan-in. The reason this will be helpful is that the depth of a boolean circuit required to add $N$ integers of $N$ bits is asymptotically the same as for $2$ integers of $N$ bits, so increasing the fan-in will allow us to more efficiently convert the arithmetic circuit into a boolean circuit.
Claim 3.
There is a polynomial-time algorithm that takes as input an arithmetic circuit using $\max$ and ++ gates of depth $d$ and fan-in $2$ and outputs an equivalent circuit of depth $O(d/\log d)$ using $\max$ and ++ gates of fan-in at most $O(d)$.
{claimproof}
To simplify notation we assume that $d$ is a power of $2$. Suppose without loss of generality that the given circuit is arranged in layers, where layers alternate between $\max$ and ++ gates. Suppose the layers are numbered $1,\ldots,d$ and gates at layer $i$ have inputs in layer i1i-1. We will conserve the gates in one out of each $\log d$ layers, call these layers the selected layers.
Consider now a gate $g$ at a selected layer. Its value depends on the values of at most $2^{\log d}=d$ gates in the previous selected layer. Consider then the binary tree rooted at $g$ with the $d$ gates from the previous selected layer as leaves. We can replace this tree with a constant-depth circuit of fan-in $O(d)$. In particular, our new tree will have a $\max$ gate as output, which takes as input $O(d)$ ++ gates, which each take as input some subset of the $d$ gates of the previous selected layer. This can be done inductively: for each gate $g^{\prime}$ in the path from the previous selected layer to $g$ we construct such a bounded-depth circuit: if $g^{\prime}$ is a $\max$ gate we take the circuits constructed for its inputs and merge their outputs; while if $g^{\prime}$ is a ++ gate we apply the distributive property $\max\{a,b\}+\max\{c,d\}=\max\{a+c,a+d,b+c,b+d\}$ to maintain the property that the output is a $\max$ gate. Doing this for all gates in selected layers allows us to eliminate intermediate layers and obtain the desired depth bound.
We now recall (Theorem˜2.1) that one can perform additions and maximum calculations of $N$ numbers on $N$ bits with depth $O(\log n)$ boolean circuits. Putting this together with the two claims we obtain the proof of the statement. In particular, we initially constructed a bounded fan-in arithmetic circuit of depth $O(\log|\phi|)$ and then after applying the second claim and Theorem˜2.1 converted it to an equivalent boolean circuit of depth $O(\log|\phi|)$. We add to the end a boolean comparator that checks if the output is the binary encoding of at least the target value $t$. The whole circuit now has depth $O(\log|\phi|)$, which as we explained can be assumed to be at most $\varepsilon m$, so the supposed algorithm for circuit satisfiability can be used to solve MaxSAT on the given instance.
5 Modulators to Logarithmic Pathwidth and 2-SAT Backdoors
In this section we show an equivalence between the following two hypotheses.
Theorem 5.1.
The following are equivalent:
  1. 1.
    (2SatB-SETH is false) There is an algorithm that takes as input a CNF formula $\phi$ and a strong 2-SAT backdoor of size $b$, and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{b}|\phi|^{O(1)}$.
  2. 2.
    ($\log\!\textrm{pw}$M-SETH is false) There is an algorithm that takes as input a CNF formula $\phi$, a modulator $M$ of size $m$, and a path decomposition of the primal graph of $\phi-M$ of width $O(\log|\phi|)$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
Before we proceed, let us give some intuition. One can think of both problems that we are dealing with here as reachability problems, in the following sense: if we are given an assignment to a 2-SAT backdoor, we can verify if it can be extended to a satisfying assignment by solving 2-SAT, which is known to be complete for the class NL, that is, equivalent to reachability. On the other hand, if we have an instance of logarithmic pathwidth, the intuition of the class EPNL defined in [31] (see also [36]) is that solving such an instance corresponds exactly to deciding if a non-deterministic machine with space $O(\log n)$ will accept a certain input, that is, this problem is again complete for NL. This intuition leads us to believe that modulators to logarithmic pathwidth and 2-SAT backdoors are equivalent.
There is, however, one complication in the above reasoning. What we have established is that once someone supplies us with an assignment to the 2-SAT backdoor, checking if this assignment is correct has the same complexity as checking whether an assignment to the logarithmic pathwidth modulator is correct. However, what we really care about is whether an algorithm which avoids enumerating all assignments in one case can be used to avoid enumerating the assignments in the other. Hence, an argument which assumes that the assignment is known cannot be used as a black box. What we need is a reduction that transfers the answer to all possible assignments for a backdoor to the possible assignments to a modulator, without going through the assignments (ideally running in polynomial time).
Given this, it is natural to revisit the reasons for which both 2-SAT and SAT on graphs of logarithmic pathwidth are essentially equivalent to reachability. We then run into the following obstacle: while SAT on graphs of logarithmic pathwidth easily reduces to and from reachability, 2-SAT is actually only easily seen to be equivalent to the complement of reachability. The reason 2-SAT ends up being complete for NL is the famous Immerman–Szelepcsényi theorem, which established that NL=coNL [29].
We are therefore in a situation where, if we attempt to go through reachability, one problem we are dealing with has a quantification of the form $\exists\exists$ (for logarithmic pathwidth modulators the question is if there exists an assignment to the modulator so that the resulting graph has an $s\to t$ path) and the other has the form $\exists\forall$ (does there exist an assignment to the backdoor so that in the graph representing the 2-SAT implications there is no cycle creating a contradiction). What we need is to change the inner quantifier without knowing the value of the variables of the first quantifier block. We are thus led to the following definition of two auxiliary problems:
Definition 5.2.
Suppose we are given a DAG $G=(V,E)$, a designated source $s$ and sink $t$ of $G$, a set $M$ of $m$ boolean variables $x_{1},\ldots,x_{m}$, and an annotation function which assigns to some of the arcs of $G$ a literal (that is, a variable or its negation) from $M$.
The Ann-s-t-Reach problem is the following: decide if there exists an assignment to $M$ such that if we only keep in the graph arcs with no annotation or with an annotation set to True by the assignment, then there is a directed path from $s$ to $t$.
The Ann-s-t-Non-Reach problem is the following: decide if there exists an assignment to $M$ such that if we keep the same arcs as above, then there is no directed path from $s$ to $t$.
Intuitively, Ann-s-t-Reach allows us to encode $2^{m}$ instances of reachability simultaneously: for each assignment to the variables of $M$ we obtain a distinct graph. It is important to note here that Ann-s-t-Reach and Ann-s-t-Non-Reach are not complementary problems. In both problems the outer quantification is existential: we ask if there exists an assignment to $M$ such that, in one case a path exists, and in the other no path exists.
Our strategy is now the following: we will show that Ann-s-t-Reach is equivalent to SAT parameterized by a logarithmic pathwidth modulator (intuitively this is easy, because both problems have an $\exists\exists$ quantification); we will show that SAT parameterized by a 2-SAT backdoor is equivalent to Ann-s-t-Non-Reach (again, both problems have an $\exists\forall$ quantification). Finally, we will revisit the proof of the Immerman–Szelepcsényi theorem and show that the key arguments still go through. In particular, the Immerman–Szelepcsényi theorem can be seen as a reduction from reachability to its complement. What we need to show is that this reduction can still be executed if arcs are annotated, as in our problems, in a way that preserves the answer for all assignments to $M$.
Proof 5.3 (Proof of Theorem˜5.1).
We split the proof into three parts. In Lemma˜5.4 we show an equivalence between 2-SAT backdoors and Ann-s-t-Non-Reach; in Lemma˜5.10 we show an equivalence between SAT with a modulator to logarithmic pathwidth and Ann-s-t-Reach; and in Lemma˜5.12 we show that Ann-s-t-Reach and Ann-s-t-Non-Reach have the same complexity by revisiting the proof of the Immerman–Szelepcsényi theorem.
5.1 2-SAT and Non-Reachability
Lemma 5.4.
There is an algorithm solving Ann-s-t-Non-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$, for some $\varepsilon>0$ if and only if the 2SatB-SETH is false.
Proof $5.$
The proof is broken down into Lemma˜5.6 and Lemma˜5.8.
Lemma 5.6.
If there is an algorithm falsifying 2SatB-SETH, then there is an algorithm solving Ann-s-t-Non-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$.
Proof 5.7.
We use a standard reduction of non-reachability to 2-SAT, except we have to take into account the arc annotations. Suppose we are given a Ann-s-t-Non-Reach instance. Let $G=(V,E)$ be the given DAG on $n$ vertices. We construct a CNF formula with variables the variables of $M$, given with the Ann-s-t-Non-Reach instance, to which we add a variable $x_{v}$ for each $v\in V$. We will set $M$ to be the backdoor set. Now, for each arc $(v_{1},v_{2})\in E$ we add the clause $(\neg x_{v_{1}}\lor x_{v_{2}})$ to the formula if the arc is not annotated, otherwise we add the clause $(\neg x_{v_{1}}\lor x_{v_{2}}\lor\neg\ell)$ if the arc is annotated with the literal $\ell$. We also add the unit clauses $(x_{s})$ and $(\neg x_{t})$. This completes the construction.
We now claim that if the formula we constructed is satisfiable, using the same assignment to $M$ yields a digraph where there is no path from $s$ to $t$. Indeed, suppose that for the same assignment there is a path from $s$ to $t$, meaning that the assignment sets to True all the literals used as annotations to the edges of the path. This implies that in the corresponding clauses, the backdoor literals we have added are set to False by the satisfying assignment. But then this would create a chain of implications given $x_{s}\to x_{t}$, which in conjunction with the two unit clauses would falsify the formula.
For the converse direction, suppose that there is an assignment to $M$ that makes $t$ unreachable from $s$. We use the same assignment in the formula and extend it by setting $x_{v}$ to True if and only if $v$ is reachable from $s$ under the same assignment to $M$. It is not hard to see that this assignment will validate the formula: the only possibility of failure is for a clause representing the arc $(v_{1},v_{2})$, where $v_{1}$ is reachable from $s$ and $v_{2}$ is not. In such a case, the arc must be annotated by a literal that is set to False by the assignment to $M$, so the corresponding clause is satisfied.
Finally, we observe that the construction is polynomial and the size of the backdoor is exactly $m$, as each clause contains at most $2$ of the new variables.
Lemma 5.8.
If there is an algorithm solving Ann-s-t-Non-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$, for some $\varepsilon>0$, then the 2SatB-SETH is false.
Proof 5.9.
We want to reduce 2-SAT to Ann-s-t-Non-Reach while using the annotation to encode the backdoor variables. We are given a CNF formula $\phi$ with a set $B=\{y_{1},\ldots,y_{b}\}$ of $b$ backdoor variables and $n$ non-backdoor variables $X=\{x_{1},\ldots,x_{n}\}$. We assume without loss of generality that each clause contains exactly $2$ non-backdoor variables: indeed, no clause can contain three or more such variables (otherwise $B$ would not be a strong backdoor), and if $C$ contains fewer than $2$ such variables, we can replace it with the pair of clauses $(C\lor x_{i}),(C\lor\neg x_{i})$, for some $x_{i}$ that does not appear in $C$.
We now describe a basic implication digraph: for each variable $x_{i}\in X$ we construct two vertices $x_{i},\neg x_{i}$. For each clause $C$, let $\ell_{1},\ell_{2}$ be the literals of $C$ made up of variables of $X$. We construct an arc from $\neg\ell_{1}$ to $\ell_{2}$ and an arc from $\neg\ell_{2}$ to $\ell_{1}$. If $C$ has size $2$ we do not annotate these arcs. Otherwise, if $C$ contains $b^{\prime}\leq b$ other literals, we subdivide these arcs so that they become directed paths of length $b^{\prime}$ and annotate each arc of the resulting path with the negation of a distinct literal of $C$ involving a variable of $B$. The intuitive meaning of this construction is that we have a path from $\neg\ell_{1}$ to $\ell_{2}$ and from $\neg\ell_{2}$ to $\ell_{1}$ if and only if all the backdoor literals of $C$ are set to False.
The digraph above is not necessarily a DAG, but we can construct an equivalent DAG as follows: take $2n$ copies of the graph above, number them $1,\ldots,2n$, for each vertex $\ell$ in copy $i$ add non-annotated arcs to the same vertex in all later copies, and for each edge $(\ell_{1},\ell_{2})$ and each $i\in[2n-1]$ add an arc from the vertex $\ell_{1}$ in copy $i$ to the vertex $\ell_{2}$ in copy $i+1$ maintaining the annotation of the arc. We now have the property that there is a path from $\ell_{1}$ to $\ell_{2}$ in the original graph if and only if there is a path from $\ell_{1}$ in the first copy to $\ell_{2}$ in the last copy of the vertex set in the resulting DAG. This equivalence continues to hold taking into account the arc annotations. Call the DAG we have constructed the implication DAG and denote it by $D$.
Now our construction is as follows: start with two vertices $s,t$ and for each $i\in[n]$ construct $2$ copies of the implication DAG, call them $D_{i,0},D_{i,1}$ for $i\in[n]$. For $i\in[n]$ put an arc from $s$ to the first vertex $x_{i}$ in $D_{i,0}$ from the last $\neg x_{i}$ in $D_{i,0}$ to the first $\neg x_{i}$ in $D_{i,1}$ and from the last $x_{i}$ in $D_{i,1}$ to $t$. We set $M=B$ and this completes the construction which can clearly be carried out in polynomial time. What remains is to argue for correctness.
Suppose now that the original CNF is satisfiable. We want to show that by using the same assignment to $M$ we eliminate all $s\to t$ paths. Indeed, suppose we find such a path, which for some $i\in[n]$ must go through $D_{i,0},D_{i,1}$. In particular, this means we have a path from $x_{i}$ to $\neg x_{i}$ and a path from $\neg x_{i}$ to $x_{i}$ in $D$ while respecting the annotations. However, recall that for each clause we have added a path of annotated arcs containing the negations of all backdoor literals of the clause, meaning that if we can follow an arc, the assignment to $M$ must be setting all the backdoor literals to False. We conclude that the assignment cannot be satisfying to the formula, as we can infer the implications $(x_{i}\to\neg x_{i})$ and $(\neg x_{i}\to x_{i})$.
For the converse direction, suppose there is an assignment to $B$ which eliminates all $s\to t$ paths. We claim the same assignment can be extended to a satisfying assignment of the original formula. Suppose for contradiction that this is not the case and suppose that we are working with a maximal counter-example, that is, for the given $n,b$ we have started our reduction from a formula $\phi$ that satisfies the following: (i) $\phi$ is unsatisfiable (ii) there is an assignment to $B$ that eliminates all $s\to t$ paths (iii) every other $\phi^{\prime}$ with the same backdoor size $b$ that satisfies (i) and (ii) either has $n^{\prime}>n$ non-backdoor variables or the Ann-s-t-Non-Reach instance constructed for $\phi^{\prime}$ has at most as many arcs as that for $\phi$. In other words, $\phi$ is the counter-example with smallest $n$ and among these has the maximum number of arcs in the resulting Ann-s-t-Non-Reach instance. Assuming these we will reach a contradiction, which will show that a counter-example cannot exist.
We will attempt to find a good assignment to $\phi$. Consider an $i\in[n]$ and since there is no path from $s$ to $t$ it must be the case that either there is no path from (the first) $x_{i}$ to (the last) $\neg x_{i}$ in $D$, or that there is no path from (the first) $\neg x_{i}$ to (the last) $x_{i}$ in $D$, or both.
Suppose that there is no path from $x_{i}$ to $\neg x_{i}$, but that there is a path from $\neg x_{i}$ to $x_{i}$, for some $i\in[n]$. We then set $x_{i}$ to True. Symmetrically, suppose that there is no path from $\neg x_{i}$ to $x_{i}$, but that there is a path from $x_{i}$ to $\neg x_{i}$. Then we set $x_{i}$ to False. Do this for all variables that satisfy one of these conditions.
We claim that this assignment so far satisfies all clauses containing a variable that has been set. Indeed, suppose $x_{i}$ has been set to True but there is a clause such that all its backdoor literals are false and its other non-backdoor literal is $\ell$. The interesting case is if $C$ contains $x_{i}$ negated (otherwise $C$ is satisfied by $x_{i}$), so we have the directed paths $x_{i}\to\ell$ and $\neg\ell\to\neg x_{i}$, which can be followed under the current assignment to $M$. Then, there is a path $\neg\ell\to\neg x_{i}\to x_{i}\to\ell$, which implies that the variable involved in $\ell$ has been set to a value that satisfies $C$. The reasoning is the same if $x_{i}$ has been set to False.
At this point we have an assignment to $B$ and an assignment to $X^{\prime}\subseteq X$ which together satisfy all clauses involving variables of $X^{\prime}$. If $X^{\prime}\neq\emptyset$ then we are done, because we can give the assignment we have calculated to $X^{\prime}$ and remove from $\phi$ all clauses containing a variable from $X^{\prime}$. The new formula $\phi^{\prime}$ would be a smaller counter-example, contradicting the selection of $\phi$. Indeed, the Ann-s-t-Non-Reach instance for $\phi^{\prime}$ is a sub-graph of the one for $\phi$, so it cannot contain any $s\to t$ path, while $\phi^{\prime}$ must be unsatisfiable, because if it were satisfiable, using the assignment we calculated for $X^{\prime}$ would satisfy all of $\phi$.
We then must have $X^{\prime}=\emptyset$. This implies that for all $i\in[n]$ we have neither a path from $x_{i}$ to $\neg x_{i}$ nor from $\neg x_{i}$ to $x_{i}$ in $D$. We claim that setting any $x_{i}$ to True and simplifying the formula also yields a smaller counter-example. Indeed, it will be more convenient to do this by adding the clause $(x_{i}\lor x_{i})$ to $\phi$, obtaining an Ann-s-t-Non-Reach instance with more arcs (this clause cannot have already existed, as then we would have had $X^{\prime}\neq\emptyset$). The new formula is clearly also unsatisfiable (we added a clause to $\phi$ which was assumed unsatisfiable). Furthermore, there is still no $s\to t$ path in the Ann-s-t-Non-Reach instance for $\phi^{\prime}$. To see this, suppose that there is now such a path $s\to x_{j}\to\neg x_{j}\to\neg x_{j}\to x_{j}\to t$. Previously, neither the $x_{j}\to\neg x_{j}$ path (in $D_{j,0}$) nor the $\neg x_{j}\to x_{j}$ path (in $D_{j,1}$) existed, so both paths must be using the arc $\neg x_{i}\to x_{i}$ which is the only newly added arc to $D$. So we already had paths $x_{i}\to\neg x_{j}$ and $\neg x_{j}\to\neg x_{i}$, which means we already had a path $x_{i}\to\neg x_{i}$, contradicting the assumption that $X^{\prime}=\emptyset$. We conclude that $\phi^{\prime}$ is a counter-example with the same number of variables but strictly more arcs in the corresponding Ann-s-t-Non-Reach instance, contradicting the selection of $\phi$. Hence, no counter-example exists.
5.2 Log-Pathwidth and Reachability
Lemma 5.10.
There is an algorithm solving Ann-s-t-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$, for some $\varepsilon>0$ if and only if the $\log\!\textrm{pw}$M-SETH is false.
Proof 5.11.
For one direction, assume that there is an algorithm solving Ann-s-t-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$ and we are given a CNF formula $\phi$ and a modulator $M$ of size $m$ such that removing $M$ from $\phi$ results in a graph of pathwidth $O(\log|\phi|)$. Suppose we are given a path decomposition of the resulting graph, and the bags are numbered $B_{1},\ldots,B_{r}$. Recall that each clause of $\phi$ induces a clique on the primal graph and by standard properties of tree decompositions for each clause $C$ there is a bag that contains all variables of $C$ (outside of $M$). Repeating bags as necessary we assume that there is an injective mapping from the clauses to the bags so that each clause is mapped to a bag that contains all its variables outside of $M$. Note that if a clause contains only variables of $M$, then it can be mapped to any bag.
We construct an Ann-s-t-Reach instance as follows. For each $i\in[r]$ and for each assignment $\sigma$ to the variables of $B_{i}$ we construct two vertices $v_{i,\sigma,1},v_{i,\sigma,2}$. For each $i\in[r-1]$, for each assignment $\sigma$ to $B_{i}$ and $\sigma^{\prime}$ to $B_{i+1}$ such that $\sigma,\sigma^{\prime}$ agree on $B_{i}\cap B_{i+1}$, we add an arc $(v_{i,\sigma,2},v_{i+1,\sigma^{\prime},1})$ without annotation.
For each $i\in[r]$ and assignment $\sigma$ to $B_{i}$ we do the following: if $\sigma$ satisfies the (at most one) clause $C$ mapped to $B_{i}$ or if no clause is mapped to $B_{i}$, then add an arc $(v_{i,\sigma,1},v_{i,\sigma,2})$ without annotation. If there is a clause $C$ mapped to $B_{i}$ but $C$ is not satisfied by $\sigma$, we add an arc $(v_{i,\sigma,1},v_{i,\sigma,2})$ annotated with each literal of $C$ from $M$. In particular, if $C$ contains no literals from $M$ we do not add an arc $(v_{i,\sigma,1},v_{i,\sigma,2})$, while if it contains several, we add several parallel arcs (we can then subdivide them to avoid constructing a multi-graph), each annotated with a distinct literal. We add a vertex $s$ and arcs $(s,v_{1,\sigma,1})$ for all $\sigma$, as well as a vertex $t$ and arcs $(v_{r,\sigma,2},t)$. We retain the set $M$ between the two instances.
It is now not hard to see that there is a one-to-one correspondence between satisfying assignments to $\phi$ and $s\to t$ paths. Fix an assignment to $M$. If there is an $s\to t$ path, it must use for each $i$ an arc $(v_{i,\sigma,1},v_{i,\sigma,2})$, which allows us to infer an assignment $\sigma$ for $B_{i}$. This assignment is consistent because we only added arcs between bags for consistent pairs of assignments. It is satisfying because if a clause $C$ is mapped to $B_{i}$ we can only use the arc $(v_{i,\sigma,1},v_{i,\sigma,2})$ if $\sigma$ satisfies the clause or if the assignment to $M$ agrees with one of the literals of the clause; furthermore, all clauses are mapped to some bag. For the converse direction, if there is an assignment we construct a path by using in each bag the arc $(v_{i,\sigma,1},v_{i,\sigma,2})$ corresponding to the assignment $\sigma$ used for $B_{i}$.
Let us now move to the other half of the lemma. Suppose there is an algorithm deciding satisfiability with a modulator $M$ to pathwidth $O(\log|\phi|)$ and we want to use this algorithm to solve Ann-s-t-Reach. We construct a formula which retains the set $M$ of annotation variables as the modulator. Given a digraph $G=(V,E)$ with $n$ vertices we assume without loss of generality that $n$ is a power of $2$ (otherwise add some dummy vertices). We construct a formula with $n\log n$ extra variables, called $x_{i,j}$, $i\in[n],j\in[\log n]$. The intended meaning is that the $\log n$ variables $x_{i,1},\ldots,x_{i,\log n}$ encode the index of the $i$-th vertex we would visit in an $s\to t$ path. To enforce this meaning we assume the vertices of $G$ are numbered from 0 to n1n-1 and with each vertex we associate a distinct truth assignment to $\log n$ variables by reading these variables as a binary number. We add clauses enforcing the variables $x_{1,j},j\in[\log n]$ to take the value corresponding to $s$ and the variables $x_{n,j},j\in[\log n]$ to take the value corresponding to $t$. For each $i\in[n-1]$ we add at most $n^{2}$ clauses to the formula as follows: for each $(u,v)\not\in E$ such that $u\neq v$, we add a clause ensuring that if the $x_{i,j}$ variables encode $u$ and the $x_{i+1,j}$ variables encode $v$, then the formula is not satisfied; for each $(u,v)\in E$ which is annotated with a literal $\ell$ we add the same clause (which is falsified if $x_{i,j},x_{i+1,j}$ encode $u,v$), but we add to this clause the literal $\ell$. If $(u,v)\in E$ is not annotated or if $u=v$, we do not add a clause. It is not hard to see that, if we remove the variables of $M$, this formula has pathwidth $O(\log n)$, because we can make n1n-1 bags, each containing $x_{i,j},x_{i+1,j},j\in[\log n]$ and this covers all edges.
To argue for correctness, if there is an $s\to t$ path in the original instance for some assignment to $M$ we use the same assignment in $\phi$ and extend it to a satisfying assignment as follows: if the path visits $r\leq n$ vertices in total, we set for each $i\in[r]$ the assignment to $x_{i,j},j\in[\log n]$ to encode the $i$-th vertex of the path; all remaining groups encode $t$. This satisfies the formula. For the converse direction, we can infer from a satisfying assignment to the formula a path of length at most n1n-1 by interpreting the assignment to $x_{i,j},j\in[\log n]$ as the encoding of the $i$-th vertex of the path, and the clauses we have added ensure that the path is valid, that is we are only followed to stay in the same vertex, follow an arc without annotation, or follow an arc whose annotation is set to True by the assignment to $M$.
5.3 Reachability and Non-Reachability
Lemma 5.12.
For all $\varepsilon>0$, there is an algorithm solving Ann-s-t-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$ if and only if there is an algorithm solving Ann-s-t-Non-Reach in time $(2-\varepsilon)^{m}n^{O(1)}$.
Proof 5.13.
We describe a polynomial-time algorithm which takes as input a DAG $G_{1}=(V_{1},E_{1})$, with $|V_{1}|=n$, a source $s$, a sink $t$, with the arcs annotated with variables from a set $M$. The algorithm produces a new DAG $G_{2}=(V_{2},E_{2})$, a source $s_{2}$, a sink $t_{2}$, with arcs annotated from the same set $M$. It guarantees the following: for each assignment to $M$, $G_{2}$ has an $s_{2}\to t_{2}$ path respecting the assignment if and only if $G_{1}$ does not have any $s\to t$ path respecting the assignment. Observe that if we achieve this, we immediately obtain reductions from Ann-s-t-Reach to Ann-s-t-Non-Reach in both directions, implying the lemma.
We will reuse the ideas of the famous Immerman–Szelepcsényi theorem, so the reader may think of $G_{2}$ as a graph that encodes the working of a non-deterministic machine which is verifying that $G_{1}$ has no $s\to t$ path. The construction is of course made more complicated because we also need to take into account the arc annotations of $G_{1}$.
Before we begin, let us slightly edit $G_{1}$. Take $n=|V_{1}|$ copies of $V_{1}$, constructing for each $v\in V_{1}$ the vertices $v^{1},\ldots,v^{n}$. For each $i\in[n-1]$ and $v\in V_{1}$ we add the arc $(v^{i},v^{i+1})$ without annotation; for each arc $(u,v)$ and each $i\in[n-1]$ we add the arc $(u^{i},v^{i+1})$, maintaining the annotation of $(u,v)$. Finally, for the vertex $s^{n-1}$ and all $v\in V_{i}\setminus\{t\}$ we add the non-annotated arc $(s^{n-1},v^{n})$. This transformation has achieved the following: for each assignment to $M$, an $s\to t$ path exists in $G_{1}$ if and only if an $s^{1}\to t^{n}$ path exists in the new graph; there is a path from $s_{1}$ to $v^{n}$ for all $v\in V_{1}\setminus\{t\}$. As a result the question of whether an $s\to t$ path exists in $G_{1}$ is equivalent to asking whether the number of vertices reachable from $s^{1}$ in $\{v^{n}\ |\ v\in V_{1}\}$ is equal to $n$. Call the new graph $G_{1}^{\prime}$ and we want to construct $G_{2}$ such that for all assignments to $M$, there is a path $s_{2}\to t_{2}$ in $G_{2}$ if and only if n1n-1 vertices of $\{v^{n}\ |\ v\in V_{1}\}$ are reachable from $s^{1}$ in $G_{1}^{\prime}$. In the remainder, when we refer to the $i$-th layer of $G_{1}^{\prime}$ we will mean the set $\{v^{i}\ |v\in V_{1}\}$. Observe that $G_{1}^{\prime}$ has the property that all arcs go from a layer $i$ to a layer $i+1$.
We begin the construction of $G_{2}$ with $n^{2}$ vertices $s_{i,j}$ for $i,j\in[n]$, plus a source $s_{2}$ and a sink $t_{2}$. The intended meaning is that (for all assignments to $M$) $s_{i,j}$ will be reachable from $s_{2}$ if and only if there are exactly $j$ vertices of $G_{1}^{\prime}$ in layer $i$ which are reachable from $s^{1}$. We add a non-annotated arc $(s_{n,n-1},t_{2})$ and a non-annotated arc $(s_{2},s_{1,1})$. For each $i,j\in[n]$ we want to construct a checker gadget $C_{i,j}$ such that there will be an arc from $s_{i,j}$ to $C_{i,j}$ and arcs from $C_{i,j}$ to $s_{i+1,j^{\prime}}$ for some $j^{\prime}\in[n]$. We want to achieve the following property for all assignments $\sigma$ to $M$. Suppose that under some assignment $\sigma$, there are exactly $j$ vertices of layer $i$ of $G_{1}^{\prime}$ reachable from $s^{1}$. Then, there will be a path from $s_{i,j}$ to $s_{i+1,j^{\prime}}$ through $C_{i,j}$ if and only if the number of vertices of layer $i+1$ of $G_{1}^{\prime}$ reachable from $s^{1}$ is $j^{\prime}$. If we achieve this property we are done, as any path starts from $s_{1,1}$ and it is the case that exactly $1$ vertex of layer $1$ is reachable from $s^{1}$ ($s^{1}$ itself), so we can proceed by induction and we see that there is a path to $s_{n,n-1}$ (hence to $t_{2}$) if and only if n1n-1 vertices of the last layer of $G_{1}^{\prime}$ are reachable. Hence, there will be a $s_{2}\to t_{2}$ path in $G_{2}$ if and only if there is no $s_{1}\to t_{1}$ path in $G_{1}$.
Let us now describe the construction of the checker $C_{i,j}$. For each $\alpha,\beta\in\{0,\ldots,n\}$ we construct a vertex $x_{i,j,\alpha,\beta}$. We place an incoming non-annotated arc to $x_{i,j,0,0}$ from $s_{i,j}$. The intended meaning of these vertices is the following: assuming that $j$ is correct (that is, there are exactly $j$ vertices of layer $i$ reachable from $s^{1}$ in $G_{1}^{\prime}$), there should exist a path from $s_{i,j}$ to $x_{i,j,\alpha,\beta}$ if and only if among the $\alpha$ first vertices of layer $i+1$ of $G_{1}^{\prime}$ there are exactly $\beta$ which are reachable from $s^{1}$. Clearly, the property holds for $x_{i,j,0,0}$. Given this intended meaning we place a non-annotated arc from $x_{i,j,n,\beta}$ to $s_{i+1,\beta}$ for all $\beta\in[n]$. Now if we establish that the vertices $x_{i,j,\alpha,\beta}$ satisfy the property we stated, then the checker gadget functions correctly. What remains then is to add some gadgets that imply this property.
For each $\alpha\in\{0,\ldots,n-1\}$ and $\beta\in\{0,\ldots,n-1\}$ we will construct a gadget called $\mathrm{Adj}_{i,j,\alpha,\beta}$ and a gadget called $\mathrm{NonAdj}_{i,j,\alpha,\beta}$, each with a single input and output vertex. The intended meaning is that a path should exist from the input to the output of the $\mathrm{Adj}$ gadget if and only if the $(\alpha+1)$-th vertex of layer $i+1$ of $G_{1}^{\prime}$ is reachable from $s^{1}$; such a path should exist in the $\mathrm{NonAdj}$ gadget if and only if the same vertex is not reachable from $s^{1}$. Given this meaning we place a non-annotated arc from $x_{i,j,\alpha,\beta}$ to the input of $Adj_{i,j,\alpha,\beta}$ and from the output of that gadget to $x_{i,j,\alpha+1,\beta+1}$; and an annotated arc from $x_{i,j,\alpha,\beta}$ to the input of $NonAdj_{i,j,\alpha,\beta}$ and from the output of that gadget to $x_{i,j,\alpha+1,\beta}$. What remains is to describe these gadgets and prove they work as expected.
The $\mathrm{Adj}_{i,j,\alpha,\beta}$ gadget is conceptually simple. Suppose that $v$ is the $(\alpha+1)$-th vertex of $V_{1}$ and we are testing if $v^{i+1}$ is reachable from $s^{1}$ in $G_{1}^{\prime}$. The gadget is simply a copy of the first $i+1$ layers of $G_{1}^{\prime}$, with input vertex $s^{1}$ and output vertex $v^{i+1}$. Clearly, there is a path from the input to the output of the gadget if and only if $v^{i+1}$ is reachable from $s^{1}$ in $G_{1}^{\prime}$.
The $\mathrm{NonAdj}_{i,j,\alpha,\beta}$ gadget is more complicated as it relies on the key insight of the Immerman–Szelepcsényi theorem which is that we can certify that a vertex $v$ is not reachable from $s$ at distance $i+1$ if we know how many vertices are reachable at distance $i$, because we can list all such vertices, verify that their cardinality is correct, and verify that none of them have an arc to $v$. For each $\gamma\in\{0,\ldots,n\}$ and $\delta\in\{0,\ldots,j\}$ we construct a vertex $y_{i,j,\alpha,\beta,\gamma,\delta}$, whose intended role is the following: assuming $j$ is correct, there is a path from the input of the gadget to $y_{i,j,\alpha,\beta,\gamma,\delta}$ if and only if among the first $\gamma$ vertices of the $i$-th layer of $G_{i}$ there are at least $\delta$ vertices which are reachable from $s^{1}$ but which do not have an arc that leads to $v$ under the current assignment, where again $v$ is the $(\alpha+1)$-th vertex of $V_{1}$. The input vertex of the gadget is $y_{i,j,\alpha,\beta,0,0}$ and the output is $y_{i,j,\alpha,\beta,n,j}$. For all $\gamma\in\{0,\ldots,n-1\}$ and $\delta\in\{0,\ldots,j\}$ we add a non-annotated arc from $y_{i,j,\alpha,\beta,\gamma,\delta}$ to $y_{i,j,\alpha,\beta,\gamma+1,\delta}$. The intended meaning is that if among the first $\gamma$ vertices of the $i$-th layer there are $\delta$ satisfying the property above, then there certainly are at least $\delta$ vertices satisfying this property among the first $\gamma+1$ vertices. What remains is to add a gadget so that we can transition from $y_{i,j,\alpha,\beta,\gamma,\delta}$ to $y_{i,j,\alpha,\beta,\gamma+1,\delta+1}$ if and only if the $(\gamma+1)$-th vertex for the $i$-th layer, call it $u^{i}$, satisfies the property. If there is a non-annotated arc from $u^{i}$ to $v^{i+1}$, we do nothing. If there is no such arc or the arc is annotated, we need to check if $u^{i}$ is reachable from $s^{1}$. We therefore add an $\mathrm{Adj}$ gadget checking this, and add a non-annotated arc from $y_{i,j,\alpha,\beta,\gamma,\delta}$ to the input of that gadget. From the output of that gadget we add an arc to $y_{i,j,\alpha,\beta,\gamma+1,\delta+1}$ annotated as follows: if $(u^{i},v^{i+1})$ is not an arc of $G_{1}^{\prime}$, then we add no annotation; otherwise $(u^{i},v^{i+1})$ is annotated with a literal $\ell$, and we use the annotation $\neg\ell$ for the new arc. This completes the construction.
Correctness now follows from the description of the gadgets. It is worth noting that a vertex $u^{i}$ can be used as part of the certificate that $v^{i+1}$ is not reachable only when the (possible) arc $(u^{i},v^{i+1})$ cannot be used either because it does not exist or because of its annotation. Since we have added the same annotation negated to the output arc of the $\mathrm{NonAdj}$ gadget, this correctly checks if $u^{i}$ is a valid part of the certificate that $v^{i+1}$ is not reachable. The rest of the proof follows along the lines of the classical Immerman–Szelepcsényi theorem [29].
6 Circuit-SAT and Horn Backdoors
Our goal in this section is to prove the following theorem:
Theorem 6.1.
The following statements are equivalent:
  1. 1.
    (Circuit-SETH is false) There exists $\varepsilon>0$ and an algorithm that takes as input a Boolean circuit of size $s$ on $n$ inputs and decides if the circuit is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
  2. 2.
    (HornB-SETH is false) There exists $\varepsilon>0$ and an algorithm that takes as input a CNF formula $\phi$ together with a strong Horn backdoor of size $b$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{b}|\phi|^{O(1)}$.
  3. 3.
    (W[P]-SETH is false, weak version) There exist $\varepsilon>0,k_{0}>0$ and an algorithm which for all $k>k_{0}$ takes as input a monotone Boolean circuit of size $s$ and decides if the circuit can be satisfied by setting at most $k$ inputs to True in time $O(s^{(1-\varepsilon)k})$.
  4. 4.
    (W[P]-SETH is false, strong version) There exists $\varepsilon>0$ and an algorithm which takes as input a Boolean circuit of size $s$ on $n$ inputs and an integer $k$ and decides if the circuit can be satisfied by setting exactly $k$ of its inputs to True in time $n^{(1-\varepsilon)k}s^{O(1)}$.
Before we go on let us give some intuition. Similarly to Section˜5, it will be helpful to recall some basic facts from computational complexity theory. For the first two statements, we are dealing with the satisfiability problem for circuits and SAT with a Horn backdoor. Intuitively, it is natural to expect these two problems to be equivalent because once we fix either the input to a given circuit or the assignment to the Horn backdoor, deciding if the circuit is satisfied or if the backdoor assignment can be extended to a satisfying assignment are problems with the same classical complexity, that is, they are P-complete [44]. As in Section˜5 this intuition is not in itself sufficient to imply the equivalence, because what we want is not to claim that once we have an assignment, checking it is of the same complexity in both cases, but rather that if we can avoid checking all assignments in one case, we can avoid it in both. Nevertheless, the classical results on the P-completeness of these two problems are indeed the main ingredient we need, and our proof mostly consists of verifying that we can extend them to the new setting. We then go on to consider the prototypical W[P]-complete problem of the third statement. It is natural to expect that since the first and third statement refer to circuit satisfiability problems, without any particular restriction on the circuits, it should be possible to transform one to the other, and indeed we are able to do so using standard techniques, similar to what we did in Theorem˜4.1.
Proof 6.2 (Proof of Theorem˜6.1).
The proof is given in the four lemmas below which show the implications $1\Rightarrow 2$ (Lemma˜6.3), $2\Rightarrow 1$ (Lemma˜6.5), $3\Rightarrow 1$ (Lemma˜6.7), $1\Rightarrow 4$ (Lemma˜6.9).
We note that the implication $4\Rightarrow 3$ is easy to see. First, statement 3 talks about monotone circuits, while statement 4 about general circuits. Second, if we have an algorithm running in time $n^{(1-\varepsilon)k}s^{O(1)}$ we can see that $n\leq s$ (as the input gates are part of the circuit) so the running time is at most $O(s^{(1-\varepsilon)k+c})$ for some constant $c$. We can now set $k_{0}=\frac{2c}{\varepsilon}$ and we then have $\varepsilon k>\frac{\varepsilon k}{2}+c$ if $k>k_{0}$, therefore the running time is at most $O(s^{(1-\frac{\varepsilon}{2})k})$.
6.1 Circuit-SAT and Horn Backdoors
Lemma 6.3.
Statement 1 of Theorem˜6.1 implies statement 2.
Proof 6.4.
Suppose that we have an algorithm solving Circuit SAT in time $(2-\varepsilon)^{n}s^{O(1)}$, for some fixed $\varepsilon>0$. We are given a CNF formula $\phi$ and a strong Horn backdoor of size $b$ of $\phi$. Our plan is to construct a Boolean circuit $C$ satisfying the following:
  1. 1.
    $C$ is satisfiable if and only if $\phi$ is.
  2. 2.
    $C$ has $b$ inputs and can be constructed in time $|\phi|^{O(1)}$.
Clearly, if $C$ can be constructed in time polynomial in $|\phi|$, then its size is also polynomial in $|\phi|$, so the supposed algorithm for Circuit SAT gives us the promised algorithm for SAT with a Horn backdoor.
The high-level idea now is to implement a circuit which takes as input the values of the $b$ backdoor variables of $\phi$ and then runs the standard polynomial-time algorithm to check satisfiability of Horn formulas. Recall that this algorithm starts with the all-False assignment to all (non-backdoor) variables, and then iterates, at each step setting a variable to True if this is forced. This will result in either a satisfying assignment or a contradiction after at most $n$ iterations, where $n$ is the number of (non-backdoor) variables.
Suppose that the variables of $\phi$ that belong to the backdoor are $y_{1},\ldots,y_{b}$ and the remaining variables are $x_{1},\ldots,x_{n}$. Assume $\phi$ has $m$ clauses, numbered $c_{1},\ldots,c_{m}$.
We construct a circuit with the following gates:
  1. 1.
    $b$ input gates, identified with the variables $y_{i},i\in[b]$.
  2. 2.
    For each non-backdoor variable $x_{i},i\in[n]$, we construct $n+2$ disjunction ($\lor$) gates $x_{i,t},t\in\{0,\ldots,n+1\}$. The intuitive meaning is that $x_{i,t}$ should evaluate to True if the Horn satisfiability algorithm would set variable $x_{i}$ to True after $t$ iterations of the main loop. One of the inputs of each $x_{i,t}$ is the gate $x_{i,t-1}$, for all $i,t\in[n+1]$. We specify the remaining inputs below.
  3. 3.
    For each clause $c_{j},j\in[m]$, we construct a disjunction ($\lor$) gate using the variables $x_{i,n+1}$ and the input gates as inputs. In particular, if $x_{i}$ appears positive in $c_{j}$ we connect $x_{i,n+1}$ to the gate representing $c_{j}$, and if it appears negative, we connect $x_{i,n+1}$ to a negation ($\neg$) gate whose output is sent to the gate representing $c_{j}$, and similarly for input gates. We route the output of all gates representing clauses to a $\land$ gate which is the output of the whole circuit. Intuitively, the circuit contains a copy of $\phi$, with the $x_{i}$ variables replaced by the gates $x_{i,n+1}$.
To complete the construction we describe the gates we need to add so that the $x_{i,t}$ gates play their intuitive role. For each variable $x_{i},i\in[n]$, for each $t\in[n+1]$ and each clause $c_{j},j\in[m]$, such that $x_{i}$ appears positive in $c_{j}$ we add a conjunction ($\land$) gate $c_{i,j,t}$ whose output is sent to $x_{i,t}$. The inputs to $c_{i,j,t}$ are all gates $x_{i^{\prime},t-1}$ such that $x_{i^{\prime}}$ appears (negative) in $c_{j}$, all gates $y_{i^{\prime\prime}}$ such that $y_{i^{\prime\prime}}$ appears negative in $c_{j}$, and the negation of each gate $y_{i^{\prime\prime}}$, such that $y_{i^{\prime\prime}}$ appears positive in $c_{j}$. Finally, set all $x_{i,0}$ gates, for $i\in[n]$ to False and simplify the circuit accordingly. This completes the construction, which can clearly be carried out in polynomial time. We have described a circuit with unbounded fan-in, but it is easy to transform it into a circuit of bounded fan-in without significantly increasing its size, in polynomial time. Note also that we have used the fact that if $x_{i}$ appears positive in $c_{j}$ every other non-backdoor variable $x_{i^{\prime}}$ must appear negative in $c_{j}$; this follows from the fact that if two variables appear positive in a clause, at least one of them must be in the backdoor [42].
It is easy to see that if the circuit is satisfiable then so is the formula, because we can extract an assignment from the values of the input gates and the gates $x_{i,n+1}$. This assignment satisfies the formula since we have added a copy of $\phi$ at the end of the circuit, with inputs the $y_{i}$ and $x_{i,n+1}$ gates.
For the converse direction, we first make some easy observations: for each $i\in[n]$ and $t\in[n+1]$ we have that if $x_{i,t-1}$ evaluates to True, then $x_{i,t}$ evaluates to True, since $x_{i,t-1}$ is an input to $x_{i,t}$. As a result, as $t$ increases, the set of gates $x_{i,t}$ which evaluate to True can only monotonically increase. Furthermore, if for some $t\in[n+1]$ we have for all $i\in[n]$ that $x_{i,t-1}$ and $x_{i,t}$ evaluate to the same value, then for all $t^{\prime}>t$ and for all $i$ we have $x_{i,t^{\prime}}$ evaluates to the same value as $x_{i,t}$. This is because the $c_{i,j,t}$ gates and the $c_{i,j,t+1}$ gates will evaluate to the same values, forcing $x_{i,t+1}$ to have the same value as $x_{i,t}$, and we can proceed by induction. It now follows that $x_{i,n}$ evaluates to the same value as $x_{i,n+1}$ for all $i\in[n]$. Indeed, there are $n+2$ possible values of $t$ (time steps), at each step the set of True gates can only monotonically increase, so we have at most $n+1$ distinct sets of True gates. As a result, by pigeonhole principle, there exists $t$ such that $x_{i,t}$ evaluates to the same value as $x_{i,t+1}$ for all $i\in[n]$, which implies the same for $t=n$.
Suppose now that $\phi$ is satisfiable and fix a satisfying assignment. Using the same assignment restricted to the backdoor variables as input to our circuit we claim that we obtain a satisfying assignment to the circuit. To see this, we claim that if for some $t\in[n+1]$ we have that $x_{i,t}$ evaluates to True, then the satisfying assignment to $\phi$ must also set $x_{i}$ to True. For the sake of contradiction, suppose that $t$ is minimum such that $x_{i,t}$ evaluates to True but the satisfying assignment sets $x_{i}$ to False. Since $x_{i,t}$ is True and $t$ is minimum, one of the $c_{i,j,t}$ gates must evaluate to True. But this means that the satisfying assignment sets all literals of the clause $c_{j}$ to False, except $x_{i}$, because by the minimality of $t$ we know that for each $x_{i^{\prime},t-1}$ which evaluates to True, $x_{i^{\prime}}$ is True in the satisfying assignment. Hence, the satisfying assignment must set $x_{i}$ to True.
We can now conclude that if $x_{i,n+1}$ evaluates to True, then the satisfying assignment to $\phi$ sets $x_{i}$ to True. Suppose then for contradiction that the circuit outputs False, which implies that there is a gate added in the third step, representing $c_{j}$, which evaluates to False. It must be the case that $c_{j}$ has as input an $x_{i,n+1}$ which evaluates to False while $x_{i}$ is set to True in the satisfying assignment, because in all other cases the inputs to $c_{j}$ would be the same as those dictated by the satisfying assignment to $\phi$ and $c_{j}$ would evaluate to True. Furthermore, $x_{i}$ must appear positive in $c_{j}$ and is therefore the only non-backdoor variable which appears positive in $c_{j}$. We now claim that the gate $c_{i,j,n+1}$ must evaluate to True, reaching a contradiction. Indeed, because $x_{i^{\prime},n}$ and $x_{i^{\prime},n+1}$ evaluate to the same value for all $i^{\prime}\in[n]$, all non-backdoor variables of $c_{j}$ except $x_{i}$ (which all appear negative in $c_{j}$) are set to True in $x_{i,n+1}$, hence in $x_{i,n}$, and backdoor variables are shared between the gates $c_{j},c_{i,j,n+1}$, so all inputs to $c_{i,j,n+1}$ evaluate to True, contradiction. Hence, we have an assignment satisfying the circuit.
Lemma 6.5.
Statement 2 of Theorem˜6.1 implies statement 1.
Proof 6.6.
Suppose we have an algorithm solving SAT, given a Horn backdoor of size $b$, in time $(2-\varepsilon)^{b}|\phi|^{O(1)}$. We are now given a circuit $C$ with $n$ inputs and size $s$. Our plan is to construct from $C$ a CNF formula $\phi$ and a Horn backdoor such that:
  • $\phi$ is satisfiable if and only if $C$ is.
  • The backdoor has size $n$ and $\phi$ can be constructed in time $s^{O(1)}$.
Clearly, if $\phi$ can be constructed in time $s^{O(1)}$, it also has size polynomial in $s$, so the supposed algorithm for SAT with a Horn backdoor would give us an algorithm for Circuit-SAT.
The high-level idea is now to revisit the standard proof that Horn-SAT is P-complete, which relies on a reduction from the Circuit Value Problem. More precisely, assume that the given circuit is represented by a DAG that contains only $\land,\lor$, and $\neg$ gates of maximum fan-in $2$. We first transform the circuit (in polynomial time) so that $\neg$ gates only appear directly connected to the inputs. Since the circuit is a DAG, we can assume that the gates are topologically sorted. We construct $\phi$ as follows:
  1. 1.
    For each input gate of $C$ we construct a variable $x_{i}$, $i\in[n]$. These variables will form the Horn backdoor.
  2. 2.
    For each input, $\land$, and $\lor$ gate $c$ of $C$, we construct two variables: $c^{p}$ and $c^{n}$. Intuitively, in circuit assignments that set $c$ to True, $c^{p}$ must be set to True and $c^{n}$ to False, while the opposite must happen for circuit assignments that set $c$ to False. Thus, for each such gate we add the clause $(\neg c^{p}\lor\neg c^{n})$.
  3. 3.
    For each input gate $c$ corresponding to variable $x_{i}$ we add the clauses $(x_{i}\leftrightarrow c^{p})$ and $(x_{i}\leftrightarrow\neg c^{n})$.
  4. 4.
    For each negation gate $c$, whose input corresponds to the variable $x_{i}$, we add the clauses $(x_{i}\leftrightarrow\neg c^{p})$ and $(x_{i}\leftrightarrow c^{n})$.
  5. 5.
    For each conjunction gate $c$, whose inputs are gates $c_{1},c_{2}$ we add the clauses $(\neg c_{1}^{p}\lor\neg c_{2}^{p}\lor c^{p}),(\neg c_{1}^{n}\lor c^{n}),(\neg c_{2}^{n}\lor c^{n}),(\neg c^{p}\lor c_{1}^{p}),(\neg c^{p}\lor c_{2}^{p})$
  6. 6.
    For each disjunction gate $c$, whose inputs are gates $c_{1},c_{2}$ we add the clauses $(\neg c_{1}^{n}\lor\neg c_{2}^{n}\lor c^{n}),(\neg c_{1}^{p}\lor c^{p}),(\neg c_{2}^{p}\lor c^{p}),(\neg c^{n}\lor c_{1}^{n}),(\neg c^{n}\lor c_{2}^{n})$
  7. 7.
    For the unique output gate $c$ of the circuit we add the unit clauses $c^{p}$ and $\neg c^{n}$.
The construction can be performed in polynomial time. Furthermore, all clauses except those of the third and fourth steps are Horn clauses. The clauses of the third and fourth steps all have size $2$ and contain a variable from the backdoor, so we have a valid strong Horn backdoor of size $n$, as promised. It is now not hard to see that if $C$ has a satisfying assignment, then we can obtain a satisfying assignment to $\phi$ by giving to the variables of the backdoor the values of the satisfying assignment to $C$ and then for each gate $c$ of the circuit setting $c^{p},c^{n}$ to True, False respectively if $c$ is set to True and to False, True respectively otherwise.
For the converse direction, suppose we have a satisfying assignment to $\phi$. We claim that using the assignment to the backdoor variables as input to $C$ satisfies the circuit. To see this, we prove by induction that for each gate $c$, exactly one of $c^{p},c^{n}$ must be set to True in the satisfying assignment, and $c^{p}$ is set to True if and only if gate $c$ has value True in the circuit $C$ for the given input assignment. Note that it is clear that at most one of $c^{p},c^{n}$ is set to True, due to the clauses of the second step. Furthermore, for the base case we observe that input and negation gates satisfy the property, due to the clauses of the third and fourth steps. Consider then a gate $c$ with inputs $c_{1},c_{2}$, which are assumed to satisfy the property. We have four cases:
  1. 1.
    $c$ is a conjunction that has value True in $C$. Then $c_{1},c_{2}$ must evaluate to True in $C$, since these gates satisfy the property we have $c_{1}^{p},c_{2}^{p}$ set to True, so the clause $(\neg c_{1}^{p}\lor\neg c_{2}^{p}\lor c^{p})$ ensures that $c^{p}$ is True (and $c^{n}$ is False) as desired.
  2. 2.
    $c$ is a conjunction that has value False in $C$. Then one of $c_{1},c_{2}$ must evaluate to False in $C$, say without loss of generality $c_{1}$, and since these gates satisfy the property we have $c_{1}^{n}$ set to True. So, the clause $(\neg c_{1}^{n}\lor c^{n})$ ensures that $c^{n}$ is True (and $c^{p}$ is False) as desired.
  3. 3.
    $c$ is a disjunction that has value True in $C$. Then one of $c_{1},c_{2}$ must evaluate to True in $C$, say without loss of generality $c_{1}$, and since these gates satisfy the property we have $c_{1}^{p}$ set to True. So, the clause $(\neg c_{1}^{p}\lor c^{p})$ ensures that $c^{p}$ is True (and $c^{n}$ is False) as desired.
  4. 4.
    $c$ is a disjunction that has value False in $C$. Then $c_{1},c_{2}$ must evaluate to False in $C$, since these gates satisfy the property we have $c_{1}^{n},c_{2}^{n}$ set to True, so the clause $(\neg c_{1}^{n}\lor\neg c_{2}^{n}\lor c^{n})$ ensures that $c^{n}$ is True (and $c^{p}$ is False) as desired.
Therefore, $\phi$ is satisfiable if and only if $C$ is.
6.2 Circuit-SAT and Weighted Circuit-SAT
Lemma 6.7.
Statement 3 of Theorem˜6.1 implies statement 1.
Proof 6.8.
Suppose we have an algorithm which takes as input a monotone boolean circuit $C^{\prime}$ of size $S$ and $N$ inputs and an integer $k$ and decides if $C^{\prime}$ has a satisfying assignment setting $k$ inputs to True in time $O(S^{(1-\varepsilon)k})$, for some $\varepsilon>0$, assuming $k>k_{0}$ for some fixed $k_{0}$. We will use this supposed algorithm to solve Circuit-SAT. We are given a circuit $C$ with $n$ inputs and size $s$. We will assume that $s<2^{\frac{\varepsilon}{2k_{0}}n}$, because otherwise $2^{n}\leq s^{\frac{2k_{0}}{\varepsilon}}=s^{O(1)}$, so we immediately have a polynomial time algorithm for such cases by trying out all input assignments.
We construct a monotone circuit $C^{\prime}$ and an integer $k=k_{0}$ with the following properties:
  1. 1.
    $C^{\prime}$ has a satisfying assignment setting $k$ of its inputs to True if and only if $C$ is satisfiable.
  2. 2.
    $C^{\prime}$ can be constructed in time $O(2^{n/k}s)$.
  3. 3.
    $C^{\prime}$ has $N=O(2^{n/k})$ inputs and size $S=O(Ns)=O(2^{\frac{n}{k}+\frac{\varepsilon n}{2k}})$.
If we achieve the above, then we obtain the lemma as follows: we can decide satisfiability of $C$ by running the reduction and then the supposed algorithm on $C^{\prime}$. The running time is dominated by the second step which takes time $O(S^{(1-\varepsilon)k})=O(2^{(1-\varepsilon)(1+\frac{\varepsilon}{2})n})=O(2^{(1-\varepsilon^{\prime})n})$ for some appropriate $\varepsilon^{\prime}$, as desired.
Let us then describe the construction of $C^{\prime}$ in which we start from $C$ and add some gadgets. We assume that after running standard transformations $C$ only contains negation gates at the first layer, that is, every input of a negation is an input to $C$. Partition the inputs of $C$ arbitrarily into $k$ groups of (almost) equal size. Each group contains at most $\lceil\frac{n}{k}\rceil$ inputs of $C$. Consider a group $I_{i}$ of input gates of $C$, for $i\in[k]$ and enumerate all $2^{|I_{i}|}=O(2^{n/k})$ possible assignments to the inputs of $I_{i}$. For each such assignment $\sigma$ we construct a new input gate $g_{i,\sigma}$. We convert the gates of $I_{i}$ from input gates to disjunction ($\lor$) gates and for each assignment $\sigma$ we connect $g_{i,\sigma}$ to all gates of $I_{i}$ such that $\sigma$ assigns them value True. Similarly, for each negation gate whose input is in $I_{i}$ we convert it into a disjunction gate and for each $\sigma$ we connect $g_{i,\sigma}$ to all gates which were previously negations such that $\sigma$ was assigning to their input the value False. Since we have converted all negations to disjunctions the circuit is now monotone.
At this point it is not hard to see that the new circuit has $N=O(k2^{n/k})=O(2^{n/k})$ inputs (as $k$ is a fixed constant), its size is at most $O(Ns)$ (in the worst case we connected each new input to each old gate) and that a satisfying assignment to $C$ exists if and only if there exists a satisfying assignment to $C^{\prime}$ that sets exactly one new input gate from each of the $k$ groups to True and all others to False.
We now add some gadgets to ensure that a satisfying assignment of weight $k$ must set one input from each group to True. For each $i\in[k]$ add a disjunction ($\lor$) gate $t_{i}$ and connect all $g_{i,\sigma}$ to $t_{i}$. Add a conjunction gate that will be the new output of the circuit and feed it as input all $t_{i}$, for $i\in[k]$ as well as the output gate of $C$. This completes the construction. The total size is still $S=O(Ns)$ and the construction can be performed in linear time in the size of the output. We can also transform the circuit into an equivalent bounded fan-in circuit without increasing the size by more than a constant factor. With the extra gates we added in the last step we ensure that at least one input gate from each group is set to True, so since we are looking for an assignment of weight exactly $k$, exactly one input is set to True in each group.
Lemma 6.9.
Statement 1 of Theorem˜6.1 implies statement 4.
Proof 6.10.
Suppose there is an algorithm that takes as input a boolean circuit $C^{\prime}$ of size $S$ with $N$ inputs and decides if $C^{\prime}$ is satisfiable in time $2^{(1-\varepsilon)N}S^{c}$, for some constants $\varepsilon>0,c>0$. We are given a circuit $C$ with $n$ inputs and size $s$ as well as an integer $k$ and we want to decide if $C$ has a satisfying assignment of weight exactly $k$ in time $n^{(1-\varepsilon^{\prime})k}s^{O(1)}$. We will do this by constructing a circuit $C^{\prime}$ with the following properties:
  1. 1.
    $C^{\prime}$ is satisfiable if and only if $C$ has a satisfying assignment of weight $k$.
  2. 2.
    $C^{\prime}$ has at most $N=k\log n+k$ input gates and can be constructed in time $O(sk^{2}n^{2})$.
If we achieve the above we obtain the desired algorithm because we can produce $C^{\prime}$ and run the supposed Circuit-SAT algorithm on it. The running time is dominated by the second step which takes $2^{(1-\varepsilon)N}S^{c}$. We have $2^{(1-\varepsilon)N}\leq n^{(1-\varepsilon)k}2^{k}$ and $S^{c}=O(s^{c}(kn)^{2c})$. Observe now that we can assume that $n$ and $k$ are sufficiently large. In particular, if $k<8c/\varepsilon$, then since without loss of generality $s>n$, the trivial $O(n^{k}s)$ time algorithm which checks all assignments of weight $k$ would run in time $s^{O(1)}$. But if $n,k$ are sufficiently large, then $2^{k}(kn)^{2c}<n^{\frac{\varepsilon k}{2}}$ and the total running time is at most $n^{(1-\frac{\varepsilon}{2})k}s^{O(1)}$ as desired.
We construct $C^{\prime}$ by adding some gadgets to $C$. First, number the input gates of $C$ as $c_{1},\ldots,c_{n}$. We transform them into disjunction gates. For each $i\in[n]$ we construct $k$ new gates $c_{i}^{1},\ldots,c_{i}^{k}$ and feed their output into $c_{i}$. The intuitive meaning is that $c_{i}^{j}$ will be True if $c_{i}$ is the $j$-th input set to True in a satisfying assignment of weight $k$. Each $c_{i}^{j}$, for $i\in[n],j\in[k]$ is a conjunction gate.
We now construct $k\lceil\log n\rceil\leq k\log n+k$ new input gates, partitioned into $k$ groups $I_{1},\ldots,I_{k}$, each containing $\lceil\log n\rceil$ input gates. The intuitive meaning is that the assignment to the gates of $I_{j}$ should encode the index of the $j$-th input of $C$ set to True in a satisfying assignment. For each $j\in[k]$ and each assignment $\sigma$ to the input gates of $I_{j}$ we interpret $\sigma$ as a binary number in $[n]$ and suppose that this number is $i$. We then connect the input gates of $I_{j}$ to $c_{i}^{j}$ via appropriate negations such that $c_{i}^{j}$ evaluates to True if and only if the gates of $I_{j}$ are given assignment $\sigma$.
At this point it is not hard to see that $C$ is satisfiable with an assignment of weight $k$ if and only if $C^{\prime}$ is satisfiable via an assignment which for each distinct $j,j^{\prime}\in[k]$ gives distinct assignments to $I_{j},I_{j^{\prime}}$. What remains is to add some gadgets so that $I_{j},I_{j^{\prime}}$ cannot have the same assignment, so that satisfying assignments to $C^{\prime}$ encode satisfying assignments of weight exactly $k$ to $C$. This is easily achieved by adding for each pair of distinct $j,j^{\prime}\in[k]$ a disjunction gate, which takes its input from $2\lceil\log n\rceil$ conjunction gates, where each conjunction gate checks if the assignments to $I_{j},I_{j^{\prime}}$ differ in a corresponding bit by giving False to the bit of $I_{j}$ and True to the bit of $I_{j^{\prime}}$ or vice-versa.
Let us analyze the properties of this construction. The number of inputs of $C^{\prime}$ is at most $k\lceil\log n\rceil\leq k\log n+k$. The number of wires we added is at most $O(kn\log n)$ for the wires incident on $c_{i}^{j}$ gates, and $O(k^{2}\log n)$ for the gadget that ensures we have an assignment of weight $k$. The construction can be performed in linear time in the size of the output and the total size is at most $O(sk^{2}n^{2})$.
7 Applications
In this section we list five characteristic example of our equivalence classes. In each case we highlight a problem and a lower bound on its time complexity such that breaking the bound is equivalent to falsifying the hypotheses of the corresponding class.
7.1 SETH-equivalence for Coloring
Theorem 7.1.
For all $q\geq 3$, there is an $\varepsilon>0$ and an algorithm solving $q$-Coloring in time $(q-\varepsilon)^{m}n^{O(1)}$, where $m$ is the size of a given modulator to constant tree-depth, if and only if the SETH is false. The same holds if $m$ is the size of a given $(\sigma,\delta)$-hub.
Proof 7.2.
It was already established in [19] that if the SETH is true, then there is no algorithm for $q$-Coloring with the stated running time, even for parameter the size of a given $(\sigma,\delta)$-hub. Therefore, if the SETH is true, no such algorithm exists for parameter size of modulator to constant tree-depth.
We establish that if the SETH is false, then we do get such a $q$-Coloring algorithm for parameter modulator to constant tree-depth. We are given an $n$-vertex graph $G$ and a modulator $M$ of size $m$ so that GMG-M has tree-depth $c$. We will use Theorem˜3.1 and reduce this instance to an instance of $k$-SAT with a modulator to constant tree-depth. According to Theorem˜3.1, if the SETH is false, we can solve such an instance is time $2^{(1-\varepsilon)m^{\prime}}|\phi|^{O(1)}$, where $m^{\prime}$ is the size of the new modulator, for some $\varepsilon$. In particular, we will make sure that $m^{\prime}\leq(1+\frac{\varepsilon}{2})m\log q+O(1)$, and that $\phi$ has size polynomial in $n$. Then, the running time of the whole procedure will be at most $2^{(1-\varepsilon)(1+\frac{\varepsilon}{2})m\log q}n^{O(1)}=q^{(1-\varepsilon^{\prime})m}n^{O(1)}$ for some appropriate $\varepsilon^{\prime}$.
To begin our construction, we consider every vertex of GMG-M. For each such $v$, we define $q$ variables $x_{v}^{1},\ldots,x_{v}^{q}$, indicating informally that $v$ takes color $i\in[q]$. We add the clause $(x_{v}^{1}\lor\ldots\lor x_{v}^{q})$. For each $uv\in E$ with $u,v\not\in M$ we add for each $i\in[q]$ the clause $(\neg x_{v}^{i}\lor\neg x_{u}^{i})$. It is not hard to see that the graph we have constructed so far has tree-depth at most $cq$. All the other variables we will construct will belong in the new modulator.
Consider now the vertices of $M$ and for some integer $\gamma$ to be defined later partition them into groups of size at most $\gamma$ as equitably as possible. We have then at most $\tau=\lceil\frac{m}{\gamma}\rceil\leq\frac{m}{\gamma}+1$ groups, $M_{1},\ldots,M_{\tau}$. For some integer $\rho$, also to be defined later, we replace each $M_{i}$ with a set of $\rho$ new variables $Y_{i}$. We select the integers $\gamma,\rho$ so that we have the following:
$2^{(1-\frac{\varepsilon}{2})\rho}\leq q^{\gamma}\leq 2^{\rho}$
Since $q^{\gamma}\leq 2^{\rho}$, we can map injectively all proper colorings of $G[M_{i}]$ to assignments of $Y_{i}$. For each assignment that is not an image of a proper coloring we add a clause ensuring that this assignment cannot be used in a satisfying assignment. Furthermore, for each $uv\in E$ such that $u,v$ belong in distinct sets $M_{i},M_{j}$ we consider all pairs of colorings of $M_{i},M_{j}$ that set $u,v$ to the same color. For each such pair, we add a clause ensuring that $Y_{i},Y_{j}$ cannot simultaneously take the assignments which are the images of these colorings. Finally, for each $u\in M_{i}$ and $v\not\in M$ such that $uv\in E$, for each color $j\in[q]$ we add for each coloring of $M_{i}$ that colors $u$ with $j$ a clause that contains $\neg x_{v}^{j}$ and literals which are all falsified if we select for $Y_{i}$ the image of this coloring of $M_{i}$.
This completes the construction and it is not hard to see that there is a one-to-one correspondence between proper colorings of $G$ and satisfying assignments of the CNF formula we have constructed. To bound the size of the new modulator we observe that we have $\frac{m}{\gamma}+1$ groups of $\rho$ variables, but $(1-\frac{\varepsilon}{2})\rho\leq\gamma\log q$, therefore, $\rho\leq(1+\frac{\varepsilon}{2})\gamma\log q$, so the total number of variables of the modulator is at most $(\frac{m}{\gamma}+1)\rho\leq(1+\frac{\varepsilon}{2})m\log q+O(1)$.
Finally, to argue that we can select appropriate $\gamma,\rho$, we observe that it is sufficient that $2^{\rho}>2^{(1-\frac{\varepsilon}{2})\rho}q$, because then there is an integer power of $q$ between $2^{\rho}$ and $2^{(1-\frac{\varepsilon}{2})\rho}$. This gives $\rho>\frac{4\log q}{\varepsilon}$, so selecting $\rho$ to be any integer at least this large also allows us to select an appropriate $\gamma$.
7.2 MaxSAT-equivalence for Max Cut
Theorem 7.3.
There is an $\varepsilon>0$ and an algorithm which takes as input an $n$-vertex graph $G$ and a $(\sigma,\delta)$-hub of size $h$ and computes a maximum cut of $G$ in time $(2-\varepsilon)^{h}n^{O(1)}$ if and only if the MaxSat-SETH is false.
Proof 7.4.
One direction is easy: given a Max Cut instance $G=(V,E)$, we can construct a CNF formula by constructing a variable $x_{v}$ for each vertex $v$ and for each edge $uv$ the clauses $(x_{u}\lor x_{v})$ and $(\neg x_{u}\lor\neg x_{v})$. The primal graph of this formula is simply $G$, so the $(\sigma,\delta)$-hub is preserved. It is now not hard to see that the maximum number of clauses that it is possible to satisfy in the new instance is exactly $|E|+\textrm{max-cut}(G)$, so an algorithm as described in the second statement of Theorem˜3.5 gives the desired algorithm for Max Cut.
For the converse direction, we are given a $k$-CNF formula $\phi$ with $n$ variables and $m$ clauses and a target value $t$ and are asked if we can satisfy at least $t$ clauses. We reuse the reduction of [38], except with smaller weights, so that we obtain a $(\sigma,\delta)$-hub. Let us give the relevant details. We will first describe a reduction to edge-weighted Max Cut and then remove the weights.
For each variable of $\phi$, $x_{1},\ldots,x_{n}$ we construct a vertex with the same name. We also construct a vertex $x_{0}$. For each clause $C_{j}$ of $\phi$ we construct a cycle of length $4k+1$, one of whose vertices is $x_{0}$, while the others are new vertices. Each edge of this cycle has weight $8k$. Number the vertices of the cycle $p_{j,1},\ldots,p_{j,4k}$, starting from a neighbor of $x_{0}$. Suppose that $C_{j}$ contains literals $\ell_{1},\ell_{2},\ldots,\ell_{k^{\prime}}$ for $k^{\prime}\leq k$. For each $i\in[k^{\prime}]$ if $\ell_{i}$ is a positive appearance of a variable $x_{r}$, then we connect $p_{j,4i+2},p_{j,4i+3}$ to $x_{r}$, otherwise if it is a negative appearance of $x_{r}$, then we connect $p_{j,4i+1},p_{j,4i+2}$ to $x_{r}$. These edges have weight $1$. Furthermore, we attach $k-k^{\prime}$ leaves to $p_{j,1}$ with edges of weight $1$. We now claim that there is an assignment satisfying $t$ clauses if and only if there is a cut of weight at least $32k^{2}m+km+t$.
For the forward direction, if there is such an assignment, place all the vertices representing False variables on the same side as $x_{0}$ in the cut and all the vertices representing True variables on the other side. For each unsatisfied clause, pick a bipartition of the corresponding odd cycle so that $4k$ of the $4k+1$ edges of the cycle are cut and the uncut edge is incident on $x_{0}$. This contributes $32k^{2}$ from the edges of the cycle. We also obtain at least $k^{\prime}$ from the edges connecting the cycle to the variable vertices and $k-k^{\prime}$ from the attached leaves by placing each leaf to the opposite side of its neighbor, so we get $32k^{2}+k$ for each unsatisfied clause. For each satisfied clause $C_{j}$, let $\ell_{i}$ be the first literal that is set to True by the assignment. For $i^{\prime}<i$ we set $p_{j,4i^{\prime}+2},p_{j,4i^{\prime}+4}$ on the same side as $x_{0}$ and $p_{j,4i^{\prime}+1},p_{j,4i^{\prime}+3}$ on the other side; while for $i^{\prime}>i$ we use the opposite partition. We now have that the vertex preceding $p_{j,4i+1}$ and the vertex after $p_{j,4i+4}$ in the cycle are both on the side of $x_{0}$. For the four remaining vertices, if $\ell_{i}$ is a variable that appears positive, we set $p_{j,4i+2},p_{j,4i+3}$ on the same side as $x_{0}$, cutting both edges connecting these vertices to the corresponding variable vertex, and $p_{j,4i+1},p_{j,4i+4}$ on the other side; while if $\ell_{i}$ is a variable that appears negative, we set $p_{j,4i+1},p_{j,4i+2},p_{j,4i+4}$ on the side opposite $x_{0}$ and $p_{j,4i+3}$ on the side of $x_{0}$, again cutting both edges to the corresponding variable vertex. We now observe that for a satisfied clause we have cut $32k^{2}+k+1$ edges, obtaining the bound on the total size of the cut.
For the converse direction we observe that an optimal cut must cut $4k$ of the $4k+1$ edges of the odd cycle constructed for each clause. Indeed, it is not possible to cut more (since the cycle is odd), while if we leave two edges uncut, we may replace the solution by a solution that cuts $4k$ of the edges. This will add $8k$ to the total weight, more than making up for the at most $2k$ lost edges connecting the cycle to the rest of the graph. Thus, for each cycle we obtain $32k^{2}$ from the edges inside the cycle. Furthermore, we can obtain at most $k+1$ from the remaining edges incident on the cycle. To see this, note that we have added $k^{\prime}$ pairs of edges (one pair for each literal of the clause) which are adjacent to consecutive vertices in the cycle. If we cut both edges in a pair, an edge of the cycle is uncut, so there is at most one pair of cut edges and from each other pair we cut at most one edge. Taking into account the $k-k^{\prime}$ incident on leaves (which are always cut), we have that each cycle contributes at most $32k^{2}+k+1$ to the cut. Hence, there must be at least $t$ cycles contributing exactly this much, which we claim correspond to satisfied clauses by the assignment we extract by setting to False variables whose corresponding vertices are on the same side as $x_{0}$. Indeed, take a cycle representing $C_{j}$ and suppose that $C_{j}$ is not satisfied by this assignment. However, there exists a literal $\ell_{i}$ in $C_{j}$ such that we cut both edges added for $\ell_{i}$. If $\ell_{i}$ is positive, then $p_{j,4i+2},p_{j,4i+3}$ have a common neighbor. If $C_{j}$ is not satisfied, this neighbor must be on the same side as $x_{0}$, so $p_{j,4i+2},p_{j,4i+3}$ are on the other side. It is now not hard to see via a parity argument that two edges of the cycle must be uncut, contradiction. The reasoning is similar if $\ell_{i}$ is negative, as then $p_{j,4i+1},p_{j,4i+2}$ must have a common neighbor, which must be on the side not containing $x_{0}$ (otherwise $C_{j}$ is satisfied), so $p_{j,4i+1},p_{j,4i+2}$ are on the same side as $x_{0}$, and again two edges are uncut. Hence, a cut of the specified size leads to an assignment satisfying at least $t$ clauses.
In order to remove weights, for each edge $uv$ of weight $w$ we construct $w$ parallel paths of length $3$ between $uv$ and increase the target by $2w$. It is not hard to see that any decision on $u,v$ can be extended in a way that cuts at least $2w$ of the new edges, and it is possible to cut all $3w$ new edges if and only if $u,v$ are on different sides.
After the transformation we have that the variable vertices plus $x_{0}$ form a hub of size $n+1$ such that its removal leaves components with $O(k^{2})$ vertices. Furthermore, each component is connected to at most $k$ vertices of the hub. We therefore have a $(\sigma,\delta)$-hub, for $\sigma=O(k^{2})$ and $\delta=k$.
7.3 Coloring and Pathwidth Modulators
Theorem 7.5.
For all $q\geq 3$ we have the following. There is an $\varepsilon>0$ such that for all $c>0$ there is an algorithm that takes as input an $n$-vertex graph $G$ and a modulator $M$ of size $m$ such that GMG-M has pathwidth at most $c$ and decides if $G$ is $q$-colorable in time $(q-\varepsilon)^{m}n^{O(1)}$ if and only if the LD-C-SETH is false. The same is true if we replace “pathwidth cc” with “tree-depth $c\log n$” in the preceding statement.
Proof 7.6.
Clearly, an algorithm parameterized by modulator to tree-depth $c\log n$ implies the existence of an algorithm parameterized by modulator to pathwidth $c$. We show two statements: if the LD-C-SETH is false, then there is an algorithm for parameter modulator to tree-depth $c\log n$; and if there is an algorithm for parameter modulator to pathwidth $c$, then the LD-C-SETH is false.
The first claim readily follows from the proof of Theorem˜7.1. Recall that in that proof we showed how to transform an instance $G$ of $q$-Coloring with a modulator $M$ to tree-depth $c$ to an instance of SAT with a modulator of size essentially $|M|\log q$, while ensuring that in the new instance the removal of the modulator leaves a graph of tree-depth at most $cq$. We run the same reduction, except now the tree-depth is at most $cq\log n$. One of the statements of Theorem˜4.1 states that we can avoid brute-force for SAT parameterized by a modulator to logarithmic tree-depth if and only if the LD-C-SETH is false.
For the second claim we need to reduce SAT with parameter modulator to pathwidth $c$ to $q$-Coloring. Suppose for some $q\geq 3$ there exists $\varepsilon>0$ such that for all $c^{\prime}$ there is a $q$-Coloring algorithm which takes as input a modulator to pathwidth $c^{\prime}$ of size $m^{\prime}$ and solves $q$-Coloring in time $q^{(1-\varepsilon)m^{\prime}}n^{O(1)}$.
We are given a CNF formula $\phi$ and a modulator $M$ of size $m$ such that $\phi-M$ has pathwidth $c$. Using Section˜A.2 we can assume that $\phi$ is a 3-CNF formula. Select integers $\gamma,\rho$ such that we have the following:
$q^{(1-\frac{\varepsilon}{2})\gamma}<2^{\rho}<q^{\gamma}$
For such integers to exist it is sufficient to have $q^{\gamma}>2q^{(1-\frac{\varepsilon}{2})\gamma}$, which gives $\gamma>\frac{2}{\varepsilon\log q}$. Selecting $\gamma=\lceil\frac{2}{\varepsilon\log q}\rceil$ then guarantees that an appropriate integer $\rho$ exists.
We will first reduce our 3-CNF formula to an instance of CSP with alphabet $[q]$ and arity at most $3\gamma$. In this problem we are given a list of variables which take values in $[q]$ and a list of constraints, each of which involves at most $3\gamma$ variables. Each constraint lists the assignment combinations to the involved variables which are acceptable. The goal is to select an assignment so that we have selected an acceptable combination in each constraint. Observe that $q$-Coloring is a special case of this problem where the arity is $2$ and all constraints state that the endpoints of an edge must receive distinct colors.
We now partition $M$ into groups of variables of size $\rho$ as equitably as possible, constructing at most $\frac{m}{\rho}+1$ groups. For each group we construct in our CSP instance a group of $\gamma$ variables. We construct an injective mapping from boolean assignments to a group of variables of the original instance to an $q$-ary assignment to the corresponding group of variables of the new instance. Since $2^{\rho}<q^{\gamma}$, this is always possible. For each variable of $\phi$ outside of $M$ we construct a variable whose acceptable values are only $\{1,2\}$ (representing True, False). For each clause $C$ of $\phi$ we construct a constraint with maximum arity $3\gamma$: if a variable of $C$ is part of a group $M_{i}$ of variables of $M$, we involve all the $\gamma$ variables representing this group in the constraint; otherwise we involve the single variable that represents the boolean variable outside of $M$ appearing in $C$. We now list all combinations of assignments to the at most $3\gamma$ involved variables and we set as unacceptable those combinations which are not an image of boolean assignments that satisfy the clause. This completes the construction and it is not hard to see that the answer is preserved. Furthermore, if we consider the primal graph of the CSP instance, the graph is unchanged once we remove the (new) modulator. The new modulator contains at most $(\frac{m}{\rho}+1)\gamma$ variables. But $(1-\frac{\varepsilon}{2})\gamma\log q<\rho$ so the size of the new modulator is at most $m^{\prime}\leq(1-\frac{\varepsilon}{2})\frac{m}{\log q}+O(1)$. If we had an algorithm solving a CSP instance $\psi$ with bounded arity and alphabet $[q]$ in time $q^{(1-\varepsilon)m^{\prime}}|\psi|^{O(1)}$, where $m^{\prime}$ is the modulator to constant pathwidth, this would give the desired algorithm for deciding our original CNF. We will therefore show how to obtain this from the supposed $q$-Coloring algorithm.
We first reduce the arity of the CSP instance to $2$. For each constraint $C$ involving more than $2$ and at most $3\gamma$ variables, let $\mathcal{S}_{C}$ be the set of acceptable assignments, which has cardinality at most $q^{3\gamma}=O(1)$. We construct $2|\mathcal{S}_{C}|+1$ new variables and connect them in an odd cycle with constraints of arity $2$, ensuring that all the variables which are involved in a constraint take distinct values, that $|\mathcal{S}_{C}|$ variables, which are pairwise non-adjacent, have possible values $\{1,2,3\}$, and that all the others have possible values $\{1,2\}$. The intuition here is that we have formed an odd cycle and we can use colors $\{1,2,3\}$ to color it. Since the cycle is odd, we must use color $3$ somewhere, and if we use color $3$ once we can 2-color the rest. The idea then is to associate each of the $|\mathcal{S}_{C}|$ variables that have three possible values with a satisfying assignment to $C$ and add constraints that ensure that if this variable takes color $3$ then the variables of $C$ must agree with the corresponding assignment. For this, for each of the $|\mathcal{S}_{C}|$ variables of the cycle that represents some assignment $\sigma$ we add constraints of arity $2$ involving this variable and each variable of $C$. The constraints state that if the variable of the cycle has value $1$ or $2$, then everything is acceptable, otherwise the other variable must agree with $\sigma$. We do this exhaustively and it is not hard to see that the answer is preserved. We have to argue that the pathwidth of the primal graph (after removing the modulator) is almost preserved. Indeed, for each constraint $C$, find a bag that contains all its variables (outside of $M$), and after this bag add a sequence of copies of it, where we add a path decomposition of the cycle we have constructed. Repeating this will not add to the pathwidth more than a small additive constant.
Finally, we want to reduce the CSP of arity $2$ and alphabet $q$ to $q$-Coloring. We will reduce to List Coloring where all lists are subsets of $[q]$. Adding a clique of size $q$ and appropriately connecting to the graph to simulate the lists will then complete the reduction. To reduce to List Coloring we can use the weak edge gadgets of [35]. For each variable of the CSP instance we construct a vertex. For each constraint over $u,v$ we consider every assignment to $u,v$ that is ruled out by this constraint. A weak edge is a gadget that we can add between $u,v$ that will rule out the corresponding coloring, but will always be colorable if we pick any other color combination for $u,v$. In fact, a weak edge is implemented using a path of length $3$ between $u,v$ (with appropriate lists) so using weak edges instead of edges does not affect the pathwidth of the graph by more than an additive constant. By adding weak edges for each forbidden combination of assignments we obtain an equivalent $q$-Coloring instance. The size of the modulator has been preserved and we obtain the theorem.
7.4 k-Neighborhood Cut
In this section we consider the following problem:
Definition 7.7.
In the $k$-Neighborhood-Cut problem we are given as input a DAG $G$ with two designated vertices $s,t$. The question is whether it is possible to select $k$ vertices of $G$, other than $s,t$, such that deleting all the out-neighbors of the selected vertices results in a graph where no path from $s$ to $t$ exists.
Before we proceed, let us note that it is trivial that $k$-Neighborhood-Cut admits an algorithm running in time $n^{k+O(1)}$: simply try out all solutions and for each solution check if there is a path from $s$ to $t$ avoiding the out-neighborhoods of selected vertices. Furthermore, it is not hard to see that the problem is W[2]-hard parameterized by $k$. Indeed, take an instance of $k$-Dominating Set $G=(V,E)$ and construct a DAG that consists of two copies of $V$, call them $V_{1},V_{2}$, such that there is an arc from each $u\in V_{1}$ to each $v\in V_{2}$ if $u=v$ or $uv\in E$. Add a vertex $s$ with arcs to all of $V_{2}$ and a vertex $t$ with arcs from all of $V_{2}$. Clearly, selecting a dominating set in $V_{1}$ is necessary and sufficient to eliminate all paths from $s$ to $t$. From the results of Patrascu and Williams [46], the above reduction implies that $k$-Neighborhood-Cut cannot be solved in time $n^{k-\varepsilon}$, assuming the SETH.
In this section we show the same lower bound but assuming the weaker assumption that SAT cannot be solved faster than brute-force for instances with a 2-SAT backdoor. More interestingly, we show that this is precisely the right assumption for this problem, in the sense that a better algorithm can be obtained if and only if the assumption is false.
Theorem 7.8.
There exist $\varepsilon>0,k\geq 2$ such that $k$-Neighborhood-Cut can be solved in time $n^{k-\varepsilon}$ if and only if the 2SatB-SETH is false.
Proof 7.9.
We will show reductions in both directions from Ann-s-t-Non-Reach which was established to be equivalent to 2SatB-SETH in Section˜5. As a reminder, in Ann-s-t-Non-Reach we are given a DAG with its arcs annotated from a set of $m$ boolean variables. We are asked to select an assignment to the variables so that in the DAG resulting from keeping only non-annotated arcs or arcs whose annotation is set to True by the assignment there is no $s\to t$ path, for two given vertices $s,t$. We have shown that 2SatB-SETH is equivalent to assuming that it is impossible to solve Ann-s-t-Non-Reach without considering essentially all assignments to the $m$ boolean variables.
First, suppose we have an instance $G=(V,E)$ of Ann-s-t-Non-Reach on a DAG with $n$ vertices and a set $M$ of $m$ boolean variables. We construct a DAG $G^{\prime}$ as an instance of $k$-Neighborhood-Cut as follows. We begin with the same DAG $G$. We then partition the variables into $k$ sets, call them $M_{1},\ldots,M_{k}$, each of size at most $\frac{m}{k}+1$. For each set $M_{i}$ we enumerate all assignments $\sigma$, and for each assignment we construct a vertex $x_{i,\sigma}$. We construct for each $i\in[k]$ a vertex $y_{i}$ and add arcs $sy_{i}$, $y_{i}t$ as well as all arcs $x_{i,\sigma}y_{i}$. For each arc $e$ of $G$ that is annotated with a literal from a variable of $M_{i}$, we subdivide the arc and call the new vertex $z_{e}$. We add an arc to $z_{e}$ from each $x_{i,\sigma}$ such that $\sigma$ sets the literal of $e$ to False. This completes the construction, which can be carried out in time at most $2^{m/2}n^{O(1)}$ for all $k\geq 2$. The new graph has $N$ vertices with $N\leq 2^{m/k}+n^{O(1)}$.
We now observe that the reduction above preserves the solution. In particular, because of the vertices $y_{i}$ we are forced to select exactly one $x_{i,\sigma}$ for each $i\in[k]$, so there is a one-to-one correspondence between assignments to $M$ and potential solutions in the new instance. Furthermore, an annotated arc can be used if and only if we have selected the encoding of an assignment that sets the annotation to True.
Suppose that there is an algorithm for $k$-Neighborhood-Cut running in time $N^{k-\varepsilon}$ for some fixed $k\geq 2,\varepsilon>0$. We can now solve Ann-s-t-Non-Reach by running the reduction described above and then solving $k$-Neighborhood-Cut. The running time will be dominated by the second part, which would take $N^{k-\varepsilon}=2^{\frac{m}{k}(k-\varepsilon)}n^{O(1)}=2^{(1-\frac{\varepsilon}{k})m}n^{O(1)}$, which would falsify the 2SatB-SETH by the results of Section˜5.
For the converse direction, suppose we are given an instance $G=(V,E)$ of $k$-Neighborhood-Cut on $n$ vertices. We will reduce it to Ann-s-t-Non-Reach as follows. We start with the same DAG and construct a set $M$ of $k\log n$ variables (assume without loss of generality that $n$ is a power of $2$, otherwise adding dummy vertices to $G$). Intuitively, the variables of $M$ are partitioned into $k$ sets $M_{1},\ldots,M_{k}$, where the assignment of $M_{i}$ is meant to encode the selection of the $i$-th vertex of the solution to $k$-Neighborhood-Cut.
We now edit $G$ as follows. We replace every vertex $u\in V$ with two vertices $u_{in},u_{out}$. Each arc $uv$ is replaced with the arc $u_{out},v_{in}$. We set $s_{out}$ as the start vertex and $t_{in}$ as the destination vertex. For each $u\in V$ we further add $k+1$ vertices $u_{0},u_{1},\ldots,u_{k}$. The idea is that we should be able to go from $u_{in}$ to $u_{i}$ if and only if the first $i$ vertices of the solution do not have any arc to $u$ in the original instance. We add an arc $u_{in}u_{0}$ and an arc $u_{k}u_{out}$.
Let us now describe the gadget we add between $u_{i-1}$ and $u_{i}$ for all $i\in[k]$. Intuitively, we should be able to go from $u_{i-1}$ to $u_{i}$ if and only if the assignment to $M_{i}$ does not encode a vertex that dominates $u$. For each $v\in V$ such that $vu\not\in E$ and $v\neq u$ we therefore construct a path from $u_{i-1}$ to $u_{i}$ of length $\log n$ going through new vertices. The arcs of the path are annotated with literals from $M_{i}$ ensuring that we can only follow all the arcs if the assignment to $M_{i}$ encodes vertex $v$.
This completes the construction. Observe that each arc of $G$ has been replaced by $k$ collections of at most $n$ parallel paths, each of length $\log n$, so the new graph has at most $n^{5}$ vertices. We observe that the answer is preserved. Indeed, a solution $S$ of $k$-Neighborhood-Cut can be translated to an assignment to $M$ in the natural way and we can see that if $u\in S$ dominates $v$, then it is impossible to go from $v_{in}$ to $v_{out}$ with the current assignment. In the other direction, if we have an assignment, we can extract a set $S$ in $G$ by interpreting the assignments to the $k$ groups as the indices of the selected vertices.
Suppose now that there is an algorithm solving Ann-s-t-Non-Reach in time $2^{(1-\varepsilon)m}N^{d}$, for some fixed $d$. The running time of executing the reduction and then the algorithm would be at most $2^{(1-\varepsilon)k\log n}n^{5d}=n^{k-\varepsilon k+5d}$. This implies that when $k\geq\frac{5d}{\varepsilon}+1$ the running time is at most $n^{k-\varepsilon}$, therefore there exist $k,\varepsilon$ for which we can obtain a fast enough algorithm for $k$-Neighborhood-Cut.
7.5 Degenerate Deletion
In this section we consider the problem of editing a graph by deleting vertices so that the result becomes $r$-degenerate. Recall that a graph is $r$-degenerate if all of its induced subgraphs contain a vertex of degree at most $r$. Equivalently, a graph is $r$-degenerate if there exists an ordering of its vertices $v_{1},\ldots,v_{n}$ such that each $v_{i}$ has at most $r$ neighbors among vertices with higher index.
The problem we are interested in is whether we can delete $k$ vertices of a given graph so that the graph becomes $r$-degenerate (where $k$ can be thought of as the parameter and $r\geq 2$ is an absolute constant). This problem was shown to be W[P]-complete by Mathieson [41]. By reusing essentially the same chain of reductions, we are able to show the following equivalence between this problem and Circuit-SETH.
Theorem 7.10.
For all $r\geq 2$ we have the following. There exists $\epsilon>0,k_{0}>0$ and an algorithm which for all $k>k_{0}$, if given as input an $n$-vertex graph, decides if $G$ can be made $r$-degenerate by deleting $k$ vertices in time $n^{(1-\varepsilon)k}$ if and only if the Circuit-SETH is false.
Proof 7.11.
For the direction that Circuit-SETH implies hardness for our problem we essentially reuse the reduction of [41], which relies on a previous reduction of [48], and observe that these reductions from the standard W[P]-complete problem we dealt with in Theorem˜6.1 actually preserve the value of $k$ exactly and only increase the size of the instance by a constant factor, therefore the lower bound is preserved as well. For the other direction we will perform a direct reduction to Circuit-SAT.
Let us give the relevant details. For the first direction, we start with an instance of monotone weight-$k$ circuit satisfiability. By Theorem˜6.1, if we can solve such an instance in time $s^{(1-\varepsilon)k}$ for monotone circuits of $n$ inputs and size $s$, assuming $k$ is a large enough constant, we will falsify the Circuit-SETH. As shown by Szeider [48], there is a simple reduction from this problem to the Cyclic Monotone Circuit Activation problem. In the latter problem we are given a monotone boolean circuit which is a general digraph (not a DAG) and are asked to select $k$ initial gates to activate (set to True). From that point on at each round a gate becomes True if at least one of its in-neighbors is True for disjunction ($\lor$) gates and if all its in-neighbors are True, for conjunction ($\land$) gates. The question is whether there is a selection of $k$ initial gates that eventually makes everything True.
Szeider observed that we can reduce weight-$k$ monotone circuit satisfiability to this problem as follows: take $k+1$ copies of the given circuit, identify their inputs, and then replace the inputs with conjunctions whose inputs are the $k+1$ outputs. If there is a weight-$k$ assignment to the original circuit we use the same activation set and it activates all outputs, therefore eventually all gates. If on the other hand there is a way to activate everything, because we have $k+1$ copies, there is at least one copy in which the output was activated using only the input gates, so from this we can extract a satisfying assignment to the original circuit. Mathieson [41] presents a reduction from this problem which, for each $r\geq 2$ produces an equivalent instance of $r$-Degenerate Deletion, without modifying $k$, by replacing each gate with a constant-size gadget. Executing these reductions starting from a circuit of size $s$ we obtain a graph on $N=O(s)$ vertices. If the problem could be solved in time $N^{(1-\varepsilon)k}$, we would have an algorithm deciding weight-$k$ satisfiability of the original instance in time $O(s^{(1-\varepsilon)k})$, disproving the Circuit-SETH.
For the converse direction, suppose that Circuit-SETH is false and that we have an instance of the $r$-Degenerate Deletion problem with budget $k$ on a graph with $n$ vertices. Assume without loss of generality that $n$ is a power of $2$ (otherwise add some dummy vertices). We will reduce this to satisfiability of a circuit with $k\log n$ inputs and size $n^{O(1)}$. Essentially, this circuit encodes the workings of a polynomial-time algorithm which checks if a solution is correct. The solution is encoded by the $k\log n$ inputs which are meant to be seen as $k$ groups of $\log n$ bits, encoding the indices of the selected vertices to delete. The algorithm we are trying to encode can be seen as follows: at each step we mark some active vertices (initially this set is the set of deleted vertices); then if a vertex has at most $r$ unmarked neighbors, we mark it; continue this exhaustively for $n$ rounds and in the end check if all vertices are marked. This algorithm can easily be encoded by a circuit of $n$ layers, each consisting of $n$ gates, where vertex $j$ of layer $i$ encodes whether vertex $j$ is marked after $i$ rounds.
If we had the supposed algorithm for checking circuit satisfiability, it would run in time $2^{(1-\varepsilon)k\log n}n^{O(1)}=n^{(1-\varepsilon)k+O(1)}$. Setting $k_{0}$ sufficiently large that $\epsilon k_{0}/2$ is larger than the constant additive term in the exponent gives the desired algorithm.
8 Independent Set and Classes of Perfect Graphs
In this section we present some results that seek to clarify the structural relations between fine-grained questions on the complexity of Independent Set parameterized by vertex deletion distance to a base class. Let us recall the general context. Fix a hereditary (induced-subgraph closed) class $\mathcal{C}$, such that Independent Set is polynomial-time solvable on graphs from $\mathcal{C}$. We will now consider instances from the class $\mathcal{C}+kv$, that is, the class of graphs which can be obtained by a graph in $\mathcal{C}$ by adding at most $k$ new vertices (and arbitrarily many edges incident on the new vertices). We parameterize by $k$ and ask what is the complexity of the best FPT algorithm. Recall again that the problem is clearly solvable in $2^{k}n^{O(1)}$ by branching on the $k$ extra vertices, assuming that these are given in the input [^4] We will always make this assumption, as we are concerned with the question of exploiting, rather than finding, this set of vertices.. Furthermore, this simple algorithm also applies to Weighted Independent Set, as long as this problem is also polynomial-time solvable on $\mathcal{C}$.
As a case study, we will consider four standard graph classes $\mathcal{C}$: Cluster graphs, that is, graphs where every connected component is a clique; Block graphs, that is, graphs where every $2$-vertex-connected component is a clique; Cographs, that is, $P_{4}$-free graphs; and interval graphs. Weighted Independent Set is known to be in P for all these classes (even if input weights are encoded in binary). Furthermore, cluster graphs are clearly a subclass of all the other classes: for block graphs this follows from the definition; for cographs, this follows because cluster graphs are $P_{3}$-free; and for interval graphs this follows because cliques are interval graphs and interval graphs are closed under disjoint union. However, observe that the other three classes are incomparable: any tree is a block graph, but a $K_{1,3}$ where each edge is sub-divided once is neither a cograph nor an interval graph; a complete bipartite graph $K_{n,n}$ is a cograph, but not a block graph nor an interval graph; and a disjoint union of a diamond ($K_{4}-e$) and a $P_{4}$ is an interval graph but not a block graph nor a cograph.
We can therefore see that graph theory already gives us some basic information about the relative difficulties of solving Independent Set on $\mathcal{C}+kv$ graphs, where $\mathcal{C}$ is one of the classes above, as it indicates that the easiest case should be when $\mathcal{C}$ is the class of cluster graphs. Nevertheless, this tells us nothing about the relationship between the other classes and it also ceases to apply if we want to compare the complexity of Weighted Independent Set on cluster graphs with that of Independent Set on the other classes.
Our point of departure is a result which recently appeared in [19] and establishes the following:
Theorem 8.1 ([19]).
If there exist $\varepsilon>0$ and an algorithm which takes as input a graph $G=(V,E)$ and a set $M\subseteq V$ such that GMG-M is a disjoint union of cliques, and solves Independent Set on $G$ in time $(2-\varepsilon)^{|M|}|V|^{O(1)}$, then the SETH is false.
We note that the result of [19] is stated in terms of the more general $q$-Colorable Deletion problem, where we want to delete the minimum number of vertices to make a graph $q$-colorable; Independent Set is just the case $q=1$. Furthermore, even though this is not explicitly stated, one can easily verify that in the case $q=1$ the modulator constructed in the reduction of [19] is indeed a modulator to a cluster graph, as all gadgets used for $q=1$ are cliques. We give a self-contained exposition of this further below.
It would therefore appear that, in some sense, the case is closed. All the problems we have promised to consider are solvable in time $2^{k}n^{O(1)}$; but according to Theorem˜8.1 even the easiest among them is not solvable in time $(2-\varepsilon)^{k}n^{O(1)}$. Hence, if one believes the SETH, there is nothing else to investigate here and all these problems have the same complexity, despite the fact that from the graph-theoretic point of view, interval graphs (for instance) are clearly a much wider class of graphs than cluster graphs. This seems rather paradoxical.
The results we present in this section will somewhat resolve this tension and confirm the graph-theory based intuition that these problems are not of the same difficulty. Concretely, we give a hierarchical ranking of these problems (taking also into account their weighted versions) which will allow us to compare their difficulties by examining which case is reducible to which other, using the main complexity hypothesis we have discussed as our guiding framework. In summary, we will show the following: (see also Figure˜2)
  1. 1.
    As expected, cluster graphs are the easiest of the considered cases, even if we allow (binary) weights. We show (Theorem˜8.8) that refuting the LD-C-SETH would be sufficient to achieve a faster than brute-force algorithm for this, placing it somewhere below the middle class of Figure˜1. For the case of unary weights, we also show that the problem is at least as hard as the MaxSat-SETH (Theorem˜8.6), placing both weighted variants of this problem between the second and third classes (from the bottom) of Figure˜1.
  2. 2.
    On the other hand, we show that for both cographs and block graphs, obtaining a faster than brute-force algorithm seems harder, as we are able to establish that refuting the LD-C-SETH is both necessary and sufficient for this. This places these problems in the third class of Figure˜1. Equivalence continues to hold even if we allow unary weights.
  3. 3.
    Finally, for the case of interval graphs with unary weights we obtain an equivalence with the $\log\!\textrm{pw}$M-SETH, establishing that obtaining a faster than brute-force algorithm for this problem is exactly as hard as refuting this conjecture. So, this problem is placed squarely inside the fourth class (from the bottom) of Figure˜1.
Our results therefore point to a clear ranking of these questions, albeit not with respect to their time complexity, but with respect to their structural complexity. It is also important to note that this investigation goes much further than just confirming the graph-theory intuition that cluster graphs are easier, as there is no obvious graph-theoretic reason why (unary) Weighted Independent Set on cographs should be reducible to (and probably easier than) the same problem on interval graphs (recall that these classes are incomparable).
An interpretation of our results is to conclude that some problems seem to be strictly more difficult than others, even when all the questions we are dealing with (are likely to) have the same time and space complexity. For instance, (unary) Weighted Independent Set parameterized by a modulator to an interval graph appears harder than (that is, appears not reducible to) the same problem parameterized by a modulator to a cluster graph. Our argument for this is that if the contrary were true, two of the classes of Figure˜1 (the ones containing the LD-C-SETH and the $\log\!\textrm{pw}$M-SETH) would collapse. Given the informal intuition that these classes can be thought of as rough analogues of NC1 and NL, we would find this rather remarkable, so we interpret this as evidence that the interval graph question is structurally harder.
Our investigation is then part of a first foray into the structural complexity of fine-grained parameterized questions. As with classical complexity theory, one of the great advantages of constructing a framework around equivalence classes is that this allows us to draw connections and make comparisons between questions which are at first glance unrelated. Indeed, in Section˜7 we consider the complexity of Chromatic Number on graphs which are $k$ vertices away from having constant pathwidth. At first glance this type of question seems incomparable to those of this section, since we are talking about a different problem on a different class of graphs. Nevertheless, because we show this problem to be equivalent to the LD-C-SETH we immediately obtain a relation between this question and the questions of this section. In particular, Chromatic Number parameterized by a modulator to a bounded-pathwidth graph is easier than Weighted Independent Set on interval graphs, again unless the classes containing the LD-C-SETH and $\log\!\textrm{pw}$M-SETH collapse. We believe that many more such characterizations await to be discovered.
Techniques
Let us also briefly sketch some of the main ideas we use. We start with the case of cluster graphs, where for the lower bound we rely on a slight modification of the reduction of [19] (for which we give a self-contained exposition), while for the upper bound we rely on classical results from circuit complexity (Theorem˜2.1). The intuition that the LD-C-SETH morally corresponds to NC1 is crucial here. For block graphs and cographs, the lower bounds rely on reductions from the LD-C-SETH where we can assume that the input circuit has fan-out $1$, and is therefore acyclic (thanks to the depth bound). Block graphs and cographs both possess a tree-like structure that allows us to represent the circuit. In turn, this structure has at most logarithmic tree-depth, enabling a reduction to MaxSAT that gives the upper bounds via Theorem˜4.11.
The most technically challenging of the questions we tackle is the case of weighted interval graphs. The upper bound, given via a reduction to the $\log\!\textrm{pw}$M-SETH is not too difficult, as it consists of an encoding of the direct DP algorithm for this problem. However, for the converse reduction we need a gadget that represents all the assignments of a formula of pathwidth $p$ via an interval graph which for each bag contains $2^{p}$ intersecting intervals (one for each assignment). The difficulty is then to ensure that an independent set selection in one bag correctly propagates to the rest of the graph.
8.1 Cluster Graphs
In this section we present our results for Weighted Independent Set parameterized by cluster vertex deletion, that is, the size of a modulator to a cluster graph. For the lower bound we present a construction that is a slight modification of that used in [19] and which we will reuse later. For the upper bound we construct a circuit with one input for each vertex of the modulator $M$. The idea is that this circuit should output the weight of the best independent set consistent with this choice in the modulator, while still having depth $\varepsilon M$. For this, it is crucial that we can compute the sum of $N$ integers with $N$ bits in depth $O(\log N)$ (this is a standard but non-obvious fact from circuit complexity), as this allows us to sum up the weight contributions from all the cliques of the cluster graph.
8.1.1 Lower Bound
Before we proceed to the lower bound, we recall a gadget construction from [19] which we will repeatedly reuse (with minor modifications). We therefore give here all the relevant details, for the sake of completeness.
Modulator Gadget.
In several reductions in this section we will need to reduce a satisfiability problem, where we are given a boolean formula or circuit with a special set of variables $M$, into an instance of Independent Set where $M$ is represented by a special set of vertices $M^{\prime}$ (a modulator). One obstacle in such constructions is that, even though all $2^{|M|}$ boolean assignments to the original instance are a priori equally plausible, in the new instance, solutions which place more vertices of $M^{\prime}$ in the independent set have higher objective value, making it complicated to achieve a one-to-one correspondence between solutions. We therefore use a trick, going back to [15], which represents $M$ by a modulator of slightly larger size (say $|M^{\prime}|=(1+\varepsilon)|M|$), which allows us to map assignments of $M$ to (roughly) bisections of $M^{\prime}$, that is, independent sets which select roughly half the vertices. By adding appropriate simple gadgets we can ensure that any reasonable solution will indeed take an approximate bisection of $M^{\prime}$ in the independent set, and therefore will encode an assignment to $M$, while keeping the objective value in the new instance predictable.
We start with a simple proposition which states that, for all $\varepsilon$, we can make the extra cost of this gadget at most $(1+\varepsilon)$.
Proposition 8.2.
For all $\varepsilon>0$ there exists an integer $\rho$ and an odd integer $\gamma$ such that $2^{\rho}\leq{\gamma\choose(\gamma-1)/2}$ and $\gamma\leq(1+\varepsilon)\rho$.
Proof 8.3.
First, suppose that $\rho$ is sufficiently large so that there exists an odd integer $\gamma$ in the interval $(1+\varepsilon/2)\rho,(1+\varepsilon)\rho]$ (for instance it suffices to have $\rho>10/\varepsilon$). We will set $\gamma$ to be such an integer, so $(1+\varepsilon)\rho\geq\gamma\geq(1+\varepsilon/2)\rho$.
We now observe that ${\gamma\choose(\gamma-1)/2}\geq\frac{2^{\gamma}}{\gamma+1}$ (because $\sum_{i=0}^{\gamma}{\gamma\choose i}=2^{\gamma}$ and $\gamma\choose i$ is maximized for $i=\frac{\gamma-1}{2}$). Since our goal is to have ${\gamma\choose(\gamma-1)/2}\geq 2^{\rho}$ it will suffice to have $\frac{2^{\gamma}}{\gamma+1}\geq 2^{\rho}$. For this, it suffices to have $2^{\varepsilon\rho/2}\geq(1+\varepsilon)\rho+1$, which always holds for sufficiently large $\rho$.
We now describe the modulator gadget we will use in several reductions of this section. Fix some $\varepsilon>0$ and to ease notation we will switch the informal roles of $M,M^{\prime}$ given above. So, given a set of boolean variables $M^{\prime}$, we will construct a graph $G_{M}$ consisting of a set of vertices $M$ and a disjoint union of cliques (that is, $M$ is a modulator to a cluster graph), as well as a target integer $T$.
More precisely, recall that Proposition˜8.2 ensures that there exist integers $\rho,\gamma$, with $\gamma$ odd, so that $\gamma\leq(1+\varepsilon)\rho$ and $2^{\rho}\leq{\gamma\choose(\gamma-1)/2}$. We partition the variables of $M^{\prime}$ into groups of size at most $\rho$, obtaining at most $\lceil\frac{|M^{\prime}|}{\rho}\rceil$ groups. For each group of variables we construct in our new modulator $M$ a group of $\gamma$ vertices. Therefore, $|M|\leq(1+\varepsilon)|M^{\prime}|+O(1)$.
For each such group of vertices we do the following: construct $\gamma$ identical cliques of size $\gamma\choose(\gamma+1)/2$ where for each clique each vertex is adjacent to a distinct set of vertices of the group of size $\frac{\gamma+1}{2}$. Finally, we set the target independent set size $T$ to be equal to $(\frac{\gamma-1}{2}+\gamma)\lceil\frac{|M^{\prime}|}{\rho}\rceil$. In other words, the maximum independent set is expected to select, for each group, one vertex from each of the $\gamma$ cliques, plus $\frac{\gamma-1}{2}$ vertices of $M$.
Proposition 8.4.
Our construction $G_{M}$ has the following properties:
  1. 1.
    $|M|\leq(1+\varepsilon)|M^{\prime}|+O(1)$.
  2. 2.
    For each group of $\rho$ variables of $M^{\prime}$ and the corresponding group of $\gamma$ vertices of $M$ we can construct an injective function mapping each truth assignment of the $\rho$ boolean variables to a distinct set of $\frac{\gamma-1}{2}$ vertices.
  3. 3.
    All independent sets of $G_{M}$ have size at most $T$.
  4. 4.
    Any independent set that has size $T$ must contain exactly $\frac{\gamma-1}{2}$ vertices from each group of $M$.
  5. 5.
    Any independent set of vertices that contains $\frac{\gamma-1}{2}$ vertices from each group of $M$ can be extended to an independent set of size $T$.
Proof 8.5.
We have already argued for the first property. For the second property, recall that $2^{\rho}\leq{\gamma\choose\frac{\gamma-1}{2}}$, so the number of sets of size $\frac{\gamma-1}{2}$ suffices to construct an injective mapping. For the third and fourth properties, observe that in a group an independent set that contains none of the clique vertices can have at most $\gamma$ vertices (from $M$); while an independent set that contains some vertices of the cliques may contain at most $\frac{\gamma-1}{2}$ vertices of $M$ from this group (otherwise the neighborhood of a selected vertex from the clique, which has size $\frac{\gamma+1}{2}$ would intersect with the selected vertices in $M$). Hence, no set can have size more than $T$ and to attain this size we must be selecting exactly $\frac{\gamma-1}{2}$ vertices of $M$ from each group. For the last property, for each selection of size $\frac{\gamma-1}{2}$ for a group, each clique contains a vertex whose neighborhood is the complement of this selection, so we can select this vertex in every clique.
Theorem 8.6.
Suppose there exist $\varepsilon>0$ and an algorithm which takes as input a vertex-weighted graph $G=(V,E)$ with unary weights and a modulator $M\subseteq V$ such that GMG-M is a cluster graph and computes the maximum weight independent set of $G$ in time $(2-\varepsilon)^{|M|}|V|^{O(1)}$. Then, the MaxSat-SETH is false.
Proof 8.7.
Shifting notation slightly, fix an $\varepsilon>0$ for which the supposed algorithm for Weighted Independent Set exists and runs in time $2^{(1-\varepsilon)|M|}|V|^{O(1)}$. We will use it to obtain a fast algorithm for Max-SAT. Suppose that we are given a CNF formula on $n$ variables and $m$ clauses and a target $t$ and are asked if it is possible to satisfy at least $t$ clauses. We will show how to do this in time $(2-\varepsilon^{\prime})^{n}(n+m)^{O(1)}$.
Let $M^{\prime}$ be the set of variables of the input formula. We execute the modulator gadget construction described above to obtain a graph $G_{M}$ with the properties of Proposition˜8.4. Assign to each vertex of this construction weight $m+1$. This part of the construction represents the variables of the input formula.
To represent the clauses, we consider each clause $c$ in turn. We construct for each such clause a clique which contains, for each literal $\ell$ of $c$, $2^{\rho-1}$ vertices. Let $x$ be the variable of the literal $\ell$. Then, each of these $2^{\rho-1}$ vertices corresponds to an assignment to the variables of the group containing $x$ that sets $\ell$ to True. Recall that each such assignment $\sigma$ is mapped injectively to a set of $\frac{\gamma-1}{2}$ vertices of $M$. We make the clique vertex representing $\sigma$ adjacent to all the other ($\frac{\gamma+1}{2}$) vertices of $M$ in this group. All vertices constructed in this part have weight $1$.
This completes the construction and we claim that the new instance has an independent set of weight $T(m+1)+t$ if and only if it is possible to satisfy at least $t$ clauses in the original instance, where $T$ is the integer of Proposition˜8.4. For one direction, if some assignment to the formula satisfies $t$ clauses, we use the injective mapping from assignments to sets of vertices of $M$ to obtain a set of $T$ vertices of the first part of the construction (this is always possible by the last property of Proposition˜8.4), which therefore contributes a weight of $T(m+1)$. We now claim that for each satisfied clause, there is a vertex in the corresponding clique that we can add to the set. Indeed, each such clause $c$ must contain a True literal $\ell$, so we have selected an assignment mapped to a set of $\frac{\gamma-1}{2}$ vertices of $M$ which are disjoint from the neighborhood of a vertex of the clique representing $\ell$.
For the converse direction, observe that the maximum independent set of the new instance must select $T$ vertices in the first part. Indeed, since the second part consists of $m$ cliques, any solution that contains at most T1T-1 vertices of the first part has weight at most (T1)(m+1)+m(T-1)(m+1)+m, while a solution that selects $T$ vertices of the first part and nothing else always exists and has weight $T(m+1)$, which is strictly larger. But then, by the fourth property of Proposition˜8.4 we must have selected exactly $\frac{\gamma-1}{2}$ vertices of $M$ from each group. From this selection we extract an assignment to the variables by selecting for each group the pre-image of the set of $\frac{\gamma-1}{2}$ vertices of $M$ placed in the independent set under our injective mapping (if no such pre-image exists, give some arbitrary truth assignment). We claim this assignment satisfies $t$ clauses. Indeed, there must be $t$ cliques constructed in the second part where the independent set takes a vertex, say a vertex representing literal $\ell$ and assignment $\sigma$ to the variables of the group of the variable of $\ell$. However, since the selected vertex has $\frac{\gamma+1}{2}$ neighbors in the corresponding group of $M$, the vertices of $M$ we have selected must be the image of assignment $\sigma$, so we have indeed given an assignment setting $\ell$ to True.
Finally, observe that $M$ is a modulator to a cluster graph, the reduction takes polynomial time, and weights are polynomially bounded. Executing the supposed algorithm on the new instance takes time $2^{(1-\varepsilon)|M|}(n+m)^{O(1)}=2^{(1-\varepsilon^{2})|M^{\prime}|}(n+m)^{O(1)}=(2-\varepsilon^{\prime})^{n}(n+m)^{O(1)}$, for some $\varepsilon^{\prime}>0$.
8.1.2 Upper Bound
Relying heavily on Theorem˜2.1 we have the following:
Theorem 8.8.
If the LD-C-SETH is false, then we have the following: there exist $\varepsilon>0$ and an algorithm which takes as input a graph $G=(V,E)$ and a modulator $M\subseteq V$ such that GMG-M is a cluster graph and computes the maximum independent set of $G$ in time $(2-\varepsilon)^{|M|}|V|^{O(1)}$. The same holds if the vertices of $G$ have weights bounded by $2^{|V|^{O(1)}}$ and we want to compute the independent set of maximum weight.
Proof 8.9.
Suppose the LD-C-SETH is false, so there exist $\varepsilon>0$ and an algorithm that takes as input a bounded fan-in boolean circuit on $N$ inputs with depth at most $\varepsilon N$ and total size $s$ and decides if the circuit is satisfiable in time $(2-\varepsilon)^{N}s^{O(1)}$. We will use this algorithm to solve Weighted Independent Set. We are given a graph $G=(V,E)$ and a modulator $M$ as in the statement of the lemma. Let $|M|=m$ and $|V\setminus M|=n$.
Observe that it is easy to solve Weighted Independent Set of the given instance in time $2^{m}n^{O(1)}$: guess the vertices of $M$ to take into the set, remove all their neighbors from the graph, and then select the maximum weight vertex from each cluster. We can therefore assume without loss of generality that for all $\delta>0$ we have $n<2^{\delta m}$, as otherwise, $2^{m}<n^{1/\delta}$ and we can solve the given instance in polynomial time.
We now construct a circuit with $N=m$ inputs, one for each vertex of $M$, where an input is meant to be set to True if the corresponding vertex is placed in the independent set. For each $v\in V\setminus M$ we construct a gate that decides if $v$ is available by checking if its neighbors in $M$ have been selected. This gate is just a conjunction ($\land$) of at most $m$ negated inputs, so implementing this with bounded fan-in gates gives depth at most $O(\log m)$.
Let $W$ be the smallest power of $2$ such that the sum of all input weights is at most $2^{W}$, that is, the sum of all input weights can be encoded with $W$ bits. We have $W=(n+m)^{O(1)}$, because each input weight can be encoded with $(n+m)^{O(1)}$ bits.
For each $v\in V$ we construct a group of $W$ gates, hard-coded to 0 or $1$, that encode the weight of $v$. We take the conjunction of each such bit with the output of the gate that states whether $v$ is available (for $v\not\in M$) or with the corresponding input gate (for $v\in M$). We now have $n+m$ blocks of $W$ gates each, encoding the weight that each vertex may contribute to the solution without selecting the two endpoints of any edge incident on $M$. What remains is to add some machinery to ensure we also do not select two vertices from a clique and that we select the maximum weight vertex from each clique.
Consider a clique $C$ of GMG-M and let $C=\{v_{1},\ldots,v_{c}\}$. We construct a Max circuit which, given as input the $c$ blocks of $W$ gates encoding the weights of vertices from $C$ compatible with a choice for $M$ outputs the weight of the heaviest available vertex of $C$. As stated in Theorem˜2.1 this can be done with depth $O(\log W)=O(\log(n+m))$.
Suppose that GMG-M contains $r$ clique components $C_{1},\ldots,C_{r}$ and for each such component we have constructed a circuit with $W$ outputs that encodes the weight of the heaviest available vertex. We now add a circuit that computes the sum of these numbers as well as the (at most) $m$ weights of the vertices of the modulator (taking into account that the weight of a vertex whose input is set to 0 becomes 0). By Theorem˜2.1 we construct a circuit with depth $O(\log W)=O(\log(n+m))$. Finally, given some desired target value, we add to the end a circuit checking that the calculated sum is at least as large as the target; this adds another $O(\log W)=O(\log(n+m))$ to the depth.
What remains is to argue that the depth of the constructed circuit is at most $\varepsilon N=\varepsilon m$. However, we can see that the total depth is at most $O(\log(n+m))$. We have $\log(n+m)\leq\log n+\log m\leq\delta m$ for all $\delta>0$. This follows because we can assume that $m$ is sufficiently large so that $\log m$ is insignificant compared to $m$; and because as we argued $n<2^{\delta m}$ for all $\delta>0$.
8.2 Cographs
In this section we deal with cographs, which as mentioned are exactly the graphs which exclude $P_{4}$ as an induced subgraph. We will use an equivalent characterization of this class of graphs, which states that $G$ is a cograph if it can be constructed from singleton vertices using repeated applications of the union and join operations (the join operation takes the union of two graph $G_{1},G_{2}$ and adds all edges between them). It is known that, given a cograph $G$, one can construct in polynomial time a binary tree proving that $G$ can be constructed using these operations [13].
Our lower bound reduces the LD-C-SETH to Independent Set parameterized by cograph modulator size. The two main intuitions we rely on are that circuits with $N$ inputs and depth $\varepsilon N$ can, without loss of generality, be considered acyclic; and that the binary tree construction representing a cograph can encode a sufficient amount of information to represent an acyclic (tree-like) circuit. For the upper bound, we rely on the fact that this tree-like representation has logarithmic tree-depth.
8.2.1 Lower Bound
Theorem 8.10.
Suppose there exist $\varepsilon>0$ and an algorithm which takes as input a graph $G=(V,E)$ and a modulator $M\subseteq V$ such that GMG-M is a cograph and computes the maximum independent set of $G$ in time $(2-\varepsilon)^{|M|}|V|^{O(1)}$. Then, the LD-C-SETH is false.
Proof 8.11.
Slightly switching notation, we are given a bounded fan-in boolean circuit with $N$ inputs, size $s$, and depth $\varepsilon N$ and want to decide if it is satisfiable in time $2^{(1-\varepsilon)N}s^{O(1)}$, for some $\varepsilon>0$. We first show that we can assume without loss of generality that the circuit is actually a formula, that is, all non-input gates have fan-out $1$.
Indeed, suppose that we have an algorithm for this special case running in time $2^{(1-\varepsilon)N}s^{c}$, for some fixed constant $c$ and for circuits of depth at most $\varepsilon N$. We will show that if we are given a depth $\varepsilon^{\prime}N$ circuit (for some $\varepsilon^{\prime}<\varepsilon$ to be set later) where some gates have fan-out more than $1$ we can decide its satisfiability in time $2^{(1-\varepsilon^{\prime})N}$. More precisely, in this case we replicate each gate with large fan-out that is at distance $d$ from the output with $2^{d}$ gates of fan-out $1$, increasing the size of the circuit by at most a factor of $2^{\varepsilon^{\prime}N}$. The new circuit has fan-out $1$ and depth $\varepsilon^{\prime}N<\varepsilon N$, so the supposed algorithm solves it in $2^{(1-\varepsilon)N}2^{\varepsilon^{\prime}cN}s^{c}$. It suffices to have $\varepsilon-\varepsilon^{\prime}c>\varepsilon^{\prime}$ or equivalently to set $\varepsilon^{\prime}<\frac{\varepsilon}{c+1}$ to obtain the desired algorithm for the case of circuits with larger fan-out.
We can therefore view the input circuit as a rooted tree, with the output gate at the root, and each leaf labeled with one of the $N$ input variables or its negation (which of course may appear on many leaves). Notice that by moving the negations as far away from the root as possible we can ensure that all internal gates are conjunctions or disjunctions.
We will now produce an instance of Independent Set. Suppose the hypothesized algorithm for Independent Set runs in time $2^{(1-\delta)|M|}|V|^{O(1)}$.
For the first part of our construction, we use the modulator gadget of Proposition˜8.4 to represent the $N$ input variables using a modulator $M$ of size at most $(1+\delta)N+O(1)$. Let $T$ be the target independent set size in this part and recall that we have now partitioned $N$ into groups of $\rho$ variables, each of which is represented by $\gamma$ vertices of $M$ such that each assignment is encoded by a distinct set of $\frac{\gamma-1}{2}$ vertices.
For the second part, for each gate $g$ of our circuit we will consider the sub-tree rooted at this gate and produce a pair $(G_{g},T_{g})$ such that (i) $G_{g}$ is a cograph (ii) all independent sets of $G_{g}$ have size at most $T_{g}$ (iii) it is possible to place $T_{g}$ vertices of $G_{g}$ into an independent set if and only if our selection in the modulator part encodes an assignment that sets gate $g$ to True. These properties will be established by induction.
We start with the leaves, which are labeled with literals using input variables. For such a gate labeled with a literal $\ell$ using variable $x$ we construct a clique on $2^{\rho-1}$ vertices, where each vertex corresponds to an assignment to the $\rho$ variables of the group of $x$ that sets $\ell$ to True. Take each such vertex that represents an assignment $\sigma$ and make it adjacent to the $\frac{\gamma+1}{2}$ vertices of the group in $M$ which are the complement of the encoding of $\sigma$. The target size for this graph is $1$. We see that (i) is satisfied as cliques are cographs (ii) is satisfied as the maximum independent set size of a clique is $1$ and (iii) is satisfied because if we have selected an assignment to the group containing $x$ that falsifies $\ell$, then all vertices of the clique have a neighbor in the independent set; while in the opposite case, one vertex of the clique is available.
We now proceed inductively. For an AND gate $g$ with children $g_{1},g_{2}$, suppose we have constructed the pairs $(G_{g_{1}},T_{g_{1}})$ and $(G_{g_{2}},T_{g_{2}})$. We construct the pair $(G_{g_{1}}\cup G_{g_{2}},T_{g_{1}}+T_{g_{2}})$. We have that (i) the new graph is a cograph, as it is the disjoint union of two cographs (ii) its maximum independent set is at most $T_{g_{1}}+T_{g_{2}}$ (iii) an independent set can have this many vertices only if it has $T_{g_{1}}$ vertices from $G_{g_{1}}$ and $T_{g_{2}}$ vertices from $G_{g_{2}}$, which implies that $g_{1},g_{2}$ evaluate to True, therefore so does $g$; conversely, if the assignments sets $g$ to True, then $g_{1},g_{2}$ are set to True, so independent sets of sizes $T_{g_{1}},T_{g_{2}}$ exist by induction.
For an OR gate $g$ with children $g_{1},g_{2}$, suppose we have constructed the pairs $(G_{g_{1}},T_{g_{1}})$ and $(G_{g_{2}},T_{g_{2}})$, where without loss of generality $T_{g_{1}}\leq T_{g_{2}}$. We set $G_{g}$ to be the join of $G_{g_{1}}\cup(T_{g_{2}}-T_{g_{1}})K_{1}$ with $G_{g_{2}}$ and $T_{g}=T_{g_{2}}$. In other words, we add $T_{g_{2}}-T_{g_{1}}$ isolated vertices to $G_{g_{1}}$, take the join of the two graphs, and set the target to the higher of the targets of the two children. We have that (i) the new graph is a cograph, as it is obtained from cographs via union and join operations (ii) no independent set can have size more than $T_{g_{2}}$, since any independent set can only take vertices from one side of the join operation (iii) if an independent set of the new graph has size $T_{g_{2}}$ either $G_{g_{2}}$ has an independent set of this size (so $g_{2}$ evaluates to True, therefore so does $g$) or $G_{g_{1}}$ has an independent set of size $T_{g_{1}}$ (and similarly $g$ evaluates to True); conversely if $g$ evaluates to True, then either $g_{1}$ does too (so we can form an independent set of size $T_{g_{1}}$ and augment it with the $T_{g_{2}}-T_{g_{1}}$ new vertices) or $g_{2}$ does (so we have an independent set of size $T_{g_{2}}$).
Once we arrive at the root $r$ we have constructed a cograph $G_{r}$ and a target size $T_{r}$. We set the total target size to be $T+T_{r}$ and our instance is made up of $G_{r}$ and the modulator gadget construction which consists of $M$ and some disjoint cliques. Therefore, removing $M$ indeed gives us a cograph and the whole construction can be performed in polynomial time. Correctness is not hard to see: because $G_{r}$ cannot have an independent set of size more than $T_{r}$ and the modulator part cannot have an independent set of size $T$ (Proposition˜8.4) we must take exactly $T$ vertices from the modulator part and $T_{r}$ vertices from the cograph. The $T$ vertices from the modulator part force us to encode an assignment to the $N$ inputs, which must set the output gate to True (otherwise we cannot attain $T_{r}$ vertices in the cograph).
Running the hypothetical algorithm on this instance takes time $2^{(1-\delta)(1+\delta)N}s^{O(1)}=2^{(1-\delta^{2})N}s^{O(1)}$. So, for $\varepsilon=\delta^{2}$, if we start with a circuit of fan-out $1$ and depth at most $\varepsilon N$ we obtain the desired algorithm.
8.2.2 Upper Bound
Theorem 8.12.
If the LD-C-SETH is false, then we have the following: there exist $\varepsilon>0$ and an algorithm which takes as input a graph $G=(V,E)$ and a modulator $M\subseteq V$ such that GMG-M is a cograph and computes the maximum independent set of $G$ in time $(2-\varepsilon)^{|M|}|V|^{O(1)}$. The same holds if the vertices of $G$ have weights bounded by $|V|^{O(1)}$ and we want to compute the independent set of maximum weight.
Proof 8.13.
Suppose the LD-C-SETH is false, so by Theorem˜4.11 there is an algorithm that takes as input an integer $t$, a CNF formula $\phi$, a modulator $M^{\prime}$, and a tree-depth decomposition of the primal graph of $\phi-M^{\prime}$ of depth $O(\log|\phi|)$, and decides if at least $t$ clauses of $\phi$ can be satisfied, in time $(2-\varepsilon)^{|M^{\prime}|}|\phi|^{O(1)}$. We will use this algorithm to solve Independent Set. We are given a graph $G=(V,E)$ and a modulator $M$ as in the statement of the lemma. Let $|M|=m$ and $|V\setminus M|=n$.
Recall that for any cograph $H$ one can construct in polynomial time a rooted binary tree showing how $H$ can be constructed from isolated vertices using Union and Join operations. In particular, there exists a rooted binary tree where each internal node has two children and is labeled Union or Join and where we can inductively associate with every node of the tree an induced subgraph of $H$ as follows: leaves correspond to single vertices of $H$; for an internal node labeled Union (resp. Join) the corresponding graph is constructed by taking the disjoint union (resp. the join, that is, the union where we add all possible edges between the two parts) of the graphs corresponding to the children. In such a tree the root corresponds to $H$ and there is a one-to-one correspondence between vertices of $H$ and leaves of the tree.
We construct a CNF formula for $G$ as follows. First, for each vertex $v\in V$ we construct a variable $x_{v}$ (indicating whether $v$ is in the maximum independent set). Then, we construct the binary tree representing the cograph GMG-M, which has at most $2n$ nodes. For each node $t$ of this tree we construct a variable $a_{t}$, informally indicating whether the corresponding sub-tree is “active”, that is whether the optimal independent set contains any vertices from this sub-tree. We now construct the following clauses:
  1. 1.
    For each $v_{1}v_{2}\in E$ such that $\{v_{1},v_{2}\}\cap M\neq\emptyset$ we add the clause $(\neg x_{v_{1}}\lor\neg x_{v_{2}})$.
  2. 2.
    For each leaf node representing the vertex $v\in V\setminus M$ we add the clause $(\neg x_{v}\lor a_{v})$.
  3. 3.
    For each non-root node $t$ of the binary tree, if $p$ is the parent of $t$ in the tree we add the clause $(\neg a_{t}\lor a_{p})$.
  4. 4.
    For each Join node $t$ with children $t_{1},t_{2}$ we add the clause $(\neg a_{t_{1}}\lor\neg a_{t_{2}})$.
We set $M^{\prime}=\{x_{v}\ |\ v\in M\}$. We have now constructed a CNF formula $\phi$ with the following properties:
  1. 1.
    For all $\ell>0$, $G$ has an independent set of size $\ell$ if and only if $\phi$ has a satisfying assignment setting $\ell$ variables from $\{x_{v}\ |\ v\in V\}$ to True.
  2. 2.
    $\phi-M^{\prime}$ has tree-depth $O(\log n)$, $|\phi|=(n+m)^{O(1)}$, and $|M^{\prime}|=|M|=m$.
To see the second property, observe that $\phi$ is a 2-CNF formula, and once we remove $M^{\prime}$ the primal graph has treewidth $2$ (it is a tree, except the two children of each Join node are adjacent). Therefore, this graph has tree-depth $O(\log n)$. For the first property, if we have an independent set $S$ of $G$ of size $\ell$, we set the corresponding variables according to their informal meaning, that is, $x_{v}$ is True if and only if $v\in S$; and $a_{t}$ is True if and only if $S$ contains a vertex mapped to a leaf that is a descendant of $t$. It is not hard to see that this satisfies $\phi$. For the converse direction, we extract a subset $S$ of $V(G)$ from a satisfying assignment of $\phi$ by selecting all the vertices $v$ for which $x_{v}$ is True. It is clear that $S$ cannot contain both endpoints of an edge that is incident on $M$, by the clauses of step 1. Suppose then that $S$ contains both endpoints of an edge of GMG-M and let $t$ be the Join node of the tree whose corresponding subgraph contains this edge and which is as far from the root as possible. Suppose that the two children of $t$ are $t_{1}$ and $t_{2}$, and by the selection of $t$ each contains one endpoint of the offending edge. We claim that in this case we must have $a_{t_{1}},a_{t_{2}}$ both set to True, thus falsifying a clause constructed in step 4. To see this, observe that in a satisfying assignment of $\phi$, whenever we have $x_{v}$ set to True, we must have $a_{t}$ set to True for all nodes $t$ such that $v$ is mapped to a descendant of $t$. This can be seen by induction: clauses of step 2 ensure that it holds for leaves; and clauses of step 3 ensure that it holds for all higher levels of the tree.
To conclude the proof we have to argue that the satisfiability algorithm implied by the assumption that the LD-C-SETH is false and Theorem˜4.11 allows us to solve the resulting instance. To see this, we first observe that the algorithm of Theorem˜4.11 can handle instances where clauses have (polynomially-bounded) weights (for instance, by repeating each clause a number of times equal to its weight). We can therefore assign to all clauses of our constructed instance a high enough weight (polynomial in the construction) ensuring that all clauses need to be satisfied. Then, for each vertex $v$, we have constructed a variable $x_{v}$. We simply add to the instance the clause $x_{v}$ with weight equal to the weight of the corresponding vertex in the input graph.
8.3 Block Graphs
A graph is a block graph if every $2$-vertex-connected component is a clique [5]. Block graphs are a well-studied special subclass of chordal graphs which also includes cluster graphs and trees as subclasses. Clearly, Weighted Independent Set can be solved in polynomial time on block graphs. In this section we investigate the complexity of solving this problem parameterized by the size of a modulator to a block graph. Interestingly, even though block graphs are incomparable to cographs, the general tenor of our results is the same for the two problems, namely: we show that beating brute-force is equivalent to refuting the LD-C-SETH (using again the intuition that we can focus on acyclic circuits and that block graphs contain a tree-like structure in the form of the block decomposition).
8.3.1 Lower Bound
Theorem 8.14.
Suppose there exist $\varepsilon>0$ and an algorithm which takes as input a graph $G=(V,E)$ and a modulator $M\subseteq V$ such that GMG-M is a block graph and computes the maximum independent set of $G$ in time $(2-\varepsilon)^{|M|}|V|^{O(1)}$. Then, the LD-C-SETH is false.
Proof 8.15.
Similarly to Theorem˜8.10, we will assume that we are given a boolean circuit with $N$ inputs, size $s$, bounded fan-in, fan-out $1$, and dept at most $\varepsilon N$ and we want to decide its satisfiability in time $2^{(1-\varepsilon)N}s^{O(1)}$, for some $\varepsilon>0$. Suppose that an algorithm solves Independent Set in time $2^{(1-\delta)|M|}|V|^{O(1)}$, given a modulator $M$ to a block graph.
We construct a graph from the given circuit as follows. To represent the $N$ variables we (again) use the construction of Proposition˜8.4 to obtain a set $M$ of at most $(1+\delta)N+O(1)$ vertices. The variables $N$ are partitioned into groups of size $\rho$ and the groups are matched to disjoint groups of $\gamma$ vertices. Furthermore, assignments to each group are injectively mapped to sets of $\frac{\gamma-1}{2}$ vertices inside the corresponding set of size $\gamma$.
We now inductively build a block graph from the circuit. For each gate $g$ we build a graph representing the circuit rooted at $g$. If there are $s(g)$ gates in total in the sub-tree rooted at $g$, then this graph will have maximum independent set of size exactly $s(g)$. Furthermore, it will contain a special output vertex $o_{g}$. We will have the property that it is possible to find an independent set of size $s(g)$ without selecting $o_{g}$ if and only if $g$ evaluates to True if we select the input assignment encoded by the independent set we selected in the modulator $M$.
For an input gate labeled with the literal $\ell$, using variable $x$, we construct a clique of size $2^{\rho-1}+1$ (so the maximum independent set indeed has size $s(g)=1$). Label one of the vertices of the clique as the output vertex. Each other vertex is labeled with an assignment $\sigma$ to the variables of the group of $x$ that sets $\ell$ to True. For each such vertex representing assignment $\sigma$, we make it adjacent to the $\frac{\gamma+1}{2}$ vertices of the corresponding group in $M$ which are the complement of the encoding of $\sigma$. We can now see that we can form an independent set of size $1$ while avoiding the output vertex if and only if we select one of the vertices representing an assignment, which can be done if and only if we encode an assignment to the input that set $\ell$ to True.
For an OR gate $g$ with two children $g_{1},g_{2}$ we construct a clique of size $3$. One vertex of this clique is adjacent to $o_{g_{1}}$, one to $o_{g_{2}}$, and one is the output $o_{g}$ of this gate. Since we added a clique to the graphs constructed for the sub-trees rooted at $g_{1},g_{2}$, the maximum independent set size is at most $1+s(g_{1})+s(g_{2})=s(g)$. If it is possible to find an independent set of size $s(g)$ avoiding $o_{g}$, then we must take one of the other vertices of the new clique, say the one adjacent to $o_{g_{1}}$; then, since we must have taken an independent set of size $s(g_{1})$ on the sub-tree for $g_{1}$, $g_{1}$ must evaluate to True, and so does $g$. Conversely, if $g$ evaluates to True, then one of its children (say $g_{1}$) does, so we can find an independent set that avoids $o_{g_{1}}$; we then add the corresponding vertex of the new clique in the independent set and avoid $o_{g}$.
For an AND gate $g$ we construct two adjacent vertices: one is labeled $o_{g}$ and is the output, while the other is adjacent to the outputs of the children of $g$. It is not hard to see that the desired properties are satisfied inductively in this case as well.
Finally, for the output gate we remove the output vertex that the above construction would have added. We set the target independent set size to be $T+s$, where $s$ is the number of gates of the circuit. Correctness is not hard to see, because by induction we cannot take more than $s$ vertices in the block graph part, while by Proposition˜8.4 we cannot take more than $T$ in the modulator part; hence we take exactly $T$ in the modulator, ensuring that we can only take $s$ vertices in the block graph part if we satisfy the output gate (as we have removed its output vertex). To see that removing $M$ results in a block graph, observe that in the modulator gadget we get a disjoint union of cliques (which is a block graph), while to represent the circuit we have a rooted tree of cliques, so each $2$-vertex-connected component is a clique. The remaining calculations are identical to those of Theorem˜8.10.
8.3.2 Upper Bound
Theorem 8.16.
If the LD-C-SETH is false, then we have the following: there exist $\varepsilon>0$ and an algorithm which takes as input a graph $G=(V,E)$ and a modulator $M\subseteq V$ such that GMG-M is a block graph and computes the maximum independent set of $G$ in time $(2-\varepsilon)^{|M|}|V|^{O(1)}$. The same holds if the vertices of $G$ have weights bounded by $|V|^{O(1)}$ and we want to compute the independent set of maximum weight.
Proof 8.17.
Suppose the LD-C-SETH is false, so there is an algorithm that takes as input a target integer $t$, a CNF formula $\phi$, a modulator $M^{\prime}$, and a tree-depth decomposition of the primal graph of $\phi-M^{\prime}$ of width $O(\log|\phi|)$, and decides if we can satisfy at least $t$ clauses of $\phi$, in time $(2-\varepsilon)^{|M^{\prime}|}|\phi|^{O(1)}$. We will use this algorithm to solve Independent Set. We are given a graph $G=(V,E)$ and a modulator $M$ as in the statement of the lemma. Let $|M|=m$ and $|V\setminus M|=n$.
We first formulate a claim stating that Independent Set on block graphs can be encoded by a CNF formula of tree-depth $O(\log n)$.
Claim 4.
There is a polynomial time algorithm which takes as input a block graph on $n$ vertices and produces a CNF formula $\phi$ of size $n^{O(1)}$ and a tree-depth decomposition of $\phi$ of width $O(\log n)$ as well as a special set of variables $X$ which are in one-to-one correspondence with the vertices of the graph. The algorithm guarantees that for all sets of vertices $S$ of the given graph, $S$ is independent if and only if $\phi$ has a satisfying assignment that sets the variables of $X$ corresponding to $S$ to True.
Before we give the proof of the claim let us explain why it implies the lemma, using a strategy that is essentially identical to the proof of Theorem˜8.12. We apply ˜4 to the graph GMG-M and obtain a formula $\phi$. We extend this formula with a variable $x_{v}$ for each $v\in M$ and add clauses ensuring that for each edge incident on at least one vertex of $M$, the two endpoints of the edge cannot be selected (this can be done as $\phi$ has a set of variables $X$ representing the vertices of GMG-M). We set $M^{\prime}$ to be the set of new variables and therefore $\phi-M^{\prime}$ has tree-depth $O(\log n)$. We have that the new formula has a satisfying assignment setting $\ell$ variables representing vertices to True if and only if $G$ has an independent set of size $\ell$. To reduce this maximization version to the MaxSAT case of Theorem˜4.11 we again give very high weight to necessary clauses and introduce clauses $(x_{v})$, with weight equal to the weight of vertex $v$. We omit the details as they are identical to the proof of Theorem˜8.12.
Let us then proceed to the proof of the claim.
{claimproof}
[Proof of ˜4]
Recall that a block is a maximal $2$-vertex-connected component in a graph. In any graph, if $B_{1},B_{2}$ are blocks we have that $|B_{1}\cap B_{2}|\leq 1$, because if $B_{1},B_{2}$ share two vertices, then $B_{1}\cup B_{2}$ is also $2$-vertex-connected, contradicting maximality. Furthermore, for any two blocks $B_{1},B_{2}$ we have at most one edge connecting them. Indeed, if we have two edges with distinct endpoints connecting $B_{1},B_{2}$, then $B_{1}\cup B_{2}$ is $2$-vertex-connected; while if two edges incident on $x\in B_{1}$ have their endpoints in $B_{2}$, then $B_{2}\cup\{x\}$ is $2$-vertex-connected.
Given an input block graph $H$, we construct a graph $H^{\prime}$ that contains a vertex for each block of $H$ and has an edge between $B_{1},B_{2}$ if $B_{1},B_{2}$ are adjacent or share a vertex. We can see that $H^{\prime}$ is acyclic, as otherwise one of the blocks would not be maximal. We will use this fact to show that the treewidth of our formula is $O(1)$. Because for all graphs $G$ we have $\textrm{td}(G)=O(\textrm{tw}(G)\log n)$, this will imply the claim.
We begin our formula by constructing, for each vertex $v$ a variable $x_{v}$ (indicating whether $v$ is in the independent set). These variables will form the set $X$. We now prove that we can build a CNF formula such that assignments to $X$ can be extended to satisfying assignments if and only if the set of True variables of $X$ represents an independent set.
Consider each node of $H^{\prime}$, which represents a clique $C=\{v_{1},\ldots,v_{c}\}$ with $c=|C|$ of $H$. We introduce c1c-1 new variables $y_{1},\ldots,y_{c-1}$ whose informal meaning is that if $y_{i}$ True then all the variables $x_{v_{1}},x_{v_{2}},\ldots,x_{v_{i}}$ are False. We add the clause $(\neg y_{1}\lor\neg x_{v_{1}})$ and for all $i\in\{2,\ldots,c-1\}$ the clauses $(\neg y_{i}\lor y_{i-1})$ and $(\neg y_{i}\lor\neg x_{v_{i}})$. Furthermore, for all $i\in\{2,\ldots,c-1\}$ we add the clause $(\neg x_{v_{i}}\lor y_{i-1})$. We now claim that the new clauses can only be satisfied if and only if at most $1$ of the $x_{v_{i}}$ variables is set to True. To see this, observe that if we set $x_{v_{i}}$ and $x_{v_{j}}$ to True, for $j>i$, we must set $y_{j-1}$ to True; then we must set $y_{j-2}$ to True, and proceeding in this way $y_{i}$ to True; therefore $x_{v_{i}}$ must be set to False, contradiction. It is easy to see, however, that setting a single $x_{v_{i}}$ to True can always be extended to a satisfying assignment. In other words, the clauses we have constructed so far ensure that any assignment to $X$ that sets to True to variables representing vertices of the same clique will fail to satisfy the formula. To conclude the construction we consider any pair of vertices $u,v$ of $H$ such that $u,v$ are adjacent but belong in distinct 2-connected components. For any two such vertices we add the clause $(\neg x_{u}\lor\neg x_{v})$.
It is now not hard to see that the formula we have constructed so far is correct in the sense that satisfying assignments are always extensions of assignments that encode independent sets of $H$. Let us then argue that we have constant treewidth. We consider each vertex of $H^{\prime}$ separately. Consider the node of $H^{\prime}$ that represents the clique $C=\{v_{1},\ldots,v_{c}\}$ and recall that we used the variables $x_{v_{1}},\ldots,x_{v_{c}}$ and the c1c-1 new variables $y_{1},\ldots,y_{c-1}$. Because our clauses have a path-like structure (each $y_{i}$ is involved in clauses with $y_{i-1}$ and $x_{v_{i}}$) we can produce a tree (in fact path) decomposition with at most $3$ variables per bag.
We now have a tree decomposition for the part of the formula representing each vertex of $H^{\prime}$. To transform this into a decomposition of the whole formula, we consider any two $h_{1},h_{2}$ which are adjacent in $H^{\prime}$. If these correspond to blocks that share a vertex $u$, we select a bag in each decomposition that contains $u$ and make the two bags adjacent; if $h_{1},h_{2}$ represent blocks connected by an edge with endpoints $u,v$, then we construct a new bag that contains $u,v$ and make it adjacent to a bag that contains $u$ and a bag that contains $v$. It is now not hard to see that because $H^{\prime}$ is acyclic, performing these operations will merge the tree decompositions for the vertices of $H$ into a single tree decomposition for the whole formula.
Having established ˜4 we can conclude that a fast algorithm falsifying the $\log\!\textrm{pw}$M-SETH would imply a fast algorithm for Independent Set using a modulator to a block graph.
8.4 Weighted Interval Graphs
In this section we consider the Weighted Independent Set problem on graphs which are “close” to being interval graphs. Our goal is to prove the following:
Theorem 8.18.
The following two statements are equivalent:
  1. 1.
    There exist $\varepsilon>0$ and an algorithm which takes as input a vertex-weighted graph $G=(V,E)$ with unary weights and a modulator $M\subseteq V$ such that GMG-M is an interval graph and computes the maximum weight independent set of $G$ in time $(2-\varepsilon)^{|M|}|V|^{O(1)}$.
  2. 2.
    The $\log\!\textrm{pw}$M-SETH is false.
Before we proceed, let us give some intuition. Recall that the $\log\!\textrm{pw}$M-SETH informally corresponds to NL, for which complete problems typically have a Reachability flavor. What we are claiming here is that Weighted Independent Set on graphs which are close to interval graphs is a problem of the same type. Informally, this happens because the natural algorithm for this problem would proceed as follows: first, guess which vertices of $M$ are in an optimal solution; then, solve the problem on the remaining instance, which is an interval graph. However, solving Weighted Independent Set on interval graphs can naturally be done via dynamic programming, which in turn can be seen as a form of a reachability question.
Let us begin with the easier direction of the proof of Theorem˜8.18.
Lemma 8.19.
The second statement of Theorem˜8.18 implies the first.
Proof 8.20.
By Theorem˜5.1, we can assume that there exists an algorithm which takes as input a CNF formula $\phi$, a modulator $M^{\prime}$ of size $m^{\prime}$, and a path decomposition of the primal graph of $\phi-M^{\prime}$ of width $O(\log|\phi|)$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m^{\prime}}|\phi|^{O(1)}$, for some $\varepsilon>0$.
Suppose now we are given a weighted graph $G=(V,E)$, a modulator $M\subseteq V$ of size $m$ such that GMG-M is an interval graph, and a target weight $T$. We are asked if $G$ has an independent set of weight at least $T$. Let n=|V|mn=|V|-m. Without loss of generality, assume that the sum of all weights is at most $n$ (this can be achieved by adding isolated vertices of weight 0 to the graph so that |V|m|V|-m becomes larger than the largest weight; this can be performed in polynomial time, since weights are given in unary) and that $n$ is a power of $2$ (again, by adding isolated 0-weight vertices as necessary). Suppose also that we have an interval representation of GMG-M, which assigns an interval $s_{v},t_{v}]$ to each $v\in V\setminus M^{\prime}$ such that $s_{v},t_{v}]\cap[s_{u},t_{u}]\neq\emptyset$ if and only if $uv\in E$. Such a representation can be found in linear time [12]. Again, without loss of generality we may assume that the endpoints of all intervals are distinct, so we have $2n$ endpoints in total, and they correspond to the set $\{1,\ldots,2n\}$.
We will construct a CNF formula $\phi$ and a modulator $M^{\prime}$ with the following properties:
  1. 1.
    $\phi$ is satisfiable if and only if $G$ has a weighted independent set of weight at least $T$.
  2. 2.
    $\phi$ has size polynomial in $n+m$ and $|M|=|M^{\prime}|=m$.
  3. 3.
    $\phi-M^{\prime}$ has pathwidth $O(\log n)$.
  4. 4.
    $\phi,M^{\prime}$ and a path decomposition of $\phi-M^{\prime}$ of width $O(\log n)$ can be constructed in time polynomial in $n+m$.
Clearly, if we can achieve all the above we obtain the lemma, as in such a case we could invoke the algorithm that breaks the $\log\!\textrm{pw}$M-SETH, solving the resulting instance in $(2-\varepsilon)^{m}(n+m)^{O(1)}$.
Assume that the vertices of $M$ are numbered $\{1,\ldots,m\}$ and the vertices of $V\setminus M$ are numbered $\{0,\ldots,n-1\}$. We construct $\phi$ as follows:
  1. 1.
    For each $v\in M$ we construct a variable $x_{v}$, indicating whether $v$ is in the optimal independent set. Let $M^{\prime}$ be the set of these variables.
  2. 2.
    For each $i\in\{1,\ldots,2n\}$ we construct a variable $a_{i}$ indicating whether the optimal solution has selected an interval that contains $i$.
  3. 3.
    For each $i\in\{1,\ldots,2n\}$ we construct $\log n$ variables $y_{i,1},\ldots,y_{i,\log n}$ which are supposed to be the binary representation of the index of the (at most one) vertex of the independent set whose interval contains $i$.
  4. 4.
    For each $i\in\{1,\ldots,2n+m\}$ we construct $\log n$ variables $z_{i,1},\ldots,z_{i,\log n}$. For $j\in\{1,\ldots,2n\}$ these variables are supposed to store the weight of the optimal independent set if we consider only vertices whose intervals intersect $\{1,\ldots,j\}$. For $j>2n$ these variables will represent the weight of the optimal independent set restricted to GMG-M and the first j2nj-2n vertices of $M$.
  5. 5.
    For every edge $uv\in E$ such that $u,v\in M$ we construct the clause $(\neg x_{u}\lor\neg x_{v})$.
  6. 6.
    For every $i\in\{1,\ldots,2n\}$ and every vertex $u\not\in M$ such that $i$ does not belong in the interval associated with $u$ we construct a clause which initially contains only $\neg a_{i}$. We add to this clause literals using the variables $y_{i,1},\ldots,y_{i,\log n}$ which are falsified if and only if these variables encode the index of $u$.
  7. 7.
    For every edge $uv\in E$ such that $u\in M$ and $v\not\in M$, for each $i\in\{1,\ldots,2n\}$ such that $i$ belongs in the interval associated with $v$, we add the clause $(\neg a_{i}\lor\neg x_{u})$; we then add to this clause the variables $y_{i,1},\ldots,y_{i,\log n}$ properly negated so that these literals are all false if and only if their assignment corresponds to the binary representation of the index of $v$.
  8. 8.
    For every $i\in\{1,\ldots,2n-1\}$ and for any two distinct vertices $u,v\not\in M$ such that $i$ and $i+1$ both belong in both intervals associated with $u,v$ we construct a clause which initially contains only $(\neg a_{i}\lor\neg a_{i+1})$. We then add to this clause literals using the variables $y_{i,1},\ldots,y_{i,\log n}$ which are falsified if these variables encode the index of $u$ and literals using the variables $y_{i+1,1},\ldots,y_{i+1,\log n}$ which are falsified if these variables encode the index of $v$.
  9. 9.
    For every $i\in\{1,\ldots,2n-1\}$ and for any vertex $u\not\in M$ such that $i$ and $i+1$ both belong in the interval associated with $u$ we construct a clause which initially contains $(\neg a_{i}\lor a_{i+1})$. We then add to this clause literals using the variables $y_{i,1},\ldots,y_{i,\log n}$ which are falsified if these variables encode the index of $u$.
  10. 10.
    For every $i\in\{1,\ldots,2n\}$ which is the starting endpoint of the interval associated with a vertex $u\in V\setminus M$ we do the following: Construct (but do not yet add to the formula) a base clause which initially contains $\neg a_{i}$ and add to this clause literals using $y_{i,1},\ldots,y_{i,\log n}$ which are falsified if and only if these variables encode the index of $u$. Now, for each assignment to the variables $z_{i,1},\ldots,z_{i,\log n}$ and $z_{i-1,1},\ldots,z_{i-1,\log n}$ such that the difference of the binary numbers encoded by these two sets of variables is not equal to the weight of $u$ we construct a copy of the base clause and add literals falsified by these assignments. (For $i=1$ we assume that $z_{i-1,1},\ldots,z_{i-1,\log n}$ are all set to False, that is, they encode the number 0.) In this way we add at most $n^{2}$ clauses for each $i,u$.
  11. 11.
    For every $i\in\{1,\ldots,2n-1\}$ and for each $j\in\{1,\ldots,\log n\}$ we add clauses ensuring that $z_{i,j}=z_{i+1,j}$. If $i$ is a starting endpoint of an interval we add to all these clauses the literal $a_{i}$.
  12. 12.
    For every $i\in\{1,\ldots,m-1\}$ and for each $j\in\{1,\ldots,\log n\}$ add a clause ensuring that $z_{2n+i,j}=z_{2n+i+1,j}$. Add the literal $x_{i}$ to each such clause.
  13. 13.
    For every $i\in\{1,\ldots,m-1\}$ and for each assignment to the variables $z_{2n+i,1},\ldots,z_{2n+i,\log n}$ and $z_{2n+i-1,1},\ldots,z_{2n+i-1,\log n}$ such that the difference of the binary numbers encoded by these two sets of variables is not equal to the weight of the $i$-th vertex of $M$, add a clause that contains $\neg x_{i}$ and is falsified exactly by this combination of assignments to the $z$ variables.
  14. 14.
    Finally, for each integer strictly smaller than the target weight $T$ add a clause that is falsified if $z_{2n+m,1},\ldots,z_{2n+m,\log n}$ encodes this integer.
This completes the construction, which can clearly be carried out in polynomial time. Furthermore, it is not hard to produce a path decomposition of $\phi-M^{\prime}$ by constructing 2n+m12n+m-1 bags and including in the $i$-th bag for all $j\in\{1,\ldots,\log n\}$ the variables $a_{i},a_{i+1},y_{i,j},y_{i+1,j},z_{i,j},z_{i+1,j}$ (if they exist). This decomposition has width $O(\log n)$.
What remains then is to argue for correctness. If $G$ has an independent set $S$ of the desired weight we give assignments to the variables corresponding to their informal meanings: set $x_{v}$ to True if and only if $v\in M\cap S$; set $a_{i}$ to True if and only if $S$ contains a vertex $u\not\in M$ whose interval representation contains $i$ and in this case set $y_{i,1},\ldots,y_{1,\log n}$ to encode the index of $u$; set $z_{i,1},\ldots,z_{i,\log n}$ to encode the weight of $S$ restricted to the vertices of $V\setminus M$ whose intervals intersect $\{1,\ldots,i\}$ if $i\leq 2n$ or to the weight of $S$ restricted to $V\setminus M$ and the first i2ni-2n vertices of $M$ if $i>2n$. It is not hard to see that this assignment satisfies all the clauses.
For the converse direction, we extract a set $S\subseteq V$ of vertices from a satisfying assignment as follows: for the vertices $v\in M$ we place $v$ in $S$ if and only if the assignment sets $x_{v}$ to True; for $v\not\in M$ we place $v$ in $S$ if and only if there exists $i\in\{1,\ldots,2n\}$ which belongs in the interval of $v$ such that $a_{i}$ is True and $y_{i,1},\ldots,y_{i,\log n}$ encodes the index of $v$. We claim that the set $S$ we construct in this way is independent. Indeed, it is easy to see it cannot contain the two endpoints of an edge contained in $M$, because of the clauses of step 5. For each $v\in S\setminus M$ there is an $i$ such that $a_{i}$ is True and $y_{i,1},\ldots,y_{i,\log n}$ encodes the index of $v$. For the sake of contradiction (to $S$ being independent) select a $v\in S\setminus M$ that minimizes $i$ and has a neighbor in $S$. Because of the clauses of step 6 we observe that the interval of $v$ must contain $i$. If the neighbor of $v$ in $S$ were also in $M$, this would falsify a clause added in step 7, so this cannot happen. Because of the clauses of step 9 we observe that if $i+1$ also belongs to the interval of $v$, then $a_{i+1}$ is set to True. In this case, because of the clauses of steps 8 and 6, $y_{i+1,1},\ldots,y_{i+1,\log n}$ must also encode the index of $v$. Hence, for all $i^{\prime}>i$ which belong in the interval of $v$, the assignment encodes the index of $v$. We conclude that $S$ must be independent.
Given that $S$ is independent, it is not hard to see that the $z_{i,j}$ variables must be set in a way that calculates its total weight, and because of the clauses of the last step we know that this weight must be at least $T$ for the assignment to be satisfying.
Let us now proceed to the more challenging direction, where we need to reduce SAT parameterized by a modulator to logarithmic pathwidth to Weighted Independent Set parameterized by a modulator to an interval graph. To represent the modulator we will again rely on the gadget of Proposition˜8.4.
The main part of the reduction will encode a path decomposition of the primal graph of a CNF formula into an interval graph whose size is exponential in the width of the decomposition. To simplify presentation, we will require the given decomposition to obey a couple of simple (and easy to attain) structural properties.
Definition 8.21.
A path decomposition is called uniformized if all bags have the same size and the symmetric difference between any two consecutive non-identical bags consists of exactly two vertices.
Proposition 8.22.
There exists a polynomial time algorithm which, given a path decomposition of a graph, produces a uniformized decomposition of the same width.
Proof 8.23.
Repeat the following exhaustively: if there exist two bags of different sizes, then find two consecutive bags $B,B^{\prime}$ of distinct sizes (which must exist) and add an element of the larger bag to the smaller one. In the end of this process all bags have the same size. Then, if two consecutive bags $B,B^{\prime}$ have symmetric difference consisting of more than two elements, insert between them a sequence of bags that exchanges the elements of $B\setminus B^{\prime}$ with elements of $B^{\prime}\setminus B$ one by one. In the end of this process the decomposition is uniformized.
Lemma 8.24.
There is an algorithm which takes as input an $n$-variable CNF formula $\phi$ and a width $p$ uniformized path decomposition of its primal graph. The algorithm runs in time $2^{O(p)}n^{O(1)}$. Its output consists of: (i) an instace of Weighted Independent Set on an interval graph $G=(V,E)$ with target value $T$ (ii) an annotation function which assigns to each vertex of $G$ a partial truth assignment to some variables of $\phi$. The output satisfies the following:
  1. 1.
    For all truth assignments to the variables of $\phi$ there exists an independent set in $G$ that has weight $T$ and consists exactly of the vertices whose annotation is consistent with this assignment.
  2. 2.
    All independent sets of $G$ which contain vertices whose annotations are inconsistent on some variable have weight strictly less than $T$.
  3. 3.
    All independent sets of $G$ have weight at most $T$.
  4. 4.
    There exists an injective mapping which, given as input a bag of the decomposition and a truth assignment to its variables, returns a vertex of $G$ annotated with this assignment.
  5. 5.
    All vertices are given weights in $\{1,2,3\}$.
Proof 8.25.
Suppose the bags of the given decomposition are numbered $B_{1},\ldots,B_{t}$. Suppose also, in order to ease notation, that each bag has exactly $p$ (rather than $p+1$) variables.
We process the bags one by one in this order and will produce groups of vertices (and their associated intervals) called “bundles”. Each bundle produced for $B_{i}$ will contain $2^{p}$ vertices, such that for each assignment to the variables of $B_{i}$ exactly one of the vertices of the bundle will be annotated with this assignment (this will allow us to satisfy requirement 4). The intervals of a bundle will all share a common point (and hence form a clique). We will ensure that all intervals of each bundle have distinct endpoints and define the rank of a vertex $v$ in a bundle as the number of vertices that have strictly smaller right endpoint than $v$ in this bundle (so the rank is an integer in $\{0,\ldots,2^{p}-1\}$).
The first bundle we construct is a special Start bundle, corresponding to the variables of $B_{1}$: for each $i\in\{0,\ldots,2^{p}-1\}$ we construct a vertex represented by the interval $i,2^{p}-1+i]$ and annotated by the binary representation of $i$ (seen as a truth assignment to the variables of $B_{i}$). These vertices have weight $2$. Furthermore, for each $i\in\{0,\ldots,2^{p}-1\}$, we construct i1i-1 singleton intervals $0,0],[1,1],\ldots,[i-1,i-1]$, each annotated with the assignment $i$ and each with weight $1$. We refer the reader to Figure˜3.
We have now constructed our first bundle. Before we proceed to the construction of the remaining bundles, let us explain the invariant that this construction intends to maintain. Recall that in an independent set of intervals there is a natural ordering of the vertices. We want to ensure that for each vertex $v$ of the $j$-th bundle, the maximum weight achievable by any independent set whose last vertex is $v$ should be exactly $2j+\textrm{rank}(v)$. Furthermore, this weight should only be achievable if and only if we select vertices whose annotation is consistent with a global truth assignment. This invariant holds for the Start bundle ($j=1$) as for each vertex $v$ the best independent set that has $v$ as the last vertex has weight $2+\textrm{rank}(v)$.
We now observe that, having constructed $j$ bundles, it is easy to add a Do Nothing bundle which maintains the invariant: for each $i\in\{0,\ldots,2^{p}-1\}$ construct a weight-$2$ interval that has rank $i$ in the new bundle, is annotated with the same truth assignment as the interval of rank $i$ in the previous bundle, and intersects the interval of rank $i+1$ of the previous bundle, but not the interval of rank $i$. Clearly, the only way to maintain the invariant here is to select an interval with the same annotation as in the previous bundle, so this simple construction allows us to propagate an assignment (we again refer to Figure˜3).
The Do Nothing bundle is actually not necessary for our construction, but it allows us to more easily explain the rest. Having now constructed a bundle for the first bag, we want to proceed to the next bag of the decomposition, which has replaced one of the variables of the current bag (say $x$) with a new one. We start with an easy case. Suppose that the intervals of the current bundle are annotated such that for all $i\in\{0,\ldots,2^{p-1}-1\}$ we have that the intervals with ranks $2i$ and $2i+1$ are annotated with assignments that agree on all variables except $x$. Intuitively, this roughly corresponds to a situtation where, if we read the assignments as binary numbers and order them by rank, $x$ is the least significant bit. In this case we construct a New Variable bundle, meant to allow us to replace $x$ with a new variable. This is done by starting with a Do Nothing bundle, and then for each $i\in\{0,\ldots,2^{p-1}-1\}$ cutting the interval of rank $2i$ in the new bundle into two non-intersecting intervals of weight $1$. The starting point of the second of the two intervals is the same as that of the interval of rank $2i+1$ in the new bundle (again see Figure˜3). Cutting up the interval offers two new possibilities: if we had previously selected the interval of rank $2i+1$, we can now select the interval of rank $2i$, but which has weight $1$ (so the invariant is satisfied); conversely, if we had selected the interval of rank $2i$, we can select the left part of the cut interval and also the new interval of rank $2i+1$ (so the rank and the weight are increased by $1$ and the invariant is satisfied). Observe, however, that any selection that produces annotations that disagree on a variable still leads to a sub-optimal solution.
What remains to explain is how to handle the general case, where the assignments are not conveniently ranked so that the variable we are forgetting is the only disgreement between the intervals of rank $2i$ and $2i+1$, for all $i$. For this we introduce a Flip bundle, which allows us to switch the ranks of two assignments which are consecutive in the ranking. If we manage to do this repeatedly while maintaining the invariant, we will be able to produce any permutation of the assignments within at most $(2^{p})^{2}$ steps, hence also an assignment satisfying the conditions needed for a New Variable bundle.
A Flip bundle is then constructed as follows: start with a Do Nothing bundle, but make the interval of rank $i$ end a little later and the interval of rank $i+1$ end a little sooner, so that the ranks of all other intervals remain unchanged but the ranks of these two are exchanged. Assign weight $1$ to the interval that was shortened and $3$ to the interval that was lengthened. It is not hard to see that the only way to maintain an optimal solution that satisfies the invariant is to make a selection that is consistent in the new bundle. In particular, a solution which previously had the interval of rank $i+1$ still cannot take the interval which (now) has rank $i+1$ in the new bundle, so it must take the interval that has rank $i$ and weight $1$ (so both weight and rank are decreased by $1$); conversely a solution that had the interval of rank $i$ must select the interval that now has rank $i+1$ in the new bundle (and weight $3$) as the interval that now has rank $i$ has weight $1$.
We are now ready to describe the full construction. After the Start bundle, for each new bag we introduce a sequence of Flip bundles, ensuring that we have a ranking of the assignments so that the variable to be forgotten is the only point of disagreement between assignments of rank $2i$ and $2i+1$, for all $i\in\{0,\ldots,2^{p-1}-1\}$. We then add a New Variable bundle, and continue in the same way. Finally, for the last bag we construct a bundle that is the reverse of the Start bundle, ensuring that a solution with rank $i$ can collect additional weight of $2^{p}-1-i$. We have constructed at most $N\leq 4^{p}t$ bundles and we set $T=2N+2^{p}-1$. We can now verify by induction, using the invariant we mentioned, that all the promised properties hold.
Figure 3: Example of the construction of Lemma˜8.24. Gray intervals have weight $1$, solid black intervals weight $2$, and empty intervals weight $3$. The intervals are annotated with the truth assignment of their line, for the active variables of their bundle ($x_{1},x_{2}$ for the first two bundles, $x_{1},x_{3}$ for the next two, and $x_{4},x_{3}$ for the last one). The optimal solution is supposed to collect weight 22 per bundle, plus an extra amount equal to the rank of the selected interval in the last bundle. The first new variable gadget allows us to switch between $00$ and $01$ (similarly $10$ and $11$) without cost, as we forget variable $x_{2}$. The flip gadget switches the ranks of $01$ and $10$, allowing us to use a new variable gadget to forget $x_{1}$.
Armed with the previous lemma, we are now ready to prove the second part of Theorem˜8.18:
Lemma 8.26.
The first statement of Theorem˜8.18 implies the second.
Proof 8.27.
Suppose there exists an algorithm which takes as input a vertex-weighted graph $G=(V,E)$ with unary weights and a modulator $M$ such that GMG-M is an interval graph, and computes a maximum weight independent set in time $2^{(1-\varepsilon)|M|}|V|^{O(1)}$. We will use this algorithm to refute the $\log\!\textrm{pw}$M-SETH.
We are given as input an $n$-variable formula $\phi$, which by Section˜A.2 may be assumed to be in 3-CNF (hence has total size $O(n^{3})$). We are also given a modulator $M^{\prime}$ and a path decomposition of the primal graph of $\phi-M^{\prime}$ of width $O(\log n)$. The goal is to decide satisfiability of $\phi$ in time $(2-\varepsilon^{\prime})^{|M^{\prime}|}n^{O(1)}$. In the remainder we will assume without loss of generality that all clauses of $\phi$ involve a variable outside $M^{\prime}$; if this is not the case we can branch on the satisfying assignments of the corresponding clause (which are at most $7$) and if we can do this repeatedly we obtain an algorithm faster than brute-force. So, the interesting case is when no such clauses exist.
We will construct a graph $G=(V,E)$, a target weight $T$, and a modulator $M$ satisfying the following:
  1. 1.
    $G$ has an independent set of weight $T$ if and only if $\phi$ is satisfiable.
  2. 2.
    $|M|\leq(1+\varepsilon)|M^{\prime}|+O(1)$.
  3. 3.
    $G$ and $M$ can be constructed in time $n^{O(1)}$.
If we achieve the above, we obtain the lemma, as executing the reduction and then the supposed algorithm for computing a maximum weight independent set would run in time at most $2^{(1-\varepsilon^{2})|M^{\prime}|}n^{O(1)}$.
Our construction consists of two parts: a part that encodes the modulator $M^{\prime}$ and a part that represents the remaining formula using Lemma˜8.24.
For the first part, we use the gadget of Proposition˜8.4 and this encodes the formula modulator $M^{\prime}$ into a graph modulator $M$ of the desired size. Let $T_{1}$ be the maximum independent set size in the graph constructed so far, as given by Proposition˜8.4.
For the second part of the construction we invoke Proposition˜8.22, so we can assume the given path decomposition of $\phi-M^{\prime}$ is uniformized, and then invoke Lemma˜8.24 to obtain an interval graph of size polynomial in $n$ and a target weight $T_{2}$. Recall also that all vertices of the interval graph are annotated with assignments to some variables of $\phi-M^{\prime}$ and that we have an injective mapping which, given a bag and an assignment to its variables returns a vertex annotated with this assignment. What remains now is to encode the clauses of $\phi$. Let $m$ be the number of clauses of $\phi$. Before we proceed, we multiply all the weights of the interval graph by $m+1$, making the target weight $T_{2}(m+1)$.
Assume without loss of generality that for each clause of $\phi$ there exists a distinct bag of the given decomposition that contains all the variables of the clause (except those of $M^{\prime}$). Such a bag exists, since these variables form a clique; we can make the mapping from clauses to bags injective by simply repeating bags in the decomposition as needed. Consider now a clause $c$ whose variables appear in a bag $B$. We have a mapping that gives us a vertex of the graph for each assignment to the variables of $B$. For each such assignment that satisfies $c$, we find the corresponding vertex and increase its weight by $1$. For each assignment which does not satisfy $c$, let $v$ be the interval returned by the mapping for this assignment. We will replace $v$ with several copies of the same interval, all of which will have weight equal to the weight of $v$ plus $1$. These intervals will be connected to $M$ in a way that ensures that they can be selected in the independent set only if the assignment we have encoded for $M^{\prime}$ satisfies $c$. In particular, $c$ contains at most $2$ variables from $M^{\prime}$, which belong in two (perhaps distinct) groups of size $\rho$. Consider all assignments to the $2\rho$ variables of these groups and recall that we have mapped each such assignment to a set of $\frac{\gamma-1}{2}$ vertices of each group. For each combination of assignments that satisfies $c$ we construct a copy of $v$ adjacent to the $\frac{\gamma+1}{2}$ vertices of each group which are not in the independent set when we have selected these assignments.
We now assign to all vertices of the first part weight $T_{2}(m+1)+m+1$ and set the total target weight to $T=T_{1}(T_{2}(m+1)+m+1)+T_{2}(m+1)+m$. What remains is to argue for correctness. For the easy direction: if $\phi$ is satisfiable, then we select vertices in each group of $M$ according to the mapping we have decided from truth assignments to sets of size $\frac{\gamma-1}{2}$ and complete the selection as guaranteed by Proposition˜8.4. This gives us a weight of $T_{1}(T_{2}(m+1)+m+1)$. We then move on to the second part and select all intervals annotated with an assignment consistent with our global truth assignment. According to Lemma˜8.24 this allows us to obtain a weight of $T_{2}(m+1)$. Observe, however, that in some cases we have constructed several copies of an existing interval $v$. In case these copies are annotated in a way consistent with the truth assignment, we know that the assignment satisfies the clause represented by these copies using variables of the modulator. Therefore, there must exist a copy that we can add to the solution. To see that we attain the desired total weight observe that for each clause of $\phi$ we have increased the weight of a selected interval by $1$.
For the converse direction, we have made vertices of the first part sufficiently expensive to ensure that $T_{1}$ of them must be selected and therefore that the selection encodes a truth assignment to the modulator. In particular, the total weight we can obtain in the second part is at most $T_{2}(m+1)+m$, because by Lemma˜8.24 all independent sets in the interval graph have initial weight at most $T_{2}$ (and we have multiplied these weights by $m+1$) and we have added $1$ more for each clause to each potential solution. Therefore, vertices of the first part are sufficiently expensive to ensure we have selected $T_{1}$ of them, and in a similar way vertices of the second part are sufficiently expensive to ensure we have selected $T_{2}$ of them, which according to Lemma˜8.24 ensures we can extract a global consistent truth assignment. Observe now that for each clause we can gain an extra weight of $1$, so to reach the target weight we must have an assignment that satisfies all clauses.
Appendix A Complexity Hypotheses
A.1 List of Hypotheses
We list here the complexity hypotheses that appear in this paper.
  • SAT-related
    1. 1.
      SETH: [30] For all $\varepsilon>0$ there exists $k$ such that there is no algorithm solving $k$-SAT for $n$-variable formulas $\phi$ in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$, where $n$ is the number of variables.
    2. 2.
      MaxSat-SETH: The same as the previous statement but for Max-SAT, that is, the problem of deciding if there is an assignment satisfying at least $t$ clauses, where $t$ is given in the input.
  • Circuits
    1. 3.
      Circuit-SETH: For all $\varepsilon>0$ there is no algorithm deciding if an $n$-input circuit of size $s$ is satisfiable in time $(2-\varepsilon)^{n}s^{O(1)}$.
    2. 4.
      LD-C-SETH: The same as the previous statement, but restricted to circuits of depth at most $\varepsilon n$.
  • Backdoors
    1. 5.
      2SatB-SETH: For all $\varepsilon>0$ there is no algorithm deciding if a CNF formula $\phi$ with a given strong 2-SAT backdoor of size $b$ is satisfiable in time $(2-\varepsilon)^{b}|\phi|^{O(1)}$.
    2. 6.
      HornB-SETH: The same as the previous statement, but for a strong Horn backdoor.
  • Modulators
    1. 7.
      tdM-SETH: For all $\varepsilon>0$ there exist $c,k$ such that there is no algorithm that takes as input a $k$-CNF formula $\phi$ and a modulator $M$ of size $m$ such that $\phi-M$ has tree-depth at most $c$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
    2. 8.
      pwM-SETH: For all $\varepsilon>0$ there exists $c$ such that there is no algorithm that takes as input a CNF formula $\phi$ and a modulator $M$ of size $m$ such that $\phi-M$ has pathwidth at most $c$ and decides if $\phi$ is satisfiable in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
    3. 9.
      twM-SETH: The same as the previous statement but for treewidth at most $c$.
    4. 10.
      $\log\!\textrm{td}$M-SETH: The same as the previous statement but for tree-depth at most $c\log|\phi|$.
    5. 11.
      $\log\!\textrm{pw}$M-SETH: The same as the previous statement but for pathwidth at most $c\log|\phi|$.
    6. 12.
      $(\sigma,\delta)H$-MaxSat-SETH: For all $\varepsilon>0$ there exist $k,\sigma,\delta$ such that there is no algorithm which takes as input a $k$-CNF $\phi$ and a $(\sigma,\delta)$-hub $M$ of size $m$ and computes an assignment satisfying the maximum number of clauses in time $(2-\varepsilon)^{m}|\phi|^{O(1)}$.
    7. 13.
      $(\sigma,\delta)H$-SETH: The same as the previous statement but for deciding if the input formula is satisfiable.
  • Weighted SAT
    1. 14.
      W[SAT]-SETH: For all $\varepsilon>0$ there is no algorithm which takes as input a (general) Boolean formula $\phi$ of size $s$ on $n$ variables and an integer $k$ and decides if $\phi$ is satisfiable by an assignment that sets exactly $k$ variables to True in time $n^{(1-\varepsilon)k}s^{O(1)}$.
    2. 15.
      W[P]-SETH (weak version): Same as the previous statement but for Boolean circuits.
    3. 16.
      W[P]-SETH (strong version): For all $\varepsilon>0,k_{0}>0$, there is no algorithm which for all $k>k_{0}$ takes as input a Boolean circuit of size $s$ and decides if the circuit has a weight-$k$ satisfying assignment in time $s^{(1-\varepsilon)k}$.
  • Others
    1. 17.
      Max-3-SAT-SETH: The same as MaxSat-SETH but for CNFs of maximum arity $3$.
    2. 18.
      NC-SETH: The same as LD-C-SETH but for circuits of depth at most $\log^{O(1)}n$.
    3. 19.
      pw-SETH: For all $\varepsilon>0$ there is no algorithm that solves 3-SAT in time $(2-\varepsilon)^{\textrm{pw}}n^{O(1)}$ for formulas on $n$ variables of primal pathwidth pw.
A.2 Relations between Classes of Hypotheses
Many of the relations between the hypotheses we consider can be inferred directly from the definitions or from the relations between graph widths. In particular, we have $\textrm{tw}\leq\textrm{pw}\leq\textrm{td}\leq\textrm{tw}\log n$ for any graph on $n$ vertices. This gives $1\Rightarrow 7\Rightarrow 8\Rightarrow 9\Rightarrow 10\Rightarrow 11\Rightarrow 13\Rightarrow 12$. Furthermore, it is clear that deciding if a general circuit is satisfiable is at least as hard as a circuit with depth $\varepsilon n$, which is at least as hard as a CNF formula, so $1\Rightarrow 4\Rightarrow 3$. We have $14\Rightarrow 15$ as formulas are a special case of circuits. Also $2\Rightarrow 12$ as the set of all variables is a $(\sigma,\delta)$-hub.
It is worth noting that the arity $k$ appears explicitly in statement 7, but not in statements 8 to 11. This is because of the following observation, which allows us to decrease the arity once the parameter we are dealing with allows the construction of long paths.
{observation}
There is a polynomial time algorithm which takes as input a CNF formula $\phi$ and a modulator $M$ and produces an equisatisfiable 3-CNF formula $\psi$ by adding variables to $\phi$ such that $M$ remains a modulator of $\psi$ and we have $\textrm{pw}(\psi-M)\leq\textrm{pw}(\phi-M)+2$, $\textrm{tw}(\psi-M)\leq\textrm{tw}(\phi-M)+2$, $\textrm{td}(\psi-M)\leq\textrm{td}(\phi-M)+O(\log|\phi|)$.
Proof A.1.
We first state an easy claim.
Claim 5.
Let $G=(V,E)$ be a graph and $\mathcal{K}\subseteq 2^{V}$ be a set of cliques of $G$. Let $G^{\prime}$ be a graph obtained from $G$ by adding for each $K\in\mathcal{K}$ a path going through at most $\ell$ new vertices and connecting some of the vertices of the path to vertices of $K$. Then $\textrm{pw}(G^{\prime})\leq\textrm{pw}(G)+2$, $\textrm{tw}(G^{\prime})\leq\textrm{tw}(G)+2$, and $\textrm{td}(G^{\prime})\leq\textrm{td}(G)+\log(\ell+1)$.
{claimproof}
Assume we start with a tree or path decomposition of $G$ or a rooted tree proving the tree-depth of $G$. For path and tree decompositions we know that there is a bag $B$ that contains the clique $K$. Take a path decomposition of the path that we connected to $K$ and add to all bags the contents of bag $B$ (the width has now been increased by $2$). In the case of treewidth, attach one bag of this decomposition to $B$; in the case of pathwidth insert this decomposition after $B$. For tree-depth, let $k\in K$ be the vertex that is furthest from the root. We observe that all vertices of $K\setminus\{k\}$ must be ancestors of $k$. We insert below $k$ a tree of depth at most $\log(\ell+1)$ containing the path we connected to $k$. Here we are using the fact that a path of $\ell$ vertices has tree-depth at most $\log(\ell+1)$ (which can easily be seen by induction).
We apply the following standard trick exhaustively: as long as $\phi$ contains a clause of arity $r>3$, say $C=(\ell_{1}\lor\ell_{2}\lor\ldots\lor\ell_{r})$, introduce r3r-3 new variables $z_{1},\ldots,z_{r-3}$ and replace $C$ with the clauses $(\ell_{1}\lor\ell_{2}\lor z_{1})\land(\neg z_{1}\lor\ell_{3}\lor z_{2})\land(\neg z_{2}\lor\ell_{4}\lor z_{3})\land\ldots\land(\neg z_{r-3}\lor\ell_{r-1}\lor\ell_{r})$. Observe now that the variables of $C$ form a clique in the primal graph, while the new variables induce a path whose neighborhood is contained in this clique. We can therefore apply the claim, with $\mathcal{K}$ being the set of all clauses on which we applied this transformation.
We note here two further easy observations justifying the organization of hypotheses we have made in Figure˜1.
Theorem A.2.
(MaxSat-SETH $\Rightarrow$ LD-C-SETH) If there is an $\varepsilon>0$ and an algorithm that decides satisfiability of $n$-input circuits of size $s$ and depth at most $\varepsilon n$ in time $(2-\varepsilon)^{n}s^{O(1)}$, then there is an $\varepsilon^{\prime}>0$ and an algorithm that given as input an $n$-variable CNF formula $\phi$ finds the assignment satisfying the maximum number of clauses in time $(2-\varepsilon^{\prime})^{n}|\phi|^{O(1)}$.
Proof A.3.
Construct a circuit with $n$ inputs and a gate for each clause appropriately connected to the inputs through negations. This circuit can be made to have fan-in $2$ and depth $\log n$. For each target $t$ we now want to add some layers to the circuit so that it is satisfiable if and only if we can satisfy at least $t$ clauses of the original instance. Running the assumed algorithm for each $t$ allows us to compute the maximum number of satisfied clauses, while self-reducibility allows us to find an assignment that satisfies this many clauses.
[…]
Theorem A.4.
(2SatB-SETH $\Rightarrow$ Circuit-SETH) If there is an $\varepsilon>0$ and an algorithm that decides satisfiability of $n$-input circuits of size $s$ in time $(2-\varepsilon)^{n}s^{O(1)}$, then there is an $\varepsilon^{\prime}>0$ and an algorithm that given as input a CNF formula $\phi$ and a strong 2-SAT backdoor of size $b$ decides if $\phi$ is satisfiable in time $(2-\varepsilon^{\prime})^{b}|\phi|^{O(1)}$.
Proof A.5.
The idea is to construct a circuit implementing the standard polynomial-time algorithm for 2-SAT. Recall that one way to decide satisfiability for 2-SAT is to run resolution exhaustively, where resolution is the rule that, given two clauses $C,C^{\prime}$ that contain two opposite literals $x,\neg x$, adds to the formula a new clause containing all the other literals of $C,C^{\prime}$. It is known that a 2-CNF formula is satisfiable if and only if performing this exhaustively does not produce the empty formula.
We are given a CNF $\phi$ and a backdoor $B$ of size $b$. Let $x_{1},\ldots,x_{n}$ be the non-backdoor variables of $\phi$. We construct a circuit with $b$ inputs. Observe that there are at most $4n^{2}$ clauses we can construct with the variables $x_{1},\ldots,x_{n}$. For each $t\in[0,4n^{2}]$ and clause $C$ that can be constructed from these variables we will have in our circuit a gate $x_{t,C}$ whose meaning is that the algorithm has inferred $C$ after $t$ steps. The gates $x_{0,C}$ can be computed from the inputs: $x_{0,C}$ is the disjunction over all clauses of $\phi$ that contain $C$ of the conjunction of the negation of all their other literals. It is now easy to compute $x_{i+1,C}$ from layer $x_{i,C}$ and the output of the circuit is $x_{4n^{2},\emptyset}$. The size of the circuit is polynomial in $\phi$.